### v1

# Transversal Hamilton cycles in digraph collections

Abstract
Given a collection $\mathcal{D}=\{D_{1},D_{2},\ldots,D_{m}\}$ of digraphs on the common vertex set $V$, an $m$-edge digraph $H$ with vertices in $V$ is transversal in $\mathcal{D}$ if there exists a bijection $\varphi:E(H)\rightarrow[m]$ such that $e\in E(D_{\varphi(e)})$ for all $e\in E(H)$. Ghouila-Houri proved that any $n$-vertex digraph with minimum semi-degree at least $\frac{n}{2}$ contains a directed Hamilton cycle. In this paper, we provide a transversal generalization of Ghouila-Houri’s theorem, thereby solving a problem proposed by Chakraborti, Kim, Lee and Seo [6]. Our proof utilizes the absorption method for transversals, the regularity method for digraph collections, as well as the transversal blow-up lemma [9] and the related machinery. As an application, when $n$ is sufficiently large, our result implies the transversal version of Dirac’s theorem, which was proved by Joos and Kim [21].
1 Introduction
A transversal $X$ over a collection =$\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects represents an object that intersects every $F_{i}$. Here, the object can be any mathematical object, such as sets, spaces, set systems, matroids, and so forth. Several classical theorems, which guarantee the existence of certain objects, have been restated in the context of transversals, illustrating the feasibility or infeasibility of obtaining such an object as a transversal over a certain collection. To name a few, transversals are extensively studied for Carathéodory’s theorem [4], Helly’s theorem [22], Erdős-Ko-Rado theorem [2], Rota’s basis conjecture [20, 24], and others.
Following this trend, there has been ample research studying transversal generalizations of classical results in graph theory. These generalizations are not only interesting in their own right but also often provide a strengthening of the original results, reflecting the robustness of graph properties to some extent. Transversals over a collection of graphs have been implicitly used in the literature under the name of rainbow coloring and explicitly defined in [21] as follows.
Definition 1.1.
For a given collection $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/hypergraphs/digraphs with the same vertex set $V$, an $m$-edge graph/hypergraph/digraph $H$ with vertices in $V$ is transversal in $\mathcal{G}$ if there exists a bijection $\phi:E(H)\to[m]$ such that $e\in E(G_{\phi(e)})$ for all $e\in E(H).$
By interpreting each $G_{i}$ as the set of edges colored with color $i$, the function $\phi$ is often referred to as a coloring. We say that the coloring $\phi$ is rainbow if it is injective, and $H$ is a rainbow subgraph if there exists a rainbow coloring on it. Many classical results in extremal graph theory have been extended to such transversal settings, exhibiting interesting phenomena. The central question in studying transversal generalizations can be stated as follows.
Problem A ([21]).
Let $H$ be a graph/hypergraph/digraph with $m$ edges, $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$ be a collection of graphs/hypergraphs/digraphs on a common vertes set of size $n$. Which property $\mathcal{P}$ imposed on $\mathcal{G}$ will guarantee the existence of a transversal copy of $H$?
Since every object in the family $\mathcal{G}$ can be the same, a necessary condition for a positive answer to the above question is that every $n$-vertex graph/hypergraph/directed graph satisfying the property $\mathcal{P}$ contains a copy of $H$. However, this alone is not always sufficient. For example, Aharoni, DeVos, de la Maza, Montejano, and Šámal [1] proved that if $\mathcal{G}=\{G_{1},G_{2},G_{3}\}$ is a collection of graphs on a common vertex set of size $n$ and $|E(G_{i})|>(\frac{26-2\sqrt{7}}{81})n^{2}$ for $i\in[3]$, then $\mathcal{G}$ contains a rainbow triangle. Moreover, the constant $\frac{26-2\sqrt{7}}{81}$ is optimal, while Mantel’s theorem states that any $n$-vertex graph with more than $\lfloor\frac{n^{2}}{4}\rfloor$ edges must contain a triangle.
In 2020, Aharoni [1] conjectured that Dirac’s theorem can be extended to a transversal version. Cheng, Wang and Zhao [12] solved this conjecture asymptotically, and it was completely confirmed by Joos and Kim [21].
Theorem 1.2 ([21]).
Suppose $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of graphs with the same vertex set of size $n$. If $\delta(G_{i})\geq\frac{n}{2}$ for all $i\in[n]$, then $\mathcal{G}$ contains a transversal Hamilton cycle.
There are numerous established results on transversal problems in graph collections. For more results on transversals, we refer the reader to the survey [26].
A natural next step is to study rainbow analogues in digraph collections. Cheng, Han, Wang and Wang [8] studied the transversal version of Hajnal-Szemerédi theorem in digraph collections, which asymptotically generalized the corresponding results by Czygrinow-DeBiasio-Kierstead-Molla [15] and Treglown [30]. It is well-known that every tournament contains a Hamilton path, and every strongly connected tournament contains a Hamilton cycle. Chakraborti, Kim, Lee and Seo [6] proved the existence of a transversal directed Hamilton path and cycle in a family of tournaments. They [7] also proved that in a digraph collection consisting of tournaments, there exist transversal Hamilton cycles of all possible orientations possibly except the consistently oriented one. These results generalized the classical theorem of Thomason [28]. Babiński, Grzesik and Prorok [3] studied the transversal Mantel’s theorem in digraph and oriented graph collections. Gerbner, Grzesik, Palmer and Prorok [18] determined the minimum number of edges in each color which guarantees the existence any rainbow directed star.
1.1 Main result
In this paper, we study the transversal problem in digraph collections. Let $D$ be a digraph with no loops and at most one edge in each direction between every pair of vertices. Denote the set of vertices of $D$ by $V(D)$. Let $\delta^{+}(D)$ and $\delta^{-}(D)$ be the minimum out-degree and minimum in-degree of $D$, respectively. Let $\delta^{0}(D)=\min\{\delta^{+}(D),{\delta^{-}}(D)\}$ be the minimum semi-degree of $D.$ Let $\mathcal{D}=\{D_{1},D_{2},\dots,D_{m}\}$ be a collection of not necessarily distinct digraphs with common vertex set $V$. Define $\delta^{0}(\mathcal{D}):=\min\{\delta^{0}(D_{i}):i\in[m]\}$.
In 1960, Ghouila-Houri [19] generalized the Dirac’s theorem to digraphs.
Theorem 1.3 ([19]).
Every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains a directed Hamilton cycle.
Chakraborti, Kim, Lee, Seo [6] suggested to consider a transversal version of Ghouila-Houri’s theorem would be interesting. Our main result solves this problem.
Theorem 1.4.
Let $n$ be a sufficiently large integer. Suppose $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set of size $n$. If $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
In an undirected graph, one may replace each undirected edge by two arcs with opposite directions. Thus, for sufficiently large $n$, Theorem 1.2 is a direct consequence of Theorem 1.4. Furthermore, by letting $D_{1}=D_{2}=\cdots=D_{n}$ in Theorem 1.4, we get Theorem 1.3 immediately when $n$ is large.
1.2 Notation and Organization
Notation. Throughout the paper, we use standard graph theory notation and terminology. Let $D$ be a digraph with vertex set $V(D)$ and edge set $E(D)$, and sometimes we write $e(D):=|E(D)|$. For $x,y\in V(D)$, denote by $xy$ the directed edge from $x$ to $y$ and we always write $xy\in D$. Denote by $P_{n}$ and $C_{n}$ the directed path and directed cycle on $n$ vertices, respectively.
We use $H\subseteq D$ to denote that $H$ is a sub-digraph of $D$. For $S\subseteq V(D)$, we write $D[S]$ for the sub-digraph of $D$ induced by $S$ and $D-S$ for $D[V(D)\setminus S]$. For two not necessarily disjoint vertex subsets $A,B\subseteq V(D)$, let $D[A,B]$ be the sub-digraph of $D$ with vertex set $A\cup B$ and edge set $\{xy\in E(D):x\in A,y\in B\}$ and let $e_{D}(A,B):=|E(D[A,B])|$. Denote $D^{\pm}[A,B]=D[A,B]\cup D[B,A]$. When $A=B$, we often abbreviate $D[A,B]$ and $e_{D}(A,B)$ as $D[A]$ and $e_{D}(A)$ respectively.
For a digraph $D$ and $v\in V(D)$, we denote the set of out-neighbours and in-neighbours of $v$ by $N_{D}^{+}(v)$ and $N_{D}^{-}(v)$, respectively, and we write $d_{D}^{+}(v):=|N_{D}^{+}(v)|$ and $d_{D}^{-}(v):=|N_{D}^{-}(v)|$ for the out- and indegree of $v$ in $D$, respectively. For $S\subseteq V(D)$, we write $d_{D}^{+}(v,S):=|N_{D}^{+}(v)\cap S|$ and $d_{D}^{-}(v,S):=|N_{D}^{-}(v)\cap S|$. We often omit subscripts when they are clear from the context.
Let $\mathcal{D}=\{D_{i}:i\in[m]\}$ be a collection of digraphs on a common vertex set $V$ and $H$ be a rainbow digraph inside $\mathcal{D}$. For a vertex $v\in V$, let $N_{i}^{+}(v):=N_{D_{i}}^{+}(v)$, $N_{i}^{-}(v):=N_{D_{i}}^{-}(v)$, $d_{i}^{+}(v):=N_{i}^{+}(v)$ and $d_{i}^{-}(v):=N_{i}^{-}(v)$. For any two vertex subsets $X,\,Y\subseteq V$, define $\mathcal{D}[X]:=\{D_{i}[X]:i\in[m]\}$, $\mathcal{D}[X,Y]:=\{D_{i}[X,Y]:i\in[m]\}$ and $\mathcal{D}^{\pm}[X,Y]:=\{D_{i}^{\pm}[X,Y]:i\in[m]\}$. If $E(D)=\emptyset$, then we simply write $D=\emptyset$. Similarly, if there is no edge in $D_{i}$ for all $i\in[m]$, then we write $\mathcal{D}=\emptyset$. Denote by ${\rm col}(H)$ the set of colors appearing in $H$. Two rainbow graphs $H_{1}$ and $H_{2}$ are said to be disjoint if $V(H_{1})\cap V(H_{2})=\emptyset$ and ${\rm col}(H_{1})\cap{\rm col}(H_{2})=\emptyset$.
The notation above extends to undirected graphs in the obvious ways. For $n\in\mathbb{N}$, we denote the set $\{1,2,\ldots,n\}$ by $[n]$. For any two constants $\alpha,\beta\in(0,1)$, we write $\alpha\ll\beta$ if there exists a function $f=f(\beta)$ such that the subsequent arguments hold for all $0<\alpha\leq f.$
Organization. We conclude this section with a sketch of the proof of Theorem 1.4. In section 2, we introduce two key techniques that we will need later, i.e., the regularity lemma in digraph collections and a stability result for transversal perfect matchings in bipartite graph collections. For readability, we postpone its proofs to the appendix. The proof of Theorem 1.4 will be divided into “stable case” and “extremal case”. Section 3 introduces notions of extremality and stability for digraphs and digraph collections; we prove several results about absorption that will be used in the proof of the stable case. In Section 5, we show that there is always a transversal directed Hamilton cycle in stable case. Section 6 deals with the extremal case. We finish our paper with some concluding remarks in the final section.
1.3 Discussion of the strategy
Joos and Kim [21] proved the transversal version of Dirac’s theorem by introducing the auxiliary digraph technique, along with some rotations. However, this technique is not feasible in digraph collections because we cannot extend an existing directed cycle into a longer cycle by using the same rotation. In [10], the first author and Staden further developed the absorption method for graph collections, combined with the application of the transversal blow-up lemma, they finally proved a stability version of Joos and Kim’s result [21]. The absorption technique used in [10] originates from [12] and [8], based on the common absorption technique introduced first in [25], which was showed to be extremely powerful in the solutions of various transversal graph embedding problems. In this paper, we adapt the method developed in [10] and generalize their stability from graph collections to digraph collections. By analyzing the extremal cases, we proved a tight bound for the transversal Ghouila-Houri’s theorem. We believe our method could be used to demonstrate stabilities or tight bounds for other transversal graph embedding problem (see the survey [26]).
1.4 Sketch of the proof of Theorem 1.4
Our proof utilizes the regularity blow-up method in digraph collections, and is divided into two parts: stable case and the extremal case.
Stable case. We first characterize the extremality of a single digraph, which forms the basis of our extremal case distinction. Let $0<\frac{1}{n}\ll\epsilon\ll 1$ and $D$ be an $n$-vertex digraph with minimum semi-degree at least $(\frac{1}{2}-\epsilon)n$. If $D$ contains no directed Hamilton cycle, then $D$ contains a sub-digraph that is close to one of the digraphs in Figure 1. Based on Lemma 3.3, there exists a characteristic partition of $D$ in each case, and we say that $D$ is $(\epsilon,{\rm EC}k)$-extremal for some $k\in[3]$. Moreover, the digraph $D$ can be $(\epsilon,{\rm EC}k)$-extremal for exactly one $k$.
[…]
  • $\mathcal{D}$ is strongly stable if $\mathcal{D}$ contains many digraphs each of which is not $(\epsilon,{\rm EC}k)$-extremal for any $k\in[3]$,
  • $\mathcal{D}$ is weakly stable if for almost all colors $i\in[n]$, each $D_{i}$ is $(\epsilon,{\rm EC}k)$-extremal for some $k\in[3]$, but their vertex partitions are not similar,
  • $\mathcal{D}$ is stable if it is either strongly stable or weakly stable.
In the stable case, we are to show that if $\mathcal{D}$ is stable, then it contains a transversal directed Hamilton cycle.
Step 1. Build an absorbing cycle when $\mathcal{D}$ is stable. (Section 3)
We build an “directed absorbing rainbow cycle” $C$ for $\mathcal{D}$ with the property that $C$ is very small and there is a set $A$ consisting of some color-vertex-vertex triples $(c,u,v)$ such that whenever $A_{0}\subseteq A$ is sufficiently small compared to $|C|$, $C$ can absorb all of its elements. Additionally, we construct an additional auxiliary set $\mathcal{Q}$ consisting of color-color-vertex-vertex tuples $(c,c^{\prime},v,v^{\prime})$ to absorb those colors and vertices which $C$ cannot. Delete all colors and vertices in $C$ and $\mathcal{Q}$ from $\mathcal{D}$.
Step 2. Use the regularity-blow-up method for transversals in digraph collections to cover almost all vertices and colors with long directed paths.
We first prove the regularity lemma for digraph collections (Section 2.1), and then apply it to $\mathcal{D}$, yielding a reduced digraph collection $\mathcal{R}$ that inherits the minimum degree condition and stability of $\mathcal{D}$. By considering the characteristic bipartite graph of each digraph, we get a bipartite graph collection $\mathcal{B_{R}}$ for the reduced digraph collection $\mathcal{R}$. Notice that $\mathcal{R}$ has a disjoint rainbow directed cycle partition if and only if $\mathcal{B_{R}}$ has a transversal perfect matching. We establish a stability result for transversal perfect matching in bipartite graph collections (Section 2.2), ensuring that the disjoint rainbow directed cycle partition of $\mathcal{R}$ exists. Using the blow-up method, we obtain a set of almost spanning disjoint rainbow directed paths within each directed cycle of $\mathcal{R}$, which covers almost all vertices outside $C$ and $\mathcal{Q}$ (Section 4).
Step 3. Connect the obtained directed paths and cover remaining vertices via the absorbing cycle.
In this step, we apply the absorbing property of $C$ as well as the auxiliary set $\mathcal{Q}$ to connect all the directed paths and cover the remaining vertices, ultimately form a transversal directed Hamilton cycle.
[…]
Firstly, expand $A_{1}\cup B_{1}$ into an equitable partition $A\cup B$ of $V$. For a vertex $v\in A$, there are three bad cases (similar for vertices in $B$): $v$ lies in almost all $B_{i}$ for $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$; $v$ lies in almost all $L_{i}$ for $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$; $v\in A_{i}$ for many $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and $v\in B_{j}$ for many $j\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$. Move vertices in the first case from $A$ to $B$, and do the same operation for such vertices in $B$. Denote the set of vertices in the other two bad cases by $V_{\rm bad}$. After this, the number of vertices in $A$ and $B$ may be unbalanced, but their difference is still small. Without loss of generality, assume that $|A|\leq|B|$. In order to apply the transversal blow-up lemma to find long rainbow directed paths inside $\mathcal{D}$, the following four steps are required.
Step 1. Balance the number of vertices in $A$ and $B$.
If there exists a set of disjoint rainbow paths in $\mathcal{D}[B]$ such that after deleting their vertices, the number of vertices in $A$ and $B$ is balanced, then we are done. Otherwise, we will show that $\mathcal{D}$ must contains a transversal directed Hamilton cycle (see Lemma 5.14).
Step 2. Cover vertices in $V_{\rm bad}$.
Using colors in $\mathcal{C}_{2}$, we are to find a series of disjoint rainbow directed paths $P_{3}$ such that their centers are all vertices in $V_{\rm bad}$ and endpoints are unused vertices in $(A\cup B)\setminus V_{\rm bad}$.
Step 3. Deal with colors in $\mathcal{C}_{\rm bad}\cup\mathcal{C}_{3}$.
Choose a maximal rainbow matching $M$ using colors in $\mathcal{C}_{\rm bad}$ and avoiding vertices used in Steps 1-2. It is routine to verify that for each $j\in(\mathcal{C}_{\rm bad}\cup\mathcal{C}_{3})\setminus{\rm col}(M)$, the graph $D_{j}$ is EC1-extremal.
Step 4. Deal with colors in $(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}\cup\mathcal{C}_{3})\setminus{\rm col}(M)$.
We greedily select two rainbow directed paths using colors from $(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}\cup\mathcal{C}_{3})\setminus{\rm col}(M)$ while avoiding the vertices used in the previous three steps.
Based on the minimum degree condition and the characteristic partition of extremal digraphs, one may use colors in $\mathcal{C}_{2}$ to connect all rainbow paths obtained in the above four steps into a single short rainbow directed path, say $P$, such that $V_{\rm bad}\subseteq V(P)$, $\mathcal{C}_{1}\cup\mathcal{C}_{3}\cup\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|A\setminus V(P)|$ and $|B\setminus V(P)|$ are nearly equal. By applying the transversal blow-up lemma and some structural analysis, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle.
2 Preliminaries
2.1 Regularity for digraph collections
We use the following version of the regularity lemma for digraph collections, which is obtained by applying the degree version of the weak regularity lemma to the 4-graph of $\mathcal{D}$. The proofs of Lemma 2.2 and Lemma 2.4 are postponed to the appendix. We begin by defining regularity for digraph collections.
Definition 2.1.
Suppose that $\mathcal{D}=\{D_{c}:c\in\mathcal{C}\}$ is a collection of bipartite digraphs on a common vertex partition $V_{1}\cup V_{2}$. We say that $\mathcal{D}$ is $(\epsilon,d)$-regular if whenever $V_{i}^{\prime}\subseteq V_{i}$ with $|V_{i}^{\prime}|\geq\epsilon|V_{i}|$ for $i\in[2]$ and $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ with $|\mathcal{C}^{\prime}|\geq\epsilon|\mathcal{C}|$ we have
$\left|\frac{\sum_{c\in\mathcal{C}^{\prime}}e_{D_{c}}(V_{1}^{\prime},V_{2}^{\prime})}{|\mathcal{C}^{\prime}||V_{1}^{\prime}||V_{2}^{\prime}|}-\frac{\sum_{c\in\mathcal{C}}e_{D_{c}}(V_{1},V_{2})}{|\mathcal{C}||V_{1}||V_{2}|}\right|<\epsilon$
and $\sum_{c\in\mathcal{C}}e_{D_{c}}(V_{1},V_{2})\geq d|\mathcal{C}||V_{1}||V_{2}|$.
Lemma 2.2 (Regularity lemma for digraph collections).
For every integer $L_{0}\geq 1$ and every $\epsilon,\delta>0$, there is an $n_{0}=n_{0}(\epsilon,\delta,L_{0})$ such that for every $d\in[0,1)$ and every digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on vertex set $V$ of size $n\geq n_{0}$ with $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$, there exists a partition of $V$ into $V_{0},V_{1},\ldots,V_{L}$, of $\mathcal{C}$ into $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ and a spanning sub-digraph $D_{c}^{\prime}$ of $D_{c}$ for each $c\in\mathcal{C}$ such that the following properties hold:
  1. (i)
    $L_{0}<L,M<n_{0}$ and $|V_{0}|+|\mathcal{C}_{0}|\leq\epsilon n$,
  2. (ii)
    $|V_{1}|=\cdots=|V_{L}|=|\mathcal{C}_{1}|=\cdots=|\mathcal{C}_{M}|=:m$,
  3. (iii)
    $\sum_{c\in\mathcal{C}}d_{D_{c}^{\prime}}^{+}(v)\geq\sum_{c\in\mathcal{C}}d_{D_{c}}^{+}(v)-(\frac{24d}{\delta^{3}}+\epsilon)n^{2}$ and $\sum_{c\in\mathcal{C}}d_{D_{c}^{\prime}}^{-}(v)\geq\sum_{c\in\mathcal{C}}d_{D_{c}}^{-}(v)-(\frac{24d}{\delta^{3}}+\epsilon)n^{2}$ for all $v\in V$, and $e(D_{c}^{\prime})\geq e(D_{c})-(\frac{24d}{\delta^{3}}+\epsilon)n^{2}$ for all $c\in\mathcal{C}$,
  4. (iv)
    if for each $c\in\mathcal{C}$, the digraph $D_{c}^{\prime}$ has an edge with both end vertices in a single cluster $V_{i}$ for some $i\in[L]$, then $c\in\mathcal{C}_{0}$,
  5. (v)
    for all triples $\{(h,i),j\}\in\binom{[L]}{2}\times[M]$, we have either $D_{c}^{\prime}[V_{h},V_{i}]=\emptyset$ for all $c\in\mathcal{C}_{j}$, or $\mathcal{D}_{hi,j}^{\prime}:=\{D_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon,d)$-regular.
The sets $V_{i}$ are called vertex clusters and the sets $\mathcal{C}_{i}$ are called color clusters, while $V_{0}$ and $\mathcal{C}_{0}$ are the exceptional vertex and color sets respectively. Now, we define the reduced digraph collection.
Definition 2.3 (Reduced digraph collection).
Given a digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on $V$ and parameters $\epsilon>0$, $d\in[0,1)$ and $L_{0}\geq 1$, the reduced digraph collection $\mathcal{R}=\mathcal{R}(\epsilon,d,L_{0})$ of $\mathcal{D}$ is defined as follows. Apply Lemma 2.2 to $\mathcal{D}$ with parameters $\epsilon,\delta,d,L_{0}$ to obtain $\mathcal{D}^{\prime}$ and a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$, and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}$, where $V_{0},\mathcal{C}_{0}$ are the exceptional sets and $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ are the color clusters. Then :=$\mathcal{R}:=\{R_{1},\ldots,R_{M}\}$ is a digraph collection of $M$ digraphs each on the same vertex set $[L]$, where, for $\{(h,i),j\}\in\binom{[L]}{2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}$ is $(\epsilon,d)$-regular.
The next lemma states that clusters inherit a minimum semi-degree bound in the reduced digraph collection.
Lemma 2.4 (Degree inheritance).
Suppose $L_{0}\geq 1$ and $0<\frac{1}{n}\ll\epsilon\leq d\ll\delta,\gamma,p\leq 1$. Let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a digraph collection on vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq(p+\gamma)n$ and $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$. Let $\mathcal{R}:=\mathcal{R}(\epsilon,d,L_{0})$ be the reduced digraph collection of $\mathcal{D}$ consisting of $M$ digraphs on a common vertex set $[L]$. Then
  1. (i)
    for every $i\in[L]$, there are at least $(1-d^{\frac{1}{4}})M$ colors $j\in[M]$ for which $d_{R_{j}}^{+}(i)\geq(p+\frac{\gamma}{2})L$ and $d_{R_{j}}^{-}(i)\geq(p+\frac{\gamma}{2})L$,
  2. (ii)
    for every $j\in[M]$, there are at least $(1-d^{\frac{1}{4}})L$ vertices $i\in[L]$ for which $d_{R_{j}}^{+}(i)\geq(p+\frac{\gamma}{2})L$ and $d_{R_{j}}^{-}(i)\geq(p+\frac{\gamma}{2})L$.
Proof.
(i) Choose a vertex $v\in V\setminus V_{0}$. By Lemma 2.2 (i) and (iii), we have
$\sum_{c\in\mathcal{C}}d_{D_{c}^{\prime}-V_{0}}^{+}(v)\geq\sum_{c\in\mathcal{C}}d_{D_{c}}^{+}(v)-\left(\frac{24d}{\delta^{3}}+\epsilon\right)n^{2}-\epsilon n|\mathcal{C}|\geq\sum_{c\in\mathcal{C}}d_{D_{c}}^{+}(v)-\frac{26dn^{2}}{\delta^{3}}.$
Let $\mathcal{D}_{v}:=\left\{c\in\mathcal{C}\setminus\mathcal{C}_{0}:d_{D_{c}^{\prime}-V_{0}}^{+}(v)\geq d_{D_{c}}^{+}(v)-\sqrt{d}n\right\}$. Then
$\sum_{c\in\mathcal{C}\setminus\mathcal{C}_{0}}d_{D_{c}^{\prime}-V_{0}}^{+}(v)\leq\sum_{c\in\mathcal{C}\setminus\mathcal{C}_{0}}d_{D_{c}}^{+}(v)-\sqrt{d}n|\mathcal{C}\setminus(\mathcal{C}_{0}\cup\mathcal{D}_{v})|.$
Therefore, $|\mathcal{C}\setminus(\mathcal{C}_{0}\cup\mathcal{D}_{v})|\leq\frac{d^{\frac{1}{3}}n}{2}$ since $d\ll\delta$. It follows that $|\mathcal{D}_{v}|\geq|\mathcal{C}|-\frac{d^{\frac{1}{3}}n}{2}-\epsilon n\geq|\mathcal{C}|-d^{\frac{1}{3}}n$. Recall that $mM\leq|\mathcal{C}|\leq mM+\epsilon n$ and $mL\leq n\leq mL+\epsilon n$. Together with $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$, one has $\delta L\leq 2(1-\epsilon)M$. Hence the number of color clusters $\mathcal{C}_{j}$ containing at least one color of $\mathcal{D}_{v}$ is at least
$\frac{|\mathcal{D}_{v}|}{m}\geq M-\frac{d^{\frac{1}{3}}n}{m}\geq M-\frac{d^{\frac{1}{3}}L}{1-\epsilon}\geq M-\frac{2d^{\frac{1}{3}}M}{\delta}\geq\left(1-\frac{d^{\frac{1}{4}}}{2}\right)M.$
Now let $i\in[L]$ and $v\in V_{i}$. For each cluster $\mathcal{C}_{j}$ as above, choose an arbitrary color $c_{j}\in\mathcal{C}_{j}\cap\mathcal{D}_{v}$. Thus, the number of clusters $V_{h}$ containing some vertices in $N_{D_{c_{j}}^{\prime}}^{+}(v)$ is at least
$\frac{d_{D_{c_{j}}}^{+}(v)-\sqrt{d}n}{m}\geq\frac{(p+\gamma-\sqrt{d})n}{m}\geq\left(p+\frac{\gamma}{2}\right)L.$
Lemma 2.2 (v) implies that $i$ is adjacent to each such $V_{h}$ in $R_{j}$. Therefore, for every $i\in[L]$, $d_{R_{j}}^{+}(i)\geq(p+\frac{\gamma}{2})L$ for at least $(1-\frac{d^{\frac{1}{4}}}{2})M$ colors $j$. The proofs of other cases are similar and we omit it. ∎
2.2 Transversal perfect matching in bipartite graph collections
Let $D$ be a digraph with vertex set $V$. We define $B_{D}$ to be the characteristic bipartite graph of $D$ with bipartition $V_{1}\cup V_{2}$ where $V_{1}=V_{2}=V$, for each $u\in V_{1}$ and $v\in V_{2}$, $uv\in E(B_{D})$ if and only if $uv\in E(D)$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$. Define the characteristic bipartite graph collection of $\mathcal{D}$ to be $\mathcal{B_{D}}=\{B_{D_{1}},\ldots,B_{D_{n}}\}$.
A matching in a (di)graph $G$ is a collection of vertex-disjoint edges $M\subseteq E(G)$. We say that $M$ is a perfect matching if $V(M)=V(G)$. It is routine to check that $\mathcal{B_{D}}$ contains a transversal perfect matching if and only if the vertex set of $\mathcal{D}$ can be covered by a set of disjoint rainbow cycles inside $\mathcal{D}$. Therefore, in the following of this section, we are to find a transversal perfect matching in a bipartite graph collection.
Joos and Kim [21] established a minimum degree condition that guarantees the existence of transversal perfect matchings in a graph collection. Bradshaw [5] proved the analogue result in bipartite graph collections, which can be stated as follows.
Theorem 2.5 ([5]).
Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a collection of bipartite graphs on a common vertex bipartition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. If for each $i\in[n]$ we have the following hold:
  • (i)
    $d_{G_{i}}(v)>\frac{n}{2}$ for each $v\in V_{1}$,
  • (ii)
    $d_{G_{i}}(v)\geq\frac{n}{2}$ for each $v\in V_{2}$,
then $\mathcal{G}$ contains a transversal perfect matching.
In this subsection, we establish a stability result for transversal perfect matching in bipartite graph collections, that is, the minimum degree required to guarantee a transversal perfect matching can be below that of Theorem 2.5, as long as the bipartite graph collection $\{G_{1},\ldots,G_{n}\}$ is far in edit distance from several extremal cases. For this, we first define the extremality for a single bipartite graph.
Definition 2.6 (nice, extremal).
Let $\epsilon>0$ and $G$ be a balanced bipartite graph on vertex set $V_{1}\cup V_{2}$ of size $2n$. We say that
  • $G$ is $\epsilon$-nice if for any two sets $A\subseteq V_{1}$ and $B\subseteq V_{2}$ of size at least $(\frac{1}{2}-\epsilon)n$ we have $e_{G}(A,B)\geq\epsilon n^{2}$;
  • $G$ is $\epsilon$-extremal if it is not $\epsilon^{5}$-nice.
The following lemma is the basis for our proof.
Lemma 2.7.
Suppose that $0<\frac{1}{n}\ll d\ll\epsilon\leq 1$. Let $G$ be a balanced bipartite graph on a vertex set $V=V_{1}\cup V_{2}$ of size $2n$ with $d_{G}(x)\geq(\frac{1}{2}-\epsilon^{5})n$ for all but at most $dn$ vertices $x\in V$ which is $\epsilon$-extremal. Then there is a characteristic partition $(A_{1},B_{1},A_{2},B_{2},C_{1},C_{2})$ of $G$ such that the following hold:
  • (i)
    $A_{i},B_{i},C_{i}\subseteq V_{i}$, $|A_{i}|=|B_{i}|=(\frac{1}{2}-\epsilon)n$ and $|C_{i}|=2\epsilon n$ for $i\in[2]$,
  • (ii)
    $d_{G}(v,X_{i})\geq(\frac{1}{2}-2\epsilon)n$ for $v\in X_{3-i}$ and $X\in\{A,B\}$, $i\in[2]$,
  • (iii)
    either $e_{G}(A_{1},B_{2})\leq\epsilon n^{2}$ or $e_{G}(A_{2},B_{1})\leq\epsilon n^{2}$.
Definition 2.8 (crossing, cross graph).
Let $10<\frac{1}{n},\epsilon,\delta<$ 1 where $n\in\mathbb{N}$ and let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a balanced bipartite graph collection on a common vertex set $V$ of size $2n$. Given $i,j\in[n]$ such that $G_{i}$ and $G_{j}$ are both $\epsilon$-extremal, assume that the characteristic partition corresponding to $G_{\ell}$ is $(A_{1}^{\ell},B_{1}^{\ell},C_{1}^{\ell},A_{2}^{\ell},B_{2}^{\ell},C_{2}^{\ell})$ for $\ell\in\{i,j\}$. We say that they are $\delta$-crossing if $|A^{i}_{1}\Delta A^{j}_{1}|\geq\delta n$ and $|A^{i}_{1}\Delta B^{j}_{1}|\geq\delta n$.
We define the cross graph $C_{\mathcal{G}}^{\epsilon,\delta}$ to be the graph with vertex set $[n]$ where $i$ is adjacent to $j$ if and only if $G_{i}$ and $G_{j}$ are both $\epsilon$-extremal and $\delta$-crossing.
Definition 2.9 (strongly stable, weakly stable [^1] To avoid introducing too many terms, we have similarly defined the concepts of “nice, extremal, crossing, cross graph, strongly stable, weakly stable, stable” in the context of digraph collections. We should note that, unless explicitly stated otherwise, those concepts in bipartite graph collections are only discussed in Section 2.2 and the appendix.).
Let $10<\gamma,\alpha,\epsilon,\delta<$ 1. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. We say that
  • $\mathcal{G}$ is $(\gamma,\alpha)$-strongly stable if $G_{i}$ is $\alpha$-nice for at least $\gamma n$ colors $i\in[n]$;
  • $\mathcal{G}$ is $(\epsilon,\delta)$-weakly stable if $e(C_{\mathcal{G}}^{\epsilon,\delta})\geq\delta n^{2}$;
  • $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable if it is either $(\gamma,\alpha)$-strongly stable or $(\epsilon,\delta)$-weakly stable.
The following is a stability result for transversal perfect matching in bipartite graph collections.
Theorem 2.10.
Let $0<\frac{1}{n}\ll d\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1$. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. If the following hold:
  • $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable,
  • for every $i\in[n]$, $d_{G_{i}}(x)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ vertices $x\in V_{1}\cup V_{2}$,
  • for every $x\in V_{1}\cup V_{2}$, $d_{G_{i}}(x)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ colors $i\in[n]$,
then $\mathcal{G}$ contains a transversal perfect matching.
For convenience, we postpone the proofs of Lemma 2.7 and Theorem 2.10 to the appendix.
2.3 Probabilistic tool
We use the following version of Chernoff’s bound.
Lemma 2.11.
Let $X$ be a random variable with binomial or hypergeometric distribution, and let $0<\epsilon<\frac{3}{2}$. Then
$\mathbb{P}[|X-\mathbb{E}(X)|\geq\epsilon\mathbb{E}(X)]\leq 2e^{-\frac{\epsilon^{2}}{3}\mathbb{E}(X)}.$
3 Absorbing
Denote by $V^{k}$ the set of all $k$-tuples of distinct elements of $V$. A pair $H=(V,E)$ where $V$ is a set and $E\subseteq V^{k}$ is called a directed $k$-graph, while if we allow $E$ to be a multiset, it is a directed multi-$k$-graph. We use the same notation for directed $k$-graphs as for graphs. The following lemma was obtained by Cheng and Staden [10], which is the key tool that we will use to build an absorbing structure.
Lemma 3.1 ([10]).
Suppose that $k,C,n\in\mathbb{N}$, $0\leq\frac{1}{n}\ll\gamma\ll\epsilon\ll\frac{1}{k},\frac{1}{C}\leq 1$, $m\in[n^{C}]$ and $t:=\gamma n$. Let $\mathbf{H}=\{H_{1},\ldots,H_{t}\}$ be a collection of directed $k$-graphs and let $\mathbf{Z}=\{Z_{1},\ldots,Z_{m}\}$ be a collection of directed multi-$k$-graphs all defined on a common vertex set $V$ of size $n$. Suppose that $|E(H_{i})|\geq\epsilon n^{k}$ for all $i\in[t]$, and for each $j\in[m]$ we have $|E(Z_{j})\cap E(H_{i})|\geq\epsilon n^{k}$ for at least $\epsilon t$ indices $i\in[t]$. Then there is a rainbow matching $M$ in $\mathbf{H}$ of size at least $(1-\frac{\epsilon^{2}}{4})t$ and $|E(Z_{j})\cap E(M)|\geq\frac{\epsilon^{2}t}{4}$ for each $j\in[m]$.
We now define the extremality for a single digraph.
Definition 3.2 (nice, extremal).
Let $\epsilon>0$ and $D$ be a digraph on vertex set $V$ of size $n$. We say that
  • $D$ is $\epsilon$-nice if for any two sets $A,B\subseteq V$ of size at least $(\frac{1}{2}-\epsilon)n$, we have $e_{D}(A,B)\geq\epsilon n^{2}$;
  • $D$ is $\epsilon$-extremal if it is not $\epsilon$-nice.
Cheng, Wang and Yan [13] characterized the structure of $\epsilon$-extremal digraphs, which forms the basis of our extremal case distinction.
Lemma 3.3 ([13]).
Suppose that $0<\frac{1}{n}\ll d\ll\mu\ll\epsilon\leq 1$. Let $D$ be an $\epsilon$-extremal digraph on vertex set $V$ of size $n$ with $\min\{d^{+}(x),d^{-}(x)\}\geq(\frac{1}{2}-\mu)n$ for all but at most $dn$ vertices $x\in V$. Then exactly one of the following holds:
  1. (i)
[…]
  2. (ii)
[…]
  3. (iii)
[…]
Figure 1: Extremal digraphs EC1, EC2, and EC3. The gray shaded elliptical indicates that the digraph induced by this vertex set is complete, the gray shaded arrow between two vertex sets indicates that the induced digraph by them is complete in this direction.
[…]
Definition 3.4 (directed absorbing path and absorbing cycle).
Given any two not necessarily distinct vertices $u,v\in V$ and a rainbow directed path $P=v_{1}v_{2}v_{3}v_{4}$ with $u,v\not\in V(P)$, we call $P$ a Type-I directed $c$-absorbing path of $(v,u)$ if $c\in L({v_{2}v})$ and ${\rm col}({v_{2}v_{3}})\in L({uv_{3}})$ (see Figure 2), a Type-II directed $c$-absorbing path of $(v,u)$ if $c\in L({vv_{3}})$ and ${\rm col}({v_{2}v_{3}})\in L({v_{2}u})$ (see Figure 3).
Given $\delta,\delta^{\prime},\gamma,\gamma^{\prime}>0$ and ${\rm K}\in\{{\rm I,II}\}$, a rainbow directed cycle $C=v_{1}v_{2}\ldots v_{t}v_{1}$ is a Type-K directed absorbing cycle with parameters $(\delta,\delta^{\prime},\gamma,\gamma^{\prime})$ if $t\leq\gamma n$ and there exists a color set $\mathcal{C}$ of size at least $\delta n$ such that
  • given any color $c\in\mathcal{C}$ and any $D_{c}$-good vertex $v$, for all but at most $\delta^{\prime}n$ vertices $u\in V$, there are at least $\gamma^{\prime}n$ disjoint Type-K directed $c$-absorbing paths of $(v,u)$ inside $C$,
  • given any color $c\in\mathcal{C}$, for all but at most $\delta^{\prime}n$ $D_{c}$-good vertices $v$, there are at least $\gamma^{\prime}n$ disjoint Type-K directed $c$-absorbing paths of $(v,v)$ inside $C$.
If there is no ambiguity, we typically refer to the Type-I and Type-II directed $c$-absorbing path (cycle) simply as directed $c$-absorbing path (cycle).
Figure 2: A Type-I directed $c$-absorbing path of $(v,v)$ and a Type-I directed $c$-absorbing path of $(v,u)$ with $v\neq u$.
Figure 3: A Type-II directed $c$-absorbing path of $(v,v)$ and a Type-II directed $c$-absorbing path of $(v,u)$ with $v\neq u$.
Definition 3.5 (strongly stable).
Let $10<\gamma,\alpha<$ 1. Suppose that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$. We say that $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable if $D_{i}$ is $\alpha$-nice for at least $\gamma n$ colors $i\in[n]$.
The following lemma finds a directed absorbing cycle when $\mathcal{D}$ is strongly stable.
Lemma 3.6.
Let $1<\frac{1}{n}\ll\lambda,\mu\ll\gamma,\alpha\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$. If $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, then there exists a Type-I directed absorbing cycle with parameters $(1,0,\lambda,\lambda^{2})$.
Proof.
Since $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, one may assume that $D_{1},\ldots,D_{\gamma n}$ are $\alpha$-nice and $6|\lambda n$. We divide the color set $[\frac{\lambda n}{2}]$ into consecutive sets $\{1,2,3\},\{4,5,6\},\ldots,\{\frac{\lambda n}{2}-2,\frac{\lambda n}{2}-1,\frac{\lambda n}{2}\}$. For each $i\in I:=[\frac{\lambda n}{6}]$, define $\mathcal{F}_{i}$ to be a directed $4$-graph with vertex $V$ and edge set
$\left\{(v_{1},v_{2},v_{3},v_{4})\in V^{4}:3(i-1)+\ell\in L({v_{\ell}v_{\ell+1}})\ \text{for each}\ \ell\in[3]\right\}.$
The minimum semi-degree condition implies that $e(\mathcal{F}_{i})\geq n\left(\frac{1}{2}-\mu\right)^{3}n^{3}\geq\frac{n^{4}}{9}$.
Given a color $c\in[n]$ and two not necessarily distinct vertices $v,v^{\prime}\in V$ (which are not required to be $D_{c}$-good), let $Z_{i}(c,vv^{\prime})$ be the collection of $(v_{1},v_{2},v_{3},v_{4})$ for which $v_{1}v_{2}v_{3}v_{4}$ is a Type-I directed $c$-absorbing path of $(v,v^{\prime})$. Since $D_{3i+2}$ is $\alpha$-nice, there are at least $\alpha n^{2}$ ways to choose $v_{2},v_{3}$ with $v_{2}\in N_{c}^{-}(v),v_{3}\in N_{{3i+2}}^{+}(v^{\prime})$ and ${v_{2}v_{3}}\in E(D_{3i+2})$. When $v_{2},v_{3}$ are fixed, there are at least $\left(\frac{1}{2}-\mu\right)^{2}n^{2}$ ways to choose $v_{1}$ and $v_{4}$. Thus $|Z_{i}(c,vv^{\prime})|\geq\alpha\left(\frac{1}{2}-\mu\right)^{2}n^{4}\geq\frac{\alpha n^{4}}{5}$.
Applying Lemma 3.1 with
$t:=|I|,\ \epsilon:=\frac{\alpha}{5},\ \mathbf{H}:=\{\mathcal{F}_{i}:i\in I\},\ \text{and}$
$\mathbf{Z}:=\left\{Z(c,vv^{\prime}):=\bigcup_{i\in I}Z_{i}(c,vv^{\prime}):c\in[n],v,v^{\prime}\in V\right\},$
we obtain a rainbow matching $M$ inside $\mathbf{H}$ of size at least $(1-\frac{\alpha^{2}}{100})\frac{\lambda n}{6}$ such that $|E(Z(c,vv^{\prime}))\cap E(M)|\geq\frac{\alpha^{2}\lambda n}{600}$ for all $c\in[n]$ and $v,v^{\prime}\in V$. That is, there is a set $I^{\prime}\subseteq I$ with $|I^{\prime}|\geq(1-\frac{\alpha^{2}}{100})\frac{\lambda n}{6}$ such that there is a rainbow directed path $P^{i}=v_{1}^{i}v_{2}^{i}v_{3}^{i}v_{4}^{i}$ with $(v_{1}^{i},v_{2}^{i},v_{3}^{i},v_{4}^{i})\in\mathcal{F}_{i}$ for each $i\in I^{\prime}$, and for every $c\in[n]$ and $v,v^{\prime}\in V$, there are at least $\frac{\alpha^{2}\lambda n}{600}$ directed $c$-absorbing paths $P^{i}$ of $(v,v^{\prime})$.
By relabeling indices, one may assume $I^{\prime}=:[s]$ with $s=(1-\frac{\alpha^{2}}{100})\frac{\lambda n}{6}$. We are to connect the directed paths $P^{1},\ldots,P^{s}$ one by one into a rainbow directed cycle $C$. Firstly, we connect $P^{1}$ and $P^{2}$ into a single rainbow directed path $P^{1}xyP^{2}$. For this, choose distinct unused colors $c_{1},c_{2}\in[\lambda n+1,n]$ and $c_{3}\in[\frac{\lambda n}{2}+1,\lambda n]$. Let $U:=V\setminus V(\bigcup_{i\in[s]}P^{i})$. Clearly, $|U|\geq n-\lambda n$. It follows from $\delta^{0}(\mathcal{D})\geq(\frac{1-\mu}{2})n$ that
$d_{{c_{1}}}^{+}(v_{4}^{1},U),\,d_{{c_{2}}}^{-}(v_{1}^{2},U)\geq\left(\frac{1-\mu}{2}-\lambda\right)n\geq\left(\frac{1}{2}-\alpha\right)n.$ (1)
Notice that $D_{c_{3}}$ is $\alpha$-nice. Therefore, $|E(D_{c_{3}}[N_{{c_{1}}}^{+}(v_{4}^{1},U),N_{{c_{2}}}^{-}(v_{1}^{2},U)])|\geq\alpha n^{2}$. Hence, there is an edge ${xy}\in E(D_{c_{3}})$ with $x\in N_{{c_{1}}}^{+}(v_{4}^{1},U)$ and $y\in N_{{c_{2}}}^{-}(v_{1}^{2},U)$ such that $P^{1}xyP^{2}$ is a rainbow directed path with color set ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{2},c_{3}\}$.
We apply the above process for each pair $(P^{i},P^{i+1})$ for each $i\in[s]$ with $P^{s+1}:=P^{1}$. In each process, we need two unused colors from $[\lambda n+1,n]$ and one unused color from $[\frac{\lambda n}{2}+1,\lambda n]$. This is feasible since $2s\leq\frac{\lambda n}{3}$. The total number of used vertices at each step is at most $6s\leq{\lambda n}$, so we have the same degree bounds as (1). It is straightforward to verify that the resulting directed cycle $C$ is a Type-I directed absorbing cycle with parameters $(1,0,\frac{6s}{n},\frac{\alpha^{2}\lambda}{600})$ and hence with parameters $(1,0,\lambda,\lambda^{2})$ since $\lambda\ll\alpha$. ∎
The extremality of a digraph collection also depends on the relationships between the digraphs, so we give the following definitions.
Definition 3.7 (crossing, cross graph, weakly stable).
Let $10<\frac{1}{n},\epsilon,\delta<$ 1 with $n\in\mathbb{N}$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$. Choose $i,j\in[n]$ such that $D_{i}$ is $(\epsilon,{\rm EC}t)$-extremal and $D_{j}$ is $(\epsilon,{\rm EC}s)$-extremal with $t,s\in[3]$. We say that $D_{i}$ and $D_{j}$ are $\delta$-crossing if one of the following holds:
  1. (A1)
    $t,s\in\{1,2\}$, $|A_{i}\Delta A_{j}|\geq\delta n$ and $|A_{i}\Delta B_{j}|\geq\delta n$;
  2. (A2)
    $t=s=3$, either $|W_{i}^{1}\Delta W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\Delta W_{j}^{3}|\geq\delta n$, or $|W_{i}^{2}\Delta W_{j}^{2}|\geq\delta n$ and $|W_{i}^{2}\Delta W_{j}^{4}|\geq\delta n$;
  3. (A3)
    $t\in\{1,2\},s=3$, either $|A_{i}\Delta W_{j}^{1}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{3}|\geq\delta n$, or $|A_{i}\Delta W_{j}^{2}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{4}|\geq\delta n$.
For $k\in[3]$, define the cross graph $C_{\mathcal{D},k}^{\epsilon,\delta}$ to be the graph with vertex set $[n]$ where $i$ and $j$ are adjacent if and only if $D_{i},D_{j}$ are satisfying (Ak).
We say that $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable if $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$.
Observation 3.8.
Assume that $0<\epsilon\leq\frac{\delta}{8}$.
  • If $D_{i}$ and $D_{j}$ are satisfying (A1), then $|X_{i}\cap Y_{j}|\geq\frac{\delta n}{4}$, where $X,Y\in\{A,B\}$.
  • If $D_{i}$ and $D_{j}$ are satisfying (A2), then $|W_{i}^{x}\cap W_{j}^{y}|\geq\frac{\delta n}{4}$, where either $x,y\in\{1,3\}$ or $x,y\in\{2,4\}$.
  • If $D_{i}$ and $D_{j}$ are satisfying (A3), then $|X_{i}\cap W_{j}^{y}|\geq\frac{\delta n}{4}$, where $X\in\{A,B\}$ and either $y\in\{1,3\}$ or $y\in\{2,4\}$.
Our following lemma finds a directed absorbing cycle when $\mathcal{D}$ is weakly stable.
Lemma 3.9.
Let $10<\frac{1}{n}\ll\lambda,\mu,\epsilon\ll\delta<$ 1 and $\delta\ll\min\{\zeta,\frac{1}{2}-\zeta-\epsilon\}$ (where $\zeta$ is defined in Lemma 3.3). Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$. If $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, then there exists a directed absorbing cycle with parameters $(\frac{\delta}{3},\sqrt{\epsilon},\lambda,\lambda^{2})$.
Proof.
Since $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, one has $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$. In particular, if $k=2$, then we consider only $|W_{i}^{1}\Delta W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\Delta W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},2}^{\epsilon,\delta})$; if $k=3$, then we consider only $|A_{i}\Delta W_{j}^{1}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},3}^{\epsilon,\delta})$ (since the other case can be discussed similarly).
For each $i\in[\lfloor\frac{n}{3}\rfloor]$, define $\mathcal{F}_{i}$ to be a directed $4$-graph with vertex set $V$ and edge set
$\{(v_{1},v_{2},v_{3},v_{4})\in V^{4}:3(i-1)+\ell\in L(v_{\ell}v_{\ell+1})\ \text{for}\ \ell\in[3]\}.$
If follows from $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$ that $e(\mathcal{F}_{i})\geq n\left(\left(\frac{1}{2}-\mu\right)n\right)^{3}\geq\frac{n^{4}}{9}$.
We first give the following claim, which requires only the minimum semi-degree condition.
Claim 3.10.
Assume that $ij\in E(C_{\mathcal{D},k}^{\epsilon,\delta})$ for some $k\in[3]$. Let $u$ be a $D_{j}$-good vertex and $v$ be a $D_{i}$-good vertex with $j+1<3n$. For ${\rm K}\in\{{\rm I,II}\}$, let $Z_{j}^{\rm K}(i,vu)$ be the collection of $(v_{1},v_{2},v_{3},v_{4})\in\mathcal{F}_{\frac{j+1}{3}}$ where $v_{1}v_{2}v_{3}v_{4}$ is a Type-K directed $i$-absorbing path of $(v,u)$. Then $|Z_{j}^{\rm K}(i,vu)|\geq 2^{-7}\delta n^{4}$.
Proof of Claim 3.10.
Fix any such $i,j,u,v$. We proceed by considering the following cases.
Case 1. $k=1$.
In this case, we prove $|Z_{j}^{\rm I}(i,vu)|\geq 2^{-7}\delta n^{4}$. By Lemma 3.3 (i) and (ii), we have either $d_{i}^{-}(v,A_{i})\geq(\frac{1}{2}-2\epsilon)n$ or $d_{i}^{-}(v,B_{i})\geq(\frac{1}{2}-2\epsilon)n$. Without loss of generality, we assume that the former case holds since the latter case can be proved similarly. Note that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then $|A_{i}\cap A_{j}|\geq\frac{\delta n}{4}$ and $|A_{i}\cap B_{j}|\geq\frac{\delta n}{4}$. Therefore,
$d_{i}^{-}(v,A_{j})\geq|A_{i}\cap A_{j}|-(|A_{i}|-d_{i}^{-}(v,A_{i}))\geq\frac{\delta n}{4}-\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\ll\delta).$
Similarly, we have $d_{i}^{-}(v,B_{j})\geq\frac{\delta n}{5}$.
[…]
$|N_{j}^{+}(u,Z_{j})\cap N_{j}^{+}(x,Z_{j})|\geq|N_{j}^{+}(u,Z_{j})|+|N_{j}^{+}(x,Z_{j})|-|Z_{j}|\geq\left(\frac{1}{2}-3\epsilon\right)n.$
Any choice of $y\in N_{j}^{+}(u,Z_{j})\cap N_{j}^{+}(x,Z_{j}),\ x^{\prime}\in N_{{j-1}}^{-}(x)$ and $y^{\prime}\in N_{{j+1}}^{+}(y)$ yields a directed $i$-absorbing path $x^{\prime}xyy^{\prime}$ of $(v,u)$ in $\mathcal{F}_{\frac{j+1}{3}}$. The number of such paths is therefore at least $\frac{\delta n}{5}(\frac{1}{2}-3\epsilon)n\left(\frac{1}{2}-\mu\right)^{2}n^{2}\geq 2^{-7}\delta n^{4}$.
Case 2. $k=2$.
In this case, we prove $|Z_{j}^{\rm II}(i,vu)|\geq 2^{-7}\delta^{2}n^{4}$. By Lemma 3.3 (iii), we have either $d_{D_{i}}^{+}(v,W_{i}^{1})\geq(\frac{1}{2}-3\epsilon)n$ or $d_{D_{i}}^{+}(v,W_{i}^{3})\geq(\frac{1}{2}-3\epsilon)n$. Without loss of generality, we assume that the former case holds since the latter case can be proved similarly. Note that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then $|W_{i}^{1}\cap W_{j}^{1}|\geq\frac{\delta n}{4}$ and $|W_{i}^{1}\cap W_{j}^{3}|\geq\frac{\delta n}{4}$. Therefore,
$d_{i}^{+}(v,W_{j}^{1})\geq|W_{i}^{1}\cap W_{j}^{1}|-\left(|W_{i}^{1}|-d_{i}^{+}(v,W_{i}^{1})\right)\geq\frac{\delta n}{4}-2\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\ll\delta).$
Similarly, we have $d_{i}^{+}(v,W_{j}^{3})\geq\frac{\delta n}{5}$.
Since $u$ is $D_{j}$-good, one has $u\in Y_{j}$ for some $Y\in\{W^{1},W^{3}\}$. Let $Z=W^{4}$ if $Y=W^{1}$, and let $Z=W^{2}$ if $Y=W^{3}$. Note that $|N_{j}^{-}(w,Z_{j})|\geq(\frac{1}{2}-3\epsilon)n$ for each vertex $w\in Y_{j}$. Choose $x\in N_{i}^{+}(v,Y_{j})$. Then
$|N_{j}^{-}(u,Z_{j})\cap N_{j}^{-}(x,Z_{j})|\geq|N_{j}^{-}(u,Z_{j})|+|N_{j}^{-}(x,Z_{j})|-|Z_{j}|\geq\left(\frac{1}{2}-5\epsilon\right)n.$
Any choice of $y\in N_{j}^{-}(u,Z_{j})\cap N_{j}^{-}(x,Z_{j}),\ x^{\prime}\in N_{{j+1}}^{+}(x)$ and $y^{\prime}\in N_{{j-1}}^{-}(y)$ yields a Type-II directed $i$-absorbing path $y^{\prime}yxx^{\prime}$ of $(v,u)$ in $\mathcal{F}_{\frac{j+1}{3}}$. Therefore, the number of such paths is at least $\frac{\delta n}{5}(\frac{1}{2}-5\epsilon)n\left(\frac{1}{2}-\mu\right)^{2}n^{2}\geq 2^{-7}\delta n^{4}$.
Case 3. $k=3$.
This case can be proved by a similar discussion, so we omit the proof. ∎
For each vertex $v$, let
[…]
Since $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$, there is a subgraph $H$ of $C_{\mathcal{D},k}^{\epsilon,\delta}$ such that $|V(H)|\geq\delta n$ and $\delta(H)\geq\delta n$. For each $i\in V(H)$, define
$T_{i}:=\left\{u\in V:|N_{H}(i)\setminus\mathcal{C}_{u}|\geq\frac{\delta n}{2}\right\}.$
Notice that $|L_{i}|=2\epsilon n$ when $D_{i}$ is $\epsilon$-extremal. Hence $|T_{i}|\frac{\delta n}{2}\leq 2\epsilon n^{2}$, that is, $|T_{i}|\leq\frac{\sqrt{\epsilon}}{2}n$ since $\epsilon\ll\delta$.
Let $\overline{T_{i}}:=V\setminus T_{i}$. Then $|\overline{T_{i}}|\geq(1-\frac{\sqrt{\epsilon}}{2})n$. For each $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|N_{H}(i)\cap\mathcal{C}_{u}|\geq\frac{\delta n}{2}$. Now we independently and randomly select vertices from $V(H)$ with probability $\kappa:=\frac{\lambda}{14}$ to obtain a color set $\mathcal{U}$. By using Chernoff’s bound, with high probability the following hold:
  1. (*)
    $\frac{\kappa|V(H)|}{2}\leq t:=|\mathcal{U}|\leq 2\kappa|V(H)|$,
  2. (**∗ ∗)
    for every $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|N_{H}(i,\mathcal{U})\cap\mathcal{C}_{u}|\geq\frac{\delta\kappa n}{4}$.
Fix such a color set $\mathcal{U}$ and one may assume $\mathcal{U}=\{3j-1:j\in[t]\}$. Let $\overline{\mathcal{U}}:=V(H)\cap[3t+1,n]$. It is easy to see that $\mathcal{U}\cap\overline{\mathcal{U}}=\emptyset$, $|\overline{\mathcal{U}}|\geq\frac{\delta n}{2}$ and $\delta(H[\overline{\mathcal{U}}])\geq\frac{\delta n}{2}$.
Given $i\in\overline{\mathcal{U}}$ and $u\in\overline{T_{i}}$, choose $j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})$. Then $\frac{j+1}{3}\in[t]$ and the number of such choices of $j$ is at least $\frac{\delta\kappa n}{4}\geq\frac{\delta t}{8}$ by (**∗ ∗). Let $v$ be a $D_{i}$-good vertex. It follows from Claim 3.10 that $|Z_{j}^{\rm K}(i,vu)|\geq 2^{-7}\delta n^{4}$ for some ${\rm K}\in\{{\rm I,II}\}$.
By (*) and the fact that $\lambda\ll\delta$, one has $\frac{t}{n}\leq 2\kappa=\frac{\lambda}{7}\ll 2^{-7}\delta$. Applying Lemma 3.1 with
$\gamma:=\frac{\lambda}{7},\ \epsilon:=2^{-7}\delta,\ \mathbf{H}:=\{\mathcal{F}_{j}:j\in[t]\}\ \text{and}$
[…]
we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(1-2^{-16}\delta^{2})t$ such that $|E(Z(i,vu))\cap E(M)|\geq 2^{-16}\delta^{2}t$ for all $Z(i,vu)\in\mathbf{Z}$. That is, there is a subset $I\subseteq[t]$ with $|I|\geq(1-2^{-16}\delta^{2})t$ such that there is a Type-K directed path $P^{j}=v_{1}^{j}v_{2}^{j}v_{3}^{j}v_{4}^{j}\in\mathcal{F}_{j}$ for each $j\in I$, and for every $i\in\overline{\mathcal{U}}$, $u\in\overline{T_{i}}$ and a $D_{i}$-good vertex $v$, there are at least $2^{-16}\delta^{2}t$ Type-K directed $i$-absorbing paths $P^{j}$ of $(v,u)$.
Relabel indices such that $I=:[s]$ with $s\geq(1-2^{-16}\delta^{2})t$. We are to connect those disjoint rainbow directed paths $P^{1},\ldots,P^{s}$ into a directed absorbing cycle $C$. Notice that $\overline{\mathcal{U}}\cap{\rm col}(\bigcup_{i\in I}P^{i})=\emptyset$. We first connect $P^{1}$ and $P^{2}$ into a single rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xzv_{1}^{2}P^{2}$. For this, choose distinct colors $c_{1},c_{2},c_{3},c_{4}\in\overline{\mathcal{U}}$ such that $c_{1}c_{2}\in E(H)$. Clearly, there are at least $|\overline{\mathcal{U}}|\geq\frac{\delta n}{2}$ choices for $c_{1}$, and given this at least $\delta(H[\overline{\mathcal{U}}])\geq\frac{\delta n}{2}$ choices for $c_{2}$, and at least $|\overline{\mathcal{U}}|-3>\frac{\delta n}{3}$ choices for each of $c_{3},c_{4}$. Next, choose an unused $D_{c_{1}}$-good vertex $x\in N^{+}_{{c_{3}}}(v_{4}^{1})$ and an unused $D_{c_{2}}$-good vertex $y\in N^{-}_{{c_{4}}}(v_{1}^{2})$.
Claim 3.11.
One of the following holds:
  • there exists a rainbow directed path $P^{1}v_{4}^{1}xyzv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{2},c_{3},c_{4}\}$ and $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$, each of $x,y,z$ has at least $\frac{\delta n}{10}$ choices;
  • there exists a rainbow directed path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$ and $y\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})$, each of $x,y$ has at least $\frac{\delta n}{10}$ choices.
Proof of Claim 3.11.
We proceed by considering the following three cases.
Case 1. $k=1$.
[…]
$|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|-7e(M)\geq|N_{{c_{1}}}^{+}(x,Z_{c_{1}})\cap N_{{c_{2}}}^{-}(y,W_{c_{2}})|-14\kappa n\geq|Z_{c_{1}}\cap W_{c_{2}}|-2\epsilon n-12\kappa n\geq\frac{\delta n}{5},$
as desired.
Case 2. $k=2$.
[…]
If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$, then choose $x\in W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$ and $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Hence, there are at least $\delta n-7e(M)-2\epsilon n\geq\frac{\delta n}{2}$ choices for each of $x$ and $y$. Furthermore, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{3}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired.
If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|<\delta n$, then $|W_{c_{1}}^{4}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq(\frac{1}{2}-\mu-\delta-2\epsilon)n$. Choose $x\in W_{c_{1}}^{4}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$. Then $d_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\geq|W_{c_{1}}^{1}|-2\epsilon n$. Notice that either $|W_{c_{2}}^{1}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$ or $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$. We may assume the former case holds. Choose $y\in W_{c_{2}}^{1}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Then $d_{{c_{2}}}^{-}(y,W_{c_{2}}^{4})\geq|W_{c_{2}}^{4}|-2\epsilon n$. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|\geq\delta n$, then $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\frac{\delta n}{9}$, then $|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|\geq(\frac{1}{2}-\delta-3\epsilon)n$ and choose $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$. Similarly, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired. If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|<\frac{\delta n}{9}$, then $d^{-}_{{c_{4}}}(v_{1}^{2},W_{c_{2}}^{1})\geq|W_{c_{2}}^{1}|-(2\epsilon+\frac{\delta}{9})n$. Therefore,
$|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})|\geq|N_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\cap N_{{c_{4}}}^{-}(v_{1}^{2},W_{c_{2}}^{1})|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{1}|-\left(4\epsilon+\frac{\delta}{9}\right)n\geq\frac{\delta n}{9}.$
Now, we connect $P^{1}$ and $P^{2}$ by using a directed rainbow path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
Case 3. $k=3$.
[…]
Now, we consider $|Z_{c_{1}}\cap W_{c_{2}}|<\frac{\delta n}{4}$. Without loss of generality, assume that $Z=A$ and $W=W^{2}$. It follows that $|A_{c_{1}}\cap W_{c_{2}}^{4}|\geq|W_{c_{2}}^{4}|-(\frac{\delta}{4}+2\epsilon)n$. This implies $|A_{c_{1}}\cap W_{c_{2}}^{3}|\geq|C_{c_{2}}^{4}|-(\frac{\delta}{4}+2\epsilon)n\geq\frac{\delta n}{2}$ since $|C^{2}_{4}|\geq\delta n$. If $|W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq\frac{\delta n}{9}$, then choose $y\in W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})$. Hence $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|A_{c_{1}}\cap W_{c_{2}}^{4}|-3\epsilon n\geq\frac{\delta n}{5}$. If $|W_{c_{2}}^{1}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|<\frac{\delta n}{9}$, then $|W_{c_{2}}^{3}\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|W_{c_{2}}^{3}|-(\frac{\delta}{9}+2\epsilon)n$. Thus, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|A_{c_{1}}\cap W_{c_{2}}^{3}|-(3\epsilon+\frac{\delta}{9})n\geq\frac{\delta n}{9}$. Now, we connect $P^{1}$ and $P^{2}$ by using a directed rainbow path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
This completes the proof of Claim 3.11. ∎
In view of Claim 3.11, there are at least $\frac{\delta n}{10}$ choices for each of $c_{1},c_{2},c_{3},c_{4},x,y,(z)$ given any previous choices. Thus, we obtain the desired directed rainbow path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$. Applying the above process for each pair $(P^{j},P^{j+1})$ with $j\in[s]$ where $P^{s+1}:=P^{1}$. Every time we use at most four unused colors in $\overline{\mathcal{U}}$ and at most three unused vertices. Therefore, so for each choice, at most $4s\leq 4t\leq 8\kappa n\ll\frac{\delta n}{10}$ colors are forbidden and $3s\ll\frac{\delta n}{10}$ vertices are forbidden. Thus we obtain a directed rainbow cycle $C$ of length at most $7s$ which contains every $P^{j}$ with $j\in[s]$ as a segment.
By the construction of $C$, for every color $i\in\overline{\mathcal{U}}$, any $D_{i}$-good vertex $v$ and for every vertex $u\in\overline{T_{i}}$, there are at least $2^{-16}\delta^{2}t\geq 2^{-17}\kappa\delta^{3}n$ disjoint directed $i$-absorbing paths of $(v,u)$. Moreover, the number of $D_{i}$-good vertices inside $\overline{T_{i}}$ is at least $|\overline{T_{i}}|-2\epsilon n\geq(1-\frac{\sqrt{\epsilon}}{2})n-2\epsilon n\geq(1-{\sqrt{\epsilon}})n$. Thus $C$ is a directed absorbing cycle with parameters $(\frac{\delta}{3},\sqrt{\epsilon},\lambda,\lambda^{2})$. ∎
4 Stable case
In this section, by combining the results of the previous sections, we use the regularity-blow-up method in a digraph collection $\mathcal{D}$ to find a transversal directed Hamilton cycle when $\mathcal{D}$ is strongly stable or weakly stable. Let $10<\gamma,\alpha,\epsilon,\delta<$ 1. We say that $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable if it is either $(\gamma,\alpha)$-strongly stable or $(\epsilon,\delta)$-weakly table. We first show that the reduced digraph collection inherits stability.
Lemma 4.1.
Let $0<\frac{1}{n}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu,\alpha\ll\gamma,\epsilon\ll\delta\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. Let $\mathcal{R}=\mathcal{R}(\epsilon_{0},d,L_{0})$ be the reduced digraph collection of $\mathcal{G}$. If $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2},\epsilon,\frac{\delta}{2})$-stable.
Proof.
We may assume that $\frac{1}{n}\leq\frac{1}{n_{0}}$ where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is the constant from Lemma 2.2. Write $[L]$ for the common vertex set of digraphs in $\mathcal{R}$ where $L_{0}\leq L\leq n_{0}$, and $[M]$ for the set of color clusters. Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $[n]$ and a digraph collection $\mathcal{D}^{\prime}$ satisfying (i)-(v) of Lemma 2.2. Therefore, for $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}:=\{D_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular. Furthermore, $Lm\leq n\leq Mm+\epsilon_{0}n$ and $Mm\leq n\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$. Thus we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$. Given $X\subset[L]$, we write $\hat{X}:=\bigcup_{j\in X}V_{j}$. Hence $|\hat{X}|=m|X|$.
Claim 4.2.
Let $d\ll\alpha^{\prime}$ and $i\in[L]$. Assume that there are sets $A,B\subset[L]$ of size at least $(\frac{1}{2}-\alpha^{\prime})L$ such that $e_{R_{i}}(A,B)\leq\alpha^{\prime}L^{2}$. Then $e_{D_{c}}(\hat{A},\hat{B})\leq\sqrt{\alpha^{\prime}}n^{2}$ for all but at most $2\sqrt{\alpha^{\prime}}m$ colors $c\in\mathcal{C}_{i}$.
Proof of Claim 4.2.
Let $t$ be the number of colors $c\in\mathcal{C}_{i}$ such that $e_{D_{c}}(\hat{A},\hat{B})>\sqrt{\alpha^{\prime}}n^{2}$. Since $e(D_{c}^{\prime})-e(D_{c})\leq(3d+\epsilon_{0})n^{2}$ for all $c\in[n]$, we have
$\sqrt{\alpha^{\prime}}n^{2}t$ $\leq\sum_{c\in\mathcal{C}_{i}}e_{D_{c}}(\hat{A},\hat{B})\leq\sum_{c\in\mathcal{C}_{i}}\left(e_{D_{c}^{\prime}}(\hat{A},\hat{B})+(3d+\epsilon_{0})n^{2}\right)$
$\leq|\mathcal{C}_{i}|e_{R_{i}}(A,B)m^{2}+|\mathcal{C}_{i}|(3d+\epsilon_{0})n^{2}$
$\leq\alpha^{\prime}L^{2}m^{3}+4dmn^{2}\leq 2\alpha^{\prime}mn^{2}.$
This implies that $t\leq 2\sqrt{\alpha^{\prime}}m$. ∎
Case 1. $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable.
We prove that $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2})$-strongly stable. Suppose that there exists a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for all $i\in I$. By Definition 3.2, for each $i\in I$, there are two sets $A^{i},B^{i}\subseteq[L]$ of size at least $(\frac{1}{2}-\alpha^{2})L$ such that $e_{R_{i}}(A^{i},B^{i})<\alpha^{2}L^{2}$. Applying Claim 4.2 with $\alpha^{\prime}:=\alpha^{2}$, we get that there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i_{0}}$ such that $e_{D_{c}}(\hat{A^{i}},\hat{B^{i}})\leq\alpha n^{2}$. Notice that $|\hat{A^{i}}|,|\hat{B^{i}}|\geq(\frac{1}{2}-\alpha^{2})Lm>(\frac{1}{2}-\alpha)n$. Then there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i}$ such that $D_{c}$ is not $\alpha$-nice. Thus the number of colors $c\in[n]$ for which $D_{c}$ is not $\alpha$-nice is at least $(1-2\alpha)m|I|>(1-\gamma)n$ since $\alpha\ll\gamma$. This contradicts the fact that $\mathcal{D}$ is $(\gamma,\alpha)$-strong stable.
Case 2. $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable.
Suppose that $\mathcal{R}$ is not $(\frac{\gamma}{2},\alpha^{2})$-strongly stable. Then there is subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$. It suffices to show that $\mathcal{R}$ is $(\epsilon,\frac{\delta}{2})$-weakly stable. Choose $i\in I$, then $R_{i}$ is $\alpha^{2}$-extremal. Based on Lemma 2.4 (i), one has $d_{R_{i}}^{+}(j),d_{R_{i}}^{-}(j)\geq(\frac{1}{2}-2\mu)L$ for all but at most $d^{\frac{1}{4}}L$ vertices $j\in[L]$. Together with Lemma 3.3, we obtain that
[…]
Note that $|\hat{Z}_{i}^{\prime}|,|\hat{Y}_{i}^{\prime}|=(\frac{1}{2}-\alpha^{2})Lm\geq(\frac{1}{2}-2\alpha)n$. In view of Claim 4.2, we know that there is $\mathcal{B}_{i}\subseteq\mathcal{C}_{i}$ with $|\mathcal{C}_{i}\setminus\mathcal{B}_{i}|\leq 2\alpha m$ such that for all $c\in\mathcal{B}_{i}$, we have $e_{D_{c}}(\hat{Z}_{i}^{\prime},\hat{Y}_{i}^{\prime})\leq\alpha n^{2}$. It follows that $D_{c}$ is $2\alpha$-extremal for each $c\in\mathcal{B}_{i}$.
[…]
Choose $c\in\mathcal{B}_{i},c^{\prime}\in\mathcal{B}_{j}$ for some $i,j\in[L]$. Suppose that $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing. We must have $i\neq j$.
[…]
$|A_{i}^{\prime}\Delta A_{j}^{\prime}|m=|\hat{A}_{i}^{\prime}\Delta\hat{A}_{j}^{\prime}|\geq|A_{c}\Delta A_{c^{\prime}}|-|\hat{A}_{i}^{\prime}\Delta A_{c}|-|\hat{A}_{j}^{\prime}\Delta A_{c^{\prime}}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
[…]
$|{W_{i}^{3}}^{\prime}\Delta{W_{j}^{3}}^{\prime}|m=|{\hat{W_{i}^{3}}^{\prime}}\Delta{\hat{W_{j}^{3}}^{\prime}}|\geq|W_{c}^{3}\Delta W_{c^{\prime}}^{3}|-|{\hat{W_{i}^{3}}^{\prime}}\Delta W_{c}^{3}|-|{\hat{W_{j}^{3}}^{\prime}}\Delta W_{c^{\prime}}^{3}|\geq\delta n-21\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
and similarly, $|{W_{i}^{3}}^{\prime}\Delta{W_{j}^{1}}^{\prime}|\geq\frac{\delta L}{2}$. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
[…]
$|A_{i}^{\prime}\Delta{W_{j}^{3}}^{\prime}|m=|\hat{A}_{i}^{\prime}\Delta\hat{{W_{j}^{3}}^{\prime}}|\geq|A_{c}\Delta W_{c^{\prime}}^{3}|-|\hat{A}_{i}^{\prime}\Delta A_{c}|-|\hat{{W_{j}^{3}}^{\prime}}\Delta W_{c^{\prime}}^{3}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
Similarly, $|B_{i}^{\prime}\Delta{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
Since $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, we have $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$. The number of pairs $i,j\in[L]$ such that $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing is at least $\frac{\delta n^{2}}{m^{2}}\geq\frac{\delta L^{2}}{2}$. Thus $e(C_{\mathcal{R},k}^{\epsilon,\frac{\delta}{2}})\geq\frac{\delta L^{2}}{2}$, and hence $\mathcal{R}$ is $(\epsilon,\frac{\delta}{2})$-weakly stable. ∎
Lemma 4.3.
Let $0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. If $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Proof.
Choose additional parameters $n_{0},L_{0},\epsilon_{0},d,\theta,\beta,\lambda,\zeta$, where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is obtained from Lemma 2.2, and $\zeta$ is obtained from Lemma 3.3, such that
$1<\frac{1}{n}\ll\frac{1}{n_{0}}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu\ll\beta\ll\lambda\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1,$
and $\delta\ll\min\{\zeta,\frac{1}{2}-\zeta-\epsilon\}$, where the previous lemmas hold with suitable parameters.
Since $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, in view of Lemma 3.6 and Lemma 3.9, we know that $\mathcal{D}$ has a directed absorbing cycle $C$ with parameters $(\frac{\delta}{3},\sqrt{\epsilon},\lambda,\lambda^{2})$. Fix any two colors $c,c^{\prime}\in[n]$, two vertices $x,y\in V$ and ${\rm{K\in\{I,II\}}}$, we say the triple $(c,x,y)$ is absorbable if there are at least $\lambda^{2}n$ disjoint Type-K directed $c$-absorbing paths of $(x,y)$ inside $C$; we say the pair $(c,x)$ is absorbable if there are at least $\lambda^{2}n$ disjoint Type-K directed $c$-absorbing paths of $(x,x)$ inside $C$; we say $(c,c^{\prime},x,y)$ is totally absorbable if $(c,x),\ (c^{\prime},y)$ and $(c,x,y)$ are all absorbable.
By Definition 3.4, $C$ has length at most $\lambda n$, and there exists a color set $\mathcal{C}\subseteq[n]\setminus{\rm col}(C)$ of size at least $\frac{\delta n}{3}$ such that
  1. (a)
    given any color $c\in\mathcal{C}$ and any $D_{c}$-good vertex $v$, the triple $(c,v,u)$ is absorbable for all but at most $\sqrt{\epsilon}n$ vertices $u$,
  2. (b)
    given any color $c\in\mathcal{C}$, for all but at most $\sqrt{\epsilon}n$ $D_{c}$-good vertices $v$, the pair $(c,v)$ is absorbable.
Claim 4.4.
There is an integer $r\in[(2-2^{-10})\beta n,2\beta n]$ and for each $i\in[r]$, disjoint vertex pairs $(v_{i},v_{i}^{\prime})\in V^{2}$ and disjoint color pairs $(c_{i},c_{i}^{\prime})\in\mathcal{C}^{2}$ such that the family $\mathcal{Q}:=\{(c_{i},c_{i}^{\prime},v_{i},v_{i}^{\prime}):i\in[r]\}$ has the following properties:
  1. (i)
    $(c_{i},c_{i}^{\prime},v_{i},v_{i}^{\prime})$ is totally absorbable for all $i\in[r]$,
  2. (ii)
    for every pair $(u_{1},u_{2})\in V^{2}$ and $c\in[n]$, there are at least $2^{-9}\beta n$ values $i\in[r]$ such that $c\in L({v_{i}u_{1}})$ and $c_{i}^{\prime}\in L({u_{2}v_{i}^{\prime}})$ (see Figure 4).
[…]
Figure 4: Absorbing.
Proof of Claim 4.4.
For every color pair $(b_{1},b_{2})\in\mathcal{C}^{2}$, every vertex pair $(u_{1},u_{2})\in V^{2}$ and color $c\in\mathcal{C}$, let $S(b_{1},b_{2},u_{1},u_{2},c)$ be the collection of pairs $(v_{1},v_{2})\in V^{2}$ such that $c\in L({v_{1}u_{1}})$, $b_{2}\in L({u_{2}v_{2}})$ and $(b_{1},b_{2},v_{1},v_{2})$ is totally absorbable. We will prove that $|S(b_{1},b_{2},u_{1},u_{2},c)|\geq 2^{-4}n^{2}$.
[…]
Let $\{c_{i},c_{i}^{\prime}:i\in[2\beta n]\}\subseteq\mathcal{C}$ be a collection of distinct colors. For each $i\in[2\beta n]$, define $\mathcal{F}_{i}$ to be a directed $2$-graph with vertex set $V$ and edge set $\{(v_{1},v_{2})\in V^{2}:(c_{i},c_{i}^{\prime},v_{1},v_{2})\ \text{is totally absorbable}\}$ and $\mathbf{H}:=\{\mathcal{F}_{i}:i\in[2\beta n]\}$. Define the collection of directed multi-graphs:
$\mathbf{Z}:=\left\{S(u_{1},u_{2},c):=\bigcup_{i\in[2\beta n]}S(c_{i},c_{i}^{\prime},u_{1},u_{2},c):(u_{1},u_{2})\in V^{2},c\in[n]\right\}.$
For every $S:=S(u_{1},u_{2},c)\in\mathbf{Z}$ and every $i\in[2\beta n]$, we have $|E(S)\cap E(\mathcal{F}_{i})|\geq 2^{-4}n^{2}$. Applying Lemma 3.1 with $t=2\beta n$ and $\epsilon=2^{-4}$, we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(2-2^{-10})\beta n$ (and at most $2\beta n$) and $|E(S)\cap E(M)|\geq 2^{-9}\beta n$ for all $S\in\mathbf{Z}$. ∎
Let
$V_{\rm abs}:=\bigcup_{i\in[r]}\{v_{i},v_{i}^{\prime}\},\ \ U:=V\setminus(V(C)\cup V_{\rm abs}),\ \ \mathcal{C}_{\rm abs}:=\bigcup_{i\in[r]}\{c_{i},c_{i}^{\prime}\},$
$\mathcal{C}_{\rm rem}=[n]\setminus({\rm col}(C)\cup\mathcal{C}_{\rm abs}),\ \ \mathcal{J}:=\{J_{i}:i\in\mathcal{C}_{\rm rem}\}\ \text{where}\ J_{i}=D_{i}[U].$
Thus, $|[n]\setminus\mathcal{C}_{\rm rem}|=|V\setminus U|\leq\lambda n+4\beta n<2\lambda n$ since $\beta\ll\lambda$, and $\delta^{0}(\mathcal{J})\geq(\frac{1}{2}-\mu-2\lambda)n\geq(\frac{1}{2}-3\lambda)n$.
Since $\lambda\ll\gamma,\alpha,\epsilon,\delta$ and $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, it is easy to see that $\mathcal{J}$ is $(\frac{\gamma}{2},\frac{\alpha}{2},2\epsilon,\frac{\delta}{2})$-stable. Apply Lemma 2.2 to $\mathcal{J}$ with parameters $(\epsilon_{0},1,d,L_{0})$. Let $\mathcal{R}$ be the reduced digraph collection of $\mathcal{J}$. Write $[L]$ for the common vertex set of digraphs in $\mathcal{R}$ where $L_{0}\leq L\leq n_{0}$, and $[M]$ for the set of color clusters. Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $U$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}_{\rm rem}$ and a digraph collection $\mathcal{J}^{\prime}$ satisfying (i)-(v) of Lemma 2.2. Therefore, for $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have that ${hi}\in R_{j}$ if and only if $\mathcal{J}_{hi,j}^{\prime}:=\{J_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular. Furthermore, $Lm\leq|U|\leq Mm+\epsilon_{0}n$ and $Mm\leq|\mathcal{C}_{\rm rem}|\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$. Thus we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$. Based on Lemma 2.4, we obtain that, for each vertex $i\in[L]$, there are at least $(1-d^{1/4})M$ colors $j\in[M]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$; for each color $j\in[M]$, there are at least $(1-d^{1/4})L$ vertices $i\in[L]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$.
It follows from Lemma 4.1 that $\mathcal{R}$ is $(\frac{\gamma}{4},\frac{\alpha^{2}}{4},2\epsilon,\frac{\delta}{4})$-stable. It is routine to check that the bipartite graph collection corresponding to $\mathcal{R}$, say $\mathcal{B_{R}}$, is $(\frac{\gamma}{4},\frac{\alpha^{2}}{4},2\epsilon,\frac{\delta}{4})$-stable. Together with Theorem 2.10, there is a transversal perfect matching inside $\mathcal{B_{R}}$. Therefore, there exists a set of disjoint rainbow directed cycles inside $\mathcal{R}$, say $C^{1},C^{2},\ldots,C^{t}$, such that $|V(C^{1})|+\cdots+|V(C^{t})|=L$. For each rainbow directed cycle $C^{i}$ in $\mathcal{R}$, by using colors in $\bigcup_{j\in{\rm col}(C^{i})}\mathcal{C}_{j}$, we are to find a rainbow directed path $P^{i}$ in $\mathcal{J}^{\prime}$ with length at least $(1-\sqrt{\epsilon_{0}})|C^{i}|m$.
Claim 4.5.
Let $C=v_{1}v_{2}\ldots v_{s}v_{1}$ be a rainbow directed cycle in $\mathcal{R}$ with $v_{s+1}:=v_{1}$ and $v_{i}v_{i+1}\in E(R_{i})$ for each $i\in[s]$. Then $\{J_{i}^{\prime}[V_{1},V_{2},\cdots,V_{s}]:i\in\bigcup_{j\in[s]}\mathcal{C}_{j}\}$ contains a rainbow directed path with order at least $(1-\sqrt{\epsilon_{0}})sm$.
Proof of Claim 4.5.
We construct a rainbow directed path greedily. Assume that we already have a rainbow directed path $P=x_{1}\ldots x_{i}$ with $x_{j}\in V_{j_{0}}$ and $j_{0}\equiv j\pmod{s}$ for each $j\in[i]$. If $i\geq(1-\sqrt{\epsilon_{0}})sm$, then we are done. Otherwise, $|V_{j}\setminus V(P)|\geq\sqrt{\epsilon_{0}}m-1$ for each $j\in[s]$. We claim that $\max\{|N_{J^{\prime}_{c}}^{+}(x_{i})\cap(V_{i_{0}+1}\setminus V(P))|:c\in\mathcal{C}_{i_{0}}\setminus{\rm col}(P)\}\geq 2\epsilon_{0}m$. If so, then choose $c\in\mathcal{C}_{i_{0}}\setminus{\rm col}(P)$ such that $|N_{J^{\prime}_{c}}^{+}(x_{i})\cap(V_{i_{0}+1}\setminus V(P))|$ is maximum and choose $x_{i+1}\in N_{J_{c}^{\prime}}^{+}(x_{i})\cap(V_{i_{0}+1}\setminus V(P))$ such that $\max\{|N_{J_{c}^{\prime}}^{+}(x_{i+1})\cap(V_{i_{0}+2}\setminus V(P))|:c\in\mathcal{C}_{i_{0}+1}\setminus{\rm col}(P)\}$ is maximum.
Suppose that $\max\{|N_{J_{c}^{\prime}}^{+}(x_{i})\cap(V_{i_{0}+1}\setminus V(P))|:c\in\mathcal{C}_{i_{0}}\setminus{\rm col}(P)\}<2\epsilon_{0}m$. Clearly, $i\geq 2$. Let $W_{i_{0}}:=N_{J^{\prime}_{c(x_{i-1}x_{i})}}^{+}(x_{i-1})\cap(V_{i_{0}}\setminus V(P))$ and $W_{i_{0}+1}$ be a subset of $V_{i_{0}+1}\setminus V(P)$ with $|W_{i_{0}+1}|\geq\frac{\sqrt{\epsilon_{0}}}{2}m$. By the choice of each $x_{j}$ with $j\leq i$, we have $|W_{i_{0}}|\geq 2\epsilon_{0}m$ and each $y\in W_{i_{0}}$ satisfies $|N_{J_{c}^{\prime}}^{+}(y)\cap(V_{i_{0}+1}\setminus V(P))|<2\epsilon_{0}m$ for each color $c\in\mathcal{C}_{i_{0}}\setminus{\rm col}(P)$. Then
$\frac{\sum_{c\in\mathcal{C}_{i_{0}}\setminus{\rm col}(P)}e_{J^{\prime}_{c}}(W_{i_{0}},W_{i_{0}+1})}{|\mathcal{C}_{i_{0}}\setminus{\rm col}(P)||W_{i_{0}}||W_{i_{0}+1}|}\leq 2\sqrt{\epsilon_{0}}\ll d,$
which contradicts the fact that $\{J^{\prime}_{i}[V_{i_{0}},V_{i_{0}+1}]:i\in\mathcal{C}_{i_{0}}\}$ is $(\epsilon_{0},d)$-regular. ∎
Based on Claim 4.5, we know that there exists a set of disjoint rainbow directed paths $P^{i}$ in $\mathcal{J}$ with order at least $(1-\sqrt{\epsilon_{0}})|C^{i}|m$ and colors from $\bigcup_{j\in{\rm col}(C^{i})}\mathcal{C}_{j}$ for all $i\in[t]$. The number of vertices or colors not in any $P^{i}$ is at most $\sqrt{\epsilon_{0}}n+3\epsilon_{0}n\leq 2\sqrt{\epsilon_{0}}n$. We consider each vertex in $U$ but not in any $P^{i}$ to be a single path (of length 0).
Without loss of generality, assume that $P^{1},P^{2},\ldots,P^{s}$ are all the disjoint rainbow directed paths obtained in the above. Then $s\leq 2\sqrt{\epsilon_{0}}n+L<2^{-10}\beta n$. Let $x_{i},y_{i}$ be the start vertex and end vertex of $P^{i}$ for each $i\in[s]$ (where $x_{i}=y_{i}$ if $P^{i}$ has length 0). Clearly, $U=\bigcup_{i\in[s]}V(P^{i})$. It follows that $|\mathcal{C}_{\rm rem}\setminus{\rm col}(\bigcup_{i\in[s]}P^{i})|=s$.
Next, we do the following operation for each $i\in[s]$ in turn. Let $a_{i}\in\mathcal{C}_{\rm rem}$ be an arbitrary unused color. Choose an unused 4-tuple $Q_{i}:=(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})\in\mathcal{Q}$ where $j_{i}\in[r]$, $a_{i}\in L({v_{j_{i}}x_{i}})$ and $c_{j_{i}}^{\prime}\in L({y_{i}v_{j_{i}}^{\prime}})$. This is feasible since Claim 4.4 implies that there are at least $2^{-9}\beta n$ choices for $Q_{i}$, of which at most $s\leq 2^{-10}\beta n$ have been used. Now, $(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})$ is totally absorbable, so there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j_{i}}$-absorbing path of $(v_{j_{i}},v_{j_{i}}^{\prime})$ inside ${C}$. Then choose one of them, say $S_{i}:=x_{1}^{i}x_{2}^{i}x_{3}^{i}x_{4}^{i}$, whose vertices have not been previously chosen since $s<2^{-10}\beta n<2^{-10}\lambda^{2}n$, and whose colors are $b_{1}^{i},b_{2}^{i},b_{3}^{i}$ in order.
Notice that there remains $I\subseteq[r]$ such that the $(c_{j},c_{j}^{\prime},v_{j},v_{j}^{\prime})\in\mathcal{Q}$ with $j\in I$ are precisely the $4$-tuples which were not chosen to be some $Q_{i}$. For each one, there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j}$-absorbing paths of $(v_{j},v_{j})$ and disjoint rainbow directed $c_{j}^{\prime}$-absorbing paths of $(v_{j}^{\prime},v_{j}^{\prime})$ inside $C$. Since $\beta<\lambda^{2}$, there are two such rainbow directed rainbow paths $T_{i},T_{i}^{\prime}$ for each $(c_{j},v_{j})$ and $(c_{j}^{\prime},v_{j}^{\prime})$, which are disjoint and whose vertices have not previously been chosen.
At the end of this process, we have a collection $\{S_{i}:i\in[s]\},\ \{T_{i}:i\in I\},\ \{T_{i}^{\prime}:i\in I\}$ of disjoint rainbow directed paths inside $C$ satisfying
  • for each $i\in[s]$, $S_{i}$ is a directed $c_{j_{i}}$-absorbing path of $(v_{j_{i}},v_{j_{i}}^{\prime})$,
  • for each $i\in I$, $T_{i}$ is a directed $c_{i}$-absorbing path of $(v_{i},v_{i})$ and $T_{i}^{\prime}$ is a $c_{i}^{\prime}$-directed absorbing path of $(v_{i}^{\prime},v_{i}^{\prime})$.
For each $i\in[s]$, we replace $S_{i}$ by $x_{1}^{i}x_{2}^{i}v_{j_{i}}x_{i}P^{i}y_{i}v_{j_{i}}^{\prime}x_{3}^{i}x_{4}^{i}$ with colors $b_{1}^{i},c_{j_{i}},a_{i}$ followed by the colors inherited from $P^{i}$, followed by $c_{j_{i}}^{\prime},b_{2}^{i},b_{3}^{i}$. That is, we have replaced $S_{i}$ by a directed rainbow path with the same endpoints, vertices $V(S_{i})\cup V(P^{i})\cup\{v_{j_{i}},v_{j_{i}}^{\prime}\}$, and colors ${\rm col}(S_{i})\cup{\rm col}(P^{i})\cup\{c_{j_{i}},c_{j_{i}}^{\prime},a_{i}\}$. For each $i\in I$, we replace $T_{i}:=y_{1}^{i}y_{2}^{i}y_{3}^{i}y_{4}^{i}$ by $y_{1}^{i}y_{2}^{i}v_{i}y_{3}^{i}y_{4}^{i}$ where colors are inherited except ${\rm col}(y_{2}^{i}v_{i})=c_{i}$ and ${\rm col}(v_{i}y_{3}^{i})={\rm col}(y_{2}^{i}y_{3}^{i})$. We do a similar replacement of $T_{i}^{\prime}$, using new vertex $v_{i}^{\prime}$ and new color $c_{i}^{\prime}$. So we have replaced $T_{i}$ (resp. $T_{i}^{\prime}$) by a rainbow directed path with the same endpoints, vertices $V(T_{i})\cup\{v_{i}\}$ (resp. $V(T_{i}^{\prime})\cup\{v_{i}^{\prime}\}$) and colors ${\rm col}(T_{i})\cup\{c_{i}\}$ (resp. ${\rm col}(T_{i}^{\prime})\cup\{c_{i}^{\prime}\}$). Thus we have obtained a directed rainbow cycle using vertices $V(C)\cup\bigcup_{i\in[s]}V(P^{i})\cup\{v_{i},v_{i}^{\prime}:i\in[r]\}=V$ and colors ${\rm col}(C)\cup\bigcup_{i\in[s]}{\rm col}(P^{i})\cup\{c_{i},c_{i}^{\prime}:i\in[r]\}=[n]$ where each color is used at most once (and hence exactly once). That is, $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired. ∎
5 Extremal case
Choose constants $\mu,\alpha,\gamma,\epsilon,\delta$ such that
$0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1\ \text{and}\ \delta\ll\min\left\{\zeta,\frac{1}{2}-\zeta-\epsilon\right\},$
where $\zeta$ is defined in Lemma 3.3. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. By Lemma 4.3, we know $\mathcal{D}$ is not $(\gamma,\alpha)$-strongly stable. Hence there are at least $(1-\gamma)n$ digraphs in $\mathcal{D}$ such that each of them is $\alpha$-extremal. Without loss of generality, assume that $D_{i}$ is $\alpha$-extremal for each $i\in[(1-\gamma)n]$. Note that every $\alpha$-extremal digraph is also $\epsilon$-extremal, since $\alpha\ll\epsilon$. Denote
[…]
Hence $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|+|\tilde{\mathcal{C}}_{3}|\geq(1-\gamma)n$. By Lemma 3.3, we know that $D_{i}$ has a characteristic partition $(\tilde{A_{i}},\tilde{B_{i}},\tilde{L_{i}})$ for each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, or a characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each $i\in\tilde{\mathcal{C}}_{3}$.
Under the above characteristic partition, applying Lemma 4.3 yields that $\mathcal{D}$ is not $(\epsilon,\delta)$-weakly stable. For $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{1}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. For $i\in\tilde{\mathcal{C}}_{3}$, define $I^{2}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. For $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{3}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then we have
$\sum_{i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}}|I^{1}_{i}|<2\delta n^{2},\ \sum_{i\in\tilde{\mathcal{C}}_{3}}|I^{2}_{i}|<2\delta n^{2},\,\,\text{and}\,\,\sum_{i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}}|I^{3}_{i}|<\delta n^{2}.$
Hence there exists a color set, say $I^{k}$ (where $I^{k}\subseteq\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ if $k=1$ and $I^{k}\subseteq\tilde{\mathcal{C}}_{3}$ if $k\in\{2,3\}$), with size at most $\sqrt{\delta}n$ such that the following hold:
  1. (B1)
    if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\Delta\tilde{A}_{i}|<\delta n$ and $|\tilde{B}_{1}\Delta\tilde{B}_{i}|<\delta n$ for all $i\in(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$,
  2. (B2)
    if $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels, for all $k\in[4]$ and all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{2}$ we have $|W_{1}^{k}\Delta W_{i}^{k}|<\delta n$, in particular $|C_{1}^{k}\Delta C_{i}^{k}|<2\delta n$,
  3. (B3)
    if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\Delta W_{i}^{1}|<\delta n$ and either $|\tilde{A}_{1}\Delta W_{i}^{2}|<\delta n$ or $|\tilde{A}_{1}\Delta W_{i}^{4}|<\delta n$ for all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$.
Note that if ${\bf(B3)}$ holds, then for each $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$, there is an $i_{0}\in\{2,4\}$ such that
$2\delta n>|\tilde{A}_{1}\Delta W_{i}^{1}|+|\tilde{A}_{1}\Delta W_{i}^{i_{0}}|\geq|W_{i}^{1}\Delta W_{i}^{i_{0}}|>2\delta n,$
a contradiction. This implies that exactly one of $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$ holds. Based on (B1)-(B2), we proceed our proof by considering the following two theorems.
Theorem 5.1.
Assume that $\delta$ is defined in (5). Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $|\tilde{\mathcal{C}}_{3}|<2\sqrt{\delta}n$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Theorem 5.2.
Assume that $\delta$ is defined in (5). Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
In the following of this section, we prove the above two theorems respectively.
5.1 Proof of Theorem 5.1
In this subsection, we give the proof of Theorem 5.1.
Proof of Theorem 5.1.
Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. Let $\epsilon,\delta$ be defined in (5), and choose $\eta$ with $\delta\ll\eta\ll 1$. Denote $[m]:=(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$ and $\mathcal{C}_{\rm bad}:=[m+1,n]$. By adding colors to $\mathcal{C}_{\rm bad}$ if necessary we may assume $m=(1-4\sqrt{\delta})n$. Recall that $\epsilon\ll\delta$. Hence for each $i\in[2,m]$, there exists a new partition $(A_{i},B_{i},L_{i})$ of $D_{i}$ such that
  • (C1)
[…]
  • (C2)
[…]
  • (C3)
    subject to (C1)-(C2), each vertex in $V$ lies in as many as possible $A_{i}\cup B_{i}$ for $i\in[2,m]$.
Denote $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$. Expand $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$. For simplicity, we assume $Y\in\{A,B\}$ and $Z\in\{A,B\}\setminus\{Y\}$ in the following. Let $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ be a partition of $[m]$ with
[…]
Hence for $i\in[2,m]$, we have $|Y_{1}\Delta Y_{i}|\leq|\tilde{Y_{1}}\Delta\tilde{Y_{i}}|+|\tilde{L}_{i}|\leq\delta n+2{\epsilon}n<2\delta n$. Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise. Clearly, $|\hat{\mathcal{C}}|\geq(1-4\sqrt{\delta}-\eta)n$. Define
$X:=$$\{x\in V:x\not\in A_{i}\cup B_{i}\leavevmode\nobreak\ \textrm{forat least}\leavevmode\nobreak\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in\hat{\mathcal{C}}\},$
$X_{Y}:=$$\{x\in Y\setminus X:x\not\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 10\sqrt{\delta}|\hat{\mathcal{C}}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in\hat{\mathcal{C}}\}.$
Recall that $|L_{i}|\leq 2{\epsilon}n$ and $|Y_{1}\Delta Y_{i}|<2\delta n$ for each $i\in\hat{\mathcal{C}}$. Hence $6\sqrt{\delta}|\hat{\mathcal{C}}||X|\leq 2{\epsilon}n|\hat{\mathcal{C}}|$ and
$10\sqrt{\delta}|\hat{\mathcal{C}}||X_{Y}|\leq\sum_{i\in\hat{\mathcal{C}}}|Y\setminus Y_{i}|\leq\sum_{i\in\hat{\mathcal{C}}}(|Y\setminus Y_{1}|+|Y_{1}\Delta Y_{i}|)\leq(2\delta+2{\epsilon})n|\hat{\mathcal{C}}|.$
It follows that $|X|<\frac{1}{3}\sqrt{\epsilon}n$ and $|X_{Y}|<\frac{1}{4}\sqrt{\delta}n$. For $k\in[2]$, define
$X_{Y}^{k}=\left\{\begin{array}[]{ll}\{x\in X_{Y}:x\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 3\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{k}}\},&\textrm{if $|\mathcal{C}_{k}|\geq\eta n$,}\\[5.0pt] \emptyset,&\textrm{otherwise.}\end{array}\right.$
For a vertex $x\in X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})$, we know $x\in Y_{i}$ for at most $6\sqrt{\delta}n$ colors $i\in\hat{\mathcal{C}}$ and $x\in A_{i}\cup B_{i}$ for at least $(1-6\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Thus, $x\in Y\setminus Y_{i}$ for at least $|\hat{\mathcal{C}}|-6\sqrt{\delta}n$ colors $i\in\hat{\mathcal{C}}$ and $x\in Z_{i}$ for at least $(1-6\sqrt{\delta})|\hat{\mathcal{C}}|-6\sqrt{\delta}n\geq(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Hence
$|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|(|\hat{\mathcal{C}}|-6\sqrt{\delta}n)\leq\sum_{i\in\hat{\mathcal{C}}}|Y\setminus Y_{i}|\leq(2\delta+2\epsilon)n|\hat{\mathcal{C}}|.$
It follows that $|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|<4\delta n$. Move vertices in $X_{A}\setminus(\bigcup_{k\in[2]}X_{A}^{k})$ to $B$ and vertices in $X_{B}\setminus(\bigcup_{k\in[2]}X_{B}^{k})$ to $A$.
Let $X^{\prime}$ be a subset of $X$ consisting of vertices $x$ such that $d_{i}^{+}(x,Y_{i})\geq(1-3\sqrt{\delta})n$ (resp. $d_{i}^{-}(x,Y_{i})\geq(1-3\sqrt{\delta})n$) for at least $(1-3\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$ and $d_{i}^{-}(x,Z_{i})\geq\frac{5}{2}\sqrt{\delta}n$ (resp. $d_{i}^{+}(x,Z_{i})\geq\frac{5}{2}\sqrt{\delta}n$) for at least $3\sqrt{\delta}|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Delete vertices in $X^{\prime}$ from $A\cup B$. Without loss of generality, assume that $|B|-|A|=r$ with $0\leq r\leq 8{\delta}n+\sqrt{\epsilon}n$. Define $V_{\rm bad}:=\bigcup_{k\in[2]}(X^{k}_{A}\cup X^{k}_{B})\cup(X\setminus X^{\prime})$. Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+\frac{1}{2}\sqrt{\delta})n$. Moreover, each vertex in $Y\setminus V_{\rm bad}$ lies in $Y_{i}$ for at least $(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
The subsequent result can be used to connect two disjoint short rainbow directed paths into a single short rainbow directed path.
Claim 5.3 (Connecting tool).
Assume $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ are two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s},v_{1}\notin V_{\rm bad}\cup X^{\prime}$ and $s+t\leq 5\eta n$.
  1. (i)
    If $u_{s}\in Y$, $v_{1}\in Z$ and $|\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, then there are three colors $c_{1},c_{2},c_{3}\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$ and two vertices $w_{1}\in Z\setminus(V(P\cup Q)\cup V_{\rm bad})$, $w_{1}^{\prime}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}w_{1}^{\prime}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2},c_{3}\}$.
  2. (ii)
    If $u_{s},v_{1}\in Y$ and $|\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Z\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
  3. (iii)
    If $u_{s},v_{1}\in Y$ and $|\mathcal{C}_{1}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{1}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
Proof of Claim 5.3.
We only give the proof of (i), the other two statements can be proved by similar discussions, whose procedures are omitted.
Note that $u_{s},v_{1}\not\in V_{\rm bad}\cup X^{\prime}$. Then $u_{s}$ (resp. $v_{1}$) lies in $Y_{i}$ (resp. $Z_{i}$) for at least $(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Since $|\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, we obtain that $u_{s}$ (resp. $v_{1}$) lies in $Y_{i}$ (resp. $Z_{i}$) for at least $\sqrt{\delta}n$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$. Hence there are two distinct colors $c_{1},c_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$ such that $u_{s}\in Y_{c_{1}}$ and $v_{1}\in Z_{c_{2}}$. It is routine to check that
$|N_{{c_{1}}}^{+}(u_{s})\cap(Z\setminus(V_{\rm bad}\cup X^{\prime}))|\geq$ $|N_{{c_{1}}}^{+}(u_{s})\cap Z_{c_{1}}|-|Z_{1}\Delta Z_{c_{1}}|-|X\cup X_{A}\cup X_{B}|$
$\geq$ $\left(\frac{1}{2}-3\sqrt{\delta}\right)n-\left(2\delta+\sqrt{\epsilon}+\frac{1}{2}\sqrt{\delta}\right)n\geq\left(\frac{1}{2}-4\sqrt{\delta}\right)n.$
Hence there exists a vertex $w_{1}\in N_{{c_{1}}}^{+}(u_{s})\cap Z$ that avoids $V(P\cup Q)\cup V_{\rm bad}\cup X^{\prime}$. Similarly, since $w_{1}\notin V_{\rm bad}\cup X^{\prime}$, there is a color $c_{3}\in\mathcal{C}_{2}\setminus({\rm col}(P\cup Q)\cup\{c_{1},c_{2}\})$ such that $w_{1}\in B_{c_{3}}$ and
$|N_{{c_{2}}}^{-}(v_{1})\cap N_{{c_{3}}}^{+}(w_{1})\cap(Y\setminus(V_{\rm bad}\cup X^{\prime}))|\geq\left(\frac{1}{2}-10\sqrt{\delta}\right)n.$
Therefore, there is a vertex $w_{1}^{\prime}\in N_{{c_{2}}}^{-}(v_{1})\cap N_{{c_{3}}}^{+}(w_{1})\cap A$ that avoids $V(P\cup Q)\cup V_{\rm bad}\cup X^{\prime}$. It follows that $u_{1}Pu_{s}w_{1}w_{1}^{\prime}v_{1}Qv_{t}$ is a rainbow directed path inside $\mathcal{D}$ with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2},c_{3}\}$. ∎
Next, based on the sizes of $|\mathcal{C}_{1}|$ and $|\mathcal{C}_{2}|$, we distinguish our proof into the following three cases.
Case 1. $|\mathcal{C}_{1}|<\eta n$.
In this case, each vertex in $Y\setminus(X\cup X_{Y}^{2})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$. Recall that $|B|-|A|=r$ with $0\leq r\leq(8{\delta}+\sqrt{\epsilon})n$. The proof is divided into four steps.
Step 1. Balance the number of vertices in $A$ and $B$.
In this step, we claim that by moving vertices in $B\cap V_{\rm bad}$ or deleting rainbow directed paths inside $\mathcal{D}[B]$, we can make the number of remaining vertices in $B$ and $A$ differ by $\sigma$, where $\sigma=1$ if $|A|+|B|$ is odd and $\sigma=0$ otherwise. Denote $s:=|V_{\rm bad}\cap B|$. If $|B|-|A|-\sigma\leq 2s$, then move $\frac{|B|-|A|-\sigma}{2}$ vertices in $V_{\rm bad}\cap B$ to $A$, and our desired result holds. Hence, it suffices to consider $|B|-|A|-\sigma>2s$.
Now, we move all vertices in $V_{\rm bad}\cap B$ to $A$. Hence, $B\cap V_{\rm bad}=\emptyset$. Assume that $\{Q_{1},\ldots,Q_{t}\}$ is a set of disjoint maximal rainbow directed paths in $\mathcal{D}[B]$. If $|E(Q_{1})|+\cdots+|E(Q_{t})|\geq|B|-|A|-\sigma$, then there must exist a set of disjoint rainbow directed paths, say $\{Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}\}$, such that $|E(Q_{1}^{\prime})|+\cdots+|E(Q_{t^{\prime}}^{\prime})|=|B|-|A|-\sigma$. That is, $|A|-t^{\prime}=|B|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-\sigma$. Using Claim 5.3 (ii), we connect $Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}$ into a single rainbow directed path $P^{1}$ with length at most $27\delta n$, starting at $A$ and ending at $B$. (In fact, by using Claim 5.3 (ii), we only get a rainbow path with two endpoints in $B$. But we usually want to find a rainbow directed path with endpoints in different parts. Hence one may extend it by using an unused color of $\mathcal{C}_{2}$ and an unused vertex in $A$). Therefore, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|-\sigma$, as desired.
In what follows, it suffices to consider $|E(Q_{1})|+\cdots+|E(Q_{t})|<|B|-|A|-\sigma=r-2s-\sigma$. Now, we move all vertices of $Q_{1},\ldots,Q_{t}$ to $A$. Let $\mathcal{C}^{\prime}:=\mathcal{C}\setminus{\rm col}(\bigcup_{i\in[t]}Q_{i})$. Hence
$|A|$ $\leq\frac{n-|X^{\prime}|-r}{2}+s+2(r-2s-\sigma)\leq\frac{n}{2}+18{\delta}n,$
$|B|$ $\geq\frac{n-|X^{\prime}|+r}{2}-s-2(r-2s-\sigma)\geq\frac{n}{2}-18{\delta}n,$
$|\mathcal{C}^{\prime}|$ $=|\mathcal{C}|-(|E(Q_{1})|+\cdots+|E(Q_{t})|)\geq(1-9\delta)n.$
By the maximality of $\{Q_{1},\ldots,Q_{t}\}$, one has $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$. It follows from $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$ that $\left\lceil\frac{n}{2}\right\rceil\leq|A|+|X^{\prime}|\leq\frac{n}{2}+19{\delta}n$. Now, by Lemma 5.14 (which is presented in the final of this subsection), we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
In this step, if we only move vertices in $V_{\rm bad}$ from $B$ to $A$, then let $P^{1}$ be a null digraph (i.e., no vertices); otherwise, assume that $P^{1}$ is a rainbow directed path obtained in the above, whose length is at most $27\delta n$ and endpoints are $u_{1}\in A\setminus V_{\rm bad},\,v_{1}\in B\setminus V_{\rm bad}$. Hence $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|-\sigma$.
Step 2. Construct a series of disjoint rainbow directed rainbow $P_{3}$ such that all centers of them are exactly all vertices in $V_{\rm bad}\cup X^{\prime}$.
Recall that $|X_{A}^{2}\cup X_{B}^{2}|\leq|X_{A}\cup X_{B}|<\frac{1}{2}\sqrt{\delta}n$. Let $v\in X_{Y}^{2}$. If $v\in Z$, then $v$ is moved from $B$ to $A$ in Step 1, so $v\in X_{B}$. Therefore, $v\in A_{i}\cup B_{i}$ for at least $(1-6\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$ and $v\not\in B_{i}$ for at least $10\sqrt{\delta}|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$. Hence, $v\in A_{i}$ for at least $4\sqrt{\delta}|\mathcal{C}_{2}|-27\delta n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$. If $v\in Y$, then $v$ belongs to $Y_{i}$ for at least $(3\sqrt{\delta}-27\delta)n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$. Hence, using colors in $\mathcal{C}_{2}\setminus{\rm col}(P^{1})$, we can greedily choose $|X_{A}^{2}\cup X_{B}^{2}|$ disjoint rainbow directed $P_{3}$, such that their centers correspond to all vertices in $X_{A}^{2}\cup X_{B}^{2}$. Moreover, if $v\in(X_{A}^{2}\cup X_{B}^{2})\cap Y$, then the rainbow directed $P_{3}$ with center $v$ has endpoints in $Z\setminus(V_{\rm bad}\cup X^{\prime}\cup V(P^{1}))$.
For vertices in $X^{\prime}$, we know from its definition that by avoiding vertices and colors in $P^{1}$ and all rainbow directed paths chosen for vertices in $X_{A}^{2}\cup X_{B}^{2}$, there exists a set of disjoint rainbow directed $P_{3}$ with centers in $X^{\prime}$ and endpoints in different parts.
By using Claim 5.3 (i)-(ii), we connect $P^{1}$, all rainbow directed $P_{3}$ with centers in $X_{A}^{2}\cup X_{B}^{2}\cup X^{\prime}$ (as obtained above) into a single rainbow directed path. This results in a rainbow directed path $P^{2}=u_{1}\ldots v_{2}$ with $v_{2}\in B\setminus V_{\rm bad}$ and length at most
$|E(P^{1})|+4|X_{A}^{2}\cup X_{B}^{2}|+5|X^{\prime}|+6\leq 27\delta n+2\sqrt{\delta}n+5\sqrt{\epsilon}n+6.$
For vertices in $X\setminus X^{\prime}$, we consider the following claim. Notice that the position (in $A$ or $B$) of vertices in $X\setminus X^{\prime}$ does not affect our proof, so if $v\in X$ is a vertex moved from $B$ to $A$, then we only need consider $v\in A$.
Claim 5.4.
Assume $X\setminus V(P^{2})=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. If $x_{i}\in X\cap Y$, then there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{2}\setminus({\rm col}(P^{2})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},x_{i}^{4}\in Z\setminus(V(P^{2})\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$.
Proof of Claim 5.4.
Suppose there exists an $i_{0}\in[s]$ such that Claim 5.4 holds for all $i\in[i_{0}-1]$ but does not hold for $i_{0}$. Assume for convenience that $x_{i_{0}}\in X\cap A$. Denote $\mathcal{C}_{2}^{\prime}:=\mathcal{C}_{2}\setminus({\rm col}(P^{2})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i_{0}-1]\})$. Let
$\mathcal{C}_{2}^{1}:=\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{+}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P^{2})|\}\ \text{and}$
$\mathcal{C}_{2}^{2}:=\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{-}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P^{2})|\}.$
Hence either $|\mathcal{C}_{2}^{1}|\leq 3$ or $|\mathcal{C}_{2}^{2}|\leq 3$. Without loss of generality, assume that $|\mathcal{C}_{2}^{1}|\leq 3$. Thus, for each $j\in\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}$, we have
$d_{j}^{+}(x_{i_{0}},A_{j})$ $\geq d_{j}^{+}(x_{i_{0}},A)-|A\setminus A_{1}|-|A_{1}\Delta A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
$\geq\delta^{0}(D_{j})-d_{j}^{+}(x_{i_{0}},B)-|A\setminus A_{1}|-|A_{1}\Delta A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
$\geq\frac{n}{2}-\left(4(i_{0}-1)+4+|V(P^{2})|\right)-2{\epsilon}n-2\delta n-4\delta n-\sqrt{\epsilon}n$
$\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n.$
Furthermore, $|\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}|\geq(1-3\sqrt{\delta})|\mathcal{C}_{2}|$. Notice that $x_{i_{0}}\notin X^{\prime}$. Then, for at least $(1-3\sqrt{\delta})|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, $d_{j}^{-}(x_{i_{0}},B_{j})<\frac{5}{2}\sqrt{\delta}n$, and consequently, $d_{j}^{-}(x_{i_{0}},A_{j})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for the same set of colors. By (C3), we know that $x_{i_{0}}\in A_{j}\cup B_{j}$ for all but at most $6\sqrt{\delta}|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, which leads to $x_{i_{0}}\not\in X$, a contradiction. ∎
By Claim 5.4, for each $x_{i}\in(X\setminus X^{\prime})\cap Y$, there exists a rainbow directed path $Q_{x_{i}}=x_{i}^{1}x_{i}x_{i}^{2}$ with ${\rm col}(Q_{x_{i}})=\{c_{i}^{1},c_{i}^{2}\}\subseteq\mathcal{C}_{2}$ and $x_{i}^{1},x_{i}^{2}\in Z\setminus V_{\rm bad}$. Moreover, each color or endpoint in $Q_{x_{i}}$ can be replaced by an unused color or vertex (if needed). Using Claim 5.3 (i)-(ii), one may connect $P^{2}$ and all rainbow directed paths in $\{Q_{x_{i}}:x_{i}\in X\setminus X^{\prime}\}$ into a single rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{3}\in B\setminus V_{\rm bad}$, whose length is at most $1027\delta n+2\sqrt{\delta}n+5\sqrt{\epsilon}n+$ 10 .
Step 3. Select a rainbow matching with colors in $\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{3})$.
Choose a maximum rainbow matching in $\{D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]:i\in\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{3})\}$, say $\tilde{M}$. If there exists an edge with color $j$ in $\tilde{M}$ such that $D_{j}^{\pm}[A\setminus V(P^{3}\cup\tilde{M}),B\setminus V(P^{3}\cup\tilde{M})]$ contains no $2$-matching, then delete all such edges from $\tilde{M}$ and denote the resulting rainbow directed matching by $M$. Connect $P^{3}$ and all rainbow directed edges in $M$ by Claim 5.3 (i), we obtain a rainbow directed path $P^{4}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{4}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{3})|+16\sqrt{\delta}n\leq 18\sqrt{\delta}n+5\sqrt{\epsilon}n+27\delta n+10$.
[…]
Step 4. Construct two rainbow directed paths inside $\mathcal{D}[A\setminus V(P^{4})]$ and $\mathcal{D}[B\setminus V(P^{4})]$ by using unused colors in $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$.
It is straightforward to check that
$|B\setminus V(P^{4})|-|A\setminus V(P^{4})|=|B\setminus V(P^{1})|-|A\setminus V(P^{1})|=\sigma.$
By the construction of $P^{4}$, we know that if $w\in V(P^{4})$ is a vertex not in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X$, then it is possible to avoid $w$ when constructing $P^{4}$; if $i\in{\rm col}(P^{4})$ is a color not in ${\rm col}(P^{1}-A)\cup{\rm col}(M)$, then it is possible to avoid $i$ when constructing $P^{4}$. That is, we can assume that $w\not\in V(P^{4})$ unless $w\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i\not\in{\rm col}(P^{4})$ unless $i\in{\rm col}(P^{1}-A)\cup{\rm col}(M)$.
Furthermore, in the process of choosing rainbow directed paths inside $\mathcal{D}[B]$ (in Step 1), one may use colors from $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$ before other colors and avoid vertices in $X^{\prime}\cup V_{\rm bad}\cup V(M)$. This implies that either $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}\subseteq{\rm col}(P^{1})$ or ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$. Denote $\tilde{\mathcal{C}}:=(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P^{1})$.
Notice that for each color $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{4})$, every vertex $v\in A\setminus V(P^{4})$ has at least $(\frac{1}{2}-3\sqrt{\delta})n-|V(P^{4})|-|B\Delta B_{i}|$ out-neighbors (resp. in-neighbors) in $B\setminus V(P^{4})$ in $D_{i}$. Hence,
$\left|N_{i}^{+}(v)\cap N_{i}^{-}(v)\cap(B\setminus V(P^{4}))\right|\geq\left(\frac{1}{2}-45\sqrt{\delta}\right)n.$
Therefore, by replacing every pair of arcs with the same endpoints and opposite directions by an undirected edge, the digraph $D_{i}^{\pm}[A\setminus V(P^{4}),B\setminus V(P^{4})]$ can be viewed as an undirected graph $G_{i}$ with vertex set $(A\cup B)\setminus V(P^{4})$ and minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$.
Next, we give an application of the transversal blow-up lemma (see [9]) for embedding transversal undirected Hamilton paths inside very dense bipartite graph collections, which can be proved by minor modifications to the proof of [10, Lemma 6.1], and we omit the proof here.
Claim 5.5.
Assume $W,Z\in\{A,B\}$. Let $W^{*}\subseteq W\setminus V_{\rm bad}$ and $Z^{*}\subseteq Z\setminus V_{\rm bad}$, where $|W^{*}|,|Z^{*}|\geq\eta n$, $W^{*}\cap Z^{*}=\emptyset$ and $|W^{*}|-|Z^{*}|=t\in\{0,1\}$. Let $T^{*}=Z^{*}$ if $t=0$ and $T^{*}=W^{*}$ if $t=1$. Let $\mathcal{C^{*}}\subseteq\mathcal{C}$ satisfy $|\mathcal{C^{*}}|=|W^{*}|+|Z^{*}|-1$, where $\mathcal{C^{*}}\subseteq\mathcal{C}_{1}$ if $W=Z$ and $\mathcal{C^{*}}\subseteq\mathcal{C}_{2}$ if $W\neq Z$. Let $W^{-}\subseteq W^{*}$ and $T^{+}\subseteq T^{*}$ with $|W^{-}|,|T^{+}|\geq\frac{\eta n}{8}$. Then there is a transversal undirected Hamilton path in $\{G_{i}[W^{*},Z^{*}]:i\in\mathcal{C}^{*}\}$ starting at $W^{-}$ and ending at $T^{+}$.
Claim 5.6.
Let $Q=x_{1}\ldots x_{s}$ be a rainbow directed path inside $\mathcal{D}$ with $s\leq 19\sqrt{\delta}n$, $V_{\rm bad}\subseteq V(Q)$, $(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad})\setminus\tilde{\mathcal{C}}\subseteq{\rm col}(Q)$, $x_{1}\in A\setminus V_{\rm bad}$ and $x_{s}\in B\setminus V_{\rm bad}$. If $|\tilde{\mathcal{C}}|-||B\setminus V(Q)|-|A\setminus V(Q)||$ is a nonnegative even integer, then $\mathcal{D}$ has a transversal directed Hamilton cycle.
Proof of Claim 5.6.
Using colors in $\tilde{\mathcal{C}}$, one may greedily choose two disjoint rainbow directed paths $P^{5}$ and $P^{6}$ in $\mathcal{D}[A\setminus V(Q)]$ and $\mathcal{D}[B\setminus V(Q)]$ with lengths $\frac{|\tilde{\mathcal{C}}|-(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$ and $\frac{|\tilde{\mathcal{C}}|+(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$ respectively, whose endpoints are $u_{5},v_{5}\in A\setminus V_{\rm bad}$ and $u_{6},v_{6}\in B\setminus V_{\rm bad}$. (Note that if a rainbow path has length 00, then it contains one vertex.) Applying Claim 5.3 (i)-(ii) to connect $Q,P^{5}$ and $P^{6}$ in turn, we get a rainbow directed path $P$ with endpoints $x_{1}\in A\setminus V_{\rm bad}$ and $v_{6}\in B\setminus V_{\rm bad}$. Clearly, $|V(P)|\leq 19\sqrt{\delta}n+4\sqrt{\delta}n+\eta n+4$ and $|A\setminus V(P)|=|B\setminus V(P)|$.
Notice that $x_{1},v_{6}\not\in V_{\rm bad}$. Hence there exist $i_{1},i_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P)$ such that $x_{1}\in A_{i_{1}}$ and $v_{6}\in B_{i_{2}}$. Let $W^{*}:=A\setminus V(P),\,T^{*}:=B\setminus V(P)$, $\mathcal{C}^{*}:=\mathcal{C}\setminus({\rm col}(P)\cup\{i_{1},i_{2}\})$, $W^{-}:=N_{G_{i_{2}}}(v_{6})\cap W^{*}$ and $T^{+}:=N_{G_{i_{1}}}(x_{1})\cap T^{*}$. It is routine to check that $|W^{*}|=|T^{*}|$, $|\mathcal{C}^{*}|=|W^{*}|+|T^{*}|-1$,
$|W^{-}|$ $\geq d_{G_{i_{2}}}(v_{6},A)-|V(P)|\geq d_{G_{i_{2}}}(v_{6},A_{i_{2}})-|A_{1}\Delta A_{i_{2}}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|-|V(P)|$
$\geq\left(\frac{1}{2}-45\sqrt{\delta}\right)n-2\delta n-4{\delta}n-\sqrt{\epsilon}n-|V(P)|>\left(\frac{1}{2}-2\eta\right)n,$
and similarly $|T^{+}|>\left(\frac{1}{2}-2\eta\right)n$. Applying Claim 5.5 yields that $\{G_{i}[W^{*},T^{*}]:i\in\mathcal{C}^{*}\}$ contains a transversal undirected path $P^{\prime}$ starting at $v^{\prime}\in W^{-}$ and ending at $u^{\prime}\in T^{+}$. Thus, $x_{1}Pv_{6}v^{\prime}P^{\prime}u^{\prime}x_{1}$ is a transversal directed Hamilton cycle inside $\mathcal{D}$. ∎
If $n$ and $|\tilde{\mathcal{C}}|$ have the same parity, then in view of Claim 5.6, there exists a transversal directed Hamilton cycle inside $\mathcal{D}$, a contradiction. Hence, it suffices to consider that $n$ and $|\tilde{\mathcal{C}}|$ have the different parity. Therefore, either $|\tilde{\mathcal{C}}|-(|B\setminus V(P^{4})|-|A\setminus V(P^{4})|)$ is a positive odd integer or $|\tilde{\mathcal{C}}|=0$ and $n$ is odd. Based on the definition of $X^{\prime}$, move vertices in $X^{\prime}$ to $Y$ if it is covered by a rainbow directed path from $Z$ to $Y$ (in Step 2).
Claim 5.7.
  1. (i)
    If $|\tilde{\mathcal{C}}|=0$, then $D_{i}[B\setminus(X^{\prime}\cup V(P^{1}))]=\emptyset$ for all $i\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$.
  2. (ii)
    For all $i\in\tilde{\mathcal{C}}$, $D_{i}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]=\emptyset$ and $D_{i}[(B\setminus(V_{\rm bad}\cup X^{\prime}\cup V(P^{1})),A]=\emptyset$.
Proof of Claim 5.7.
(i) Suppose that there exists an $i_{0}\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$ such that $D_{i_{0}}[B\setminus(X^{\prime}\cup V(P^{1}))]\neq\emptyset$. Choose ${w_{1}w_{2}}\in E(D_{i_{0}}[B\setminus(X^{\prime}\cup V(P^{1}))])$. Recall that each of $w_{1}$ and $w_{2}$ can be chosen outside $V(P^{4})$ unless it is in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i_{0}$ can be chosen outside ${\rm col}(P^{4})$ unless $i_{0}\in{\rm col}(M)$. If $i_{0}\in{\rm col}(M)$, then let $\tilde{P}^{4}$ be a rainbow directed path obtained by deleting the edge with color $i_{0}$ from $P^{4}$ and connecting its two endpoints by Claim 5.3 (i). It is routine to check that $|B\setminus V(\tilde{P}^{4})|-|A\setminus V(\tilde{P}^{4})|=|B\setminus V({P}^{4})|-|A\setminus V({P}^{4})|$. Let $P^{4}:=\tilde{P}^{4}$ if $i_{0}\in{\rm col}(M)$. Therefore, it suffices to consider $i_{0}\not\in{\rm col}(P^{4})$.
$\bullet$ $w_{1},w_{2}\not\in V_{\rm bad}$. Then $w_{1},w_{2}\not\in V(P^{4})$. Connect $w_{1}$ with the endpoint $v_{4}$ of $P^{4}$ by Claim 5.3 (ii). Denote the resulting rainbow directed path by $Q_{1}$.
$\bullet$ $w_{1}\not\in V_{\rm bad}$ and $w_{2}\in V_{\rm bad}$. Hence $w_{1}\not\in V(P^{4})$ and there exists a rainbow subpath $u_{1}P^{4}w_{2}^{1}w_{2}^{2}w_{2}$ of $P^{4}$ with $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$, $w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$. Replace $w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$, and connect $w_{1}$ with the vertex in $N_{P^{4}}^{-}(w_{2}^{1})$ by Claim 5.3 (i). Let $Q_{2}$ be the resulting rainbow directed path. The case for $w_{1}\in V_{\rm bad}$ and $w_{2}\not\in V_{\rm bad}$ is similar.
$\bullet$ $w_{1},w_{2}\in V_{\rm bad}$. By rearranging, one may assume that $w_{1}$ and $w_{2}$ is connected by a rainbow directed path $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ in $P^{4}$, where $w_{1}^{1},w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$ and $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$. Replace $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$. Denote the resulting rainbow directed path by $Q_{3}$.
In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})||$ is a nonnegative even integer. It follows from Claim 5.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
(ii) We will only prove that $D_{i}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]=\emptyset$ for all $i\in\tilde{\mathcal{C}}$, the second statement can be proved by a similar discussion. Suppose that $D_{i_{1}}[A\setminus(V_{\rm bad}\cup X^{\prime}),B]\neq\emptyset$ for some $i_{1}\in\tilde{\mathcal{C}}$. Choose ${z_{1}z_{2}}\in E(D_{i_{1}}[A\setminus(V_{\rm bad}\cup X^{\prime}),B])$. Hence $z_{1}\notin V(P^{4})$.
$\bullet$ $z_{2}\not\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$. Then $z_{2}\not\in V(P^{4})$. Connect $z_{2}$ with the endpoint $u_{1}$ of $P^{4}$ by Claim 5.3 (i). Let $Q_{1}$ be the resulting rainbow directed path, and $\tilde{\mathcal{C}}_{1}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
$\bullet$ $z_{2}\in V(P^{1})\cap B$. Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{\prime}z_{2}P^{4}v_{4}$, where $z_{2}^{\prime}z_{2}\in E(P^{1})\cap E(D_{i_{2}})$ for some $i_{2}\in\mathcal{C}$. Clearly, $i_{2}\in\mathcal{C}_{2}$ if $z_{2}^{\prime}\in A$, and $i_{2}\in(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus\tilde{\mathcal{C}}$ if $z_{2}^{\prime}\in B$ (since ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$). Notice that $V(P^{1})\cap B\cap V_{\rm bad}=\emptyset$. Hence $z_{2}^{\prime}\not\in V_{\rm bad}$. Connect $u_{1}P^{1}z_{2}^{\prime}$ with $z_{1}z_{2}P^{4}v_{4}$ into a single rainbow directed path by Claim 5.3 (i)-(ii). Denote the resulting rainbow directed path by $Q_{2}$. Let $\tilde{\mathcal{C}}_{2}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$ if $z_{2}^{\prime}\in A$, and $\tilde{\mathcal{C}}_{2}:=(\tilde{\mathcal{C}}\setminus\{i_{1}\})\cup\{i_{2}\}$ if $z_{2}^{\prime}\in B$.
$\bullet$ $z_{2}\in V_{\rm bad}\cup X^{\prime}$. Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{1}z_{2}P^{4}v_{4}$, where $z_{2}^{1}\in A\setminus V_{\rm bad}$. Connect $u_{1}P^{4}z_{2}^{1}$ with $z_{1}z_{2}z_{2}^{2}P^{4}v_{4}$ into a single rainbow directed path by Claim 5.3 (ii). Let $Q_{3}$ be the resulting rainbow directed path and $\tilde{\mathcal{C}}_{3}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})||$ is a nonnegative even integer. It follows from Claim 5.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction. ∎
If $|\tilde{\mathcal{C}}|=0$, then $n$ is odd. In view of Claim 5.7 (i), we know $D_{i}[B\setminus(X^{\prime}\cup V(P^{1}))]=\emptyset$ for all $i\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$. Notice that $|B\setminus(X^{\prime}\cup V(P^{1}))|\geq\frac{n}{2}-28\delta n$ and $|{\rm col}(P^{1}-A)|\leq 27\delta n$. Together with Lemma 5.14 (see the final of this subsection), we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
If $|\tilde{\mathcal{C}}|\geq 1$, then based on Claim 5.7 (ii) and the fact that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we obtain $|A|,|B|\geq\frac{n}{2}+1$, a contradiction.
Case 2. $|\mathcal{C}_{2}|<\eta n$.
In this case, each vertex in $Y\setminus(X\cup X_{Y}^{1})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{1}|$ colors $i\in\mathcal{C}_{1}$. By a similar discussion as in Case 1, we know that there are $|V_{\rm bad}|$ disjoint rainbow directed $P_{3}$ with centers in $V_{\rm bad}$ and both endpoints in the same part. We now consider vertices in $X^{\prime}$. Based on the definition of $X^{\prime}$, we move vertices in $X^{\prime}$ to $Y$ if they are covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1). In order to connect those rainbow directed $P_{3}$ with centers in $X^{\prime}$ into a single rainbow directed path, we balance the number of rainbow directed paths from $A$ to $B$ and those from $B$ to $A$.
Claim 5.8.
There exists a rainbow directed path $P^{0}$ such that its endpoints are not in $V_{\rm bad}\cup X^{\prime}$, $X^{\prime}\subseteq V(P^{0})$ and ${\rm col}(P^{0})\subseteq\mathcal{C}_{1}$, and its length is at most $8\sqrt{\epsilon}n$. Furthermore, such $P^{0}$ can be chosen starting at $A$ and ending at $B$ or two endpoints in the same part.
Proof of Claim 5.8.
If $|X^{\prime}|=0$, then by the minimum semi-degree condition, our result holds obviously. Without loss of generality, assume that $|X^{\prime}|\geq 1$ and $|A\cap X^{\prime}|-|B\cap X^{\prime}|\geq 0$. We first select all disjoint non-trivial maximum rainbow directed paths, say $Q_{1},\ldots,Q_{t}$, inside $\{D_{i}[A\cap X^{\prime}]:i\in\mathcal{C}_{1}\}$. Let $t^{\prime}:=|V(Q_{1})|+\cdots+|V(Q_{t})|$ and $Q_{i}:=x_{i}\ldots y_{i}$ for each $i\in[t]$. Next, choose a maximum rainbow directed matching, denoted by $\tilde{M}$, inside $\{D_{i}[A,B]:i\in\mathcal{C}_{1}\setminus{\rm col}(Q_{1}\cup Q_{2}\cup\ldots\cup Q_{t})\}$ avoiding vertices in $\bigcup_{i\in[t]}V(Q_{i}-y_{i})$. Partition $\tilde{M}$ to $M_{1}\cup M_{2}$ such that $V(M_{2})\cap A=V(\tilde{M})\cap A\cap X^{\prime}$. Hence, each vertex in $V(M_{2})\cap A$ or each $Q_{i}$ with $y_{i}\in V(M_{2})\cap A$ can be covered by a rainbow directed path with length at most $|X^{\prime}|+2$ and both endpoints in $B\setminus(V_{\rm bad}\cup X^{\prime})$. Denote $\hat{X}:=(A\cap X^{\prime})\setminus V\left(\bigcup_{i\in[t]}(Q_{i}-y_{i})\cup\tilde{M}\right)$.
If $|\hat{X}|=0$ and $|B\cap X^{\prime}|\geq 1$, then the desired rainbow directed path can be obtained immediately by Claim 5.3 (iii) (note that if there is a rainbow directed path with some endpoint in $V_{\rm bad}$, then it can be extended by at most two vertices so that both of its endpoints remain in the corresponding part and are no longer in $V_{\rm bad}$).
Next, we consider $|\hat{X}|\geq 1$. We are to prove that $|E(M_{1})|\geq|\hat{X}|+1$. Suppose that $|E(M_{1})|\leq|\hat{X}|\leq\sqrt{\epsilon}n$. Together with the definition of $X^{\prime}$, one has $B\cap X^{\prime}\subseteq V(\tilde{M})\cap B$ and $\tilde{M}$ can be chosen such that each matching edge incident to a vertex in $B\cap X^{\prime}$ is not incident to vertices in $A\cap X^{\prime}$.
Choose $u_{1}\in\hat{X}$, $u_{2}\in B\setminus V(\tilde{M})$ and two colors $c_{1},c_{2}\in\mathcal{C}_{1}\setminus{\rm col}(\tilde{M}\cup Q_{1}\cup\ldots\cup Q_{t})$. We consider $N_{c_{1}}^{+}(u_{1})$ and $N_{c_{2}}^{-}(u_{2})$ respectively. It is easy to see that $u_{1}$ has at most $t^{\prime}-t$ out-neighbors inside $A\cap X^{\prime}$ and $N_{c_{1}}^{+}(u_{1},B)=N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))$. Hence
$\left\lceil\frac{n}{2}\right\rceil\leq|N_{c_{1}}^{+}(u_{1})|\leq|A\setminus X^{\prime}|+(t^{\prime}-t)+|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|.$ (2)
Similarly, we have
$\left\lceil\frac{n}{2}\right\rceil\leq|N_{c_{2}}^{-}(u_{2})|\leq|N_{c_{2}}^{-}(u_{2},A\cap V(\tilde{M}))|+|B|-1.$ (3)
Therefore,
$|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|+|N_{c_{2}}^{-}(u_{2},A\cap V(\tilde{M}))|\geq|\hat{X}|+|E(M_{2})|+1.$
Notice that, if there exists an edge $w_{1}w_{2}\in E(\tilde{M})$ such that $u_{1}w_{1}\in E(D_{c_{1}})$ and $w_{2}u_{2}\in E(D_{c_{2}})$, then we will find a large rainbow matching, a contradiction. Hence $|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|+|N_{c_{2}}^{-}(u_{2},A\cap V(\tilde{M}))|\leq|E(\tilde{M})|$. It follows that $|E(\tilde{M})|\geq|\hat{X}|+|E(M_{2})|+1$, which implies $|E(M_{1})|\geq|\hat{X}|+1$. Thus, together with Claim 5.3 (iii), the desired rainbow directed path $P^{0}$ can be obtained.
Finally, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|=0$. If $|E(M_{1})|\geq 1$, then we are done. Otherwise, by the maximality of $\tilde{M}$ and the definition of $X^{\prime}$, we deduce that $D_{i}[A\setminus X^{\prime},B]=\emptyset$. Using a similar argument as (2) and (3), there is a vertex in $A\setminus(V_{\rm bad}\cup X^{\prime})$ that is adjacent to some vertex in $\{x_{1},\ldots,x_{t}\}\cup((A\cap X^{\prime})\setminus V(Q_{1}\cup\ldots\cup Q_{t}))$. Together with Claim 5.3 (iii), the desired rainbow directed path can be obtained immediately. ∎
By a similar discussion as Claim 5.4, we get the following claim.
Claim 5.9.
Assume $X\setminus X^{\prime}=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. If $x_{i}\in(X\setminus X^{\prime})\cap Y$, then there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{1}\setminus({\rm col}(P^{0})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},\,x_{i}^{4}\in Y\setminus(V(P^{0})\cup V_{\rm bad}\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$.
By Claim 5.9 and a similar discussion as Steps 1-3 in Case 1, the following hold.
  1. (D1)
    In $\{D_{i}[A\setminus V(P^{0})]\cup D_{i}[B\setminus V(P^{0})]:i\in\mathcal{C}_{1}\setminus{\rm col}(P^{0})\}$, there are $|V_{\rm bad}\setminus V(P^{0})|$ disjoint rainbow directed $P_{3}$ with centers in $V_{\rm bad}\setminus V(P^{0})$ and both endpoints in the same part. Let $\mathbf{P}$ be a set consisting of those rainbow directed $P_{3}$.
  2. (D2)
[…]
  3. (D3)
    Let $q:=|\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime}|$. In the digraph collection $\{D_{i}^{\pm}[A,B]:i\in\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\}$, there exists a rainbow directed path $Q$ of length $q-1$ or $q$ that avoids vertices in $\mathbf{P}\cup P^{0}\cup M$.
  4. (D4)
    We can assume that a vertex $v\notin V(\mathbf{P}\cup M\cup Q)$ unless $v\in V_{\rm bad}$, and a color $c\notin{\rm col}(\mathbf{P})$ if $c\in\mathcal{C}_{1}$.
Next, we give the following claim.
Claim 5.10.
Assume that $Q_{1}=z_{1}\ldots z_{s}$ and $Q_{2}=z_{t}^{\prime}\ldots z_{1}^{\prime}$ are two disjoint rainbow directed paths inside $\mathcal{D}$ such that $z_{1},z_{1}^{\prime}\in A\setminus V_{\rm bad}$, $z_{s},z_{t}^{\prime}\in B\setminus V_{\rm bad}$ and $s+t<4\eta n$. If $V_{\rm bad}\subseteq V(Q_{1}\cup Q_{2})$ and $\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}\subseteq{\rm col}(Q_{1}\cup Q_{2})$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Proof of Claim 5.10.
Note that $z_{1},z_{1}^{\prime}\in A\setminus V_{\rm bad}$ and $z_{s},z_{t}^{\prime}\in B\setminus V_{\rm bad}$. There exist $i_{1},i_{2},i_{3},i_{4}\in\mathcal{C}_{1}\setminus{\rm col}(Q_{1}\cup Q_{2})$ such that $z_{1}\in A_{i_{1}}$, $z_{1}^{\prime}\in A_{i_{2}}$, $z_{s}\in B_{i_{3}}$ and $z_{t}^{\prime}\in B_{i_{4}}$. Split $\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{i_{1},i_{2},i_{3},i_{4}\})$ into two parts $\mathcal{C}_{a}\cup\mathcal{C}_{b}$, where $|\mathcal{C}_{a}|=|A\setminus V(Q_{1}\cup Q_{2})|-1$ and $|\mathcal{C}_{b}|=|B\setminus V(Q_{1}\cup Q_{2})|-1$. Let $I_{1}\cup I_{2}$ and $J_{1}\cup J_{2}$ be equitable partitions of $A\setminus V(Q_{1}\cup Q_{2})$ and $B\setminus V(Q_{1}\cup Q_{2})$ with $|I_{1}|\geq|I_{2}|$ and $|J_{1}|\geq|J_{2}|$, respectively. Define
$\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{1},I_{2}]:i\in\mathcal{C}_{a}\}\ \text{and}\ \mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{1},J_{2}]:i\in\mathcal{C}_{b}\}.$
Notice that
$|Y\setminus V(Q_{1}\cup Q_{2})|\geq\frac{n-|X^{\prime}|-r}{2}-|V(Q_{1}\cup Q_{2})|>\frac{n}{2}-5\eta n,$
and for $i\in\mathcal{C}_{a}\cup\{i_{1},i_{2}\}$, $j\in[2]$ and $z\in I_{3-j}\cup\{z_{1},z_{1}^{\prime}\}$, we have
$|N_{{i}}^{+}(z,I_{j})|$ $\geq d_{i}^{+}(z,A\setminus V(Q_{1}\cup Q_{2}))-|I_{3-j}|$
$\geq d_{i}^{+}(z,A_{i})-|A_{1}\Delta A_{i}|-|X_{A}\setminus X_{A}^{1}|-|X^{\prime}|-|V(Q_{1}\cup Q_{2})|-|I_{3-j}|$
$\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n-\left(2\delta+4\delta+4\eta+\sqrt{\epsilon}\right)n-\left\lceil\frac{n+r}{4}\right\rceil>\frac{n}{4}-5\eta n.$
Similarly, we have $|N_{{i}}^{-}(z,I_{j})|\geq\frac{n}{4}-5\eta n$. It implies that
$|N_{{i}}^{+}(z,I_{j})\cap N_{{i}}^{-}(z,I_{j})|\geq 2\left(\frac{n}{4}-5\eta n\right)-|I_{j}|\geq\frac{n}{4}-10\eta n.$
Hence, for each $i\in\mathcal{\mathcal{C}}_{a}\cup\{i_{1},i_{2}\}$, one may view $D_{i}^{\pm}[I_{1},I_{2}]$ as an undirected graph $G_{i}[I_{1},I_{2}]$ with minimum degree at least $\frac{n}{4}-10\eta n$, where two vertices are adjacent in $G_{i}[I_{1},I_{2}]$ if and only if there are two opposite arcs between them. By Claim 5.5, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with two endpoints $x_{a}\in N_{G_{i_{1}}}(z_{1},I_{1})$ and $y_{a}\in N_{G_{i_{2}}}(z_{1}^{\prime},I)$ (here $y_{a}\in I_{2}$ if and only if $|I_{1}|=|I_{2}|$).
Similarly, there exists a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with two endpoints $x_{b}\in N_{G_{i_{3}}}(z_{s},J_{1})$ and $y_{b}\in N_{G_{i_{4}}}(z_{t}^{\prime},J)$ (here $y_{b}\in J_{2}$ if and only if $|J_{1}|=|J_{2}|$). Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle $z_{1}Q_{1}z_{s}x_{b}Q_{b}y_{b}z_{t}^{\prime}Q_{2}z_{1}^{\prime}y_{a}Q_{a}x_{a}z_{1}$. ∎
Now, we proceed by considering the value of $q$.
Subcase 2.1. $q$ is even and $q\geq 2$.
Using Claim 5.3 (iii), we connect $P^{0}$ (with endpoints in the same part), all rainbow directed paths in $\mathbf{P},M$ and $Q$ (from $A$ to $B$ with length $q-1$) into a single rainbow directed path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in A\setminus(V_{\rm bad}\cup X^{\prime})$, $y\in B\setminus(V_{\rm bad}\cup X^{\prime})$. Next, choose an edge $u_{q}v_{q}$ from $D_{c}[B\setminus V(P),A\setminus V(P)]$, where $c\in(\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P)$. Consequently, $P$ and $u_{q}v_{q}$ are two disjoint rainbow directed paths satisfying all conditions in Claim 5.10. Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
Subcase 2.2. $q$ is odd.
Using Claim 5.3 (iii), we connect all rainbow paths in $\mathbf{P},M$ and $Q$ (from $B$ to $A$ with length $q$) into a single rainbow directed path $P$, whose length is at most $(3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in B\setminus V_{\rm bad}$, $y\in A\setminus V_{\rm bad}$. Consequently, $P$ and $P^{0}$ (from $A$ to $B$) are two disjoint rainbow paths satisfying all conditions in Claim 5.10. Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
Subcase 2.3. $q=0$.
If $|X^{\prime}|\geq 2$, then choose $u\in A\cap X^{\prime}$ if $|A\cap X^{\prime}|>|B\cap X^{\prime}|$, and choose $u\in B\cap X^{\prime}$ if $|A\cap X^{\prime}|=|B\cap X^{\prime}|$. Delete $u$ from $X^{\prime}$ when constructing $P^{0}$. Based on the definition of $X^{\prime}$, there exists a rainbow directed $P_{3}$ from $B$ to $A$ with center $u$ that avoids vertices in $P^{0}$. Applying Claim 5.3 (iii) again to connect $P^{0}$ (from $A$ to $B$), all rainbow directed paths in $\mathbf{P},M$ into a single rainbow path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$, with endpoints in different parts. Thus, $P$ and $P_{x}$ are two disjoint rainbow paths satisfying all conditions in Claim 5.10. Therefore, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
If $|X^{\prime}|\leq 1$, then together with the condition that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, it is easy to find two disjoint rainbow edges with opposite directions between $A$ and $B$. By Claim 5.10 and a similar discussion as above, we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
Case 3. $|\mathcal{C}_{1}|\geq\eta n$ and $|\mathcal{C}_{2}|\geq\eta n$.
In this case, each vertex in $Y\setminus V_{\rm bad}$ lies in $Y_{i}$ for at least $(1-13\sqrt{\delta})|{\mathcal{C}_{1}}\cup\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{1}\cup{\mathcal{C}_{2}}$. Denote $\bigcup_{k\in[2]}(X_{A}^{k}\cup X_{B}^{k})=:\{v_{1},v_{2},\ldots,v_{t}\}$. Notice that $t<\frac{1}{2}\sqrt{\delta}n$. By the definition of $X_{Y}^{k}$, we obtain that for each $i\in[t]$, there exist $c_{4i-3},c_{4i-2},c_{4i-1},c_{4i}\in(\mathcal{C}_{1}\cup\mathcal{C}_{2})\setminus\{c_{1},c_{2},\ldots,c_{4i-5},c_{4i-4}\}$ and $v_{i}^{1},v_{i}^{2},v_{i}^{3},v_{i}^{4}\in(A\cup B)\setminus(V_{\rm bad}\cup\{v_{\ell}^{1},v_{\ell}^{2},v_{\ell}^{3},v_{\ell}^{4}:\ell\in[i-1]\})$ such that
  • if $v_{i}\in X_{Y}^{1}$, then $c_{4i-4+j}\in\mathcal{C}_{1}$, $v_{i}^{j}\in Y\cap Y_{c_{4i-4+j}}$, ${v_{i}^{j}v_{i}}\in E(D_{c_{4i-4+j}})$ for all $j\in[2]$ and ${v_{i}v_{i}^{j}}\in E(D_{c_{4i-4+j}})$ for all $j\in\{3,4\}$,
  • if $v_{i}\in X_{Y}^{2}$, then $c_{4i-4+j}\in\mathcal{C}_{2}$, $v_{i}^{j}\in Z\cap Z_{c_{4i-4+j}}$, ${v_{i}^{j}v_{i}}\in E(D_{c_{4i-4+j}})$ for all $j\in[2]$ and ${v_{i}v_{i}^{j}}\in E(D_{c_{4i-4+j}})$ for all $j\in\{3,4\}$.
Therefore, there exists a set of disjoint rainbow directed path $\mathbf{P}^{0}=\{v_{i}^{1}v_{i}v_{i}^{2}:i\in[t]\}$ with ${\rm col}(v_{i}^{1}v_{i}v_{i}^{2})=\{c_{4i-3},c_{4i-2}\}$ for each $i\in[t]$.
Based on the definition of $X^{\prime}$, move vertices in $X^{\prime}$ to $Y$ if they are covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1). Thus, by avoiding all vertices and colors in $\mathbf{P}^{0}$, there exists a set of disjoint rainbow directed paths $\mathbf{P^{1}}=\{v^{1}vv^{2}:v\in X^{\prime}\}$ with $v^{1}\in Z\setminus(V_{\rm bad}\cup X^{\prime})$ and $v^{2}\in Y\setminus(V_{\rm bad}\cup X^{\prime})$ if $v\in Y\cap X^{\prime}$. Furthermore, at least one of the colors of ${v^{1}v}$ and ${vv^{2}}$ can be chosen from both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Without loss of generality, assume that $|A|\leq|B|$.
Connect all rainbow directed paths in $\mathbf{P}^{0}\cup\mathbf{P}^{1}$ by using colors in $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and Claim 5.3 (i)-(iii), one may obtain a rainbow directed path $P^{1}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{1}\in B\setminus V_{\rm bad}$, whose length is at most $4t+4|X^{\prime}|\leq 2\sqrt{\delta}n+4|X^{\prime}|$.
By a similar discussion as Claim 5.4, we obtain the following claim.
Claim 5.11.
Assume $X\setminus X^{\prime}=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. Then for each $i\in[s]$, there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in(\mathcal{C}_{1}\cup\mathcal{C}_{2})\setminus({\rm col}(P^{1})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},\,x_{i}^{4}\in V\setminus(V(P^{1})\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that: if $x_{i}\in Y$, then ${x_{i}^{j}x_{i}}\in E(D_{c_{i}^{j}})$ for all $j\in[2]$, and ${x_{i}x_{i}^{j}}\in E(D_{c_{i}^{j}})$ for all $j\in\{3,4\}$, and either $x_{i}^{j}\in Y$ and $c_{i}^{j}\in\mathcal{C}_{1}$ for all $j\in[4]$, or $x_{i}^{j}\in Z$ and $c_{i}^{j}\in\mathcal{C}_{2}$ for all $j\in[4]$.
By Claim 5.11, there exists a set of disjoint rainbow directed path $\mathbf{P}^{2}=\{x_{i}^{1}x_{i}x_{i}^{2}:i\in[s]\}$, each of which has colors $c_{i}^{1},c_{i}^{2}$ for $i\in[s]$. Connect all these rainbow directed paths and $P^{1}$ using Claim 5.3 (i)-(ii), and we obtain a rainbow directed path $P^{2}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{2}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{1})|+4|X\setminus X^{\prime}|+2\leq(2\sqrt{\delta}+4\sqrt{\epsilon})n+2$.
[…]
Since $u_{1},w_{3}\notin V_{\rm bad}$, there exist $i_{0},i_{1}\in\mathcal{C}_{1}\setminus{\rm col}(P^{3})$ such that $u_{1}\in A_{i_{0}}$ and $w_{3}\in B_{i_{1}}$. By the construction of $P^{3}$, one may assume that $v\not\in V(P^{3})$ unless $v\in V_{\rm bad}\cup X^{\prime}$, and $c\not\in{\rm col}(P^{3})$ unless $c\in\mathcal{C}_{\rm bad}$. Similar to the above two cases, we know that each digraph $D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{2}$ can be viewed as an undirected graph with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$; each digraph $D_{i}[A\setminus V(P^{3})]\cup D_{i}[B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{1}$ can be viewed as the union of two disjoint undirected graphs with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$. In what follows, we proceed by considering the parity of $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$.
Subcase 3.1. $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is odd.
Split $A\setminus V(P^{3})$ into $I_{1}$ and $I_{2}$ such that $|I_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|+1}{2}$, and split $B\setminus V(P^{3})$ into $J_{1}$ and $J_{2}$ such that $|J_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|+1}{2}$. Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus{\rm col}(P^{3})\}$. By Claim 5.5, there exists a transversal undirected path $Q$ inside $\mathcal{D}_{1}$ with endpoints $x\in I_{1}$ and $y\in J_{1}$. Since $x,y\not\in V_{\rm bad}$, there exist $j_{0},j_{1}\in\mathcal{C}_{1}$ such that $x\in A_{j_{0}}$ and $y\in B_{j_{1}}$.
Suppose that $|I_{2}|=|A\setminus V(P^{3})|-|I_{1}|=k_{1}$ and $|J_{2}|=|B\setminus V(P^{3})|-|J_{1}|=k_{2}$. We further split $I_{2}$ into $I_{2}^{\prime}$ and $I_{2}^{\prime\prime}$ such that $|I_{2}^{\prime}|=\lceil\frac{k_{1}}{2}\rceil$, split $J_{2}$ into $J_{2}^{\prime}$ and $J_{2}^{\prime\prime}$ such that $|J_{2}^{\prime}|=\lceil\frac{k_{2}}{2}\rceil$. Then split the color set $\mathcal{C}_{1}\setminus({\rm col}(P^{3})\cup\{j_{0},j_{1}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$. Let $\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}$ and $\mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}$. Applying Claim 5.5 again yields that there is a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with endpoints $x_{a}\in N_{{i_{0}}}^{-}(u_{1},I_{2}^{\prime}),\,y_{a}\in N_{{j_{0}}}^{+}(x,I_{2})$; and a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with endpoints $x_{b}\in N_{{i_{1}}}^{+}(w_{3},J_{2}^{\prime}),\,y_{b}\in N_{{j_{1}}}^{-}(y,J_{2})$. Thus, $u_{1}P^{3}w_{3}x_{b}Q_{b}y_{b}yQxy_{a}Q_{a}x_{a}u_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$, a contradiction.
Subcase 3.2. $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even.
In this subcase, we first prove the following claim.
Claim 5.12.
Assume that $Q_{1}=z_{1}\ldots z_{k}$ and $Q_{2}=z_{t}^{\prime}\ldots z_{1}^{\prime}$ are two disjoint rainbow directed paths inside $\mathcal{D}$ such that $z_{1},z_{1}^{\prime}\in A\setminus V_{\rm bad}$, $z_{k},z_{t}^{\prime}\in B\setminus V_{\rm bad}$ and each of them has length at most $19\sqrt{\delta}n$. If $V_{\rm bad}\subseteq V(Q_{1}\cup Q_{2})$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(Q_{1}\cup Q_{2})$ and $|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|$ is even, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Proof of Claim 5.12.
Note that $z_{1},z_{1}^{\prime}\in A\setminus V_{\rm bad}$ and $z_{k},z_{t}^{\prime}\in B\setminus V_{\rm bad}$. There exist $\ell_{1},\ell_{2},\ell_{3}\in\mathcal{C}_{1}\setminus{\rm col}(Q_{1}\cup Q_{2})$ and $\ell_{4}\in\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})$ such that $z_{1}\in A_{\ell_{1}}$, $z_{k}\in B_{\ell_{2}}$, $z_{1}^{\prime}\in A_{\ell_{4}}$ and $z_{t}^{\prime}\in B_{\ell_{3}}$.
Split $A\setminus V(Q_{1}\cup Q_{2})$ into $I_{1}$ and $I_{2}$ such that $|I_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|}{2}$, and split $B\setminus V(Q_{1}\cup Q_{2})$ into $J_{1}$ and $J_{2}$ such that $|J_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|}{2}$. Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{4}\})\}$. By Claim 5.5, there exists a transversal undirected path $Q$ inside $\mathcal{D}_{1}$ with endpoints $z_{1}^{\prime\prime}\in N_{{\ell_{4}}}^{+}(z_{1}^{\prime})\cap J_{1}$ and $y\in I_{1}$. Since $y\not\in V_{\rm bad}$, there exists $\ell_{5}\in\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3}\})$ such that $y\in A_{\ell_{5}}$.
Assume that $|I_{2}|=k_{1}$ and $|J_{2}|=k_{2}$. We further split $I_{2}$ into $I_{2}^{\prime}$ and $I_{2}^{\prime\prime}$ such that $|I_{2}^{\prime}|=\lceil\frac{k_{1}}{2}\rceil$, and split $J_{2}$ into $J_{2}^{\prime}$ and $J_{2}^{\prime\prime}$ such that $|J_{2}^{\prime}|=\lceil\frac{k_{2}}{2}\rceil$. Split the color set $\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3},\ell_{5}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$. Let $\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}$ and $\mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}$. By Claim 5.5, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with endpoints $x_{a}\in N_{{\ell_{1}}}^{-}(z_{1})\cap I_{2}^{\prime},\,y_{a}\in N_{{\ell_{5}}}^{+}(y)\cap I_{2}$; and a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with endpoints $x_{b}\in N_{{\ell_{2}}}^{+}(z_{k})\cap J_{2}^{\prime},\,y_{b}\in N_{{\ell_{3}}}^{-}(z_{t}^{\prime})\cap J_{2}$. Hence, $z_{1}Q_{1}z_{k}x_{b}Q_{b}y_{b}z_{t}^{\prime}Q_{2}z_{1}^{\prime}z_{1}^{\prime\prime}Qyy_{a}Q_{a}x_{a}z_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$. ∎
Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$ and $|A|\leq|B|$, each vertex in $A$ has an in-neighbor in $B$ in each $D_{i}$ with $i\in\mathcal{C}_{1}\cup{\rm col}(M)$. Suppose that there exists an edge $y_{2}y_{1}\in E(D_{i_{2}}[B,A])$ for some $i_{2}\in\mathcal{C}_{1}\cup{\rm col}(M)$, where $y_{1}\in A\setminus(V(P^{3})\cup V_{\rm bad})$ and $y_{2}\in B$. If $i_{2}\in{\rm col}(M)$, then assume ${uv}\in E(M[A])\cap E(D_{i_{2}})$. By Claim 5.3 (i)-(ii), we connect $P^{2}$, all rainbow edges in $M-uv$ and all rainbow edges in $M^{\prime}$ in turn. This results a rainbow directed path $\tilde{P}^{3}$ with endpoints $\tilde{u}_{1}\in A\setminus V_{\rm bad}$ and $\tilde{w}_{3}\in B\setminus V_{\rm bad}$. Therefore, $P^{3}$ can be written as $uvv^{\prime}\tilde{u}_{1}\tilde{P}^{3}\tilde{w}_{3}$, where the colors of ${vv^{\prime}}$ and ${v^{\prime}\tilde{u}_{1}}$ are in $\mathcal{C}_{2}$. Since $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even, one has $|\mathcal{C}_{2}\setminus{\rm col}(\tilde{P}^{3})|$ is even. Hence one may let $P^{3}:=\tilde{P}^{3}$ if $i_{2}\in{\rm col}(M)$ and further assume $i_{2}\not\in{\rm col}(P^{3})$.
Notice that $y_{2}\not\in V(P^{3})$ unless it is in $V_{\rm bad}\cup X^{\prime}$. If $y_{2}\not\in V_{\rm bad}\cup X^{\prime}$, then $P^{3}$ and $y_{2}y_{1}$ are two disjoint rainbow directed paths satisfying all conditions in Claim 5.12. Hence, $\mathcal{D}$ contains a transversal Hamilton cycle, a contradiction. Therefore, $y_{2}\in V_{\rm bad}\cup X^{\prime}$.
Recall that $y_{2}^{1}y_{2}y_{2}^{2}$ is the rainbow directed path that covers $y_{2}$. Assume ${y_{2}^{1}y_{2}}$ and ${y_{2}y_{2}^{2}}$ have colors $j_{1}$ and $j_{2}$ in $P^{3}$, respectively. Hence, there exists a rainbow directed $\hat{P^{3}}$ starting at $A$ and ending at $B$ such that $P^{3}$ is obtained by connecting $y_{2}^{1}y_{2}y_{2}^{2}$ and $\hat{P^{3}}$ if $y_{2}^{1}\in A$, and $P^{3}$ is obtained by connecting $\hat{P^{3}}$ and $y_{2}^{1}y_{2}y_{2}^{2}$ otherwise.
If $y\in X^{\prime}$, then by adjusting the selections of $y_{2}^{1},y_{2}^{2}$ and $j_{1},j_{2}$, one may get a rainbow directed path $P^{3}$ such that $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd. This can be reduced to Subcase 3.1.
If $y\in V_{\rm bad}$, then $|\mathcal{C}_{2}\setminus{\rm col}(\hat{P^{3}})|$ is even. Let $\hat{P}:=y_{2}^{1}y_{2}y_{1}$ if $y_{2}^{1}\in B$, and let $\hat{P}:=y^{\prime}y_{2}^{1}y_{2}y_{1}$ with $y^{\prime}y_{2}^{1}\in E(D_{c^{\prime}})$ for unused color $c^{\prime}\in\mathcal{C}_{2}$ and unused vertex $y^{\prime}\in B$ if $y_{2}^{1}\in A$. Consequently, $\hat{P^{3}}$ and $\hat{P}$ are two rainbow directed paths satisfying all conditions in Claim 5.12. Hence, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction. ∎
In order to complete the proof of Theorem 5.1, we need to show the existence of transversal directed Hamilton cycles in Step 1 of Case 1. We first state a preliminary result from [11].
Lemma 5.13 ([11]).
Let $\mathcal{C}$ be a set of colors, and $\mathcal{G}=\{G_{i}[Y,B]:i\in\mathcal{C}\}$ be a collection of bipartite graphs with the common bipartition $Y\cup B$ such that $7|Y|<|B|\leq\frac{3}{5}|\mathcal{C}|$. If $\sum_{i\in\mathcal{C}}|E(D_{i}[Y,B])|\geq t|B||\mathcal{C}|$ for some integer $t$ with $1\leq t\leq|Y|$, then $\mathcal{G}$ contains $t$ disjoint rainbow star on $5$ vertices, each with the center in $Y$ and other vertices in $B$.
Lemma 5.14.
Assume $0<\frac{1}{n}\ll\delta\ll 1$ and $0\leq\gamma\leq 3{\delta}$. Let $\mathcal{C}$ be a set of $n$ colors, and let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs with the common vertex set $V$ of size $n$ such that $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$. Let $A\cup B$ be a partition of $V$ with $|A|=\left\lceil\frac{n}{2}\right\rceil+\gamma n$, and let $\mathcal{C}^{\prime}\cup\mathcal{C}^{\prime\prime}$ be a partition of $\mathcal{C}$ with $|\mathcal{C}^{\prime\prime}|\leq\delta n$. Assume that $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$. Then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Proof.
Define $Y_{1}:=\{v\in A:d_{i}^{+}(v,B)\leq(1-\delta^{\frac{1}{4}})|B|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$, $Y_{2}:=\{v\in A:d_{i}^{-}(v,B)\leq(1-\delta^{\frac{1}{4}})|B|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$. Since $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$, it is routine to check that
$\left\lceil\frac{n}{2}\right\rceil|B||\mathcal{C}^{\prime}|$ $\leq\sum_{i\in\mathcal{C}^{\prime}}|E({D_{i}}[A,B])|$
$\leq|Y_{1}|(1-\delta^{\frac{1}{4}})|B|\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|+|Y_{1}||B|(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|+(|A|-|Y_{1}|)|B||\mathcal{C}^{\prime}|$
$=(|A|-\delta^{\frac{1}{2}}|Y_{1}|)|B||\mathcal{C}^{\prime}|.$ (4)
It implies that $|Y_{1}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$. Similarly, $|Y_{2}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$. For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp. $v\in Y_{2}\setminus Y_{1}$), we have $d_{i}^{-}(v,B)>(1-\delta^{\frac{1}{4}})|B|$ (resp. $d_{i}^{+}(v,B)>(1-\delta^{\frac{1}{4}})|B|$) for at least $(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|$ colors $i\in\mathcal{C}^{\prime}$. Notice that Claim 5.5 and Claim 5.3 (i)-(ii) hold by setting $V_{\rm bad}:=Y_{1}\cup Y_{2}$ and $\mathcal{C}_{2}=\mathcal{C}^{\prime}$.
Since $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$, we have $|E(D_{i}[A,B])|,|E(D_{i}[B,A])|\geq|B|\left\lceil\frac{n}{2}\right\rceil$ for each $i\in\mathcal{C}^{\prime}$, i.e., there are at most $|B|(\left\lceil\frac{n}{2}\right\rceil+\gamma n)-|B|\left\lceil\frac{n}{2}\right\rceil=|B|\gamma n$ non-edges from $A$ to $B$ (resp., from $B$ to $A$) in $D_{i}$. Hence
$\sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[Y_{1},B])|\geq(|Y_{1}|-\gamma n)|B||\mathcal{C}^{\prime}|\ \text{and}\ \sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[B,Y_{2}])|\geq(|Y_{2}|-\gamma n)|B||\mathcal{C}^{\prime}|.$
According to the sizes of $|Y_{1}|$ and $|Y_{2}|$, we proceed by considering the following three cases.
Case 1. $|Y_{1}|,|Y_{2}|\geq\gamma n$.
In view of Lemma 5.13, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[Y_{1},B]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[B,Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$. Recall that each vertex in $Y_{1}$ (resp. $Y_{2}$) has at least $\left\lceil\frac{n}{2}\right\rceil-(1-\delta^{\frac{1}{4}})|B|>4|Y|$ out-neighbors (resp. in-neighbors) in $A$ for at least $\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|>\delta^{\frac{1}{4}}(1-\delta)n>4|Y|$ digraphs $D_{i}$ with $i\in\mathcal{C}^{\prime}$. Therefore, by using colors in $\mathcal{C}^{\prime}$,
  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by disjoint rainbow directed $P_{3}$ with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $B$,
  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ starting at $B$ and ending at $A$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$,
  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ starting at $A$ and ending at $B$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$,
  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $A$.
For each $y\in Y_{1}\cup Y_{2}$, denote the rainbow directed $P_{3}$ with center $y$ by $P_{y}:=y^{1}yy^{2}$ with colors $c_{y}^{1}$ and $c_{y}^{2}$. Let $\mathbf{P}=\{P_{y}:y\in Y_{1}\cup Y_{2}\}$. By using Claim 5.3, one may connect all of the above rainbow directed $P_{3}$ into a single rainbow directed path $P^{1}$ starting at $A$ and ending at $B$. Then
$|V(P^{1})\cap A|-|V(P^{1})\cap B|=|Y_{1}^{\prime}\setminus Y_{2}^{\prime}|+|Y_{2}^{\prime}\setminus Y_{1}^{\prime}|+2|Y_{1}^{\prime}\cap Y_{2}^{\prime}|=|Y_{1}^{\prime}|+|Y_{2}^{\prime}|=2\gamma n.$
Hence, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|+\sigma$, where $\sigma=0$ if $n$ is odd and $\sigma=1$ otherwise.
[…]
Subcase 1.1. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have the same parity.
Greedily choose two disjoint rainbow directed paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\}$ with lengths $\lceil\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rceil$ and $\lfloor\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rfloor$ respectively, such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$. Applying Claim 5.3 (i) to connect $P^{1}$, all rainbow edges in $M$, $P_{A}$ and $P_{B}$ in turn, we get a single rainbow path $P^{2}$ with endpoints in different parts. Clearly, $|A\setminus V(P^{2})|=|B\setminus V(P^{2})|$. Together with Claim 5.5, we know $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
Subcase 1.2. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have the different parity.
Notice that $|B|=\lfloor\frac{n}{2}\rfloor-\gamma n$. Hence for each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $B$ has at least $\gamma n+1$ in-neighbors in $A$. Therefore, each vertex in $B$ has at least one in-neighbor outside $Y_{1}^{\prime}$. We first consider $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|\geq 1$. Choose an unused vertex $y\in B$ and a color $c\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$. Suppose that ${xy}\in E(D_{c}[A,B])$ with $x\notin Y_{2}^{\prime}$.
$\bullet$ $x\in A\setminus(Y_{1}\cup Y_{2})$. Then $x$ can be chosen outside $V(P^{1}\cup M)$. By Claim 5.3 (i)-(ii), one may connect $xy$ and $P^{1}$ into a single rainbow directed path $P$. Applying a similar discussion as Subcase 1.1, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
$\bullet$ $x\in(Y_{1}\cup Y_{2})\setminus Y_{1}^{\prime}$. Then replace the end vertex of $P_{x}$ by $y$. Applying a similar discussion as Subcase 1.1, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
Therefore, it suffices to consider $n$ is odd and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|=0$. For each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $A$ has at least one in-neighbor outside $Y_{1}^{\prime}$. Choose an unused vertex $y\in A\setminus(Y_{1}\cup Y_{2})$ and a color $c\in\mathcal{C}$. Then there exists a vertex $x\in A\setminus Y_{1}^{\prime}$ such that ${xy}\in E(D_{c}[A])$. By a similar discussion as above, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
Case 2. At least one of $Y_{1}$ and $Y_{2}$ has size less than $\gamma n$.
In this case, we consider only $|Y_{1}|\geq\gamma n$ and $|Y_{2}|<\gamma n$; the other two cases can be discussed similarly. Based on the proof of Case 1, we know that all vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{1}$ and endpoints in $A\setminus(Y_{1}\cup Y_{2})$. Furthermore, each vertex in $Y_{2}\setminus Y_{1}$ can be covered by a rainbow directed $P_{3}$ with center in $Y_{2}\setminus Y_{1}$ and start vertex in $A\setminus(Y_{1}\cup Y_{2})$ using colors in $\mathcal{C}^{\prime}$. (In this process, for the choice of the other endpoint of each rainbow directed $P_{3}$, we prefer using vertices in $A$ over vertices in $B$.) By using unused colors in $\mathcal{C}$ and unused vertices in $A\setminus Y$, one may extend those rainbow directed $P_{3}$ as long as possible or choose all other disjoint rainbow directed paths. Let $\mathbf{P}=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow directed paths in the above, where each $Q_{i}$ has at least one endpoint in $A\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $A$.
Subcase 2.1. $|B|\geq|A|-(s_{1}+\cdots+s_{t})-\sigma$.
In this subcase, there exists a set $\mathbf{P}^{\prime}=\{Q_{1}^{\prime},Q_{2}^{\prime},\ldots,Q_{\ell}^{\prime}\}$ such that $|B|=|A|-(s_{1}^{\prime}+\cdots+s_{\ell}^{\prime})-\sigma$, the endpoints of $Q_{i}^{\prime}$ are not in $Y$ for all $i\in[\ell]$ and $Y_{2}\cup Y_{1}^{\prime}\subseteq V(Q_{1}^{\prime}\cup\ldots\cup Q_{\ell}^{\prime})$, where $s_{i}^{\prime}:=|E(Q_{i}^{\prime}[A])|$ for each $i\in[\ell]$. By Claim 5.3 (i)-(ii), we connect all rainbow directed paths in $\mathbf{P}^{\prime}$ into a single rainbow directed path $P^{1}$, whose endpoints are in different parts. Since $|A|-|B|=2\gamma n+\sigma$, we have $\ell\leq s_{1}^{\prime}+\cdots+s_{\ell}^{\prime}=2\gamma n$. Therefore, $|E(P^{1})|\leq 8\gamma n$ and $|A\setminus V(P^{1})|-\sigma=|B\setminus V(P^{1})|$.
Recall that all vertices in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ can also be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ and both endpoints in $B$. Applying Lemma 5.3 (i) again to connect those rainbow directed paths with $P^{1}$, we get a rainbow directed path $P^{2}$ with length at most $8\gamma n+12\sqrt{\delta}n$, and with endpoints in different parts.
[…]
Based on Claim 5.3 (i), one may connect $P^{1}$ and all rainbow directed edges in $M$ into a single rainbow directed path $P^{2}$, whose endpoints are in different parts. Clearly, $|E(P^{2})|\leq 8\gamma n+4{\delta}n$. Next, we are to choose two disjoint rainbow paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)\}$ such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$ respectively, whose lengths are determined by the parity of $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$. Clearly, if $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have the same parity, then by Claim 5.5, we know $\mathcal{G}$ contains a transversal Hamilton cycle, a contradiction. Hence $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have different parity. By similar arguments as the proof of Step 4 in Case 1, we can prove our result.
Subcase 2.2. $|B|<|A|-(s_{1}+\cdots+s_{t})-\sigma$.
In this case, $\sum_{i=1}^{t}s_{i}<2\gamma n$. Let $w$ be an arbitrary vertex in $A\setminus V(\mathbf{P})$ and $c_{1},c_{2}$ be two colors in $\mathcal{C}\setminus{\rm col}(\mathbf{P})$. Clearly, in $D_{c_{1}}$ and $D_{c_{2}}$, $w$ has at least $\left\lceil\frac{n}{2}\right\rceil-(\lfloor\frac{n}{2}\rfloor-\gamma n)=\gamma n+\sigma$ out-neighbors (resp. in-neighbors) in $V(\mathbf{P})$ and it cannot be adjacent to the start vertex of $Q_{i}[A]$ (resp. from the end vertex of $Q_{i}[A]$) for all $i\in[t]$ (by the maximality of $\mathbf{P}$). For convenience, assume that $s_{0}=0$ and $V(Q_{i})\cap A:=\{v_{s_{1}+\cdots+s_{i-1}+i},\ldots,v_{s_{1}+\cdots+s_{i}+i}\}$ for each $i\in[t]$. Define
$I_{1}:=\left\{i:v_{i}\in N_{{c_{1}}}^{-}(w)\right\}\ \text{and}\ I_{2}:=\left\{i:v_{i+1}\in N_{{c_{2}}}^{+}(w)\right\}.$
Hence $I_{1},I_{2}\subseteq\bigcup_{i\in[t]}\left[\sum_{j\in[i-1]}s_{j}+i,\sum_{j\in[i]}s_{j}+i-1\right]$. Therefore, $I_{1}\cap I_{2}\neq\emptyset$, which contradicts the maximality of $\mathbf{P}$. ∎
5.2 Proof of Theorem 5.2
In this subsection, we give the proof of Theorem 5.2.
Proof of Theorem 5.2.
Recall that if $|\mathcal{C}_{3}|\geq 2\sqrt{\delta}n$, then $|\mathcal{C}_{1}|+|\mathcal{C}_{2}|<2\sqrt{\delta}n$. Denote $[m]:=\mathcal{C}_{3}\setminus I^{2}$ and $\mathcal{C}_{\rm bad}:=[m+1,n]$. By adding colors to $\mathcal{C}_{\rm bad}$ if necessary, we may assume $m=(1-4\sqrt{\delta})n$. By (B2), we have $|C_{1}^{k}\Delta C_{i}^{k}|\leq 2\delta n$ for all $k\in[4]$ and all $i\in[m]$. Expand $C_{1}^{1}\cup C_{1}^{2}\cup C_{1}^{3}\cup C_{1}^{4}$ into a partition $C^{1}\cup C^{2}\cup C^{3}\cup C^{4}$ of $V$ such that $|C^{1}|-|C^{3}|\leq 1$ and $|C^{2}|=|C^{4}|$. Hence, for $i\in[2,m]$ and $k\in[4]$, we have $|C^{k}\Delta C_{i}^{k}|\leq 2\delta n+2{\epsilon}n<3\delta n$.
Now, we define the set of bad vertices. Let $(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}$. Define
$X:=\left\{x\in V:x\in L_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 3\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
$X_{YZ}:=\left\{x\in(Y\cup Z)\setminus X:x\not\in Y_{i}\cup Z_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 10\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
$X_{YZ}^{\prime}:=\left\{x\in X_{YZ}:x\in W_{i}\leavevmode\nobreak\ \textrm{for some $W\in\{Y,Z\}$ with at least}\leavevmode\nobreak\ \frac{3}{2}\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{3}}\right\}.$
Recall that for all $i\in{\mathcal{C}}_{3}$, we have $|L_{i}|\leq 2{\epsilon}n$ and $|(Y_{1}\cup Z_{1})\Delta(Y_{i}\cup Z_{i})|<4\delta n$. Hence
$3\sqrt{\delta}|{\mathcal{C}}_{3}||X|\leq 2{\epsilon}n|{\mathcal{C}}_{3}|\ \ \text{and}\ \ 10\sqrt{\delta}|{\mathcal{C}}_{3}||X_{YZ}|\leq\sum_{i\in\mathcal{C}_{3}}|(Y\cup Z)\setminus(Y_{i}\cup Z_{i})|\leq(4\delta+2{\epsilon})n|{\mathcal{C}}_{3}|.$
It follows that $|X|\leq\sqrt{\epsilon}n$ and $|X_{YZ}|\leq\frac{9}{20}\sqrt{\delta}n<\frac{1}{2}\sqrt{\delta}n$. For a vertex $x\in X_{YZ}\setminus X_{YZ}^{\prime}$, we know $x\in Y_{i}\cup Z_{i}$ for at most $3\sqrt{\delta}n$ colors $i\in{\mathcal{C}}_{3}$ and $x\in\bigcup_{k\in[4]}C_{i}^{k}$ for at least $(1-3\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in{\mathcal{C}}_{3}$. Thus, $x\in(Y\cup Z)\setminus(Y_{i}\cup Z_{i})$ for at least $(1-4\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in\mathcal{C}_{3}$ and $x\in\bigcup_{k\in[4]}C_{i}^{k}\setminus(Y_{i}\cup Z_{i})$ for at least $(1-3\sqrt{\delta})|{\mathcal{C}}_{3}|-3\sqrt{\delta}n\geq(1-10\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in\mathcal{C}_{3}$. Hence
$|X_{YZ}\setminus X_{YZ}^{\prime}|(1-4\sqrt{\delta})|{\mathcal{C}}_{3}|\leq$ $\sum_{i\in{\mathcal{C}}_{3}}|(Y\cup Z)\setminus(Y_{i}\cup Z_{i})|$
$\leq$ $\sum_{i\in\mathcal{C}_{3}}\left(|(Y\cup Z)\setminus(Y_{1}\cup Z_{1})|+|(Y_{1}\cup Z_{1})\Delta(Y_{i}\cup Z_{i})|\right)$
$\leq$ $(4\delta+2\epsilon)n|{\mathcal{C}}_{3}|.$
It follows that $|X_{YZ}\setminus X_{YZ}^{\prime}|\leq 8\delta n$. Then move vertices in $X_{YZ}\setminus X_{YZ}^{\prime}$ to $V\setminus(Y\cup Z)$. Therefore, each vertex in $(Y\cup Z)\setminus(X\cup X_{YZ}^{\prime})$ lies in $Y_{i}\cup Z_{i}$ for at least $(1-10\sqrt{\delta})|\mathcal{C}_{3}|$ colors $i\in\mathcal{C}_{3}$.
Let $\{Y,Z\}\in\left\{\{C^{1},C^{3}\},\{C^{2},C^{4}\}\right\}$. Define
$X_{Y}:=$$\left\{x\in Y\setminus(X\cup X_{YZ}^{\prime}):x\not\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 20\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
$X_{Y}^{\prime}:=$$\left\{x\in X_{Y}:x\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 6\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{3}}\right\}.$
By a similar discussion as above, we know that $|X_{Y}|<\frac{1}{2}\sqrt{\delta}n$ and each vertex of $X_{Y}\setminus X_{Y}^{\prime}$ lies in $Z_{i}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in\mathcal{C}_{3}$. Furthermore, $|X_{Y}\setminus X_{Y}^{\prime}|\leq 4\delta n$. Then move vertices in $X_{Y}\setminus X_{Y}^{\prime}$ to $Z$. Hence $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n$.
Define $V_{\rm bad}:=X\cup(\bigcup_{(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}}X_{YZ}^{\prime})\cup(\bigcup_{k\in[4]}X_{C^{k}}^{\prime})$. Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+3\sqrt{\delta})n$. Moreover, each vertex in $C^{k}\setminus V_{\rm bad}$ lies in $C_{i}^{k}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in{\mathcal{C}}_{3}$.
The subsequent result can be used to connect two disjoint short rainbow directed paths into a single short rainbow directed path. The proof follows a similar argument to that of Claim 5.3, and thus we omit it.
Claim 5.15 (Connecting tool).
Let $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ be two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s}\in Y\setminus V_{\rm bad}$, $v_{1}\in Z\setminus V_{\rm bad}$, where $Y,Z\in\{C^{1},C^{2},C^{3},C^{4}\}$. Assume that $|\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)|>21\sqrt{\delta}n$ and $|C^{i}\setminus(V(P\cup Q)\cup V_{\rm bad})|>25\delta n$ if $w_{1}$ or $w_{1}^{\prime}$ is chosen from $C^{i}$ for some $i\in[4]$.
  1. (i)
    If $(Y,Z)\in\{(C^{1},C^{2}),(C^{3},C^{4})\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$. Similar result holds for $(Y,Z)\in\{(C^{2},C^{3}),(C^{4},C^{1})\}$.
  2. (ii)
    If $(Y,Z)\in\{(C^{1},C^{1}),(C^{3},C^{3})\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and one vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
  3. (iii)
    If $\{Y,Z\}=\{C^{2},C^{4}\}$, then there are three colors $c_{1},c_{2},c_{3}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and two vertices $w_{1}\in Z\setminus(V(P\cup Q)\cup V_{\rm bad})$, $w_{1}^{\prime}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}w_{1}^{\prime}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2},c_{3}\}$.
  4. (iv)
    If $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(1,3),(1,4),(4,3)\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in C^{2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$. Similar result holds for $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(3,1),(3,2),(2,1)\}$.
  5. (v)
    If $(Y,Z)=(C^{i},C^{i})$ with $i\in\{2,4\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and two vertices $w_{1}\in C^{i+2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
Delete vertices in $X$ from $\bigcup_{k\in[4]}C^{k}$ and choose a vertex $x\in X$. Then for some $j\in[4]$, we have $|N_{i}^{-}(x,C^{j}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$. Recall that $|X|\leq\sqrt{\epsilon}n$. We first consider $j=1$ and do the following operations.
  1. (a)
    If $|N_{i}^{+}(x,C^{1}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{3}$ with center in $x$ and both endpoints in $C^{1}$. Move such $x$ to $C^{1}$.
  2. (b)
    Notice that each vertex in $C^{1}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ in-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$. If $|N_{i}^{+}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{4}$ with center in $x$ and endpoints in $C^{4},C^{2}$ respectively. Move such $x$ to $C^{1}$.
    Suppose (a)-(b) do not hold. Then for at least $(1-400\delta)|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|C^{3}\cup C^{4}|-400\delta n-|V_{\rm bad}|$ out-neighbors in $(C^{3}\cup C^{4})\setminus V_{\rm bad}$. Notice that each vertex in $C^{3}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ out-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$.
  3. (c)
    If $|N_{i}^{-}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{3}$ or $P_{4}$ with center in $x$ and endpoints in $C^{2},C^{4}$ respectively. Move such $x$ to $C^{3}$. Similar result holds when $|N_{i}^{-}(x,C^{3}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, move such $x$ to $C^{3}$.
  4. (d)
    Suppose (a)-(c) do not hold. Then in each $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|(C^{1}\cup C^{4})\setminus V_{\rm bad}|-400\delta n-|V_{\rm bad}|$ in-neighbors inside $(C^{1}\cup C^{4})\setminus V_{\rm bad}$ for all but at most $(1-400\delta)|\mathcal{C}_{3}|$ colors $i\in\mathcal{C}_{3}$. Move such $x$ to $C^{2}$.
Applying the above operation for $j\in\{2,3,4\}$ in turn. Denote by $X^{\prime}$ the subset of $X$ consisting of vertices from item (d). Hence, for vertices in $X\setminus X^{\prime}$, there exists a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $X\setminus X^{\prime}$, each having both endpoints in $C^{k}$ with $k\in\{1,3\}$, or one endpoint in $C^{2}$ and the other in $C^{4}$. Moreover, $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n+2\sqrt{\epsilon}n$.
By the definition of $V_{\rm bad}\setminus X$, for vertices in $(V_{\rm bad}\setminus X)\cap Y$, we can find a set of disjoint rainbow directed $P_{3}$ with centers in $(V_{\rm bad}\setminus X)\cap Y$ and endpoints in $Y\setminus V_{\rm bad}$ (if $Y\in\{C^{1},C^{3}\}$) or in $(C^{2}\cup C^{4})\setminus(Y\cup V_{\rm bad})$ (if $Y\in\{C^{2},C^{4}\}$). Denote the rainbow directed path with center $x$ obtained above by $P_{x}$ for each $x\in V_{\rm bad}\setminus X^{\prime}$.
In order to use the transversal blow-up lemma inside $\{D_{i}[C^{2},C^{4}]:i\in\mathcal{C}_{3}\}$, we need to balance the sizes of $|C^{2}|$ and $|C^{4}|$. Without loss of generality, assume that $|C^{2}|-|C^{4}|=r$ with $0\leq r\leq 25{\delta}n$. We now claim that by moving vertices in $C^{2}\cap V_{\rm bad}$ or deleting rainbow directed paths inside $\mathcal{D}[C^{2}]$ and $\mathcal{D}[C^{2},C^{1}\cup C^{3}]$, we can make the number of remaining vertices in $C^{2}$ and $C^{4}$ differ by $\sigma$, where $\sigma=1$ if $|C^{2}|+|C^{4}|$ is odd and $\sigma=0$ otherwise.
Denote $s_{1}:=|X_{C^{2}C^{4}}^{\prime}\cap C^{2}|$ and $s_{2}:=|X_{C^{2}}^{\prime}|$. If $|C^{2}|-|C^{4}|-\sigma\leq s_{1}+2s_{2}$, then by moving $s_{1}^{\prime}$ vertices in $X_{C^{2}C^{4}}^{\prime}\cap C^{2}$ to $C^{1}\cup C^{3}$ and $s_{2}^{\prime}$ vertices in $X_{C^{2}}^{\prime}$ to $C^{4}$ with $s_{1}^{\prime}+2s_{2}^{\prime}=|C^{2}|-|C^{4}|-\sigma$, we can get our desired result. Hence, it suffices to consider $|C^{2}|-|C^{4}|-\sigma>s_{1}+2s_{2}$. We first move all vertices in $X_{C^{2}C^{4}}^{\prime}\cap C^{2}$ to $C^{1}\cup C^{3}$ and all vertices in $X_{C^{2}}^{\prime}$ to $C^{4}$. Then $(V_{\rm bad}\setminus X)\cap C^{2}=\emptyset$. Choose a set of disjoint maximal rainbow directed paths in $\mathcal{D}[C^{2}]\cup\mathcal{D}[C^{2},C^{1}]$, denote such a set by $\{Q_{1},\ldots,Q_{t}\}$. Assume $\{Q_{1}^{\prime},\ldots,Q_{k}^{\prime}\}$ is a set of disjoint rainbow directed $P_{3}$ in $\mathcal{D}[C^{2},C^{3}]$ with centers in $C^{2}$ and endpoints in $C^{3}$.
If $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$, then there must exist an integer $k^{\prime}\in[k]$ and a set of disjoint rainbow directed subpaths of $\{Q_{1},\ldots,Q_{t}\}$, say $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}}\}$, such that $|E(\tilde{Q}_{1})|+\cdots+|E(\tilde{Q}_{t^{\prime}})|+k^{\prime}=|C^{2}|-|C^{4}|-\sigma$. That is, $|C^{4}|-t^{\prime}=|C^{2}|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-k^{\prime}-\sigma$. If the end vertex of $\tilde{Q_{i}}$ with $i\in[t^{\prime}]$ (resp. the start vertex of $Q_{i}^{\prime}$ with $i\in[k^{\prime}]$) lies in $V_{\rm bad}\setminus X^{\prime}$, then extend it by a vertex not in $V_{\rm bad}$. Using Claim 5.15 to connect the resulting $\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}},Q_{1}^{\prime},\ldots,Q_{k^{\prime}}^{\prime}$ into a single rainbow path $P^{0}$ with endpoints in $C^{2}$ and $C^{4}$ respectively, whose length is at most $190\delta n$. Therefore, $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-\sigma$, as desired.
In each of the above cases, the remaining vertices in $C^{2}$ and $C^{4}$ are balanced. Since $|E(P^{0})|\leq 190\delta n$, by avoiding vertices in $V(P^{0})$, one may choose a set of disjoint rainbow directed paths $\mathbf{P}^{0}=\{P_{x}:x\in V_{\rm bad}\setminus(X^{\prime}\cup P^{0})\}$. For colors in $\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{0})$, we select a maximal rainbow matching $M$ inside $\mathcal{D}^{\pm}[C^{2},C^{4}]\cup\mathcal{D}[C^{1}]\cup\mathcal{D}[C^{3}]\cup(\bigcup_{i\in[4]}\mathcal{D}[C^{i},C^{i+1}])$ ensuring it avoids vertices in $V(P^{0}\cup\mathbf{P}^{0})$. If there exists a color $c\in\mathcal{C}\setminus({\rm col}(P^{0}\cup M))$, then $D_{c}^{\pm}[C^{1},C^{3}]$ is almost an undirected complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost undirected complete graphs. For colors in $\mathcal{C}\setminus({\rm col}(P^{0}\cup M))$,we can choose a set of disjoint rainbow directed paths, say $\mathbf{P}^{1}$, with centers in $C^{1}$ and endpoints inside $C^{3}$, ensuring that at most one color in $\mathcal{C}_{\rm bad}$ is remains unused.
Let $\mathbf{P}:=\mathbf{P}^{0}\cup\mathbf{P}^{1}\cup M\cup\{P^{0}\}$. By Claim 5.15 (ii) and (v), we connect all rainbow directed paths inside $\mathbf{P}$ with one endpoints in $C^{2}$ and the other in $C^{4}$ into $Q_{0}=u_{0}\ldots v_{0}$, all rainbow directed paths inside $\mathbf{P}$ with both endpoints inside $C^{1}$ and $C^{3}$ into $Q_{1}=u_{1}\ldots v_{1}$ and $Q_{2}=u_{2}\ldots v_{2}$ respectively. Since $u_{0},v_{0},u_{1},v_{2}\notin V_{\rm bad}$, there are four distinct colors $i_{1},i_{2},i_{3},i_{4}\in\mathcal{C}_{3}\setminus{\rm col}(Q_{0}\cup Q_{1}\cup Q_{3})$ such that $u_{0}\in C_{i_{1}}^{2}$, $v_{0}\in C_{i_{2}}^{4}$, $u_{1}\in C_{i_{3}}^{1}$ and $v_{2}\in C_{i_{4}}^{3}$. We proceed by considering the following two cases.
$\bullet$ $\sigma=0$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|$. We first assume that $\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})=\emptyset$. By applying Claim 5.5 inside the digraph collection $\{D_{i}[C^{1}]\cup D_{i}[C^{3}]:i\in\mathcal{C}_{3}\setminus\{i_{1},i_{2},i_{3},i_{4}\}\}$ and avoiding vertices and colors in $Q_{0}\cup Q_{1}\cup Q_{2}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, where $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$. Connecting $Q_{1},Q_{2}$ by using Claim 5.15 (iv), one gets a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.
Applying Claim 5.5 to the digraph collection $\{D_{i}^{\pm}[C^{2}\setminus V(P\cup Q_{0}),C^{4}\setminus V(P\cup Q_{0})]:i\in\mathcal{C}_{3}\setminus({\rm col}(P\cup Q_{0})\cup\{i_{1},i_{2},c_{1},c_{2}\})\}$, there exist two disjoint rainbow directed paths $P^{3}=x_{3}\ldots y_{3}$ and $P^{4}=x_{4}\ldots y_{4}$ with $x_{3}\in N_{{c_{1}}}^{+}(y_{2},C^{4}\setminus V_{\rm bad})$, $y_{3}\in N_{{i_{1}}}^{-}(u_{0},C^{4}\setminus V_{\rm bad})$, $x_{4}\in N_{{i_{2}}}^{+}(v_{0},C^{2}\setminus V_{\rm bad})$ and $y_{4}\in N_{{c_{2}}}^{-}(x_{1},C^{4}\setminus V_{\rm bad})$, respectively. Hence $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
Now, assume that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$. Together with the condition $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, there exists an edge $uv\in E(D_{c}[C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{3}\cup C^{4}])$ if $|C^{1}|+|C^{2}|\leq|C^{3}|+|C^{4}|$, and $vu\in E(D_{c}[C^{4}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{1}\cup C^{2}])$ otherwise. By a similar discussion as above and considering whether the vertex $v$ belongs to $V_{\rm bad}$ or not, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
$\bullet$ $\sigma=1$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-1$. We first consider that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$. Choose a directed edge, say ${xy}$, in $E(D_{c}[C^{3}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3}),C^{1}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3})])$. By applying Claim 5.5 to the digraph collection $\{D_{i}[C^{1}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{x\})]\cup D_{i}[C^{3}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{y\})]:i\in\mathcal{C}_{3}\setminus({\rm col}(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{i_{1},i_{2},i_{3},i_{4},c\})\}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, with $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$. By connecting $Q_{1},{xy},Q_{3}$ in turn using Claim 5.15 (iv), we obtain a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$. Similar to the previous discussion, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
Now, assume that $\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})=\emptyset$. Since $|C^{2}|\geq|C^{4}|$, either $|C^{1}|+|C^{2}|\geq|C^{3}|+|C^{4}|$ or $|C^{3}|+|C^{2}|\geq|C^{1}|+|C^{4}|$. Hence, for each $c\in\mathcal{C}_{3}$, there exists an edge $uv\in E(D_{c}[C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{1}\cup C^{2}])$ or $uv\in E(D_{c}[C^{2}\cup C^{3},C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2})])$. By a similar argument as above, and considering whether the vertex $v$ belongs to $V_{\rm bad}$ or not, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
In what follows, it suffices to consider the case that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k<|C^{2}|-|C^{4}|-\sigma=r-s_{1}-2s_{2}-\sigma$. Move all vertices inside $Q_{1},\ldots,Q_{t}$, all vertices in $(V(Q_{1}^{\prime})\cup\ldots\cup V(Q_{k}^{\prime}))\cap C^{2}$ along with their in-neighbors in $C^{3}$ to $C^{4}$. Denote the resulting partition by $\tilde{C}^{1}\cup\tilde{C}^{2}\cup\tilde{C}^{3}\cup\tilde{C}^{4}$. Let $\tilde{\mathcal{C}}:=\mathcal{C}\setminus({\rm col}(\bigcup_{i\in[t]}Q_{i})\cup{\rm col}(\bigcup_{i\in[k]}Q_{i}^{\prime}[C^{3},C^{2}]))$. Therefore,
$|\tilde{C}^{4}|\leq\frac{|C^{2}|+|C^{4}|-r}{2}+s_{2}+2(r-s_{1}-2s_{2}-\sigma)\leq\frac{|C^{2}|+|C^{4}|}{2}+50{\delta}n,$
$|\tilde{C}^{2}|\geq\frac{|C^{2}|+|C^{4}|+r}{2}-s_{1}-s_{2}-2(r-s_{1}-2s_{2}-\sigma)\geq\frac{|C^{2}|+|C^{4}|}{2}-50{\delta}n,$
$|\tilde{\mathcal{C}}|=|\mathcal{C}|-(|E(Q_{1})|+\cdots+|E(Q_{t})|)-k\geq(1-25\delta)n.$
By the maximality of $\{Q_{1},\ldots,Q_{t}\}$, we know that $D_{i}[\tilde{C}^{2}]=D_{i}[\tilde{C}^{2},\tilde{C}^{1}]=D_{i}[\tilde{C}^{3},\tilde{C}^{2}]=\emptyset$ for all color $i\in\tilde{\mathcal{C}}$. It follows from $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$ that $\left\lceil\frac{n}{2}\right\rceil\leq|\tilde{C}^{3}|+|\tilde{C}^{4}|$ and $\left\lceil\frac{n}{2}\right\rceil\leq|\tilde{C}^{1}|+|\tilde{C}^{4}|$. Therefore,
$|\tilde{C}^{1}|+|\tilde{C}^{2}|+|\tilde{C}^{3}|+|\tilde{C}^{4}|=n\leq\left\lceil\frac{n}{2}\right\rceil+\left\lceil\frac{n}{2}\right\rceil\leq|\tilde{C}^{1}|+|\tilde{C}^{3}|+2|\tilde{C}^{4}|,$
which implies that $|\tilde{C}^{2}|\leq|\tilde{C}^{4}|$. We complete our proof by the subsequent lemma (i.e., Lemma 5.16), which leads to a contradiction by showing that $\mathcal{D}$ contains a transversal directed Hamilton cycle. ∎
Lemma 5.16.
Suppose $0<\frac{1}{n}\ll\delta\ll 1$ and $0\leq\gamma<\delta$. Let $\mathcal{C}$ be a set of $n$ colors, and $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ such that $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$. Assume that
  • $C^{1}\cup C^{2}\cup C^{3}\cup C^{4}$ is a partition of $V$ with $|C^{2}|\geq 50\delta^{\frac{1}{4}}$, $\min\{|C^{3}|+|C^{4}|,|C^{1}|+|C^{4}|\}\geq\left\lceil\frac{n}{2}\right\rceil$ and $|C^{4}|=\left\lfloor\frac{|C^{2}|+|C^{4}|+1}{2}\right\rfloor+\gamma n$,
  • $\mathcal{C}^{\prime}\cup\mathcal{C}^{\prime\prime}$ is a partition of $\mathcal{C}$ with $|\mathcal{C}^{\prime\prime}|\leq\delta n$,
  • $D_{i}[C^{2}]=D_{i}[C^{2},C^{1}]=D_{i}[C^{3},C^{2}]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$.
Then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Proof.
It is easy to see that $|C^{3}|+|C^{4}|+|C^{1}|+|C^{4}|=n+2\gamma n+\sigma$ and $\max\{|C^{3}|+|C^{4}|,|C^{1}|+|C^{4}|\}\leq\left\lceil\frac{n}{2}\right\rceil+2\gamma n+\sigma$, where $\sigma=1$ if $|C^{2}|+|C^{4}|$ is odd and $\sigma=0$ otherwise. Without loss of generality, assume that $|C^{3}|+|C^{4}|=\left\lceil\frac{n}{2}\right\rceil+\gamma_{1}n$ and $|C^{1}|+|C^{4}|=\left\lceil\frac{n}{2}\right\rceil+\gamma_{2}n$ with $2\left\lceil\frac{n}{2}\right\rceil+\gamma_{1}n+\gamma_{2}n=n+2\gamma n+\sigma$. Define $Y_{1}:=\{v\in C^{4}:d_{i}^{+}(v,C^{2})\leq(1-\delta^{\frac{1}{4}})|C^{2}|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$, $Y_{2}:=\{v\in C^{4}:d_{i}^{-}(v,C^{2})\leq(1-\delta^{\frac{1}{4}})|C^{2}|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$, and $Y:=Y_{1}\cup Y_{2}$. It is routine to check that
$\left\lceil\frac{n}{2}\right\rceil|C^{2}||\mathcal{C}^{\prime}|$ $\leq\sum_{i\in\mathcal{C}^{\prime}}|E({D_{i}}[C^{2},C^{3}\cup C^{4}])|=\sum_{i\in\mathcal{C}^{\prime}}|E({D_{i}}[C^{2},C^{3}])|+\sum_{i\in\mathcal{C}^{\prime}}|E({D_{i}}[C^{2},C^{4}])|$
$\leq|C^{2}||C^{3}||\mathcal{C}^{\prime}|+|Y_{2}|(1-\delta^{\frac{1}{4}})|C^{2}|\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|+|Y_{2}||C^{2}|(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|+(|C^{4}|-|Y_{2}|)|C^{2}||\mathcal{C}^{\prime}|$
$=(|C^{4}|-\delta^{\frac{1}{2}}|Y_{2}|+|C^{3}|)|C^{2}||\mathcal{C}^{\prime}|.$ (5)
Thus, $|Y_{2}|\leq 3\sqrt{\delta}n$. Similarly, $|Y_{1}|\leq 3\sqrt{\delta}n$. Notice that Claim 5.5 and Claim 5.15 hold by setting $V_{\rm bad}:=Y$ and $\mathcal{C}_{3}:=\mathcal{C}^{\prime}$.
For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp. $v\in Y_{2}\setminus Y_{1}$), we have $d_{i}^{-}(v,C^{2})\geq(1-\delta^{\frac{1}{4}})|C^{2}|$ (resp. $d_{i}^{+}(v,C^{2})\geq(1-\delta^{\frac{1}{4}})|C^{2}|$) for at least $(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|$ colors $i\in\mathcal{C}^{\prime}$. For a vertex $y\in Y_{2}$, one has $d_{i}^{-}(y,C^{1}\cup C^{4})\geq 40\sqrt{\delta}n$ for at least ${\delta}^{\frac{1}{4}}|\mathcal{C}^{\prime}|-1$ colors $i\in\mathcal{C}^{\prime}$. Otherwise, by the definition of $Y_{2}$, there exists a color $c\in\mathcal{C}^{\prime}$ such that $d_{c}^{-}(y,C^{2})\leq(1-\delta^{\frac{1}{4}})|C^{2}|$ and $d_{c}^{-}(y,C^{1}\cup C^{4})\leq 40\sqrt{\delta}n$. Together with $|C^{2}|\geq 50{\delta}^{\frac{1}{4}}n$, one has
$d_{c}^{-}(y)\leq 40\sqrt{\delta}n+(1-\delta^{\frac{1}{4}})|C^{2}|+|C^{3}|\leq\left\lceil\frac{n}{2}\right\rceil-\delta^{\frac{1}{4}}|C^{2}|+40\sqrt{\delta}n<\left\lceil\frac{n}{2}\right\rceil,$
a contradiction. Similarly, each vertex in $Y_{1}$ has at least $40\sqrt{\delta}n$ out-neighbors inside $C^{3}\cup C^{4}$ for at least ${\delta}^{\frac{1}{4}}|\mathcal{C}^{\prime}|-1$ digraphs $D_{i}$ with $i\in\mathcal{C}^{\prime}$.
Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we have $|E(D_{i}[C^{2},C^{3}\cup C^{4}])|,|E(D_{i}[C^{1}\cup C^{4},C^{2}])|\geq|C^{2}|\left\lceil\frac{n}{2}\right\rceil$ for each $i\in\mathcal{C}^{\prime}$. This implies that for each $i\in\mathcal{C}^{\prime}$, there are at most $|C^{2}|(\left\lceil\frac{n}{2}\right\rceil+\gamma_{1}n)-|C^{2}|\left\lceil\frac{n}{2}\right\rceil=|C^{2}|\gamma_{1}n$ non-edges in $D_{i}[C^{2},C^{3}\cup C^{4}]$ and at most $|C^{2}|\gamma_{2}n$ non-edges in $D_{i}[C^{1}\cup C^{4},C^{2}]$. Therefore, Hence
$\sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[Y_{1},C^{2}])|\geq(|Y_{1}|-\gamma_{1}n)|C^{2}||\mathcal{C}^{\prime}|\ \text{and}\ \sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[C^{2},Y_{2}])|\geq(|Y_{2}|-\gamma_{2}n)|C^{2}||\mathcal{C}^{\prime}|.$
We proceed our proof by considering the values of $|Y_{1}|$ and $|Y_{2}|$.
Case 1. $|Y_{1}|\geq\gamma_{1}n,\ |Y_{2}|\geq\gamma_{2}n$.
In view of Lemma 5.13, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma_{1}n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ inside $\{D_{i}[Y_{1},C^{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma_{2}n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[C^{2},Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$. Therefore, by using colors in $\mathcal{C}^{\prime}$,
  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $C^{2}$,
  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$, starting at $C^{4}$ and ending at $C^{2}$,
  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$, starting at $C^{2}$ and ending at $C^{4}$,
  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$, $P_{4}$ or $P_{5}$ with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $C^{4}$.
For each $y\in Y$, denote the rainbow directed path with center $y$ by $P_{y}=y^{1}*y*y^{2}$ with colors in $\mathcal{C}_{3}$. Let $\mathbf{P}:=\{P_{y}:y\in Y\}$.
It is routine to verify that there exists a rainbow matching inside $D_{i}^{\pm}[C^{2}\setminus V(\mathbf{P}),C^{4}\setminus V(\mathbf{P})]\cup D_{i}[C^{1}]\cup D_{i}[C^{3}]\cup(\bigcup_{k\in[4]}D_{i}[C^{k},C^{k+1}]):i\in\mathcal{C}^{\prime\prime}\}$, say $M$. For all but at most one unused colors in $\mathcal{C}_{\rm bad}$, we can find a rainbow directed $P_{3}$ with center in $Y$ and endpoints in $Z$, where $\{Y,Z\}=\{C^{1},C^{3}\}$. If some color $c\in\mathcal{C}_{\rm bad}$ remains unused, then $D_{c}[C^{1},C^{3}]$ is nearly a complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost complete graphs. Then, by a similar argument to that in Theorem 5.14, $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Case 2. At most one of $|Y_{1}|\geq\gamma_{1}n$ and $|Y_{2}|\geq\gamma_{2}n$ holds.
We only consider the case that $|Y_{1}|\geq\gamma_{1}n$ and $|Y_{2}|<\gamma_{2}n$; other cases can be discussed similarly. Notice that vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed paths with centers in $Y_{1}$ and ending at $C^{4}\setminus(Y_{1}\cup Y_{2})$. In particular, those rainbow directed paths can be chosen such that none of them contains edges inside $\mathcal{D}[C^{4},C^{3}]$ or $\mathcal{D}[C^{1},C^{4}]$. Furthermore, vertices in $Y_{2}\setminus Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{2}\setminus Y_{1}$ and starting at $C^{4}\setminus Y$, using colors in $\mathcal{C}^{\prime}$.
In the digraph collection $\mathcal{D}[C^{4}\setminus Y]$, extend those directed rainbow paths or choose other disjoint directed rainbow directed paths into a set of disjoint maximal rainbow directed paths. Let $\mathbf{P}:=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow paths in the above, each of which has at least one endpoint in $C^{4}\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $C^{4}$. Let $\{e_{1},e_{2},\ldots,e_{r}\}$ be a set consisting of all disjoint rainbow edges in $\mathcal{D}[C^{4}\setminus Y,C^{3}]$, and let $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{k}\}$ be a set of disjoint maximal directed rainbow $P_{3}$ in $\mathcal{D}^{\pm}[C^{4}\setminus Y,C^{1}]$ with centers in $C^{4}$ and endpoints in $C^{1}$, each of which is chosen while avoiding previously used vertices and colors.
Subcase 2.1. $|C^{2}|\geq|C^{4}|-k-r-(s_{1}+\cdots+s_{t})-\sigma$.
In this subcase, there exist two integers $r^{\prime}\in[r]$, $k^{\prime}\in[k]$ and a set of disjoint rainbow directed paths $\mathbf{P}^{\prime}=\{{Q}_{1}^{\prime},{Q}_{2}^{\prime},\ldots,{Q}_{\ell}^{\prime}\}$ such that $|C^{4}\setminus V(\mathbf{P}^{\prime})|-r^{\prime}-k^{\prime}-\sigma=|C^{2}|-\ell$, where $Y_{1}^{\prime}\cup Y_{2}\subseteq V(\mathbf{P}^{\prime})$ and the endpoints of $\tilde{Q}_{i}$ are not in $Y$ for all $i\in[\ell]$. By Claim 5.15, we connect all the rainbow directed paths in $\mathbf{P}^{\prime}$ into a single rainbow path $P^{1}$, whose endpoints lie in different parts. Since $|C^{4}|-|C^{2}|=2\gamma n+\sigma$, we have $\ell\leq|E(Q_{1}^{\prime})|+\cdots+|E(Q_{\ell}^{\prime})|=2\gamma n+\sigma$. Therefore, $|E(P^{1})|\leq 6\gamma n+3\sigma$ and $|C^{4}\setminus V(P^{1})|-\sigma=|C^{2}\setminus V(P^{1})|$. By a similar discussion as Subcase 2.1 of Theorem 5.14, $\mathcal{D}$ contains a transversal directed Hamilton cycle.
Subcase 2.2. $|C^{2}|<|C^{4}|-k-r-(s_{1}+\cdots+s_{t})-\sigma$.
In this subcase, $\sum_{i=1}^{t}s_{i}+k+r<2\gamma n$. Assume that $V(Q_{i}[C^{4}]):=\{v_{s_{1}+\cdots+s_{i-1}+i},\ldots,v_{s_{1}+\cdots+s_{i}+i}\}$ for each $i\in[t]$. Let $w$ be an unused vertex in $C^{4}$ and $c_{1},c_{2}$ be two unused colors in $\mathcal{C}$. By the maximality of $\mathbf{P}$, in $D_{c_{1}}$ (resp. $D_{c_{2}}$), $w$ cannot be adjacent to the start vertices of $Q_{i}[C^{4}]$ (resp. adjacent from the end vertices of $Q_{i}[C^{4}]$) for all $i\in[t]$.
Define
$I_{1}:=\{i:v_{i}\in N_{{c_{1}}}^{-}(w)\}\ \text{and}\ I_{2}:=\{i:v_{i+1}\in N_{{c_{2}}}^{+}(w)\}.$
Hence $I_{1},I_{2}\subseteq\bigcup_{i\in[t]}\left[\sum_{j\in[i-1]}s_{j}+i,\sum_{j\in[i]}s_{j}+i-1\right]$ with $s_{0}=0$. By the maximality of $\mathbf{P}$, we know that
  • $I_{1}\cap I_{2}=\emptyset$, which implies $|I_{1}|+|I_{2}|\leq s_{1}+\cdots+s_{t}$,
  • each vertex in $N_{c_{1}}^{+}(w,C^{3})$ must be an end vertex of some $e_{i}$ for $i\in[r]$ (hence $|N_{c_{1}}^{+}(w,C^{3})|\leq r$),
  • each vertex in $N_{c_{2}}^{-}(w,C^{1})$ must be a start vertex of some $\tilde{Q}_{i}$ for $i\in[k]$ (hence $|N_{c_{2}}^{-}(w,C^{1})|\leq k$).
Therefore,
$n\leq$ $d_{D_{c_{1}}}^{+}(w)+d_{D_{c_{2}}}^{-}(w)\leq s_{1}+\cdots+s_{t}+r+|C^{3}|+|C^{2}|+k+|C^{1}|+|C^{2}|$
$\leq$ $2\gamma n-1+n-|C^{4}|+|C^{2}|=n+2\gamma n-1-(2\gamma n+\sigma)=n-1-\sigma,$
a contradiction. ∎
6 Concluding remarks
In this paper, we generalize Ghouila-Houri’s theorem [19] (Theorem 1.3) to transversal version, which solves a problem proposed by Chakraborti, Kim, Lee, Seo [6]. Thereby, the transversal version of Dirac’s theorem can be deduced, which was obtained by Joos and Kim [21].
An anti-directed cycle is a digraph in which the underlying graph forms a cycle and no pair of consecutive edges forms a directed path. DeBiasio and Molla [17] proved the the anti-directed Hamilton cycle is guaranteed to appear in a digraph $D$ if $\delta^{0}(D)\geq\frac{n}{2}+1$. DeBiasio, Kühn, Molla, Osthus and Taylor [16] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one. It would be interesting to generalize the above two results to transversal version.
An oriented graph is a digraph with no cycle of length two. Keevash, Kühn, and Osthus [23] proved the oriented version of Dirac’s theorem, which states that every $n$-vertex digraph with minimum semidegree at least $\frac{3n-4}{8}$ contains a directed Hamilton cycle. It would be interesting to consider a transversal version of this theorem.
Question 6.1.
Let $\mathcal{D}=\left\{D_{1},\ldots,D_{n}\right\}$ be a collection of oriented graphs with common vertex set $V$ of size $n.$ If $\delta^{0}(\mathcal{D})\geq\lceil\frac{3n-4}{8}\rceil$, does $\mathcal{D}$ contain a transversal directed Hamilton cycle?
We remark that our main technique used in this paper cannot be directly generalized to the above oriented case. This is because the absorbing structure we used does not exist in the above oriented graph collections since their semi-degree is much lower than $\frac{n}{2}$. To overcome this difficulty, one has to introduce new stable conditions and build an absorption structure using Lemma 3.1. It seems that new ideas are needed to prove the tight bound in Question 6.1.
Appendix A Regularity for digraph collections
In this section, we prove the regularity lemma for digraph collections (i.e., Lemma 2.2). Before proceeding with the proof, we first make some preparations, including some necessary definitions and key theorems.
A $k$-uniform hypergraph (or $k$-graph) $\mathscr{F}=(V,E)$ consists of a vertex set $V$ and an edge set $E$ which is a family of $k$-element subsets of $V$, i.e., $E\subseteq\binom{V}{k}.$ For a vertex $v\in V$, the degree of $v$ in $\mathscr{F}$, denoted by $\deg_{\mathscr{F}}(v)$, is the number of edges containing $v$. Given a $k$-graph $\mathscr{F}$, and $k$ disjoint sets $V_{1},V_{2},\ldots,V_{k}\subseteq V(\mathscr{F})$, define $\mathscr{F}[V_{1},V_{2},\ldots,V_{k}]$ to be the subhypergraph of $\mathscr{F}$ with edge set $\left\{\{v_{1},v_{2},\ldots,v_{k}\}\in E(\mathscr{F}):v_{i}\in V_{i}\ {\text{for all}}\ i\in[k]\right\}$. We call
$d_{\mathscr{F}}(V_{1},V_{2},\ldots,V_{k}):=\frac{e(\mathscr{F}[V_{1},V_{2},\ldots,V_{k}])}{|V_{1}||V_{2}|\ldots|V_{k}|}$
the density of the tuple $(V_{1},V_{2},\ldots,V_{k})$ in $\mathscr{F}$. If $d_{\mathscr{F}}(V_{1},V_{2},\ldots,V_{k})=0$, then we say that $(V_{1},V_{2},\ldots,V_{k})$ is empty in $\mathscr{F}$. Given $\epsilon>0$ and $d\in[0,1)$, we say that $(V_{1},V_{2},\ldots,V_{k})$ is $\epsilon$-regular in $\mathscr{F}$ if for all $i\in[k]$, whenever $V_{i}^{\prime}\subseteq V_{i}$ with $|V_{i}^{\prime}|\geq\epsilon|V_{i}|$ one has
$|d_{\mathscr{F}}(V_{1}^{\prime},V_{2}^{\prime},\ldots,V_{k}^{\prime})-d_{\mathscr{F}}(V_{1},V_{2},\ldots,V_{k})|\leq\epsilon$,
and the tuple is $(\epsilon,d)$-regular if it is $\epsilon$-regular and $d_{\mathscr{F}}(V_{1},V_{2},\ldots,V_{k})\geq d$. The following result was proved by Chung [14], and its proof follows the approach of the original Regularity Lemma for graphs [27].
Theorem A.1 (Weak hypergraph regularity lemma).
For all integers $k\geq 2$, $L_{0}\geq 1$, and every $\epsilon>0$ there exists $N=N(\epsilon,L_{0},k)$ such that if $\mathscr{F}$ is a $k$-graph on $n\geq N$ vertices, then $V(\mathscr{F})$ has a partition $V_{0},V_{1},\ldots,V_{L}$ such that the following hold:
  • $L_{0}\leq L\leq N$ and $|V_{0}|\leq\epsilon n$,
  • $|V_{1}|=\cdots=|V_{L}|$,
  • for all but at most $\epsilon\binom{L}{k}$ $k$-tuples $\{i_{1},\ldots,i_{k}\}\in\binom{[L]}{k}$, we have that $(V_{i_{1}},\ldots,V_{i_{k}})$ is $\epsilon$-regular in $\mathscr{F}$.
Given two partitions $V_{0},V_{1},\ldots,V_{k}$ and $U_{1},\ldots,U_{\ell}$ of a vertex set, we say that $V_{0},V_{1},\ldots,V_{k}$ refines $U_{1},\ldots,U_{\ell}$ if for all $V_{i}$ with $i\in[k]$ there is $U_{j}$ for some $j\in[\ell]$ that contains $V_{i}$. For some integer $s>0$, let $\mathscr{F}_{1},\mathscr{F}_{2},\ldots,\mathscr{F}_{s}$ be $s$ $K$-partite $k$-graphs sharing a common vertex set $V:=V_{1}\cup\ldots\cup V_{K}$. As a consequence of the proof of Theorem A.1, we can deduce the following corollary.
Corollary A.2.
For all integers $K\geq k\geq 2$, $s\geq 1$, $L_{0}\geq 1$, and every $\epsilon>0$, there exists $N=N(\epsilon,s,L_{0},K,k)$ such that if $\mathscr{F}_{1},\mathscr{F}_{2},\ldots,\mathscr{F}_{s}$ are $s$ $K$-partite $k$-graphs sharing a common vertex set $V:=V_{1}\cup\ldots\cup V_{K}$ of size $n\geq N$, then $V$ can be refined into a partition $U_{0},U_{1},\ldots,U_{L}$ satisfying the following properties:
  • $L_{0}\leq L\leq N$ and $|U_{0}|\leq\epsilon n$,
  • $|U_{1}|=\cdots=|U_{L}|$,
  • for each $i\in[s]$ and for all but at most $\epsilon\binom{L}{k}$ $k$-tuples $\{i_{1},\ldots,i_{k}\}\in\binom{[L]}{k}$, we have that $(U_{i_{1}},\ldots,U_{i_{k}})$ is $\epsilon$-regular in $\mathscr{F}_{i}$.
The degree form of the weak hypergraph regularity lemma [29] is proved in the same way as the original degree form of the regularity lemma, which in fact can be derived from Theorem A.1 via some cleaning.
Theorem A.3 (Degree form of the weak hypergraph regularity lemma, [29]).
For all integers $k\geq 2$, $L_{0}\geq 1$ and every $\epsilon>0$, there exists $N=N(\epsilon,L_{0},k)$ such that for every $d\in[0,1)$ and for every $k$-graph $\mathscr{F}$ on $n\geq N$ vertices, there exists a partition of $V(\mathscr{F})$ into $V_{0},V_{1},\ldots,V_{L}$ and a spanning subhypergraph $\mathscr{F}^{\prime}$ of $\mathscr{F}$ such that the following properties hold:
  • $L_{0}\leq L\leq N$ and $|V_{0}|\leq\epsilon n$,
  • $|V_{1}|=\cdots=|V_{L}|$,
  • ${\rm deg}_{\mathscr{F}^{\prime}}(v)\geq{\rm deg}_{\mathscr{F}}(v)-(d+\epsilon)n^{k-1}$ for all $v\in V(\mathscr{F})$,
  • every edge of $\mathscr{F}^{\prime}$ with more than one vertex in a single cluster $V_{i}$ for some $i\in[L]$ has at least one vertex in $V_{0}$,
  • for all $k$-tuples $\{i_{1},i_{2},\ldots,i_{k}\}\in\binom{[L]}{k}$, we have that $(V_{i_{1}},V_{i_{2}},\ldots,V_{i_{k}})$ is either empty or $(\epsilon,d)$-regular in $\mathscr{F}^{\prime}$.
In fact, from the proof of Theorem A.3 in [29] and Corollary A.2 we can derive the following stronger version of the conclusion.
Theorem A.4.
For all integers $K\geq k\geq 2$, $s\geq 1$, $L_{0}\geq 1$, and every $\epsilon>0$, there exists $N=N(\epsilon,s,L_{0},K,k)$ such that, for any $s$ $K$-partite $k$-graphs $\mathscr{F}_{1},\mathscr{F}_{2},\ldots,\mathscr{F}_{s}$ sharing a common vertex set $V:=V_{1}\cup\ldots\cup V_{K}$ of size $n\geq N$, we can find a refinement partition $U_{0},U_{1},\ldots,U_{L}$ of $V$, and a spanning subhypergraph $\mathscr{F}_{i}^{\prime}\subseteq\mathscr{F}_{i}$ for each $i\in[s]$ such that the following properties hold:
  • $L_{0}\leq L\leq N$ and $|U_{0}|\leq\epsilon n$,
  • $|U_{1}|=\cdots=|U_{L}|$,
  • for each $i\in[s]$ and each $v\in V$, ${\rm deg}_{\mathscr{F}^{\prime}_{i}}(v)\geq{\rm deg}_{\mathscr{F}_{i}}(v)-(d+\epsilon)n^{k-1}$,
  • for each $i\in[s]$, every edge of $\mathscr{F}_{i}^{\prime}$ with more than one vertex in a single cluster $V_{j}$ for some $j\in[L]$ has at least one vertex in $V_{0}$,
  • for each $i\in[s]$ and all $k$-tuples $\{i_{1},i_{2},\ldots,i_{k}\}\in\binom{[L]}{k}$, we have that $(U_{i_{1}},U_{i_{2}},\ldots,U_{i_{k}})$ is either empty or $(\epsilon,d)$-regular in $\mathscr{F}^{\prime}_{i}$.
Since the proof of Theorem A.4 closely follows the proof of Theorem A.3, requiring only a repeated application of the cleaning operation on $s$ hypergraphs, we omit it here for brevity.
Let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs on a common vertex set $[n]$. Based on $\mathcal{D}$, we construct the auxiliary $4$-graph $\mathscr{H}$ as follows:
$V(\mathscr{H})=[n]\cup\mathcal{C}\cup S_{1}\cup S_{2},\ \text{where}\ |S_{1}|=|S_{2}|=n,$
$E(\mathscr{H})=\left\{\{i,j,c,x\}:i,j\in[n]\ \text{with}\ ij\in E(D_{c})\ {\text{for some}}\ c\in\mathcal{C}\ \text{and}\ x\in S_{1}\ \text{if}\ i<j,\ x\in S_{2}\ \text{if}\ i>j\right\}.$
Now, we prove Lemma 2.2.
Proof of Lemma 2.2.
By increasing $L_{0}$ and decreasing $\epsilon$, $\delta$ as necessary, we may assume that $0<\frac{1}{L_{0}}\ll\epsilon\ll\delta\ll 1$. Let $L_{1}:=\frac{24L_{0}}{\delta}$. Choose further positive constants $\epsilon^{\prime},\alpha$ satisfying $\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1$. Let $n_{1}$ be derived from Theorem A.3 by applying it with the parameters $(\epsilon^{\prime},L_{1},4)$. Then, $n_{2}$ is obtained from Theorem A.4 using the parameters $(\epsilon^{\prime},2,L_{0},n_{1},3)$. Let $n_{0}:=\frac{4n_{2}}{\alpha}$. By increasing $n_{0}$, we may assume that $L_{1}\ll n_{1}\ll n_{2}\ll n_{0}$. Altogether,
$0<\frac{1}{n_{0}}\ll\frac{1}{n_{2}}\ll\frac{1}{n_{1}}\ll\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1$
Let $n\geq n_{0}$ be an integer and $\mathcal{D}$ be a digraph collection on vertex set $V:=[n]$ and with color set $\mathcal{C}$, where $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$. Let $d\in[0,1)$ and let $\mathscr{H}$ be the auxiliary $4$-graph obtained from $\mathcal{D}$ with vertex set $U:=V\cup\mathcal{C}\cup S_{1}\cup S_{2}$. Theorem A.3 implies that there is a partition $U_{0},U_{1},\ldots,U_{K}$ of $U$ and a spanning subhypergraph $\mathscr{H}_{0}$ of $\mathscr{H}$ such that
  • (a)
    $L_{1}\leq K\leq n_{1}$ and $|U_{0}|\leq\epsilon^{\prime}|U|$,
  • (b)
    $|U_{1}|=\cdots=|U_{K}|=:m^{\prime}$,
  • (c)
    ${\rm deg}_{\mathscr{H}_{0}}(v)\geq{\rm deg}_{\mathscr{H}}(v)-(d+\epsilon^{\prime})|U|^{3}$ for all $v\in U$,
  • (d)
    every edge of $\mathscr{H}_{0}$ with more than one vertex in a single cluster $U_{i}$ for some $i\in[K]$ has at least one vertex in $U_{0}$,
  • (e)
    for all $4$-tuples $\{i_{1},i_{2},i_{3},i_{4}\}\in\binom{[K]}{4}$, we have that $(U_{i_{1}},U_{i_{2}},U_{i_{3}},U_{i_{4}})$ is either empty or $(\epsilon^{\prime},d)$-regular in $\mathscr{H}_{0}$.
Delete all edges $e\in E(\mathscr{H}_{0})$ satisfying $|e\cap U_{i}|\geq 2$ for some $i\in[K]$ and $e\cap(\mathcal{C}\cap U_{0})=\emptyset$. Denote the resulting spanning subhypergraph by $\mathscr{H}^{\prime}$. Note that for every $v\in U$, the number of edges incident to $v$ that are deleted is fewer than
$\epsilon^{\prime}|U|^{3}+Km^{\prime 2}|U|\leq\epsilon^{\prime}|U|^{3}+m^{\prime}|U|^{2}\leq\epsilon^{\prime}|U|^{3}+\frac{|U|^{3}}{L_{1}}\leq 2\epsilon^{\prime}|U|^{3}.$
Combining $(c)$, we have
${\rm deg}_{\mathscr{H}^{\prime}}(v)\geq{\rm deg}_{\mathscr{H}}(v)-(d+3\epsilon^{\prime})|U|^{3}$ (6)
for all $v\in U$. Partition each cluster $U_{i}$, $i\in[K]$, into $\frac{1}{\alpha}$ subclusters of size at most $m_{0}$ so that all but at most four subclusters of $U_{i}$ have size exactly $m_{0}$ and the property that they lie entirely within one of $V$, $\mathcal{C}$, $S_{1}$ and $S_{2}$. If a subcluster does not have this property, then add it to $U_{0}$. The new exceptional set has size at most $|U_{0}|+4\alpha m^{\prime}K$, and let $V_{0}$ be its intersection with $V$, $\mathcal{C}_{0}$ be its intersection with $\mathcal{C}$. Relabel the subclusters so that those which are subsets of $V$ are $V_{1},\ldots,V_{L}$ and those which are subsets of $\mathcal{C}$ are $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$. For each $c\in\mathcal{C}$, let $D_{c}^{\prime}$ be the spanning sub-digraph of $D_{c}$ with vertex set $V=:[n]$ and edge set
$\left\{uv:u<v,\ x\in S_{1}\ \text{and}\ \{u,v,c,x\}\in\mathscr{H}^{\prime}\right\}\cup\left\{vu:u<v,\ y\in S_{2}\ \text{and}\ \{u,v,c,y\}\in\mathscr{H}^{\prime}\right\}.$
We claim that the properties of the Lemma 2.2 (i)-(iv) are satisfied.
Property (i). Note that $m^{\prime}K\leq|U|$. Since $\epsilon^{\prime}\ll\epsilon$, one has
$|V_{0}|+|\mathcal{C}_{0}|\leq|U_{0}|+4\alpha m^{\prime}K\leq(\epsilon^{\prime}+4\alpha)|U|=(\epsilon^{\prime}+4\alpha)\left(3n+\frac{n}{\delta}\right)\leq\frac{\epsilon n}{2}.$ (7)
Also,
$L+M\leq\frac{2K}{\alpha}\leq\frac{2n_{1}}{\alpha}\leq\frac{n_{0}}{2},$ (8)
proving the required upper bound for both $L$ and $M$. Furthermore, $m_{0}(L+M)\leq n+|\mathcal{C}|$ and
$\frac{L_{1}}{6}\leq\frac{3n+|\mathcal{C}|}{6m_{0}}\leq\frac{n+|\mathcal{C}|}{2m_{0}}\leq\frac{n+|\mathcal{C}|-\epsilon n}{m_{0}}\leq L+M,$
so
$L=\frac{n-|V_{0}|}{m_{0}}\geq\frac{(n-|V_{0}|)(L+M)}{n+|\mathcal{C}|}\geq\frac{(n-|V_{0}|)L_{1}}{6(n+\frac{n}{\delta})}\geq\frac{\delta L_{1}}{24}=L_{0},$ (9)
and similarly for $M$.
Property (ii). It follows by construction.
Property (iii). By the construction of $\mathscr{H}^{\prime}$, for all $v\in V$ we have
$\sum_{c\in\mathcal{C}}\left(d^{+}_{D_{c}^{\prime}}(v)+d^{-}_{D_{c}^{\prime}}(v)\right)\geq\frac{{\rm deg}_{\mathscr{H^{\prime}}}(v)}{n},$
and $e(D_{c}^{\prime})\geq\frac{{\rm deg}_{\mathscr{H}^{\prime}}(c)}{n}$ for all $c\in\mathcal{C}$. Furthermore,
$(d+3\epsilon^{\prime})(3n+|\mathcal{C}|)^{3}\leq\left(\frac{8d}{\delta^{3}}+\frac{24\epsilon^{\prime}}{\delta^{3}}\right)n^{3}\leq\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{3}.$ (10)
Notice that $\sum_{c\in\mathcal{C}}\left(d^{+}_{D_{c}}(v)+d^{-}_{D_{c}}(v)\right)=\frac{{\rm deg}_{\mathscr{H}}(v)}{n}$ for all $v\in V$. Thus,
$\sum_{c\in\mathcal{C}}d^{+}_{D_{c}^{\prime}}(v)\geq$ $\frac{{\rm deg}_{\mathscr{H^{\prime}}}(v)}{n}-\sum_{c\in\mathcal{C}}d^{-}_{D_{c}^{\prime}}(v)$
$\geq$ $\frac{{\rm deg}_{\mathscr{H}}(v)}{n}-\sum_{c\in\mathcal{C}}d^{-}_{D_{c}}(v)-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}$
$=$ $\sum_{c\in\mathcal{C}}d^{+}_{D_{c}}(v)-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}$ (11)
for all $v\in V$, and similarly for $\sum_{c\in\mathcal{C}}d^{-}_{D_{c}^{\prime}}(v)$. Additionally, observe that $e(D_{c})=\frac{{\rm deg}_{\mathscr{H}}(c)}{n}$ for all $c\in\mathcal{C}$. It follows that
$e(D_{c}^{\prime})\geq\frac{{\rm deg}_{\mathscr{H}^{\prime}}(c)}{n}\geq\frac{{\rm deg}_{\mathscr{H}}(c)}{n}-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}=e(D_{c})-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}$ (12)
for all $c\in\mathcal{C}$.
Property (iv). It follows by construction.
Property (v). Suppose that $(V_{0}\cup\mathcal{C}_{0},V_{1},\ldots,V_{L},\mathcal{C}_{1},\ldots,\mathcal{C}_{M})$ does not satisfy property (v), we are to refine it and get a desired partition. Now, define the following auxiliary $3$-graphs $\mathscr{H}_{1}$ and $\mathscr{H}_{2}$ on the common vertex set $U^{\prime}:=([n]\cup\mathcal{C})\setminus(V_{0}\cup\mathcal{C}_{0})$ and
$E(\mathscr{H}_{1})=\left\{\{u,v,c\}:uv\in D^{\prime}_{c}[V_{i},V_{j}],u,v\in V\setminus V_{0},c\in\mathcal{C}\setminus\mathcal{C}_{0},1\leq i<j\leq L\right\}$, $E(\mathscr{H}_{2})=\left\{\{u,v,c\}:uv\in D^{\prime}_{c}[V_{j},V_{i}],u,v\in V\setminus V_{0},c\in\mathcal{C}\setminus\mathcal{C}_{0},1\leq i<j\leq L\right\}$.
Notice that we can view partition $P=(V_{1},\ldots,V_{L},\mathcal{C}_{1},\ldots,\mathcal{C}_{M})$ as comprising two different partitions $P_{1}$, $P_{2}$ of hypergraphs $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ (both partitions divide $V$ and $\mathcal{C}$ in the same way, but the sets of edges among the partition sets are different). By increasing $n_{2}$, we may assume that $n_{2}\gg N(\epsilon^{\prime},2,L_{0},L+M,3)$ (which is defined in Theorem A.4). Consequently, $n_{0}\gg N(\epsilon^{\prime},2,L_{0},L+M,3)$ as well. Applying Theorem A.4, we can find a refinement partition of $P$, say $U_{0}^{\prime},U_{1}^{\prime},\ldots,U_{K^{\prime}}^{\prime}$, and spanning subhypergraphs $\mathscr{H}_{1}^{\prime}\subseteq\mathscr{H}_{1}$ and $\mathscr{H}_{2}^{\prime}\subseteq\mathscr{H}_{2}$ such that the following hold:
  • (a’)
    $L_{0}\leq K^{\prime}\leq n_{2}$ and $|U_{0}^{\prime}|\leq\epsilon^{\prime}|U^{\prime}|$,
  • (b’)
    $|U^{\prime}_{1}|=\cdots=|U^{\prime}_{K^{\prime}}|$,
  • (c’)
    for each $i\in[2]$ and each $v\in V\setminus V_{0}$, ${\rm deg}_{\mathscr{H}^{\prime}_{i}}(v)\geq{\rm deg}_{\mathscr{H}_{i}}(v)-(d+\epsilon^{\prime})|U^{\prime}|^{2}$,
  • (d’)
    for each $i\in[2]$, every edge of $\mathscr{H}_{i}^{\prime}$ with more than one vertex in a single cluster $U^{\prime}_{j}$ for some $j\in[K^{\prime}]$ has at least one vertex in $U^{\prime}_{0}$,
  • (e’)
    for each $i\in[2]$ and all $3$-tuples $\{i_{1},i_{2},i_{3}\}\in\binom{[K^{\prime}]}{3}$, we have that $(U_{i_{1}}^{\prime},U_{i_{2}}^{\prime},U_{i_{3}}^{\prime})$ is either empty or $(\epsilon^{\prime},d)$-regular in $\mathscr{H}^{\prime}_{i}$.
Following the same operation as earlier, let $V^{\prime}_{0}:=V\cap U_{0}^{\prime}$ and $\mathcal{C}^{\prime}_{0}:=\mathcal{C}\cap U_{0}^{\prime}$. Relabel the subclusters such that those which are subsets of $V$ are $V^{\prime}_{1},\ldots,V^{\prime}_{L^{\prime}}$ and those which are subsets of $\mathcal{C}$ are $\mathcal{C}_{1}^{\prime},\ldots,\mathcal{C}_{M^{\prime}}^{\prime}$. For each $c\in\mathcal{C}\setminus(\mathcal{C}_{0}\cup\mathcal{C}_{0}^{\prime})$, let $D_{c}^{\prime\prime}$ be the spanning sub-digraph of $D_{c}$ with vertex set $V$ and edge set
$\left\{uv:\{u,v,c\}\in\mathscr{H}_{1}^{\prime}\,\,\text{or}\,\,\{u,v,c\}\in\mathscr{H}_{2}^{\prime}\right\}\cup E(D_{c}^{\prime\pm}[V_{0},V\setminus(V_{0}\cup V_{0}^{\prime})]).$
Let $D_{i}^{*}:=D_{i}^{\prime}$ for $i\in\mathcal{C}_{0}\cup\mathcal{C}_{0}^{\prime}$ and $D_{i}^{*}:=D_{i}^{\prime\prime}$ for $i\in\mathcal{C}\setminus(\mathcal{C}_{0}\cup\mathcal{C}_{0}^{\prime})$. Let
$P^{*}:=(V_{0}\cup\mathcal{C}_{0}\cup V_{0}^{\prime}\cup\mathcal{C}_{0}^{\prime},V_{1}^{\prime},\ldots,V_{L^{\prime}}^{\prime},\mathcal{C}_{1}^{\prime}\ldots,\mathcal{C}_{M^{\prime}}^{\prime})$
be a partition of $V\cup\mathcal{C}$. Next, we will re-verify that $P^{*}$ and digraph collection $\mathcal{D}^{*}:=\{D_{i}^{*}:i\in\mathcal{C}\}$ satisfies (i)-(v) of Lemma 2.2. By the construction of $\mathscr{H}_{1}$ and $\mathscr{H}_{2}$, it is easy to see that $P^{*}$ and $\mathcal{D^{*}}$ satisfy the property (v). Thus, we only need to re-verify properties (i)-(iv).
Property (i). By (7), (9) and (a’), we have
$|V_{0}|+|\mathcal{C}_{0}|+|V_{0}^{\prime}|+|\mathcal{C}_{0}^{\prime}|\leq\frac{\epsilon n}{2}+\epsilon^{\prime}|U^{\prime}|\leq\epsilon n,$
and
$L_{0}\leq L\leq L^{\prime}\leq n_{2}\leq n_{0}\ \text{and}\ L_{0}\leq M\leq M^{\prime}\leq n_{2}\leq n_{0}.$
Property (ii). It follows from (b’) that $|V_{1}^{\prime}|=\cdots=|V_{L^{\prime}}^{\prime}|=|\mathcal{C}_{1}^{\prime}|=\cdots=|\mathcal{C}_{M^{\prime}}^{\prime}|$, as required.
Property (iii). For each vertex $v\in V$ and each color $c\in\mathcal{C}$, by the construction of $D_{c}^{*}$, (c’), (10), (A) and (12), we have
$e(D_{c}^{*})\geq e(D_{c})-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}-2\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}-\frac{4\epsilon^{\prime}n^{2}}{\delta}\geq e(D_{c})-\left(\frac{24d}{\delta^{3}}+\epsilon\right)n^{2}$
and
$\sum_{c\in\mathcal{C}}d^{+}_{D_{c}^{*}}(v)\geq\sum_{c\in\mathcal{C}}d^{+}_{D_{c}}(v)-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}-2\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}-\frac{2\epsilon^{\prime}n^{2}}{\delta^{2}}\geq\sum_{c\in\mathcal{C}}d^{+}_{D_{c}}(v)-\left(\frac{24d}{\delta^{3}}+\epsilon\right)n^{2}.$
Analogously, the same inequality holds for $\sum_{c\in\mathcal{C}}d^{-}_{D_{c}^{*}}(v)$.
Property (iv). Notice that $\mathscr{H}_{1}$ and $\mathscr{H}_{2}$ are $(L+M)$-partite hypergraphs. Thus, there is no edge of $\mathscr{H}_{i}^{\prime}$ with more than one vertex in a single cluster $U^{\prime}_{j}$ for some $j\in[K^{\prime}]$. This implies that $P^{*}$ and $\mathcal{D}^{*}$ satisfy (iv).
This completes the proof of Lemma 2.2. ∎
Appendix B Transversal perfect matching in bipartite graph collections
B.1 Proof of Lemma 2.7
Proof of Lemma 2.7.
We first add vertices with small degree to $C_{1}\cup C_{2}$. Let $\mu=\epsilon^{5}$. Since $G$ is $\epsilon$-extremal, there are $X\subseteq V_{1},Y\subseteq V_{2}$ each of size at least $\left(\frac{1}{2}-\mu\right)n$ such that $e_{G}(X,Y)\leq\mu n^{2}$. Thus, it is easy to see that for all but at most $\sqrt{\mu}n$ vertices $v$ in $X$ we have $d_{G}(v,Y)\leq\sqrt{\mu}n$. Similarly, for all but at most $\sqrt{\mu}n$ vertices $v$ in $Y$ we have $d_{G}(v,X)\leq\sqrt{\mu}n$. We add these exceptional vertices from $X$ and from $Y$ to $C_{1}$ and $C_{2}$, respectively. Now, the new sets $X$ and $Y$ each has size at least $\left(\frac{1}{2}-\mu\right)n-\sqrt{\mu}n\geq\left(\frac{1}{2}-2\sqrt{\mu}\right)n.$ Choose $X_{1}\subseteq X,Y_{1}\subseteq Y$ such that $|X_{1}|=|Y_{1}|=\left(\frac{1}{2}-2\sqrt{\mu}\right)n$.
Notice that $|C_{i}|\leq(d+\sqrt{\mu})n\leq 2\sqrt{\mu}n$ by $d\ll\epsilon$. Thus, we can choose $Z_{1}\subseteq V_{1}\setminus(X_{1}\cup C_{1})$ and $Z_{2}\subseteq V_{2}\setminus(Y_{1}\cup C_{2})$ such that $|Z_{1}|=|Z_{2}|=(\frac{1}{2}-2\sqrt{\mu})n$. Recall that $d_{G}(v,Y_{1})\leq\sqrt{\mu}n$ for each vertex $v\in X_{1}$. This implies that for each $v\in X_{1}$, one has
$d_{G}(v,Z_{2})\geq\left(\frac{1}{2}-\mu-5\sqrt{\mu}\right)n-|C_{2}|\geq\left(\frac{1}{2}-8\sqrt{\mu}\right)n.$
Similarly, $d_{G}(v,Z_{1})\geq\left(\frac{1}{2}-8\sqrt{\mu}\right)n$ for each $v\in Y_{1}$.
Since the number of edges between $X_{1}$ and $Z_{2}$ is at least $(\frac{1}{2}-2\sqrt{\mu})n\times\left(\frac{1}{2}-8\sqrt{\mu}\right)n$ and $|Z_{2}|=(\frac{1}{2}-2\sqrt{\mu})n$, there are at most $50\mu^{\frac{1}{4}}n$ vertices $v$ in $Z_{2}$ such that $d_{G}(v,X_{1})\leq(\frac{1}{2}-10\mu^{\frac{1}{4}})n$. Similarly, for all but at most $50\mu^{\frac{1}{4}}n$ vertices $v$ in $Z_{1}$ we have $d_{G}(v,Y_{1})\geq(\frac{1}{2}-10\mu^{\frac{1}{4}})n$. Thus, we can choose $A_{1}\subseteq X_{1}$, $B_{1}\subseteq Z_{1}$, $A_{2}\subseteq Y_{1}$, $B_{2}\subseteq Z_{2}$ such that $|A_{1}|=|B_{1}|=|A_{2}|=|B_{2}|=(\frac{1}{2}-\epsilon)n$ and $d_{G}(v,X_{i})\geq(\frac{1}{2}-2\epsilon)n$ for $v\in X_{3-i}$ and $X\in\{A,B\}$, $i\in[2]$. Let $C_{1}=V_{1}\setminus(A_{1}\cup B_{1})$ and $C_{2}=V_{2}\setminus(A_{2}\cup B_{2})$. Hence, $(A_{1},B_{1},C_{1},A_{2},B_{2},C_{2})$ is the desired partition. ∎
B.2 Proof of Theorem 2.10
In this subsection, we prove Theorem 2.10. Before proceeding with the proof, we first present some key lemmas that will be used in the proof.
Assume that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. For $\mu>0$, we say that $\mathcal{G}$ is $\mu$-nice if for every $A\subseteq V_{1}$ and $B\subseteq V_{2}$ of size $\lfloor\frac{n}{2}\rfloor$, we have $e_{\mathcal{G}}(A,B):=\sum_{i\in[n]}e_{G_{i}}(A,B)\geq\mu n^{3}$.
Lemma B.1.
Suppose that $0<\frac{1}{n}\ll\mu^{\prime}\ll\alpha\ll\gamma,\epsilon,\ll\delta\ll 1$. Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. If $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{G}$ is $\mu^{\prime}$-nice.
Proof.
For the sake of contradiction, suppose that $\mathcal{G}$ is not $\mu^{\prime}$-nice. Then there is $A\subseteq V_{1}$ of size $\lfloor\frac{n}{2}\rfloor$ and $B\subseteq V_{2}$ of size $\lfloor\frac{n}{2}\rfloor$ such that $e_{\mathcal{G}}(A,B)\leq\mu^{\prime}n^{3}$. Therefore, for all but at most $\sqrt{\mu^{\prime}}n$ colors $i\in[n]$, we have $e_{G_{i}}(A,B)<\sqrt{\mu^{\prime}}n^{2}$, implying that such $G_{i}$ is not $\sqrt{\mu^{\prime}}$-nice. This further implies that $\mathcal{G}$ is not $(1-\sqrt{\mu^{\prime}},\sqrt{\mu^{\prime}})$-strong stable and hence not $(\gamma,\alpha)$-strong stable, since $\mu^{\prime}\ll\gamma,\alpha$.
Without loss of generality, assume that for each $i\in[(1-\sqrt{\mu^{\prime}})n]$, $G_{i}$ is not $\sqrt{\mu^{\prime}}$-nice. Since $\epsilon>\mu^{\prime 1/10}$, $G_{i}$ is $\epsilon$-extremal for all $i\in[(1-\sqrt{\mu^{\prime}})n]$. By Lemma 2.7, the graph $G_{i}$ has a characteristic partition, denoted as $(A_{1}^{i},B_{1}^{i},C_{1}^{i},A_{2}^{i},B_{2}^{i},C_{2}^{i})$. It follows that, after possible relabeling $A_{1}^{i},B_{1}^{i},A_{2}^{i},B_{2}^{i}$, for all but at most $\delta^{2}n$ colors $i\in[(1-\sqrt{\mu^{\prime}})n]$ we have
$|A\Delta A_{1}^{i}|\leq\frac{\delta n}{2}\,\,\,\text{and}\,\,\,|B\Delta B_{2}^{i}|\leq\frac{\delta n}{2}.$
Otherwise, $e_{\mathcal{G}}(A,B)\geq\frac{\delta^{3}n^{3}}{8}>\mu^{\prime}n^{3}$ by $\mu^{\prime}\ll\epsilon\ll\delta$, a contradiction. Delete such colors from $[(1-\sqrt{\mu^{\prime}})n]$ and let $\mathcal{C}$ be the remaining color set. Thus, for every $i,j\in\mathcal{C}$ the graphs $G_{i}$ and $G_{j}$ are not $\delta$-crossing. Therefore,
$e(C_{\mathcal{G}}^{\epsilon,\delta})\leq\sqrt{\mu^{\prime}}n^{2}+\delta^{2}n^{2}<\delta n^{2}\ \text{since}\ \mu^{\prime}\ll\delta.$
Thus, $\mathcal{G}$ is not $(\epsilon,\delta)$-weakly stable, which contradicts the assumption that $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable. ∎
The following result shows that if a bipartite graph collection $\mathcal{G}$ is $\mu$-nice under the minimum degree constraint, then we can find a large rainbow matching in $\mathcal{G}$.
Lemma B.2.
Let $0<\frac{1}{n}\ll d\ll\mu\ll 1$. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$, and for each vertex $v\in V$ we have $d_{G_{i}}(v)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ colors $i\in[n]$. If $\mathcal{G}$ is $6\mu$-nice, then $\mathcal{G}$ contains a rainbow matching $\mathcal{M}$ with $e(\mathcal{M})\geq n-dn-2$.
Proof.
Let $\mathcal{M}$ be a maximum rainbow matching in $\mathcal{G}$, and suppose for a contradiction that $e(\mathcal{M})\leq n-dn-3$. Since for each vertex $v\in V$ we have $d_{G_{i}}(v)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ colors $i\in[n]$, there are vertices $v_{1}\in V_{1}\setminus V(\mathcal{M}),v_{2}\in V_{2}\setminus V(\mathcal{M})$ and distinct colors $c_{1},c_{2}\notin{\rm col}(\mathcal{M})$ such that $d_{G_{c_{\ell}}}(v_{\ell})\geq\left(\frac{1}{2}-\mu\right)n$ for $\ell\in\{1,2\}$. Put $N_{\ell}:=N_{G_{c_{\ell}}}(v_{\ell})$. Thus, $|N_{\ell}|\geq\left(\frac{1}{2}-\mu\right)n$. Let $\mathcal{C}:=[n]\setminus({\rm col}(\mathcal{M})\cup\{c_{1},c_{2}\})$. For $x\in V(\mathcal{M})$, let $x^{+}$ denote the unique neighbour of $x$ in $\mathcal{M}$, and for $A\subseteq V(\mathcal{M})$, let $A^{+}:=\{x^{+}:x\in A\}$.
By the maximality of $\mathcal{M}$, we have $N_{1}\cap N_{2}^{+}=\emptyset,\,N_{2}\cap N_{1}^{+}=\emptyset$ and $N_{1},N_{2}\subseteq V(\mathcal{M})$. This implies that
$e(\mathcal{M})\geq|N_{1}|+|N_{2}|\geq 2\left(\frac{1}{2}-\mu\right)n=n-2\mu n.$
Moreover, for each $c\in\mathcal{C}$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})=0$ by the maximality of $\mathcal{M}$. Thus, it suffices to show that for all $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}]\cup\mathcal{M}[N_{2},N_{2}^{+}])$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$. If so,
$e_{\mathcal{G}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}(n-2\mu n)+2\mu n^{3}<6\mu n^{3},$
a contradiction to $\mathcal{G}$ being $6\mu$-nice. Next, we consider the case $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}])$; the proof for the case $c\in{\rm col}(\mathcal{M}[N_{2},N_{2}^{+}])$ is similar and is omitted.
Let $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}])$ and $w\in N_{1}$ such that ${\rm col}(ww^{+})=c$. Now let $\mathcal{M}^{\prime}=\mathcal{M}-ww^{+}+v_{1}w$ be a rainbow matching with ${\rm col}(\mathcal{M}^{\prime})=({\rm col}(\mathcal{M})\setminus\{c\})\cup\{c_{1}\}$. Clearly, $\mathcal{M}^{\prime}$ is a maximum rainbow matching. Notice that there exists a color $c_{3}\in\mathcal{C}\setminus\{c_{1},c_{2}\}$ such that $d_{G_{c_{3}}}(w^{+})\geq\left(\frac{1}{2}-\mu\right)n$. Let $N_{3}:=N_{G_{c_{3}}}(w^{+})$. Consider the new vertex pair $(w^{+},v_{2})$ and color pair $(c_{3},c_{2})$. By the maximality of $\mathcal{M}^{\prime}$, we have $N_{3}^{+}\cap N_{2}=\emptyset$. Recall that $N_{1}^{+}\cap N_{2}=\emptyset$ and $|N_{i}|\geq\left(\frac{1}{2}-\mu\right)n$ for each $i\in[3]$. Thus,
$|N_{3}^{+}\Delta N_{1}^{+}|=|N_{3}^{+}\cup N_{1}^{+}|-|N_{3}^{+}\cap N_{1}^{+}|\leq(n-|N_{2}|)-(|N_{1}|+|N_{3}|-(n-|N_{2}|))\leq 4\mu n.$ (13)
On the other hand, for the maximality of matching $\mathcal{M}^{\prime}$, we have $e_{G_{c}}(N_{3}^{+},N_{2}^{+})=0$. Thus, together with (13), we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$, as desired. ∎
Let $n$ be a positive integer and let $d,\epsilon>0$ such that $0<\frac{1}{n}\ll d\ll\epsilon\leq 1$. Consider a bipartite graph collection $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$ such that $d_{G}(x)\geq(\frac{1}{2}-\epsilon^{5})n$ for all but at most $dn$ vertices $x\in V$. Lemma 2.7 implies that for each $i\in[n]$ if $G_{i}$ is $\epsilon$-extremal, then we can fix a characteristic partition
$(A_{1}^{i},B_{1}^{i},C_{1}^{i},A_{2}^{i},B_{2}^{i},C_{2}^{i})$ of $G_{i}$.
We define a vertex $v\in V$ is $i$-good if either $G_{i}$ is $\epsilon$-extremal and $v\in V\setminus(C^{1}_{i}\cup C^{2}_{i})$ or $G_{i}$ is not $\epsilon$-extremal and $d_{G_{i}}(v)\geq(\frac{1}{2}-\epsilon^{3})n$. Note that for each $i\in[n]$ and $j\in\{1,2\}$, there are at most $2\epsilon n$ vertices in $V_{j}$ that are not $i$-good.
Definition B.3 (absorbing edge, absorbing matching).
Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. Given two vertices $u\in V_{1}$ and $v\in V_{2}$, an edge $e=w_{1}w_{2}$ with $u,v\not\in\{w_{1},w_{2}\}$ is called a $c$-absorbing edge of $(u,v)$ if $c\in L(w_{1}v)$ and ${\rm col}(w_{1}w_{2})\in L(uw_{2})$ (see Figure 5).
Given $\delta,\delta^{\prime},\gamma,\gamma^{\prime}\geq 0$, a rainbow matching $\mathcal{M}$ is an absorbing matching with parameters $(\delta,\delta^{\prime},\gamma,\gamma^{\prime})$ if $|E(\mathcal{M})|\leq\gamma n$ and there exists a color set $\mathcal{C}$ of size at least $\delta n$ such that given any color $c\in\mathcal{C}$ and any $c$-good vertex $v\in V_{2}$, for all but at most $\delta^{\prime}n$ vertices $u\in V_{1}$, there are at least $\gamma^{\prime}n$ $c$-absorbing edges of $(u,v)$ within $\mathcal{M}$.
$u$$v$$w_{1}$$w_{2}$$c$
Figure 5: $w_{1}w_{2}$ is a $c$-absorbing edge of $(u,v)$.
The following two lemmas establish the existence of absorbing matchings when $\mathcal{G}$ is stable. We first consider the case that $\mathcal{G}$ is strongly stable.
Lemma B.4.
Let $0<\frac{1}{n}\ll d\ll\lambda,\mu\ll\gamma,\alpha\ll 1$. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$, satisfying the following conditions:
  • for every $i\in[n]$, $d_{G_{i}}(x)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ vertices $x\in V_{1}\cup V_{2}$,
  • for every $x\in V_{1}\cup V_{2}$, $d_{G_{i}}(x)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ colors $i\in[n]$.
If $\mathcal{G}$ is $(\gamma,\alpha)$-strongly stable, then there exists an absorbing matching with parameters $(1,0,\lambda,\lambda^{2})$.
Proof.
Without loss of generality, we assume that $G_{1},\ldots,G_{\gamma n}$ are $\alpha$-nice. For $i\in[\lambda n]$, define
$\mathcal{F}_{i}:=\left\{(u,v)\in V_{1}\times V_{2}:i\in L(uv)\right\}.$
The degree condition of $G_{i}$ implies that $|\mathcal{F}_{i}|=e(G_{i})\geq(n-dn)\left(\frac{1}{2}-\mu\right)n>\frac{1}{4}n^{2}.$ Fix a color $c\in[n]$ and two vertices $u\in V_{1},v\in V_{2}$ such that $v$ is $c$-good, let
$Z_{i}(c,uv):=\{(w,w^{\prime})\in\mathcal{F}_{i}:ww^{\prime}\,\,\text{is a}\,\,c\text{-absorbing edge of}\,\,(u,v)\}\ \text{and}\ Z(c,uv):=\bigcup_{i\in[\lambda n]}Z_{i}(c,uv).$
Since $G_{i}$ is $\alpha$-nice, $e(G_{i}[N_{G_{i}}(u),N_{G_{c}}(v)])\geq\frac{\alpha n^{2}}{2}$ for all but at most $dn$ indices $i\in[\gamma n]$. This means that for all but at most $dn$ indices $i\in[\gamma n]$, we have $|Z_{i}(c,uv)|\geq\frac{\alpha n^{2}}{2}$. Thus, we have $|Z(c,uv)\cap\mathcal{F}_{i}|\geq\frac{\alpha n^{2}}{2}$ for at least $(\lambda-2d)n$ indices $i\in[\lambda n]$. Applying Lemma 3.1 with
[…]
and parameters $t:=\lambda n$, $\epsilon:=\frac{\alpha}{2}$, to obtain a rainbow matching $\mathcal{M}$ inside $\mathbf{H}$ of size at least $(1-\frac{\alpha^{2}}{16})\lambda n$ such that $|E(Z(c,uv))\cap E(\mathcal{M})|\geq\frac{\alpha^{2}\lambda n}{16}$ for all $Z(c,uv)\in\mathbf{Z}$. Therefore, there exists a subset $I\subseteq[\lambda n]$ of size at least $(1-\frac{\alpha^{2}}{16})\lambda n$ such that for each $i\in I$, there is an edge $w_{i}w^{\prime}_{i}$ in $\mathcal{M}$ with $i\in L(w_{i}w^{\prime}_{i})$. For every $c\in[n]$ and $(u,v)\in V_{1}\times V_{2}$ satisfying $v$ is $c$-good, there are at least $\frac{\alpha^{2}\lambda n}{16}$ matching edges in $\mathcal{M}$ which are $c$-absorbing edges of $(u,v)$. By construction, $\mathcal{M}$ is an absorbing matching with parameters $(1,0,\lambda,\frac{\alpha^{2}\lambda}{16})$ and hence with parameters $(1,0,\lambda,\lambda^{2})$ since $\lambda\ll\alpha$. ∎
Next, we demonstrate that an absorbing matching can be found when $\mathcal{G}$ is weakly stable.
Lemma B.5.
Let $10<\frac{1}{n}\ll d\ll\lambda,\mu\ll\epsilon\ll\delta<$ 1. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$ such that the following conditions hold:
  • for every $i\in[n]$, $d_{G_{i}}(x)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ vertices $x\in V_{1}\cup V_{2}$,
  • for every $x\in V_{1}\cup V_{2}$, $d_{G_{i}}(x)\geq\left(\frac{1}{2}-\mu\right)n$ for all but at most $dn$ colors $i\in[n]$.
If $\mathcal{G}$ is $(\epsilon,\delta)$-weakly stable, then there exists an absorbing matching with parameters $(\frac{\delta}{2},\sqrt{\epsilon},\lambda,\lambda^{2})$.
Proof.
For each $i\in[n]$, we define
$\mathcal{F}_{i}:=\{(u,v)\in V_{1}\times V_{2}:i\in L(uv)\}.$
Then $|\mathcal{F}_{i}|>\frac{n^{2}}{4}$. We first prove the following claim, which only requires the degree condition on $\mathcal{G}$.
Claim B.6.
Assume $ij\in E(C_{\mathcal{G}}^{\epsilon,\delta})$. Let $u\in V_{1}$ be a $j$-good vertex and $v\in V_{2}$ be an $i$-good vertex. Define
$Z_{j}(i,uv):=\{(w,w^{\prime})\in\mathcal{F}_{j}:ww^{\prime}\,\,\text{is an}\,\,i\text{-absorbing edge of}\,\,(u,v)\}.$
Then $|Z_{j}(i,uv)|\geq 2^{-4}\delta n^{2}$.
Proof of Claim B.6.
For each $\ell\in[n]$, if $G_{\ell}$ is $\epsilon$-extremal, then by Lemma 2.7, there is a characteristic partition $(A_{1}^{\ell},B_{1}^{\ell},C_{1}^{\ell},A_{2}^{\ell},B_{2}^{\ell},C_{2}^{\ell})$ corresponding to $G_{\ell}$. Fix any such indices $i,j$ and vertices $u,v$. Since $v$ is $i$-good, we have
$d_{G_{i}}(v,A^{i}_{1})\geq\left(\frac{1}{2}-2\epsilon\right)n\,\,\,\text{or}\,\,\,d_{G_{i}}(v,B^{i}_{1})\geq\left(\frac{1}{2}-2\epsilon\right)n.$
Assume, without loss of generality, that the former case holds (the latter case is analogous). Since $G_{i}$ and $G_{j}$ are $\delta$-crossing and $\epsilon\ll\delta$, one has
$|A^{i}_{1}\cap A^{j}_{1}|\geq\frac{\delta n}{2},\ |A^{i}_{1}\cap B^{j}_{1}|\geq\frac{\delta n}{2},\ |A_{1}^{j}\cap B_{1}^{i}|\geq\frac{\delta n}{2},\,\,\text{and}\,\,|B_{1}^{i}\cap B_{1}^{j}|\geq\frac{\delta n}{2}.$
Combining $|A^{i}_{1}\cap A^{j}_{1}|\geq\frac{\delta n}{2}$ with $d_{G_{i}}(v,A^{i}_{1})\geq(\frac{1}{2}-2\epsilon)n$, we obtain
$d_{G_{i}}(v,A^{j}_{1})\geq|A^{i}_{1}\cap A^{j}_{1}|-\left(|A^{i}_{1}|-d_{G_{i}}(v,A^{i}_{1})\right)\geq\frac{\delta n}{2}-\epsilon n\geq\frac{\delta n}{3}.$
Similarly, we have $d_{G_{i}}(v,B^{j}_{1})\geq\frac{\delta n}{3}$. Recall that $u\in V_{1}$ and $u$ is $j$-good, which implies that $u\in X^{j}_{1}$ for some $X\in\{A,B\}$. Since $d_{G_{j}}(u,X_{2}^{j})\geq(\frac{1}{2}-2\epsilon)n$, there are at least $(\frac{1}{2}-2\epsilon)n$ choices for $w^{\prime}\in X_{2}^{j}$. Notice that in $G_{j}$, every vertex in $X^{j}_{2}$ is adjacent to all but at most $\epsilon n$ vertices in $X^{j}_{1}$. Fixing $w^{\prime}\in X_{2}^{j}$, and using $d_{G_{i}}(v,X^{j}_{1})\geq\frac{\delta n}{3}$, we conclude that there are at least $\frac{\delta n}{4}$ choices for $w\in X^{j}_{1}$. Thus,
$|Z_{j}(i,uv)|\geq\frac{\delta n}{4}\left(\frac{1}{2}-2\epsilon\right)n\geq 2^{-4}\delta n^{2},$
as desired. ∎
For any vertex $v\in V$, let
[…]
Since $e(C_{\mathcal{G}}^{\epsilon,\delta})\geq\delta n^{2}$, there exists a subgraph $H$ of $C_{\mathcal{G}}^{\epsilon,\delta}$ such that $|V(H)|\geq\delta n$ and $\delta(H)\geq\delta n$. For each $i\in V(H)$, define the set
$T_{i}:=\left\{v\in V:|N_{H}(i)\setminus\mathcal{C}_{v}|\geq\frac{\delta n}{2}\right\}.$
Observe that there are at most $4\epsilon n^{2}$ pairs $(u,i)\in V\times[n]$ such that $u$ is not $i$-good. This implies that $\frac{\delta n}{2}|T_{i}|\leq 4\epsilon n^{2}$. Hence, by $\epsilon\ll\delta$, we have $|T_{i}|\leq\frac{\sqrt{\epsilon}n}{2}$.
Let $\overline{T_{i}}:=V\setminus T_{i}$. Clearly, $|\overline{T_{i}}|\geq 2n-\frac{\sqrt{\epsilon}n}{2}$. For each $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|\mathcal{C}_{u}\cap N_{H}(i)|\geq\frac{\delta n}{2}$. Now we independently and randomly select vertices from $V(H)$ with probability $\kappa:=\frac{\lambda}{14}$ to obtain a set $\mathcal{U}$ of colors. A Chernoff bound implies that, with high probability,
  1. (i)
    $\frac{\kappa|V(H)|}{2}\leq t:=|\mathcal{U}|\leq 2\kappa|V(H)|$,
  2. (ii)
    for every $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})|\geq\frac{\delta\kappa n}{4}$.
Fix such a $\mathcal{U}$ and let $\overline{\mathcal{U}}:=V(H)\setminus\mathcal{U}$. By relabeling colors, one may assume $\mathcal{U}=:[t]$. Clearly,
$|\overline{\mathcal{U}}|=|V(H)|-|\mathcal{U}|\geq\frac{\delta n}{2},\,\,\text{and}\,\,\delta(H[\overline{\mathcal{U}}])\geq\delta(H)-|\mathcal{U}|\geq\frac{\delta n}{2}.$
Define the multi-graph collection
[…]
Given $i$ and $u$, the number of choices for $j$ is at least $\frac{\delta\kappa n}{4}$ by (ii). Since $ij\in E(H)$, $u$ is $j$-good and $v$ is $i$-good, Claim B.6 implies that there are at least $2^{-4}\delta n^{2}$ $i$-absorbing edges of $(u,v)$ whose ordered vertex set is in $\mathcal{F}_{j}$. Thus, $|Z(i,uv)\cap\mathcal{F}_{j}|\geq 2^{-4}\delta n^{2}$ for at least $\frac{\delta\kappa n}{4}\geq 2^{-3}t$ indices $j\in[t]$. Applying Lemma 3.1 with
$\mathbf{H}:=\{\mathcal{F}_{j}:j\in[t]\}$, $\mathbf{Z}$ and $\epsilon:=2^{-4}\delta$,
we can find a rainbow matching $\mathcal{M}$ in $\mathbf{H}$ of size at least $(1-2^{-10}\delta^{2})t$ such that $|E(Z(i,uv))\cap E(\mathcal{M})|\geq 2^{-10}\delta^{2}t$ for all $Z(i,uv)\in\mathbf{Z}$. That is, there is $I\subseteq[t]$ with $|I|\geq(1-2^{-10}\delta^{2})t$ such that for each $j\in I$ there is an edge $e_{j}\in E(\mathcal{M})$ with $j\in L(e_{j})$. For every $i\in\overline{\mathcal{U}}$, $u\in\overline{T_{i}}$ and $v\in V_{2}$ is $i$-good, there are at least $2^{-10}\delta^{2}t$ indices $j\in I$ such that $e_{j}$ is an $i$-absorbing edge of $(u,v)$. Thus, $\mathcal{M}$ is an absorbing matching with parameters $(\frac{\delta}{2},\sqrt{\epsilon},(1-2^{-10}\delta^{2})t,2^{-10}\delta^{2}t)$ and hence with $(\frac{\delta}{2},\sqrt{\epsilon},\lambda,\lambda^{2})$. ∎
Now, we are ready to prove Theorem 2.10.
Proof of Theorem 2.10.
Choose additional parameters $\beta,\lambda$ and $\mu^{\prime}$ such that
$1<\frac{1}{n}\ll d\ll\mu\ll\beta\ll\lambda\ll\mu^{\prime}\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1,$
where the previous lemmas in this section hold with these suitable parameters.
By Lemma B.4 and Lemma B.5, $\mathcal{G}$ contains an absorbing matching $\mathcal{M}$ with parameters $(\frac{\delta}{2},\sqrt{\epsilon},$ $\lambda,\lambda^{2})$. For any color $c\in[n]$ and vertices $u\in V_{1}$, $v\in V_{2}$, we define the triple $(c,u,v)$ as absorbable if there are at least $\lambda^{2}n$ disjoint $c$-absorbing edges of $(u,v)$ inside $\mathcal{M}$. By the definition of absorbing matching, one has $|\mathcal{M}|\leq\lambda n$. Furthermore, there exists a color set $\mathcal{C}\subseteq[n]\setminus{\rm col}(\mathcal{M})$ of size at least $\frac{\delta n}{2}$ such that, for any given color $c\in\mathcal{C}$ and $c$-good vertex $v\in V_{2}$, the triple $(c,u,v)$ is absorbable for all but at most $\sqrt{\epsilon}n$ vertices $u\in V_{1}$.
Claim B.7.
There exists an integer $r$ with $(2-2^{-10})\beta n\leq r\leq 2\beta n$. For each $i\in[r]$, there is a tuple $q_{i}:=(c_{i}^{1},c_{i}^{2},w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})$ where $c_{i}^{1},c_{i}^{2}\in\mathcal{C}$ are distinct colors, $w_{i}^{1},w_{i}^{2}\in V_{1}$ and $w_{i}^{1^{\prime}},w_{i}^{2^{\prime}}\in V_{2}$ are distinct vertices, satisfying the following properties:
  1. (i)
    $(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})$ and $(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})$ are absorbable for all $i\in[r]$,
  2. (ii)
    for every pair $(u,v)\in V_{1}\times V_{2}$ and every color $c\in[n]$, there are at least $2^{-20}\beta n$ indices $i\in[r]$ such that $c_{i}^{1}\in L(uw_{i}^{1^{\prime}})$, $c_{i}^{2}\in L(vw_{i}^{2})$ and $c\in L(w_{i}^{1}w_{i}^{2^{\prime}})$.
Proof of Claim B.7.
For every two distinct colors $c_{i}^{1},c_{i}^{2}\in\mathcal{C}$, and vertex pair $(u,v)\in V_{1}\times V_{2}$ such that $d_{G_{c_{i}^{1}}}(u)\geq\left(\frac{1}{2}-\mu\right)n$ and $d_{G_{c_{i}^{2}}}(v)\geq\left(\frac{1}{2}-\mu\right)n$, and for every color $c\in[n]$, let $S(c_{i}^{1},c_{i}^{2},u,v,c)$ be the collection of tuples
$(w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})\in V_{1}\times V_{2}\times V_{1}\times V_{2}$
that satisfy the following properties: $(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})$ and $(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})$ are absorbable, $c_{i}^{1}\in L(uw_{i}^{1^{\prime}})$, $c_{i}^{2}\in L(vw_{i}^{2})$, and $c\in L(w_{i}^{1}w_{i}^{2^{\prime}})$. We will prove that $|S(c_{i}^{1},c_{i}^{2},u,v,c)|\geq 2^{-7}n^{4}$. For this, we count the number of possible choices for $w_{i}^{1^{\prime}},\,w_{i}^{2^{\prime}},\,w_{i}^{1}$ and $w_{i}^{2}$ in turn.
$\bullet$ Recall that all but at most $2\epsilon n$ vertices in $V_{2}$ are $c_{i}^{1}$-good and $d_{G_{c_{i}^{1}}}(u)\geq\left(\frac{1}{2}-\mu\right)n$. Therefore, there are at least $\left(\frac{1}{2}-\mu\right)n-2\epsilon n-1\geq\frac{n}{3}$ choices for $w_{i}^{1^{\prime}}$ such that $w_{i}^{1^{\prime}}$ is $c_{i}^{1}$-good and $c_{i}^{1}\in L(uw_{i}^{1^{\prime}})$.
$\bullet$ For vertices in $V_{2}$, there are at most $2\epsilon n+dn$ vertices $w$ that are either not $c_{i}^{2}$-good or $d_{G_{c}}(w)<\left(\frac{1}{2}-\mu\right)n$. Thus, we have at least $n-2\epsilon n-dn-2\geq\frac{n}{2}$ choices for $w_{i}^{2^{\prime}}$ such that $w_{i}^{2^{\prime}}$ is $c_{i}^{2}$-good and $d_{G_{c}}(w)\geq\left(\frac{1}{2}-\mu\right)n$.
$\bullet$ Fix $w_{i}^{1^{\prime}}$ and $w_{i}^{2^{\prime}}$. Since $w_{i}^{1^{\prime}}$ is $c_{i}^{1}$-good and $c_{i}^{1}\in\mathcal{C}$, we obtain that $(c_{i}^{1},w,w_{i}^{1^{\prime}})$ is absorbable for all but at most $\sqrt{\epsilon}n$ vertices $w\in V_{1}$. Recall that $d_{G_{c}}(w_{i}^{2^{\prime}})\geq\left(\frac{1}{2}-\mu\right)n$. Therefore, there are at least $\left(\frac{1}{2}-\mu\right)n-\sqrt{\epsilon}n-1\geq\frac{n}{3}$ choices for $w_{i}^{1}$ such that $(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})$ is absorbable and $c\in L(w_{i}^{1}w_{i}^{2^{\prime}})$.
$\bullet$ Similarly, since $d_{G_{c_{i}^{2}}}(v)\geq\left(\frac{1}{2}-\mu\right)n$, $c_{i}^{2}\in\mathcal{C}$ and $w_{i}^{2^{\prime}}$ is $c_{i}^{2}$-good, there are at least $\left(\frac{1}{2}-\mu\right)n-\sqrt{\epsilon}n-2\geq\frac{n}{3}$ choices for $w_{i}^{2}$ such that $c_{i}^{2}\in L(vw_{i}^{2})$ and $(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})$ is absorbable.
Therefore, $|S(c_{i}^{1},c_{i}^{2},u,v,c)|\geq 2^{-7}n^{4}$, as desired.
Let $\{c_{i}^{1},c_{i}^{2}:i\in[2\beta n]\}\subseteq\mathcal{C}$ be a set of distinct colors. Define
$\mathcal{F}_{i}:=\{(w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})\in V_{1}\times V_{2}\times V_{1}\times V_{2}:(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})\,\,\text{and}\,\,(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})\,\,\,\text{are absorbable}\}.$
Since for each color $c\in\mathcal{C}$, there are at most $2\epsilon n$ vertices in each part that are not $c$-good. Hence, there exist at least $(n-2\epsilon n)(n-\sqrt{\epsilon}n)\geq\frac{n^{2}}{2}$ vertex pairs $(x,y)\in V_{1}\times V_{2}$ such that $(c,x,y)$ is absorbable. Thus, $|\mathcal{F}_{i}|\geq 2^{-4}n^{4}$. Let $\mathbf{H}:=\{\mathcal{F}_{i}:i\in[2\beta n]\}$ be the collection of directed 4-graphs.
For a vertex pair $(u,v)\in V_{1}\times V_{2}$, let
$I_{u,v}:=\left\{i\in[2\beta n]:d_{G_{c_{i}^{1}}}(u)\geq\left(\frac{1}{2}-\mu\right)n\,\,\text{and}\,\,d_{G_{c_{i}^{2}}}(v)\geq\left(\frac{1}{2}-\mu\right)n\right\}.$
Define the collection of directed multi-4-graphs as follows:
$\mathbf{Z}:=\left\{S(u,v,c):=\bigcup_{i\in I_{u,v}}S(c_{i}^{1},c_{i}^{2},u,v,c):(u,v)\in V_{1}\times V_{2},\,c\in[n]\right\}.$
Since for each vertex $w\in V$, there are at most $dn$ indices $i\in[n]$ such that $d_{G_{i}}(w)<\left(\frac{1}{2}-\mu\right)n$. Hence for every $S:=S(u,v,c)\in\mathbf{Z}$, we have $|E(S)\cap E(\mathcal{F}_{i})|\geq 2^{-7}n^{4}$ for all but at most $2dn$ indices $i\in[2\beta n]$. By applying Lemma 3.1 with $t:=2\beta n$ and $\epsilon:=2^{-7}$, we deduce that there is a rainbow matching $M$ in $\mathbf{H}$ of size at least $(2-2^{-10})\beta n$ (and at most $2\beta n$) and $|E(S)\cap E(M)|\geq 2^{-20}\beta n$ for all $S\in\mathbf{Z}$. ∎
Let
$V_{{\rm abs}}:=\bigcup_{i\in[r]}\{w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}}\},\ \ \mathcal{C}_{{\rm abs}}:=\bigcup_{i\in[r]}\{c_{i}^{1},c_{i}^{2}\},\ \ \mathcal{C}_{{\rm rem}}:=[n]\setminus({\rm col}(\mathcal{M})\cup\mathcal{C}_{{\rm abs}}),$
$U:=V\setminus(V(\mathcal{M})\cup V_{{\rm abs}}),\ \ \mathcal{J}:=\{J_{i}:i\in\mathcal{C}_{{\rm rem}}\}\ \text{where}\ J_{i}=G_{i}[U].$
Note that $|[n]\setminus\mathcal{C}_{{\rm rem}}|=\frac{|V\setminus U|}{2}\leq\lambda n+4\beta n$. For each vertex $v\in U$, for all but at most $dn$ colors $c\in\mathcal{C}_{\rm rem}$ we have
$d_{J_{c}}(v)\geq\left(\frac{1}{2}-\mu\right)n-\lambda n-4\beta n\geq\frac{|U|}{4}-\mu^{\prime}\frac{|U|}{12}.$
Since $\beta,\lambda,\mu^{\prime}\ll\gamma,\alpha,\epsilon,\delta$, it is easy to see that $\mathcal{J}$ is $(\frac{\gamma}{2},\frac{\alpha}{2},2\epsilon,\frac{\delta}{2})$-stable. By Lemmas B.1 and B.2, we deduce that $\mathcal{J}$ is $\mu^{\prime}$-nice and there exists a rainbow matching $\mathcal{M}_{1}$ in $\mathcal{J}$ with size at least $|\mathcal{C}_{{\rm rem}}|-dn-2$. Assume that
$(U\setminus V(\mathcal{M}_{1}))\cap V_{1}=\{u_{1},\ldots,u_{t}\},\,\,(U\setminus V(\mathcal{M}_{1}))\cap V_{2}=\{v_{1},\ldots,v_{t}\},\,\,\mathcal{C}_{\rm rem}\setminus{\rm col}(\mathcal{M}_{1})=\{c_{1},\ldots,c_{t}\}.$
Do the following for each $(u_{i},v_{i},c_{i})$ with $i=1,2,\ldots,t$ in turn. Choose an unused 6-tuple $q_{j_{i}}=(c_{j_{i}}^{1},c_{j_{i}}^{2},w_{j_{i}}^{1},w_{j_{i}}^{1^{\prime}},w_{j_{i}}^{2},w_{j_{i}}^{2^{\prime}})$ for some $j_{i}\in[r]$, such that $c_{j_{i}}^{1}\in L(u_{i}w_{j_{i}}^{1^{\prime}})$, $c_{j_{i}}^{2}\in L(v_{i}w_{j_{i}}^{2})$, and $c_{i}\in L(w_{j_{i}}^{1}w_{j_{i}}^{2^{\prime}})$. This is possible since Claim B.7 implies that there are at least $2^{-20}\beta n$ choices for $q_{j_{i}}$, of which at most $dn+2$ have been used. Thus, all vertices in $U\setminus V(\mathcal{M}_{1})$ and all colors in $\mathcal{C}_{\rm rem}\setminus{\rm col}(\mathcal{M}_{1})$ are contained in a rainbow matching of $\mathcal{G}$.
At the end of this process, there remains a set $I\subseteq[r]$ such that the $q_{i}$ with $i\in[I]$ are precisely the 6-tuples that were not chosen. Recall that $(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})$ and $(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})$ are absorbable for each $i\in I$. It follows that there are at least $\lambda^{2}n$ disjoint $c_{i}^{j}$-absorbing edges of $(w_{i}^{j},w_{i}^{j^{\prime}})$ inside $\mathcal{M}$ for $i\in I$ and $j\in[2]$. Together with $\beta\ll\lambda$, we obtain a transversal perfect matching inside $\mathcal{G}$. ∎

### v2

@@
 1 Introduction
-A transversal $X$ over a collection =$\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects represents an object that intersects every $F_{i}$.
+A transversal $X$ over a collection $\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects represents an object that intersects every $F_{i}$.
 Here, the object can be any mathematical object, such as sets, spaces, set systems, matroids, and so forth.
@@
 Definition 1.1.
-For a given collection $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/hypergraphs/digraphs with the same vertex set $V$, an $m$-edge graph/hypergraph/digraph $H$ with vertices in $V$ is transversal in $\mathcal{G}$ if there exists a bijection $\phi:E(H)\to[m]$ such that $e\in E(G_{\phi(e)})$ for all $e\in E(H).$
-By interpreting each $G_{i}$ as the set of edges colored with color $i$, the function $\phi$ is often referred to as a coloring.
-We say that the coloring $\phi$ is rainbow if it is injective, and $H$ is a rainbow subgraph if there exists a rainbow coloring on it.
+For a given collection $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/hypergraphs/digraphs with the same vertex set $V$, an $m$-edge graph/hypergraph/digraph $H$ with vertices in $V$ is transversal in $\mathcal{G}$ if there exists a bijection ${\varphi}:E(H)\to[m]$ such that $e\in E(G_{{\varphi}(e)})$ for all $e\in E(H).$
+By interpreting each $G_{i}$ as the set of edges colored with color $i$, the function $\varphi$} is often referred to as a coloring.
+We say that the coloring $\varphi$} is rainbow if it is injective, and $H$ is a rainbow subgraph if there exists a rainbow coloring on it.
 Many classical results in extremal graph theory have been extended to such transversal settings, exhibiting interesting phenomena.
@@
 Problem A ([21]).
-Let $H$ be a graph/hypergraph/digraph with $m$ edges, $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$ be a collection of graphs/hypergraphs/digraphs on a common vertes set of size $n$.
+Let $H$ be a graph/hypergraph/digraph with $m$ edges, $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$ be a collection of graphs/hypergraphs/digraphs on a common vertex set of size $n$.
 Which property $\mathcal{P}$ imposed on $\mathcal{G}$ will guarantee the existence of a transversal copy of $H$?
-Since every object in the family $\mathcal{G}$ can be the same, a necessary condition for a positive answer to the above question is that every $n$-vertex graph/hypergraph/directed graph satisfying the property $\mathcal{P}$ contains a copy of $H$.
+Since every object in the family $\mathcal{G}$ can be the same, a necessary condition for a positive answer to the above question is that every $n$-vertex graph/hypergraph/digraph satisfying the property $\mathcal{P}$ contains a copy of $H$.
 However, this alone is not always sufficient.
-For example, Aharoni, DeVos, de la Maza, Montejano, and Šámal [1] proved that if $\mathcal{G}=\{G_{1},G_{2},G_{3}\}$ is a collection of graphs on a common vertex set of size $n$ and $|E(G_{i})|>(\frac{26-2\sqrt{7}}{81})n^{2}$ for $i\in[3]$, then $\mathcal{G}$ contains a rainbow triangle.
+For example, Aharoni, DeVos, de la Maza, Montejano, and Šámal [1] proved that if $\mathcal{G}=\{G_{1},G_{2},G_{3}\}$ is a collection of graphs on a common vertex set of size $n$ and $|E(G_{i})|>(\frac{26-2\sqrt{7}}{81})n^{2}$ for all $i\in[3]$, then $\mathcal{G}$ contains a rainbow triangle.
 Moreover, the constant $\frac{26-2\sqrt{7}}{81}$ is optimal, while Mantel’s theorem states that any $n$-vertex graph with more than $\lfloor\frac{n^{2}}{4}\rfloor$ edges must contain a triangle.
@@
 There are numerous established results on transversal problems in graph collections. For more results on transversals, we refer the reader to the survey [26].
-A natural next step is to study rainbow analogues in digraph collections.
+A natural next step is to study the rainbow analogues in digraph collections.
 Cheng, Han, Wang and Wang [8] studied the transversal version of Hajnal-Szemerédi theorem in digraph collections, which asymptotically generalized the corresponding results by Czygrinow-DeBiasio-Kierstead-Molla [15] and Treglown [30].
@@
 Babiński, Grzesik and Prorok [3] studied the transversal Mantel’s theorem in digraph and oriented graph collections.
-Gerbner, Grzesik, Palmer and Prorok [18] determined the minimum number of edges in each color which guarantees the existence any rainbow directed star.
+Gerbner, Grzesik, Palmer and Prorok [18] determined the minimum number of edges in each color which guarantees the existence of any rainbow directed star.
 1.1 Main result
@@
 Denote the set of vertices of $D$ by $V(D)$.
-Let $\delta^{+}(D)$ and $\delta^{-}(D)$ be the minimum out-degree and minimum in-degree of $D$, respectively.
-Let $\delta^{0}(D)=\min\{\delta^{+}(D),{\delta^{-}}(D)\}$ be the minimum semi-degree of $D.$ Let $\mathcal{D}=\{D_{1},D_{2},\dots,D_{m}\}$ be a collection of not necessarily distinct digraphs with common vertex set $V$.
-Define $\delta^{0}(\mathcal{D}):=\min\{\delta^{0}(D_{i}):i\in[m]\}$.
-In 1960, Ghouila-Houri [19] generalized the Dirac’s theorem to digraphs.
+Define $\delta^{+}(D)$ and $\delta^{-}(D)$ as the minimum out-degree and in-degree of $D$, respectively, and let the minimum semi-degree be $\delta^{0}(D)=\min{\delta^{+}(D),\delta^{-}(D)}$.
+For a collection $\mathcal{D}=\{D_{1},\dots,D_{m}\}$ of digraphs (not necessarily distinct) on a common vertex set $V$, define $\delta^{0}(\mathcal{D})=\min\{\delta^{0}(D_{i}):i\in[m]\}$.
+In 1960, Ghouila-Houri [19] generalized Dirac’s theorem to digraphs.
 Theorem 1.3 ([19]).
-Every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains a directed Hamilton cycle.
-Chakraborti, Kim, Lee, Seo [6] suggested to consider a transversal version of Ghouila-Houri’s theorem would be interesting. Our main result solves this problem.
+For each integer $n\geq 3$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains a directed Hamilton cycle.
+Chakraborti, Kim, Lee, and Seo [6] suggested that it would be interesting to consider a transversal version of Ghouila-Houri’s theorem. Our main result solves this.
 Theorem 1.4.
@@
 Thus, for sufficiently large $n$, Theorem 1.2 is a direct consequence of Theorem 1.4.
-Furthermore, by letting $D_{1}=D_{2}=\cdots=D_{n}$ in Theorem 1.4, we get Theorem 1.3 immediately when $n$ is large.
+Furthermore, by setting $D_{1}=D_{2}=\cdots=D_{n}$ in Theorem 1.4, we immediately obtain Theorem 1.3 for all sufficiently large $n$.
 1.2 Notation and Organization
@@
 We use $H\subseteq D$ to denote that $H$ is a sub-digraph of $D$.
-For $S\subseteq V(D)$, we write $D[S]$ for the sub-digraph of $D$ induced by $S$ and $D-S$ for $D[V(D)\setminus S]$.
+For $S\subseteq V(D)$, we write $D[S]$ for the sub-digraph of $D$ induced by $S$ and DSD-S for $D[V(D)\setminus S]$.
 For two not necessarily disjoint vertex subsets $A,B\subseteq V(D)$, let $D[A,B]$ be the sub-digraph of $D$ with vertex set $A\cup B$ and edge set $\{xy\in E(D):x\in A,y\in B\}$ and let $e_{D}(A,B):=|E(D[A,B])|$.
@@
 When $A=B$, we often abbreviate $D[A,B]$ and $e_{D}(A,B)$ as $D[A]$ and $e_{D}(A)$ respectively.
-For a digraph $D$ and $v\in V(D)$, we denote the set of out-neighbours and in-neighbours of $v$ by $N_{D}^{+}(v)$ and $N_{D}^{-}(v)$, respectively, and we write $d_{D}^{+}(v):=|N_{D}^{+}(v)|$ and $d_{D}^{-}(v):=|N_{D}^{-}(v)|$ for the out- and indegree of $v$ in $D$, respectively.
+For a digraph $D$ and $v\in V(D)$, we denote the sets of out-neighbors and in-neighbors of $v$ by $N_{D}^{+}(v)$ and $N_{D}^{-}(v)$, respectively, with the corresponding out- and in-degrees given by $d_{D}^{+}(v):=|N_{D}^{+}(v)|$ and $d_{D}^{-}(v):=|N_{D}^{-}(v)|$.
 For $S\subseteq V(D)$, we write $d_{D}^{+}(v,S):=|N_{D}^{+}(v)\cap S|$ and $d_{D}^{-}(v,S):=|N_{D}^{-}(v)\cap S|$.
@@
 Let $\mathcal{D}=\{D_{i}:i\in[m]\}$ be a collection of digraphs on a common vertex set $V$ and $H$ be a rainbow digraph inside $\mathcal{D}$.
-For a vertex $v\in V$, let $N_{i}^{+}(v):=N_{D_{i}}^{+}(v)$, $N_{i}^{-}(v):=N_{D_{i}}^{-}(v)$, $d_{i}^{+}(v):=N_{i}^{+}(v)$ and $d_{i}^{-}(v):=N_{i}^{-}(v)$.
+For a vertex $v\in V$ and each $i\in[m]$, let $N_{i}^{+}(v):=N_{D_{i}}^{+}(v)$, $N_{i}^{-}(v):=N_{D_{i}}^{-}(v)$, $d_{i}^{+}(v):=|N_{i}^{+}(v)|$ and $d_{i}^{-}(v):=|N_{i}^{-}(v)|$.
 For any two vertex subsets $X,\,Y\subseteq V$, define $\mathcal{D}[X]:=\{D_{i}[X]:i\in[m]\}$, $\mathcal{D}[X,Y]:=\{D_{i}[X,Y]:i\in[m]\}$ and $\mathcal{D}^{\pm}[X,Y]:=\{D_{i}^{\pm}[X,Y]:i\in[m]\}$.
@@
 In Section 2, we introduce two key techniques that we will need later, i.e., the regularity lemma in digraph collections and a stability result for transversal perfect matchings in bipartite graph collections.
-For readability, we postpone its proofs to the appendix.
+For readability, we postpone their proofs to the appendix.
 The proof of Theorem 1.4 will be divided into “stable case” and “extremal case”.
 Section 3 introduces notions of extremality and stability for digraphs and digraph collections; we prove several results about absorption that will be used in the proof of the stable case.
-In Section 5, we show that there is always a transversal directed Hamilton cycle in stable case.
-Section 6 deals with the extremal case.
+In Section 4, we show that there is always a transversal directed Hamilton cycle in the stable case.
+Sections 5-7 deal with the extremal case.
 We finish our paper with some concluding remarks in the final section.
@@
 In [9], the first author and Staden further developed the absorption method for graph collections, combined with the application of the transversal blow-up lemma, they finally proved a stability version of Joos and Kim’s result [21].
-The absorption technique used in [10] originates from [12] and [8], based on the common absorption technique introduced first in [25], which was showed to be extremely powerful in the solutions of various transversal graph embedding problems.
+The absorption technique in [9] originates from [12] and [8], which adapted the absorption technique first introduced in [25].
+This approach has become a powerful tool for solving various transversal graph embedding problems.
 In this paper, we adapt the method developed in [9] and generalize their stability from graph collections to digraph collections.
 By analyzing the extremal cases, we proved a tight bound for the transversal Ghouila-Houri’s theorem.
-We believe our method could be used to demonstrate stabilities or tight bounds for other transversal graph embedding problem (see the survey [26]).
+We believe that our method could be used to demonstrate stabilities or tight bounds for other transversal graph embedding problems (see the survey [26]).
 1.4 Sketch of the proof of Theorem 1.4
-Our proof utilizes the regularity blow-up method in digraph collections, and is divided into two parts: stable case and the extremal case.
+Our proof utilizes the regularity blow-up method in digraph collections, and is divided into two parts: stable case and extremal case.
 Stable case.
@@
 Let $0<\frac{1}{n}\ll\epsilon\ll 1$ and $D$ be an $n$-vertex digraph with minimum semi-degree at least $(\frac{1}{2}-\epsilon)n$.
-If $D$ contains no directed Hamilton cycle, then $D$ contains a sub-digraph that is close to one of the digraphs in Figure 1.
-Based on Lemma 3.3, there exists a characteristic partition of $D$ in each case, and we say that $D$ is $(\epsilon,{\rm EC}k)$-extremal for some $k\in[3]$.
-Moreover, the digraph $D$ can be $(\epsilon,{\rm EC}k)$-extremal for exactly one $k$.
-[…]
+Suppose $D$ has no directed Hamilton cycle.
+Then $D$ contains a sub-digraph that is close to one of the digraphs in Figure 1.
+By Lemma 3.2, such a sub-digraph induces a characteristic partition of $D$; we say that $D$ is $(\epsilon,\mathrm{EC}k)$-extremal for some $k\in[3]$.
+Furthermore, this $k$ is unique.
+Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs with the same vertex set $V$ of size $n$.
+Suppose that $\delta^{0}(\mathcal{D})\geq(\frac{1}{2}-\epsilon)n$ and $\mathcal{D}$ contains no transversal directed Hamilton cycles.
+We suspect that either almost all digraphs in $\mathcal{D}$ are $(\epsilon,{\rm EC}k)$-extremal for some $k\in[2]$ (allowing $k=1$ for some digraphs and $k=2$ for others), or almost all digraphs in $\mathcal{D}$ are $(\epsilon,{\rm EC}3)$-extremal; moreover those digraphs have similar vertex partitions.
+Based on this observation, we define two kinds of stability for a digraph collection with minimum degree at least $(\frac{1}{2}-\epsilon)n$.
+We say that
   • $\mathcal{D}$ is strongly stable if $\mathcal{D}$ contains many digraphs each of which is not $(\epsilon,{\rm EC}k)$-extremal for any $k\in[3]$,
@@
 Step 1. Build an absorbing cycle when $\mathcal{D}$ is stable. (Section 3)
-We build an “directed absorbing rainbow cycle” $C$ for $\mathcal{D}$ with the property that $C$ is very small and there is a set $A$ consisting of some color-vertex-vertex triples $(c,u,v)$ such that whenever $A_{0}\subseteq A$ is sufficiently small compared to $|C|$, $C$ can absorb all of its elements.
+We build a “directed absorbing rainbow cycle” $C$ for $\mathcal{D}$ with the property that $C$ is very small and there is a set $A$ consisting of some color-vertex-vertex triples $(c,u,v)$ such that whenever $A_{0}\subseteq A$ is sufficiently small compared to $|C|$, $C$ can absorb all of its elements.
 Additionally, we construct an additional auxiliary set $\mathcal{Q}$ consisting of color-color-vertex-vertex tuples $(c,c^{\prime},v,v^{\prime})$ to absorb those colors and vertices which $C$ cannot.
 Delete all colors and vertices in $C$ and $\mathcal{Q}$ from $\mathcal{D}$.
-Step 2. Use the regularity-blow-up method for transversals in digraph collections to cover almost all vertices and colors with long directed paths.
-We first prove the regularity lemma for digraph collections (Section 2.1), and then apply it to $\mathcal{D}$, yielding a reduced digraph collection $\mathcal{R}$ that inherits the minimum degree condition and stability of $\mathcal{D}$.
+Step 2. Use the regularity-blow-up method for transversals in digraph collections to cover almost all vertices and colors with long rainbow directed paths.
+We first prove a regularity lemma for digraph collections (Section 2.1), and then apply it to $\mathcal{D}$, yielding a reduced digraph collection $\mathcal{R}$ that inherits the minimum degree condition and the stability property of $\mathcal{D}$.
 By considering the characteristic bipartite graph of each digraph, we get a bipartite graph collection $\mathcal{B_{R}}$ for the reduced digraph collection $\mathcal{R}$.
@@
 We establish a stability result for transversal perfect matching in bipartite graph collections (Section 2.2), ensuring that the disjoint rainbow directed cycle partition of $\mathcal{R}$ exists.
-Using the blow-up method, we obtain a set of almost spanning disjoint rainbow directed paths within each directed cycle of $\mathcal{R}$, which covers almost all vertices outside $C$ and $\mathcal{Q}$ (Section 4).
-Step 3. Connect the obtained directed paths and cover remaining vertices via the absorbing cycle.
-In this step, we apply the absorbing property of $C$ as well as the auxiliary set $\mathcal{Q}$ to connect all the directed paths and cover the remaining vertices, ultimately form a transversal directed Hamilton cycle.
-[…]
+Using the blow-up method, we obtain a set of almost spanning disjoint rainbow directed paths within each directed cycle of $\mathcal{R}$, which covers almost all vertices outside the absorbing cycle $C$ and the auxiliary set $\mathcal{Q}$ (Section 4).
+Step 3. Connect the obtained directed paths and cover the remaining vertices via the absorbing cycle.
+In this step, we apply the absorbing property of $C$ as well as the auxiliary set $\mathcal{Q}$ to connect all the directed rainbow paths and to cover the remaining vertices, which ultimately forms a transversal directed Hamilton cycle.
+Extremal case (Sections 5-7).
+Assume that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$ and $\mathcal{D}$ contains no transversal Hamilton cycles.
+Let $0<\frac{1}{n}\ll\epsilon\ll\delta\ll\eta\ll 1$.
+Denote $\mathcal{C}_{k}:=\{i\in[n]:D_{i}\ {\text{is ($\epsilon$,EC$k$)-extremal}}\}$ for $k\in[3]$ and $\mathcal{C}_{\rm bad}:=[n]\setminus(\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\mathcal{C}_{3})$.
+Notice that $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$ for $1\leq i<j\leq 3$.
+By “stable case”, we know that almost all digraphs in $\mathcal{D}$ are extremal and $|\mathcal{C}_{1}|+|\mathcal{C}_{2}|\leq 2\sqrt{\delta}n$ if $|\mathcal{C}_{3}|\geq 2\sqrt{\delta}n$.
+In view of Lemma 3.2, we can fix a characteristic partition $({A_{i}},{B_{i}},{L_{i}})$ for each graph $D_{i}$ with $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$, and a characteristic partition $({C_{i}^{1}},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each graph $D_{i}$ with $i\in\mathcal{C}_{3}$.
+By swapping labels, we know that $|{A_{1}}\triangle{A_{i}}|,|{B_{1}}\triangle{B_{i}}|<\delta n$ for every $i\in(\mathcal{C}_{1}\cup\mathcal{C}_{2})\setminus\{1\}$ and $|C_{1}^{k}\triangle C_{i}^{k}|<2\delta n$ for every $i\in\mathcal{C}_{3}\setminus\{1\}$.
+We proceed with our proof by considering the sizes of $|\mathcal{C}_{1}|$, $|\mathcal{C}_{2}|$ and $|\mathcal{C}_{3}|$.
+We will take the case $|\mathcal{C}_{1}|<\eta n$, $|\mathcal{C}_{2}|\geq\eta n$ and $|\mathcal{C}_{3}|\leq 2\sqrt{\delta}n$ as an example to illustrate the general idea of our proof.
 Firstly, expand $A_{1}\cup B_{1}$ into an equitable partition $A\cup B$ of $V$.
-For a vertex $v\in A$, there are three bad cases (similar for vertices in $B$): $v$ lies in almost all $B_{i}$ for $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$; $v$ lies in almost all $L_{i}$ for $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$; $v\in A_{i}$ for many $i\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and $v\in B_{j}$ for many $j\in\mathcal{C}_{1}\cup\mathcal{C}_{2}$.
-Move vertices in the first case from $A$ to $B$, and do the same operation for such vertices in $B$.
-Denote the set of vertices in the other two bad cases by $V_{\rm bad}$.
-After this, the number of vertices in $A$ and $B$ may be unbalanced, but their difference is still small.
-Without loss of generality, assume that $|A|\leq|B|$.
-In order to apply the transversal blow-up lemma to find long rainbow directed paths inside $\mathcal{D}$, the following four steps are required.
+To apply the transversal blow-up lemma (Lemma 6.2) for finding long rainbow directed paths in $\mathcal{D}$, we need to ensure most vertices in the partition $A\cup B$ satisfy uniform neighborhood properties (for example, a vertex $v\in A$ should satisfy $v\in A_{i}$ for the majority of colors $i\in\mathcal{C}_{2}$).
+We identify three vertex types in $A$ (and symmetrically in $B$) that may disrupt the desired uniformity:
+  1. Type 1.
+    $v\in A$ belongs to $L_{i}$ for many $i\in\mathcal{C}_{2}$;
+  2. Type 2.
+    $v\in A$ belongs to $A_{i}$ for many $i\in\mathcal{C}_{2}$, and $v\in B_{j}$ for many other $j\in\mathcal{C}_{2}$;
+  3. Type 3.
+    $v\in A$ belongs to nearly all $B_{i}$ for $i\in\mathcal{C}_{2}$.
+We handle these cases as follows: For Type 3 vertices in $A$, we move them to $B$ (and perform the analogous operation for Type 3 vertices in $B$).
+We define $V_{\rm bad}$ as the union of vertices of Type 1 and Type 2.
+After this reallocation, every vertex in $A\setminus V_{\rm bad}$ (resp.
+$B\setminus V_{\rm bad}$) lies in $A_{i}$ (resp.
+$B_{i}$) for the majority of colors $i\in\mathcal{C}_{2}$.
+However, the partition sizes $|A|$ and $|B|$ may become unbalanced, with their difference bounded by $9\delta n$.
+To balance them, we require the following four steps.
 Step 1. Balance the number of vertices in $A$ and $B$.
 If there exists a set of disjoint rainbow paths in $\mathcal{D}[B]$ such that after deleting their vertices, the number of vertices in $A$ and $B$ is balanced, then we are done.
-Otherwise, we will show that $\mathcal{D}$ must contains a transversal directed Hamilton cycle (see Lemma 5.14).
+Otherwise, we will show that $\mathcal{D}$ must contain a transversal directed Hamilton cycle (see Lemma 6.5).
 Step 2. Cover vertices in $V_{\rm bad}$.
-Using colors in $\mathcal{C}_{2}$, we are to find a series of disjoint rainbow directed paths $P_{3}$ such that their centers are all vertices in $V_{\rm bad}$ and endpoints are unused vertices in $(A\cup B)\setminus V_{\rm bad}$.
+Using colors in $\mathcal{C}_{2}$, we are to find a sequence of disjoint rainbow directed paths $P_{3}$ such that their centers are all vertices in $V_{\rm bad}$ and endpoints are unused vertices in $(A\cup B)\setminus V_{\rm bad}$.
 Step 3. Deal with colors in $\mathcal{C}_{\rm bad}\cup\mathcal{C}_{3}$.
@@
 Lemma 2.2 (Regularity lemma for digraph collections).
-For every integer $L_{0}\geq 1$ and every $\epsilon,\delta>0$, there is an $n_{0}=n_{0}(\epsilon,\delta,L_{0})$ such that for every $d\in[0,1)$ and every digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on vertex set $V$ of size $n\geq n_{0}$ with $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$, there exists a partition of $V$ into $V_{0},V_{1},\ldots,V_{L}$, of $\mathcal{C}$ into $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ and a spanning sub-digraph $D_{c}^{\prime}$ of $D_{c}$ for each $c\in\mathcal{C}$ such that the following properties hold:
+For every integer $L_{0}\geq 1$ and every $\epsilon,\delta>0$, there is an integer $n_{0}:=n_{0}(\epsilon,\delta,L_{0})$ such that for every $d\in[0,1)$ and every digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on vertex set $V$ of size $n\geq n_{0}$ with $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$, there exists a partition of $V$ into $V_{0},V_{1},\ldots,V_{L}$, of $\mathcal{C}$ into $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ and a spanning sub-digraph $D_{c}^{\prime}$ of $D_{c}$ for each $c\in\mathcal{C}$ satisfying the following properties:
   1. (i)
@@
 Definition 2.3 (Reduced digraph collection).
-Given a digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on $V$ and parameters $\epsilon>0$, $d\in[0,1)$ and $L_{0}\geq 1$, the reduced digraph collection $\mathcal{R}=\mathcal{R}(\epsilon,d,L_{0})$ of $\mathcal{D}$ is defined as follows.
-Apply Lemma 2.2 to $\mathcal{D}$ with parameters $\epsilon,\delta,d,L_{0}$ to obtain $\mathcal{D}^{\prime}$ and a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$, and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}$, where $V_{0},\mathcal{C}_{0}$ are the exceptional sets and $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ are the color clusters.
-Then :=$\mathcal{R}:=\{R_{1},\ldots,R_{M}\}$ is a digraph collection of $M$ digraphs each on the same vertex set $[L]$, where, for $\{(h,i),j\}\in\binom{[L]}{2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}$ is $(\epsilon,d)$-regular.
+Given a digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on $V$ and parameters $\epsilon>0$, $d\in[0,1)$ and $L_{0}\geq 1$, the reduced digraph collection $\mathcal{R}:=\mathcal{R}(\epsilon,d,L_{0})$ of $\mathcal{D}$ is defined as follows.
+Apply Lemma 2.2 to $\mathcal{D}$ with parameters $\epsilon,\delta,d,L_{0}$ to obtain $\mathcal{D}^{\prime}$, a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$, and a partition $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}$, where $V_{0},\mathcal{C}_{0}$ are the exceptional sets and $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ are the color clusters.
+Then $\mathcal{R}:=\{R_{1},\ldots,R_{M}\}$ is a digraph collection of $M$ digraphs each on the same vertex set $[L]$, where, for $\{(h,i),j\}\in\binom{[L]}{2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}$ is $(\epsilon,d)$-regular.
 The next lemma states that clusters inherit a minimum semi-degree bound in the reduced digraph collection.
@@
 For each cluster $\mathcal{C}_{j}$ as above, choose an arbitrary color $c_{j}\in\mathcal{C}_{j}\cap\mathcal{D}_{v}$.
-Thus, the number of clusters $V_{h}$ containing some vertices in $N_{D_{c_{j}}^{\prime}}^{+}(v)$ is at least
+Thus, the number of vertex clusters $V_{h}$ containing some vertices in $N_{D_{c_{j}}^{\prime}}^{+}(v)$ is at least
 $\frac{d_{D_{c_{j}}}^{+}(v)-\sqrt{d}n}{m}\geq\frac{(p+\gamma-\sqrt{d})n}{m}\geq\left(p+\frac{\gamma}{2}\right)L.$
@@
 Therefore, for every $i\in[L]$, $d_{R_{j}}^{+}(i)\geq(p+\frac{\gamma}{2})L$ for at least $(1-\frac{d^{\frac{1}{4}}}{2})M$ colors $j$.
-The proofs of other cases are similar and we omit it.
+The proofs of other cases are similar and we omit the proof.
 ∎
 2.2 Transversal perfect matching in bipartite graph collections
+In the study of digraphs, a natural approach is to consider their characteristic bipartite graphs.
+This subsection introduces the corresponding characteristic bipartite graph collection associated with a digraph system, and studies the existence of transversal perfect matchings in this bipartite system.
 Let $D$ be a digraph with vertex set $V$.
-We define $B_{D}$ to be the characteristic bipartite graph of $D$ with bipartition $V_{1}\cup V_{2}$ where $V_{1}=V_{2}=V$, for each $u\in V_{1}$ and $v\in V_{2}$, $uv\in E(B_{D})$ if and only if $uv\in E(D)$.
+We define $B_{D}$ to be the characteristic bipartite graph of $D$ with bipartition $V_{1}\cup V_{2}$ where $V_{1}=V_{2}=V$; for each $u\in V_{1}$ and $v\in V_{2}$, $uv\in E(B_{D})$ if and only if $uv\in E(D)$.
 Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$.
@@
 We say that $M$ is a perfect matching if $V(M)=V(G)$.
-It is routine to check that $\mathcal{B_{D}}$ contains a transversal perfect matching if and only if the vertex set of $\mathcal{D}$ can be covered by a set of disjoint rainbow cycles inside $\mathcal{D}$.
-Therefore, in the following of this section, we are to find a transversal perfect matching in a bipartite graph collection.
+It is straightforward to check that $\mathcal{B_{D}}$ contains a transversal perfect matching if and only if the vertex set of $\mathcal{D}$ can be covered by a set of disjoint rainbow cycles inside $\mathcal{D}$.
+Therefore, in the remainder of this section, we are to find a transversal perfect matching in a bipartite graph collection.
 Joos and Kim [21] established a minimum degree condition that guarantees the existence of transversal perfect matchings in a graph collection.
@@
 Definition 2.6 (nice, extremal).
-Let $\epsilon>0$ and $G$ be a balanced bipartite graph on vertex set $V_{1}\cup V_{2}$ of size $2n$. We say that
+Let $\epsilon>0$ and $G$ be a balanced bipartite graph on the vertex set $V_{1}\cup V_{2}$ of size $2n$. We say that
   • $G$ is $\epsilon$-nice if for any two sets $A\subseteq V_{1}$ and $B\subseteq V_{2}$ of size at least $(\frac{1}{2}-\epsilon)n$ we have $e_{G}(A,B)\geq\epsilon n^{2}$;
@@
   • (i)
-    $A_{i},B_{i},C_{i}\subseteq V_{i}$, $|A_{i}|=|B_{i}|=(\frac{1}{2}-\epsilon)n$ and $|C_{i}|=2\epsilon n$ for $i\in[2]$,
+    $A_{i},B_{i},C_{i}\subseteq V_{i}$, $|A_{i}|=|B_{i}|=(\frac{1}{2}-\epsilon)n$ and $|C_{i}|=2\epsilon n$ for each $i\in[2]$,
   • (ii)
-    $d_{G}(v,X_{i})\geq(\frac{1}{2}-2\epsilon)n$ for $v\in X_{3-i}$ and $X\in\{A,B\}$, $i\in[2]$,
+    $d_{G}(v,X_{i})\geq(\frac{1}{2}-2\epsilon)n$ for $X\in\{A,B\}$ and $v\in X_{3-i}$ with $i\in[2]$,
   • (iii)
@@
 Definition 2.8 (crossing, cross graph).
-Let $10<\frac{1}{n},\epsilon,\delta<$ 1 where $n\in\mathbb{N}$ and let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a balanced bipartite graph collection on a common vertex set $V$ of size $2n$.
+Let $0<\frac{1}{n},\epsilon,\delta<1$ where $n\in\mathbb{N}$ and let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a balanced bipartite graph collection on a common vertex set $V$ of size $2n$.
 Given $i,j\in[n]$ such that $G_{i}$ and $G_{j}$ are both $\epsilon$-extremal, assume that the characteristic partition corresponding to $G_{\ell}$ is $(A_{1}^{\ell},B_{1}^{\ell},C_{1}^{\ell},A_{2}^{\ell},B_{2}^{\ell},C_{2}^{\ell})$ for $\ell\in\{i,j\}$.
-We say that they are $\delta$-crossing if $|A^{i}_{1}\Delta A^{j}_{1}|\geq\delta n$ and $|A^{i}_{1}\Delta B^{j}_{1}|\geq\delta n$.
+We say that they are $\delta$-crossing if $|A^{i}_{1}\triangle A^{j}_{1}|\geq\delta n$ and $|A^{i}_{1}\triangle B^{j}_{1}|\geq\delta n$.
 We define the cross graph $C_{\mathcal{G}}^{\epsilon,\delta}$ to be the graph with vertex set $[n]$ where $i$ is adjacent to $j$ if and only if $G_{i}$ and $G_{j}$ are both $\epsilon$-extremal and $\delta$-crossing.
@@
 We should note that, unless explicitly stated otherwise, those concepts in bipartite graph collections are only discussed in Section 2.2 and the appendix.).
-Let $10<\gamma,\alpha,\epsilon,\delta<$ 1.
+Let $0<\gamma,\alpha,\epsilon,\delta<1$.
 Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$.
@@
 3 Absorbing
-Denote by $V^{k}$ the set of all $k$-tuples of distinct elements of $V$.
-A pair $H=(V,E)$ where $V$ is a set and $E\subseteq V^{k}$ is called a directed $k$-graph, while if we allow $E$ to be a multiset, it is a directed multi-$k$-graph.
-We use the same notation for directed $k$-graphs as for graphs.
-The following lemma was obtained by Cheng and Staden [10], which is the key tool that we will use to build an absorbing structure.
-Lemma 3.1 ([10]).
-Suppose that $k,C,n\in\mathbb{N}$, $0\leq\frac{1}{n}\ll\gamma\ll\epsilon\ll\frac{1}{k},\frac{1}{C}\leq 1$, $m\in[n^{C}]$ and $t:=\gamma n$.
-Let $\mathbf{H}=\{H_{1},\ldots,H_{t}\}$ be a collection of directed $k$-graphs and let $\mathbf{Z}=\{Z_{1},\ldots,Z_{m}\}$ be a collection of directed multi-$k$-graphs all defined on a common vertex set $V$ of size $n$.
-Suppose that $|E(H_{i})|\geq\epsilon n^{k}$ for all $i\in[t]$, and for each $j\in[m]$ we have $|E(Z_{j})\cap E(H_{i})|\geq\epsilon n^{k}$ for at least $\epsilon t$ indices $i\in[t]$.
-Then there is a rainbow matching $M$ in $\mathbf{H}$ of size at least $(1-\frac{\epsilon^{2}}{4})t$ and $|E(Z_{j})\cap E(M)|\geq\frac{\epsilon^{2}t}{4}$ for each $j\in[m]$.
-We now define the extremality for a single digraph.
+In this section, we first define the extremality for a single digraph and distinguish the characteristic partition of such digraphs.
+We then introduce our absorbing structure and prove that every digraph collection contains a directed absorbing cycle whenever either a large proportion of its members are non-extremal, or almost all of them are extremal while their characteristic partitions differ significantly.
 Definition 3.1 (nice, extremal).
@@
   1. (i)
-[…]
+    there is a partition $A\cup B\cup L$ of $V$ such that $|A|=|B|=(\frac{1}{2}-\epsilon)n$, $d^{+}(a,{A}),d^{-}(a,{A})\geq(\frac{1}{2}-2\epsilon)n$ for all $a\in{A}$, $d^{+}(b,{B}),d^{-}(b,B)\geq(\frac{1}{2}-2\epsilon)n$ for all $b\in{B}$ and either $e(A,B)\leq\epsilon n^{2}$ or $e(B,A)\leq\epsilon n^{2}$; here we say $D$ is $(\epsilon,{\rm EC}1)$-extremal,
   2. (ii)
-[…]
+    there is a partition $A\cup B\cup L$ of $V$ such that $|A|=|B|=(\frac{1}{2}-\epsilon)n$, $d^{+}(a,B),d^{-}(a,B)\geq(\frac{1}{2}-2\epsilon)n$ for all $a\in{A}$, $d^{+}(b,A),d^{-}(b,A)\geq(\frac{1}{2}-2\epsilon)n$ for all $b\in{B}$ and either $e(A)\leq\epsilon n^{2}$ or $e(B)\leq\epsilon n^{2}$; here we say $D$ is $(\epsilon,{\rm EC}2)$-extremal,
   3. (iii)
-[…]
+    there is a constant $\zeta$ with $\epsilon^{1/3}\leq\zeta\leq\frac{1}{2}-\epsilon-\epsilon^{1/3}$ and a partition $C_{1}\cup C_{2}\cup C_{3}\cup C_{4}\cup L$ of $V$ such that $|C_{1}|=|C_{3}|=\zeta n$, $|C_{2}|=|C_{4}|=(\frac{1}{2}-\zeta-\epsilon)n$, for all $a\in C_{i}$ we have $d^{+}(a,C_{i})\geq(\zeta-\epsilon)n$ if $i\in\{1,3\}$, $d^{+}(a,C_{6-i})\geq(\frac{1}{2}-\zeta-2\epsilon)n$ if $i\in\{2,4\}$, and $d^{+}(a,C_{i+1})\geq(|C_{i+1}|-\epsilon)n$ for all $i\in[4]$ (here we identify $C_{5}$ with $C_{1}$), and either $e(C_{1},C_{3})\leq\epsilon n^{2}$ or $e(C_{3},C_{1})\leq\epsilon n^{2}$, either $e(C_{2})\leq\epsilon n^{2}$ or $e(C_{4})\leq\epsilon n^{2}$; here we say $D$ is $(\epsilon,{\rm EC}3)$-extremal, (see Figure 1 for the structures of EC1, EC2 and EC3).
 […]
@@
 The gray shaded elliptical indicates that the digraph induced by this vertex set is complete, the gray shaded arrow between two vertex sets indicates that the induced digraph by them is complete in this direction.
-[…]
-Definition 3.4 (directed absorbing path and absorbing cycle).
+We call the partition obtained in Lemma 3.2 a characteristic partition of $D$.
+Assume that $0<\frac{1}{n}\ll\mu\ll\epsilon\leq 1$ and $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq(\frac{1}{2}-\mu)n$.
+In view of Lemma 3.2, whenever $D_{i}$ is $\epsilon$-extremal, one may fix a characteristic partition $(A_{i},B_{i},L_{i})$ of $D_{i}$ if it is $(\epsilon,{\rm EC}k)$-extremal for some $k\in[2]$, a characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ of $D_{i}$ if it is $(\epsilon,{\rm EC}3)$-extremal.
+For convenience, define $W_{i}^{j}:=C_{i}^{j}\cup C_{i}^{j+1}$ for $j\in[4]$ (here we identify $C_{i}^{5}$ with $C_{i}^{1}$).
+We say that a vertex $v\in V$ is $D_{i}$-good if either $D_{i}$ is not $\epsilon$-extremal or $D_{i}$ is $\epsilon$-extremal and $v\in V\setminus L_{i}$.
+For every pair of distinct vertices $x,y\in V$, we define $L({xy}):=\{i\in[n]:{xy}\in E(D_{i})\}$ to be the set of colors appearing on xy{xy}.
+Definition 3.3 (directed absorbing path and directed absorbing cycle).
 Given any two not necessarily distinct vertices $u,v\in V$ and a rainbow directed path $P=v_{1}v_{2}v_{3}v_{4}$ with $u,v\not\in V(P)$, we call $P$ a Type-I directed $c$-absorbing path of $(v,u)$ if $c\in L({v_{2}v})$ and ${\rm col}({v_{2}v_{3}})\in L({uv_{3}})$ (see Figure 2), a Type-II directed $c$-absorbing path of $(v,u)$ if $c\in L({vv_{3}})$ and ${\rm col}({v_{2}v_{3}})\in L({v_{2}u})$ (see Figure 3).
@@
   • given any color $c\in\mathcal{C}$, for all but at most $\delta^{\prime}n$ $D_{c}$-good vertices $v$, there are at least $\gamma^{\prime}n$ disjoint Type-K directed $c$-absorbing paths of $(v,v)$ inside $C$.
-If there is no ambiguity, we typically refer to the Type-I and Type-II directed $c$-absorbing path (cycle) simply as directed $c$-absorbing path (cycle).
+If there is no ambiguity, we typically refer to the Type-I and Type-II directed $c$-absorbing path (cycle) simply as a directed $c$-absorbing path (cycle).
+[…]
 Figure 2: A Type-I directed $c$-absorbing path of $(v,v)$ and a Type-I directed $c$-absorbing path of $(v,u)$ with $v\neq u$.
+[…]
 Figure 3: A Type-II directed $c$-absorbing path of $(v,v)$ and a Type-II directed $c$-absorbing path of $(v,u)$ with $v\neq u$.
+Denote by $V^{k}$ the set of all $k$-tuples of distinct elements of $V$.
+A pair $H=(V,E)$ where $V$ is a set and $E\subseteq V^{k}$ is called a directed $k$-graph, while if we allow $E$ to be a multiset, it is a directed multi-$k$-graph.
+We use the same notation for directed $k$-graphs as for graphs.
+The following lemma was obtained by Cheng and Staden [9], which is the key tool that we will use to build an absorbing structure.
+Lemma 3.4 ([9]).
+Suppose that $k,C,n\in\mathbb{N}$, $0<\frac{1}{n}\ll\gamma\ll\epsilon\ll\frac{1}{k},\frac{1}{C}\leq 1$, $m\in[n^{C}]$ and $t:=\gamma n$.
+Let $\mathbf{H}=\{H_{1},\ldots,H_{t}\}$ be a collection of directed $k$-graphs and let $\mathbf{Z}=\{Z_{1},\ldots,Z_{m}\}$ be a collection of directed multi-$k$-graphs all defined on a common vertex set $V$ of size $n$.
+Suppose that $|E(H_{i})|\geq\epsilon n^{k}$ for all $i\in[t]$, and for each $j\in[m]$ we have $|E(Z_{j})\cap E(H_{i})|\geq\epsilon n^{k}$ for at least $\epsilon t$ indices $i\in[t]$.
+Then there is a rainbow matching $M$ in $\mathbf{H}$ of size at least $(1-\frac{\epsilon^{2}}{4})t$ and $|E(Z_{j})\cap E(M)|\geq\frac{\epsilon^{2}t}{4}$ for each $j\in[m]$.
+We first consider the case where a digraph collection contains many non-extremal digraphs; such a collection is said to be strongly stable.
 Definition 3.5 (strongly stable).
-Let $10<\gamma,\alpha<$ 1.
+Let $0<\gamma,\alpha<1$.
 Suppose that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$.
 We say that $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable if $D_{i}$ is $\alpha$-nice for at least $\gamma n$ colors $i\in[n]$.
-The following lemma finds a directed absorbing cycle when $\mathcal{D}$ is strongly stable.
+By applying Lemma 3.4, we establish the following lemma, which yields a directed absorbing cycle when $\mathcal{D}$ is strongly stable.
 Lemma 3.6.
-Let $1<\frac{1}{n}\ll\lambda,\mu\ll\gamma,\alpha\ll 1$.
+Let $0<\frac{1}{n}\ll\lambda,\mu\ll\gamma,\alpha\ll 1$.
 Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$.
@@
 Since $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, one may assume that $D_{1},\ldots,D_{\gamma n}$ are $\alpha$-nice and $6|\lambda n$.
-We divide the color set $[\frac{\lambda n}{2}]$ into consecutive sets $\{1,2,3\},\{4,5,6\},\ldots,\{\frac{\lambda n}{2}-2,\frac{\lambda n}{2}-1,\frac{\lambda n}{2}\}$.
+We divide the color set $\frac{\lambda n}{2}]$ into consecutive sets $\{1,2,3\},\{4,5,6\},\ldots,\{\frac{\lambda n}{2}-2,\frac{\lambda n}{2}-1,\frac{\lambda n}{2}\}$.
 For each $i\in I:=[\frac{\lambda n}{6}]$, define $\mathcal{F}_{i}$ to be a directed $4$-graph with vertex $V$ and edge set
@@
 Clearly, $|U|\geq n-\lambda n$.
-It follows from $\delta^{0}(\mathcal{D})\geq(\frac{1-\mu}{2})n$ that
+It follows from $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$ that
 $d_{{c_{1}}}^{+}(v_{4}^{1},U),\,d_{{c_{2}}}^{-}(v_{1}^{2},U)\geq\left(\frac{1-\mu}{2}-\lambda\right)n\geq\left(\frac{1}{2}-\alpha\right)n.$ (1)
@@
 We apply the above process for each pair $(P^{i},P^{i+1})$ for each $i\in[s]$ with $P^{s+1}:=P^{1}$.
-In each process, we need two unused colors from $[\lambda n+1,n]$ and one unused color from $[\frac{\lambda n}{2}+1,\lambda n]$.
+In each process, we need two unused colors from $\lambda n+1,n]$ and one unused color from $\frac{\lambda n}{2}+1,\lambda n]$.
 This is feasible since $2s\leq\frac{\lambda n}{3}$.
@@
 ∎
-The extremality of a digraph collection also depends on the relationships between the digraphs, so we give the following definitions.
+The extremality of a digraph collection also depends crucially on the relationships between different digraphs. This leads us to introduce the following definitions.
 Definition 3.7 (crossing, cross graph, weakly stable).
-Let $10<\frac{1}{n},\epsilon,\delta<$ 1 with $n\in\mathbb{N}$.
+Let $0<\frac{1}{n}\ll\epsilon,\delta<1$ with $n\in\mathbb{N}$.
 Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$.
@@
   1. (A1)
-    $t,s\in\{1,2\}$, $|A_{i}\Delta A_{j}|\geq\delta n$ and $|A_{i}\Delta B_{j}|\geq\delta n$;
+    $t,s\in\{1,2\}$, $|A_{i}\triangle A_{j}|\geq\delta n$ and $|A_{i}\triangle B_{j}|\geq\delta n$;
   2. (A2)
-    $t=s=3$, either $|W_{i}^{1}\Delta W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\Delta W_{j}^{3}|\geq\delta n$, or $|W_{i}^{2}\Delta W_{j}^{2}|\geq\delta n$ and $|W_{i}^{2}\Delta W_{j}^{4}|\geq\delta n$;
+    $t=s=3$, either $|W_{i}^{1}\triangle W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\triangle W_{j}^{3}|\geq\delta n$, or $|W_{i}^{2}\triangle W_{j}^{2}|\geq\delta n$ and $|W_{i}^{2}\triangle W_{j}^{4}|\geq\delta n$;
   3. (A3)
-    $t\in\{1,2\},s=3$, either $|A_{i}\Delta W_{j}^{1}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{3}|\geq\delta n$, or $|A_{i}\Delta W_{j}^{2}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{4}|\geq\delta n$.
+    $t\in\{1,2\},s=3$, either $|A_{i}\triangle W_{j}^{1}|\geq\delta n$ and $|A_{i}\triangle W_{j}^{3}|\geq\delta n$, or $|A_{i}\triangle W_{j}^{2}|\geq\delta n$ and $|A_{i}\triangle W_{j}^{4}|\geq\delta n$.
 For $k\in[3]$, define the cross graph $C_{\mathcal{D},k}^{\epsilon,\delta}$ to be the graph with vertex set $[n]$ where $i$ and $j$ are adjacent if and only if $D_{i},D_{j}$ are satisfying (Ak).
 We say that $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable if $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$.
+The subsequent observation is an immediate consequence of the above definition.
 Observation 3.8.
@@
 Lemma 3.9.
-Let $10<\frac{1}{n}\ll\lambda,\mu,\epsilon\ll\delta<$ 1 and $\delta\ll\min\{\zeta,\frac{1}{2}-\zeta-\epsilon\}$ (where $\zeta$ is defined in Lemma 3.3).
+Let $0<\frac{1}{n}\ll\lambda,\mu\ll 8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1$.
 Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$.
@@
 Since $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, one has $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$.
-In particular, if $k=2$, then we consider only $|W_{i}^{1}\Delta W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\Delta W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},2}^{\epsilon,\delta})$; if $k=3$, then we consider only $|A_{i}\Delta W_{j}^{1}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},3}^{\epsilon,\delta})$ (since the other case can be discussed similarly).
+In particular, if $k=2$, then we consider only $|W_{i}^{1}\triangle W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\triangle W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},2}^{\epsilon,\delta})$; if $k=3$, then we consider only $|A_{i}\triangle W_{j}^{1}|\geq\delta n$ and $|A_{i}\triangle W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},3}^{\epsilon,\delta})$ (since the other case can be discussed similarly).
 For each $i\in[\lfloor\frac{n}{3}\rfloor]$, define $\mathcal{F}_{i}$ to be a directed $4$-graph with vertex set $V$ and edge set
-$\{(v_{1},v_{2},v_{3},v_{4})\in V^{4}:3(i-1)+\ell\in L(v_{\ell}v_{\ell+1})\ \text{for}\ \ell\in[3]\}.$
+$\left\{(v_{1},v_{2},v_{3},v_{4})\in V^{4}:3(i-1)+\ell\in L(v_{\ell}v_{\ell+1})\ \text{for}\ \ell\in[3]\right\}.$
 If follows from $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$ that $e(\mathcal{F}_{i})\geq n\left(\left(\frac{1}{2}-\mu\right)n\right)^{3}\geq\frac{n^{4}}{9}$.
-We first give the following claim, which requires only the minimum semi-degree condition.
+In order to apply Lemma 3.4, we first give the following claim, which requires only the minimum semi-degree condition.
 Claim 3.10.
@@
 Therefore,
-$d_{i}^{-}(v,A_{j})\geq|A_{i}\cap A_{j}|-(|A_{i}|-d_{i}^{-}(v,A_{i}))\geq\frac{\delta n}{4}-\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\ll\delta).$
+$d_{i}^{-}(v,A_{j})\geq|A_{i}\cap A_{j}|-(|A_{i}|-d_{i}^{-}(v,A_{i}))\geq\frac{\delta n}{4}-\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\leq\delta^{2}).$
 Similarly, we have $d_{i}^{-}(v,B_{j})\geq\frac{\delta n}{5}$.
-[…]
+Since $u$ is $D_{j}$-good, one has $u\in Y_{j}$ for some $Y\in\{A,B\}$.
+Let $Z=Y$ if $D_{j}$ is $(\epsilon,{\rm EC}1)$-extremal, and let $Z\in\{A,B\}\setminus\{Y\}$ if $D_{j}$ is $(\epsilon,{\rm EC}2)$-extremal.
+Notice that $|N_{j}^{+}(w,Z_{j})|\geq(\frac{1}{2}-2\epsilon)n$ for each vertex $w\in Y_{j}$.
+Choose $x\in N_{i}^{-}(v,Y_{j})$.
+Then
 $|N_{j}^{+}(u,Z_{j})\cap N_{j}^{+}(x,Z_{j})|\geq|N_{j}^{+}(u,Z_{j})|+|N_{j}^{+}(x,Z_{j})|-|Z_{j}|\geq\left(\frac{1}{2}-3\epsilon\right)n.$
@@
 Case 2. $k=2$.
-In this case, we prove $|Z_{j}^{\rm II}(i,vu)|\geq 2^{-7}\delta^{2}n^{4}$.
+In this case, we prove $|Z_{j}^{\rm II}(i,vu)|\geq 2^{-7}{{\delta}}n^{4}$.
 By Lemma 3.2 (iii), we have either $d_{D_{i}}^{+}(v,W_{i}^{1})\geq(\frac{1}{2}-3\epsilon)n$ or $d_{D_{i}}^{+}(v,W_{i}^{3})\geq(\frac{1}{2}-3\epsilon)n$.
@@
 Therefore,
-$d_{i}^{+}(v,W_{j}^{1})\geq|W_{i}^{1}\cap W_{j}^{1}|-\left(|W_{i}^{1}|-d_{i}^{+}(v,W_{i}^{1})\right)\geq\frac{\delta n}{4}-2\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\ll\delta).$
+$d_{i}^{+}(v,W_{j}^{1})\geq|W_{i}^{1}\cap W_{j}^{1}|-\left(|W_{i}^{1}|-d_{i}^{+}(v,W_{i}^{1})\right)\geq\frac{\delta n}{4}-2\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\leq\delta^{2}).$
 Similarly, we have $d_{i}^{+}(v,W_{j}^{3})\geq\frac{\delta n}{5}$.
@@
 For each vertex $v$, let
-[…]
+$\mathcal{C}_{v}:=\{i\in[n]:v\ \text{is}\ D_{i}\text{-good}\}.$
 Since $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$, there is a subgraph $H$ of $C_{\mathcal{D},k}^{\epsilon,\delta}$ such that $|V(H)|\geq\delta n$ and $\delta(H)\geq\delta n$.
@@
 Notice that $|L_{i}|=2\epsilon n$ when $D_{i}$ is $\epsilon$-extremal.
-Hence $|T_{i}|\frac{\delta n}{2}\leq 2\epsilon n^{2}$, that is, $|T_{i}|\leq\frac{\sqrt{\epsilon}}{2}n$ since $\epsilon\ll\delta$.
-Let $\overline{T_{i}}:=V\setminus T_{i}$.
+Hence $|T_{i}|\frac{\delta n}{2}\leq 2\epsilon n^{2}$, that is, $|T_{i}|\leq\frac{\sqrt{\epsilon}}{2}n$ since $8\epsilon^{1/2}\leq\delta$.
+For each $i\in V(H)$, let $\overline{T_{i}}:=V\setminus T_{i}$.
 Then $|\overline{T_{i}}|\geq(1-\frac{\sqrt{\epsilon}}{2})n$.
-For each $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|N_{H}(i)\cap\mathcal{C}_{u}|\geq\frac{\delta n}{2}$.
+For each $u\in\overline{T_{i}}$ with $i\in V(H)$, one has $|N_{H}(i)\cap\mathcal{C}_{u}|\geq\frac{\delta n}{2}$.
 Now we independently and randomly select vertices from $V(H)$ with probability $\kappa:=\frac{\lambda}{14}$ to obtain a color set $\mathcal{U}$.
-By using Chernoff’s bound, with high probability the following hold:
+Using Chernoff’s bound, with high probability the following hold:
   1. (*)
     $\frac{\kappa|V(H)|}{2}\leq t:=|\mathcal{U}|\leq 2\kappa|V(H)|$,
-  2. (**∗ ∗)
+  2. (**)
     for every $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|N_{H}(i,\mathcal{U})\cap\mathcal{C}_{u}|\geq\frac{\delta\kappa n}{4}$.
@@
 Given $i\in\overline{\mathcal{U}}$ and $u\in\overline{T_{i}}$, choose $j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})$.
-Then $\frac{j+1}{3}\in[t]$ and the number of such choices of $j$ is at least $\frac{\delta\kappa n}{4}\geq\frac{\delta t}{8}$ by (**∗ ∗).
+Then $\frac{j+1}{3}\in[t]$ and the number of such choices of $j$ is at least $\frac{\delta\kappa n}{4}\geq\frac{\delta t}{8}$ by (**).
 Let $v$ be a $D_{i}$-good vertex.
@@
 $\gamma:=\frac{\lambda}{7},\ \epsilon:=2^{-7}\delta,\ \mathbf{H}:=\{\mathcal{F}_{j}:j\in[t]\}\ \ \text{and}$
-[…]
+$\mathbf{Z}:=\left\{Z(i,vu):=\bigcup_{j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})}Z_{j}^{\rm K}(i,vu):i\in\overline{\mathcal{U}},\ u\in\overline{T_{i}}\ \text{and}\ v\ \text{is}\ D_{i}\text{-good}\right\},$
 we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(1-2^{-16}\delta^{2})t$ such that $|E(Z(i,vu))\cap E(M)|\geq 2^{-16}\delta^{2}t$ for all $Z(i,vu)\in\mathbf{Z}$.
@@
 Next, choose an unused $D_{c_{1}}$-good vertex $x\in N^{+}_{{c_{3}}}(v_{4}^{1})$ and an unused $D_{c_{2}}$-good vertex $y\in N^{-}_{{c_{4}}}(v_{1}^{2})$.
+The following claim considers the number of choices for $x,y$.
 Claim 3.11.
 One of the following holds:
-  • there exists a rainbow directed path $P^{1}v_{4}^{1}xyzv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{2},c_{3},c_{4}\}$ and $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$, each of $x,y,z$ has at least $\frac{\delta n}{10}$ choices;
+  • there exists a rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{2},c_{3},c_{4}\}$ and $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$, each of $x,y,z$ has at least $\frac{\delta n}{10}$ choices;
   • there exists a rainbow directed path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$ and $y\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})$, each of $x,y$ has at least $\frac{\delta n}{10}$ choices.
@@
 Case 1. $k=1$.
-[…]
+In this case, $D_{c_{1}}$ and $D_{c_{2}}$ are either $(\epsilon,{\rm EC}1)$-extremal or $(\epsilon,{\rm EC}2)$-extremal.
+Then $x\in A_{c_{1}}\cup B_{c_{1}}$ and $y\in A_{c_{2}}\cup B_{c_{2}}$.
+Clearly, each of $x,y$ has at least $\frac{n}{4}$ choices.
+Recall that $D_{c_{1}}$ and $D_{c_{2}}$ are $\delta$-crossing.
+Then, $|X_{c_{1}}\cap Y_{c_{2}}|\geq\frac{\delta n}{4}$ whenever $X,Y\in\{A,B\}$.
+By Lemma 3.2, there are $Z,W\in\{A,B\}$ such that $d_{{c_{1}}}^{+}(x,Z_{c_{1}})\geq|Z_{c_{1}}|-\epsilon n$ and $d_{{c_{2}}}^{-}(y,W_{c_{2}})\geq|W_{c_{2}}|-\epsilon n$.
+Choose $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$ and the number of choices for $z$ is at least
 $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|-7e(M)\geq|N_{{c_{1}}}^{+}(x,Z_{c_{1}})\cap N_{{c_{2}}}^{-}(y,W_{c_{2}})|-14\kappa n\geq|Z_{c_{1}}\cap W_{c_{2}}|-2\epsilon n-12\kappa n\geq\frac{\delta n}{5},$
@@
 Case 2. $k=2$.
-[…]
+In this case, both $D_{c_{1}}$ and $D_{c_{2}}$ are $(\epsilon,{\rm EC}3)$-extremal.
+Then $|W_{c_{1}}^{i}\cap W_{c_{2}}^{j}|\geq\frac{\delta n}{4}$ for $i,j\in\{1,3\}$.
+Notice that either $|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|\geq\frac{n}{5}$ or $|W_{c_{1}}^{3}\cap W_{c_{2}}^{2}|\geq\frac{n}{5}$ holds.
+Without loss of generality, assume the latter case hold.
+Choose $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$.
+Next, we consider the size of $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|$.
 If $|W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})|\geq\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|\geq\delta n$, then choose $x\in W_{c_{1}}^{2}\cap N^{+}_{{c_{3}}}(v_{4}^{1})$ and $y\in W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})$.
-Hence, there are at least $\delta n-7e(M)-2\epsilon n\geq\frac{\delta n}{2}$ choices for each of $x$ and $y$.
+Hence, there are at least $\frac{\delta n}{2}$ choices for each of $x$ and $y$.
 Furthermore, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{3}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired.
@@
 Similarly, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{2}|-4\epsilon n\geq\frac{\delta n}{5}$, as desired.
-If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|<\frac{\delta n}{9}$, then $d^{-}_{{c_{4}}}(v_{1}^{2},W_{c_{2}}^{1})\geq|W_{c_{2}}^{1}|-(2\epsilon+\frac{\delta}{9})n$.
+If $|W_{c_{1}}^{1}\cap W_{c_{2}}^{4}|<\delta n$ and $|W_{c_{2}}^{3}\cap N^{-}_{{c_{4}}}(v_{1}^{2})|<\frac{\delta n}{9}$, then $d^{-}_{{c_{4}}}(v_{1}^{2},W_{c_{2}}^{1})\geq|W_{c_{2}}^{1}|-(\mu+2\epsilon+\frac{\delta}{9})n$.
 Therefore,
-$|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})|\geq|N_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\cap N_{{c_{4}}}^{-}(v_{1}^{2},W_{c_{2}}^{1})|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{1}|-\left(4\epsilon+\frac{\delta}{9}\right)n\geq\frac{\delta n}{9}.$
-Now, we connect $P^{1}$ and $P^{2}$ by using a directed rainbow path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
+$|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{1}^{2})|\geq|N_{{c_{1}}}^{+}(x,W_{c_{1}}^{1})\cap N_{{c_{4}}}^{-}(v_{1}^{2},W_{c_{2}}^{1})|\geq|W_{c_{1}}^{1}\cap W_{c_{2}}^{1}|-\left(\mu+4\epsilon+\frac{\delta}{9}\right)n\geq\frac{\delta n}{9}.$
+Now, we connect $P^{1}$ and $P^{2}$ by using a rainbow directed path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
 Case 3. $k=3$.
-[…]
+In this case, $D_{c_{1}}$ is either $(\epsilon,{\rm EC}1)$-extremal or $(\epsilon,{\rm EC}2)$-extremal, and $D_{c_{2}}$ is $(\epsilon,{\rm EC}3)$-extremal.
+Hence $x\in A_{c_{1}}\cup B_{c_{1}}$ and $y\in W_{c_{2}}^{1}\cup W_{c_{2}}^{3}$.
+Choose $z\in N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)$.
+Note that $D_{c_{1}}$ and $D_{c_{2}}$ are $\delta$-crossing.
+Then $|X_{c_{1}}\cap Y_{c_{2}}|\geq\frac{\delta n}{4}$ whenever $X\in\{A,B\}$ and $Y\in\{W^{1},W^{3}\}$.
+By Lemma 3.2, there are $Z\in\{A,B\}$ and $W\in\{W^{2},W^{4}\}$ such that $d_{{c_{1}}}^{+}(x,Z_{c_{1}})\geq|Z_{c_{1}}|-\epsilon n$ and $d_{{c_{2}}}^{-}(y,W_{c_{2}})\geq|W_{c_{2}}|-2\epsilon n$.
+If $|Z_{c_{1}}\cap W_{c_{2}}|\geq\frac{\delta n}{4}$, then $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{2}}}^{-}(y)|\geq|Z_{c_{1}}\cap W_{c_{2}}|-3\epsilon n\geq\frac{\delta n}{5}$, as desired.
 Now, we consider $|Z_{c_{1}}\cap W_{c_{2}}|<\frac{\delta n}{4}$.
@@
 Thus, $|N_{{c_{1}}}^{+}(x)\cap N_{{c_{4}}}^{-}(v_{2}^{1})|\geq|A_{c_{1}}\cap W_{c_{2}}^{3}|-(3\epsilon+\frac{\delta}{9})n\geq\frac{\delta n}{9}$.
-Now, we connect $P^{1}$ and $P^{2}$ by using a directed rainbow path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
+Now, we connect $P^{1}$ and $P^{2}$ by using a rainbow directed path $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$ with colors ${\rm col}(P^{1}\cup P^{2})\cup\{c_{1},c_{3},c_{4}\}$, as desired.
 This completes the proof of Claim 3.11. ∎
 In view of Claim 3.11, there are at least $\frac{\delta n}{10}$ choices for each of $c_{1},c_{2},c_{3},c_{4},x,y,(z)$ given any previous choices.
-Thus, we obtain the desired directed rainbow path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$.
+Thus, we obtain the desired rainbow directed path $P^{1}v_{4}^{1}xzyv_{1}^{2}P^{2}$ or $P^{1}v_{4}^{1}xyv_{1}^{2}P^{2}$.
 Applying the above process for each pair $(P^{j},P^{j+1})$ with $j\in[s]$ where $P^{s+1}:=P^{1}$.
 Every time we use at most four unused colors in $\overline{\mathcal{U}}$ and at most three unused vertices.
-Therefore, so for each choice, at most $4s\leq 4t\leq 8\kappa n\ll\frac{\delta n}{10}$ colors are forbidden and $3s\ll\frac{\delta n}{10}$ vertices are forbidden.
-Thus we obtain a directed rainbow cycle $C$ of length at most $7s$ which contains every $P^{j}$ with $j\in[s]$ as a segment.
+Therefore, for each choice, at most $4s\leq 4t\leq 8\kappa n\ll\frac{\delta n}{10}$ colors are forbidden and $3s\ll\frac{\delta n}{10}$ vertices are forbidden.
+Thus we obtain a rainbow directed cycle $C$ of length at most $7s$ which contains every $P^{j}$ with $j\in[s]$ as a segment.
 By the construction of $C$, for every color $i\in\overline{\mathcal{U}}$, any $D_{i}$-good vertex $v$ and for every vertex $u\in\overline{T_{i}}$, there are at least $2^{-16}\delta^{2}t\geq 2^{-17}\kappa\delta^{3}n$ disjoint directed $i$-absorbing paths of $(v,u)$.
@@
 4 Stable case
-In this section, by combining the results of the previous sections, we use the regularity-blow-up method in a digraph collection $\mathcal{D}$ to find a transversal directed Hamilton cycle when $\mathcal{D}$ is strongly stable or weakly stable.
-Let $10<\gamma,\alpha,\epsilon,\delta<$ 1.
-We say that $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable if it is either $(\gamma,\alpha)$-strongly stable or $(\epsilon,\delta)$-weakly table.
+In this section, we employ the regularity-blow-up method for digraph collections $\mathcal{D}$ to prove the existence of a transversal directed Hamilton cycle when $\mathcal{D}$ is either strongly or weakly stable, combining the results from previous sections.
+Let $0<\gamma,\alpha,\epsilon,\delta<1$.
+We say that $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable if it is either $(\gamma,\alpha)$-strongly stable or $(\epsilon,\delta)$-weakly stable.
 We first show that the reduced digraph collection inherits stability.
 Lemma 4.1.
-Let $0<\frac{1}{n}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu,\alpha\ll\gamma,\epsilon\ll\delta\ll 1$.
+Let ${0<}\frac{1}{n}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu,\alpha\ll\gamma,\epsilon\leq\delta^{2}\ll 1$.
 Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$.
-Let $\mathcal{R}=\mathcal{R}(\epsilon_{0},d,L_{0})$ be the reduced digraph collection of $\mathcal{G}$.
-If $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2},\epsilon,\frac{\delta}{2})$-stable.
+Let $\mathcal{R}=\mathcal{R}(\epsilon_{0},d,L_{0})$ be the reduced digraph collection of $\mathcal{D}$.
+If $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2},\epsilon,\frac{\delta}{2})$-stable.
 Proof.
@@
 Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $[n]$ and a digraph collection $\mathcal{D}^{\prime}$ satisfying (i)-(v) of Lemma 2.2.
-Therefore, for $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}:=\{D_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular.
+Therefore, for each $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}:=\{D_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular.
 Furthermore, $Lm\leq n\leq Mm+\epsilon_{0}n$ and $Mm\leq n\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$.
@@
 Claim 4.2.
-Let $d\ll\alpha^{\prime}$ and $i\in[L]$.
+Let $d\ll\alpha^{\prime}$ and $i\in[M]$.
 Assume that there are sets $A,B\subset[L]$ of size at least $(\frac{1}{2}-\alpha^{\prime})L$ such that $e_{R_{i}}(A,B)\leq\alpha^{\prime}L^{2}$.
@@
 Let $t$ be the number of colors $c\in\mathcal{C}_{i}$ such that $e_{D_{c}}(\hat{A},\hat{B})>\sqrt{\alpha^{\prime}}n^{2}$.
-Since $e(D_{c}^{\prime})-e(D_{c})\leq(3d+\epsilon_{0})n^{2}$ for all $c\in[n]$, we have
-$\sqrt{\alpha^{\prime}}n^{2}t$ $\leq\sum_{c\in\mathcal{C}_{i}}e_{D_{c}}(\hat{A},\hat{B})\leq\sum_{c\in\mathcal{C}_{i}}\left(e_{D_{c}^{\prime}}(\hat{A},\hat{B})+(3d+\epsilon_{0})n^{2}\right)$
-$\leq|\mathcal{C}_{i}|e_{R_{i}}(A,B)m^{2}+|\mathcal{C}_{i}|(3d+\epsilon_{0})n^{2}$
-$\leq\alpha^{\prime}L^{2}m^{3}+4dmn^{2}\leq 2\alpha^{\prime}mn^{2}.$
+Since $e(D_{c})-e(D_{c}^{\prime})\leq(24d+\epsilon_{0})n^{2}$ for all $c\in[n]$, we have
+$\sqrt{\alpha^{\prime}}n^{2}t$ $\leq\sum_{c\in\mathcal{C}_{i}}e_{D_{c}}(\hat{A},\hat{B})\leq\sum_{c\in\mathcal{C}_{i}}\left(e_{D_{c}^{\prime}}(\hat{A},\hat{B})+(24d+\epsilon_{0})n^{2}\right)$
+$\leq|\mathcal{C}_{i}|e_{R_{i}}(A,B)m^{2}+|\mathcal{C}_{i}|(24d+\epsilon_{0})n^{2}$
+$\leq\alpha^{\prime}L^{2}m^{3}+25dmn^{2}\leq 2\alpha^{\prime}mn^{2}.$
 This implies that $t\leq 2\sqrt{\alpha^{\prime}}m$. ∎
+We consider the cases where $\mathcal{D}$ is strongly stable and weakly stable, respectively.
 Case 1. $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable.
 We prove that $\mathcal{R}$ is $(\frac{\gamma}{2},\alpha^{2})$-strongly stable.
-Suppose that there exists a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for all $i\in I$.
+Suppose that there exists a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$.
 By Definition 3.1, for each $i\in I$, there are two sets $A^{i},B^{i}\subseteq[L]$ of size at least $(\frac{1}{2}-\alpha^{2})L$ such that $e_{R_{i}}(A^{i},B^{i})<\alpha^{2}L^{2}$.
-Applying Claim 4.2 with $\alpha^{\prime}:=\alpha^{2}$, we get that there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i_{0}}$ such that $e_{D_{c}}(\hat{A^{i}},\hat{B^{i}})\leq\alpha n^{2}$.
+Applying Claim 4.2 with $\alpha^{\prime}:=\alpha^{2}$, we get that there are at least $(1-2\alpha)m$ colors $c\in\mathcal{C}_{i}$ such that $e_{D_{c}}(\hat{A^{i}},\hat{B^{i}})\leq\alpha n^{2}$.
 Notice that $|\hat{A^{i}}|,|\hat{B^{i}}|\geq(\frac{1}{2}-\alpha^{2})Lm>(\frac{1}{2}-\alpha)n$.
@@
 Thus the number of colors $c\in[n]$ for which $D_{c}$ is not $\alpha$-nice is at least $(1-2\alpha)m|I|>(1-\gamma)n$ since $\alpha\ll\gamma$.
-This contradicts the fact that $\mathcal{D}$ is $(\gamma,\alpha)$-strong stable.
+This contradicts the fact that $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable.
 Case 2. $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable.
 Suppose that $\mathcal{R}$ is not $(\frac{\gamma}{2},\alpha^{2})$-strongly stable.
-Then there is subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$.
+Then there is a subset $I\subseteq[L]$ with $|I|\geq(1-\frac{\gamma}{2})L$ such that $R_{i}$ is not $\alpha^{2}$-nice for any $i\in I$.
 It suffices to show that $\mathcal{R}$ is $(\epsilon,\frac{\delta}{2})$-weakly stable.
@@
 Together with Lemma 3.2, we obtain that
-[…]
-[…]
+  • if $R_{i}$ is either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal, then $R_{i}$ has a characteristic partition $(A_{i}^{\prime},B_{i}^{\prime},L_{i}^{\prime})$ such that there are $Z,Y\in\{A,B\}$ satisfying $e_{R_{i}}(Z_{i}^{\prime},Y_{i}^{\prime})\leq\alpha^{2}L^{2}$;
+  • if $R_{i}$ is $(\alpha^{2},{\rm EC}3)$-extremal, then $R_{i}$ has a characteristic partition $({C_{1}^{i}}^{\prime},{C_{2}^{i}}^{\prime},{C_{3}^{i}}^{\prime},{C_{4}^{i}}^{\prime},L_{i}^{\prime})$ such that $e_{R_{i}}(Z_{i}^{\prime},Y_{i}^{\prime})\leq\alpha^{2}L^{2}$ with $(Z_{i}^{\prime},Y_{i}^{\prime})\in\{({W_{i}^{4}}^{\prime},{W_{i}^{3}}^{\prime}),({W_{i}^{2}}^{\prime},{W_{i}^{1}}^{\prime})\}$.
 Note that $|\hat{Z}_{i}^{\prime}|,|\hat{Y}_{i}^{\prime}|=(\frac{1}{2}-\alpha^{2})Lm\geq(\frac{1}{2}-2\alpha)n$.
-In view of Claim 4.2, we know that there is $\mathcal{B}_{i}\subseteq\mathcal{C}_{i}$ with $|\mathcal{C}_{i}\setminus\mathcal{B}_{i}|\leq 2\alpha m$ such that for all $c\in\mathcal{B}_{i}$, we have $e_{D_{c}}(\hat{Z}_{i}^{\prime},\hat{Y}_{i}^{\prime})\leq\alpha n^{2}$.
+In view of Claim 4.2, we know that there is a subset $\mathcal{B}_{i}\subseteq\mathcal{C}_{i}$ with $|\mathcal{C}_{i}\setminus\mathcal{B}_{i}|\leq 2\alpha m$ such that for all $c\in\mathcal{B}_{i}$, we have $e_{D_{c}}(\hat{Z}_{i}^{\prime},\hat{Y}_{i}^{\prime})\leq\alpha n^{2}$.
 It follows that $D_{c}$ is $2\alpha$-extremal for each $c\in\mathcal{B}_{i}$.
-[…]
-[…]
-Choose $c\in\mathcal{B}_{i},c^{\prime}\in\mathcal{B}_{j}$ for some $i,j\in[L]$. Suppose that $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing. We must have $i\neq j$.
-[…]
-$|A_{i}^{\prime}\Delta A_{j}^{\prime}|m=|\hat{A}_{i}^{\prime}\Delta\hat{A}_{j}^{\prime}|\geq|A_{c}\Delta A_{c^{\prime}}|-|\hat{A}_{i}^{\prime}\Delta A_{c}|-|\hat{A}_{j}^{\prime}\Delta A_{c^{\prime}}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
+If $R_{i}$ is either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal, then each $D_{c}$ with $c\in\mathcal{B}_{i}$ admits a characteristic partition $(A_{c},B_{c},L_{c})$, and there are $W\in\{A,B\}$ and $\{Z,Y\}=\{A,B\}$ such that $e_{D_{c}}(Z_{c},W_{c})\leq\sqrt{\alpha}n^{2}$ and $e_{D_{c}}(Y_{c},W_{c})\geq(\frac{1}{4}-\sqrt{\alpha})n^{2}$.
+Hence, we must have either $|\hat{Z}_{i}^{\prime}\triangle Z_{c}|,|\hat{Y}_{i}^{\prime}\triangle Y_{c}|\leq\epsilon n$ or $|\hat{Z}_{i}^{\prime}\triangle Y_{c}|,|\hat{Y}_{i}^{\prime}\triangle Z_{c}|\leq\epsilon n$.
+So, either $|\hat{A}_{i}^{\prime}\triangle A_{c}|,|\hat{B}_{i}^{\prime}\triangle B_{c}|\leq\epsilon n$ or $|\hat{A}_{i}^{\prime}\triangle B_{c}|,|\hat{B}_{i}^{\prime}\triangle A_{c}|\leq\epsilon n$.
+That is, for all $c\in\mathcal{B}_{i}$, the characteristic partition of $D_{c}$ is almost the same as the union of clusters corresponding to the characteristic partition of $R_{i}$.
+If $R_{i}$ is $(\alpha^{2},{\rm EC}3)$-extremal, then each $D_{c}$ admits a characteristic partition $(C_{c}^{1},C_{c}^{2},C_{c}^{3},C_{c}^{4},L_{c})$ for $c\in\mathcal{B}_{i}$, and it satisfies $e_{D_{c}}(W_{c}^{2},W_{c}^{1}),e_{D_{c}}(W_{c}^{4},W_{c}^{3})\leq\sqrt{\alpha}n^{2}$, $e_{D_{c}}(C_{c}^{1}),e_{D_{c}}(C_{c}^{3})\geq|C_{c}^{1}|^{2}-\sqrt{\alpha}n^{2}$ and $e_{D_{c}}(C_{c}^{2},C_{c}^{4}),\,e_{D_{c}}(C_{c}^{4},C_{c}^{2})\geq|C_{c}^{2}|^{2}-\sqrt{\alpha}n^{2}$.
+Therefore, either $|\hat{Z}_{i}^{\prime}\triangle W_{c}^{2}|,|\hat{Y}_{i}^{\prime}\triangle W_{c}^{1}|\leq\epsilon n$ or $|\hat{Z}_{i}^{\prime}\triangle W_{c}^{4}|,|\hat{Y}_{i}^{\prime}\triangle W_{c}^{3}|\leq\epsilon n$.
+By relabeling, one may assume that $|{\hat{W_{i}^{4}}^{\prime}}\triangle W_{c}^{4}|,|{\hat{W_{i}^{3}}^{\prime}}\triangle W_{c}^{3}|\leq\epsilon n$ for each $c\in\mathcal{B}_{i}$.
+That is, for all $c\in\mathcal{B}_{i}$, the characteristic partition of $D_{c}$ is almost the same as the union of clusters corresponding to the characteristic partition of $R_{i}$.
+Choose $c\in\mathcal{B}_{i},c^{\prime}\in\mathcal{B}_{j}$ for some $i,j\in[L]$.
+Suppose that $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing.
+We must have $i\neq j$.
+We consider the following possible cases.
+$\bullet$ Suppose that $R_{i}$ and $R_{j}$ are either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal.
+If $|\hat{A}_{i}^{\prime}\triangle A_{c}|\leq\epsilon n$ and $|\hat{A}_{j}^{\prime}\triangle A_{c^{\prime}}|\leq\epsilon n$, then
+$|A_{i}^{\prime}\triangle A_{j}^{\prime}|m=|\hat{A}_{i}^{\prime}\triangle\hat{A}_{j}^{\prime}|\geq|A_{c}\triangle A_{c^{\prime}}|-|\hat{A}_{i}^{\prime}\triangle A_{c}|-|\hat{A}_{j}^{\prime}\triangle A_{c^{\prime}}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
 The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
-[…]
-$|{W_{i}^{3}}^{\prime}\Delta{W_{j}^{3}}^{\prime}|m=|{\hat{W_{i}^{3}}^{\prime}}\Delta{\hat{W_{j}^{3}}^{\prime}}|\geq|W_{c}^{3}\Delta W_{c^{\prime}}^{3}|-|{\hat{W_{i}^{3}}^{\prime}}\Delta W_{c}^{3}|-|{\hat{W_{j}^{3}}^{\prime}}\Delta W_{c^{\prime}}^{3}|\geq\delta n-21\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
-and similarly, $|{W_{i}^{3}}^{\prime}\Delta{W_{j}^{1}}^{\prime}|\geq\frac{\delta L}{2}$. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
-[…]
-$|A_{i}^{\prime}\Delta{W_{j}^{3}}^{\prime}|m=|\hat{A}_{i}^{\prime}\Delta\hat{{W_{j}^{3}}^{\prime}}|\geq|A_{c}\Delta W_{c^{\prime}}^{3}|-|\hat{A}_{i}^{\prime}\Delta A_{c}|-|\hat{{W_{j}^{3}}^{\prime}}\Delta W_{c^{\prime}}^{3}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
-Similarly, $|B_{i}^{\prime}\Delta{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
+$\bullet$ Suppose that $R_{i}$ and $R_{j}$ are $(\alpha^{2},{\rm EC}3)$-extremal.
+Recall that $|{\hat{W_{i}^{4}}^{\prime}}\triangle W_{c}^{4}|,|{\hat{W_{i}^{3}}^{\prime}}\triangle W_{c}^{3}|\leq\epsilon n$.
+Then $|\hat{W_{i}^{1}}^{\prime}\triangle W_{c}^{1}|\leq 10\epsilon n$ and $|\hat{W_{i}^{2}}^{\prime}\triangle W_{c}^{2}|\leq 10\epsilon n$.
+Since $D_{c}$ and $D_{c^{\prime}}$ are $\delta$-crossing, one may assume, without loss of generality, that $|W_{c}^{1}\triangle W_{c^{\prime}}^{1}|\geq\delta n$ and $|W_{c}^{1}\triangle W_{c^{\prime}}^{3}|\geq\delta n$.
+Hence
+$|{W_{i}^{1}}^{\prime}\triangle{W_{j}^{1}}^{\prime}|m=|{\hat{W_{i}^{1}}^{\prime}}\triangle{\hat{W_{j}^{1}}^{\prime}}|\geq|W_{c}^{1}\triangle W_{c^{\prime}}^{1}|-|{\hat{W_{i}^{1}}^{\prime}}\triangle W_{c}^{1}|-|{\hat{W_{j}^{1}}^{\prime}}\triangle W_{c^{\prime}}^{1}|\geq\delta n-21\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
+and similarly, $|{W_{i}^{1}}^{\prime}\triangle{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
+$\bullet$ Suppose that $R_{i}$ is either $(\alpha^{2},{\rm EC}1)$-extremal or $(\alpha^{2},{\rm EC}2)$-extremal, and $R_{j}$ is $(\alpha^{2},{\rm EC}3)$-extremal.
+If $|\hat{A}_{i}^{\prime}\triangle A_{c}|\leq\epsilon n$ and $|\hat{{W_{j}^{3}}^{\prime}}\triangle W_{c}^{3}|\leq\epsilon n$, then
+$|A_{i}^{\prime}\triangle{W_{j}^{3}}^{\prime}|m=|\hat{A}_{i}^{\prime}\triangle\hat{{W_{j}^{3}}^{\prime}}|\geq|A_{c}\triangle W_{c^{\prime}}^{3}|-|\hat{A}_{i}^{\prime}\triangle A_{c}|-|\hat{{W_{j}^{3}}^{\prime}}\triangle W_{c^{\prime}}^{3}|\geq\delta n-2\epsilon n\geq\frac{\delta n}{2}\geq\frac{\delta mL}{2}.$
+Similarly, $|B_{i}^{\prime}\triangle{W_{j}^{3}}^{\prime}|\geq\frac{\delta L}{2}$. The other cases are almost identical. Thus $R_{i}$ and $R_{j}$ are $\frac{\delta}{2}$-crossing.
 Since $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, we have $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$.
@@
 ∎
+For a digraph collection $\mathcal{D}$ that is either strongly stable or weakly stable, Lemma 3.6 and Lemma 3.9 guarantee the existence of a directed absorbing cycle in $\mathcal{D}$.
+We shall then employ this absorbing property to prove that $\mathcal{D}$ contains a transversal directed Hamilton cycle.
 Lemma 4.3.
-Let $0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1$.
+Let $0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1$.
 Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$.
@@
 Proof.
-Choose additional parameters $n_{0},L_{0},\epsilon_{0},d,\theta,\beta,\lambda,\zeta$, where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is obtained from Lemma 2.2, and $\zeta$ is obtained from Lemma 3.3, such that
-$1<\frac{1}{n}\ll\frac{1}{n_{0}}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu\ll\beta\ll\lambda\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1,$
-and $\delta\ll\min\{\zeta,\frac{1}{2}-\zeta-\epsilon\}$, where the previous lemmas hold with suitable parameters.
+Choose additional parameters $n_{0},L_{0},\epsilon_{0},d,\beta,\lambda$, where $n_{0}=n_{0}(\epsilon,1,L_{0})$ is obtained from Lemma 2.2, such that
+$1<\frac{1}{n}\ll\frac{1}{n_{0}}\ll\frac{1}{L_{0}}\ll\epsilon_{0}\ll d\ll\mu\ll\beta\ll\lambda\ll\alpha\ll\gamma,8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1.$
 Since $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, in view of Lemma 3.6 and Lemma 3.9, we know that $\mathcal{D}$ has a directed absorbing cycle $C$ with parameters $(\frac{\delta}{3},\sqrt{\epsilon},\lambda,\lambda^{2})$.
@@
 We will prove that $|S(b_{1},b_{2},u_{1},u_{2},c)|\geq 2^{-4}n^{2}$.
-[…]
-[…]
+For this, we first count the number of choices for $v_{1}$.
+Let $N_{1}:=\{x\in N_{{c}}^{-}(u_{1}):x\ \text{is}\ D_{b_{1}}\text{-good}\}$.
+Hence, $|N_{1}|\geq\left(\frac{1}{2}-\mu\right)n-2\epsilon n$.
+By (b), we know that all but at most $\sqrt{\epsilon}n$ vertices $v_{1}\in N_{1}$ satisfying that $(b_{1},v_{1})$ is absorbable.
+Hence there are at least $\frac{n}{4}$ choices of $v_{1}$ such that $v_{1}\in N_{1}$ and $(b_{1},v_{1})$ is absorbable.
+Now fix $v_{1}$.
+Let $N_{2}:=\{x\in N_{{b_{2}}}^{+}(u_{2}):x\ \text{is}\ D_{b_{2}}\text{-good}\}$.
+Hence, $|N_{2}|\geq\left(\frac{1}{2}-\mu\right)n-2\epsilon n$.
+By (a) and (b), we know that all but at most $2\sqrt{\epsilon}n+1$ vertices $v_{2}\in N_{2}$ such that $v_{2}\neq v_{1}$ and $(b_{1},b_{2},v_{1},v_{2})$ is totally absorbable.
+Therefore, $|S(b_{1},b_{2},u_{1},u_{2},c)|\geq 2^{-4}n^{2}$, as desired.
 Let $\{c_{i},c_{i}^{\prime}:i\in[2\beta n]\}\subseteq\mathcal{C}$ be a collection of distinct colors.
-For each $i\in[2\beta n]$, define $\mathcal{F}_{i}$ to be a directed $2$-graph with vertex set $V$ and edge set $\{(v_{1},v_{2})\in V^{2}:(c_{i},c_{i}^{\prime},v_{1},v_{2})\ \text{is totally absorbable}\}$ and $\mathbf{H}:=\{\mathcal{F}_{i}:i\in[2\beta n]\}$.
+For each $i\in[2\beta n]$, let $\mathcal{F}_{i}$ be the directed 2-graph on $V$ with edge set $\{(v_{1},v_{2})\in V^{2}:(c_{i},c_{i}^{\prime},v_{1},v_{2})\ \text{is totally absorbable}\}$, and define $\mathbf{H}:=\{\mathcal{F}_{i}:i\in[2\beta n]\}$.
 Define the collection of directed multi-graphs:
@@
 For every $S:=S(u_{1},u_{2},c)\in\mathbf{Z}$ and every $i\in[2\beta n]$, we have $|E(S)\cap E(\mathcal{F}_{i})|\geq 2^{-4}n^{2}$.
-Applying Lemma 3.1 with $t=2\beta n$ and $\epsilon=2^{-4}$, we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(2-2^{-10})\beta n$ (and at most $2\beta n$) and $|E(S)\cap E(M)|\geq 2^{-9}\beta n$ for all $S\in\mathbf{Z}$.
+Applying Lemma 3.4 with $t:=2\beta n$ and $\epsilon:=2^{-4}$, we obtain that there is a rainbow matching $M$ of $\mathbf{H}$ with size at least $(2-2^{-10})\beta n$ (and at most $2\beta n$) and $|E(S)\cap E(M)|\geq 2^{-9}\beta n$ for all $S\in\mathbf{Z}$, as desired.
 ∎
@@
 $\mathcal{C}_{\rm rem}=[n]\setminus({\rm col}(C)\cup\mathcal{C}_{\rm abs}),\ \ \mathcal{J}:=\{J_{i}:i\in\mathcal{C}_{\rm rem}\}\ \ \text{where}\ \ J_{i}=D_{i}[U].$
-Thus, $|[n]\setminus\mathcal{C}_{\rm rem}|=|V\setminus U|\leq\lambda n+4\beta n<2\lambda n$ since $\beta\ll\lambda$, and $\delta^{0}(\mathcal{J})\geq(\frac{1}{2}-\mu-2\lambda)n\geq(\frac{1}{2}-3\lambda)n$.
+Thus, $\big|[n]\setminus\mathcal{C}_{\rm rem}\big|=|V\setminus U|\leq\lambda n+4\beta n<2\lambda n$ since $\beta\ll\lambda$, and $\delta^{0}(\mathcal{J})\geq(\frac{1}{2}-\mu-2\lambda)n\geq(\frac{1}{2}-3\lambda)n$.
 Since $\lambda\ll\gamma,\alpha,\epsilon,\delta$ and $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, it is easy to see that $\mathcal{J}$ is $(\frac{\gamma}{2},\frac{\alpha}{2},2\epsilon,\frac{\delta}{2})$-stable.
-Apply Lemma 2.2 to $\mathcal{J}$ with parameters $(\epsilon_{0},1,d,L_{0})$.
-Let $\mathcal{R}$ be the reduced digraph collection of $\mathcal{J}$.
-Write $[L]$ for the common vertex set of digraphs in $\mathcal{R}$ where $L_{0}\leq L\leq n_{0}$, and $[M]$ for the set of color clusters.
+Applying Lemma 2.2 to $\mathcal{J}$ with parameters $(\epsilon_{0},1,d,L_{0})$, we obtain a reduced digraph collection $\mathcal{R}$.
+Denote by $[L]$ the common vertex set of the digraphs in $\mathcal{R}$, where $L_{0}\leq L\leq n_{0}$, and by $[M]$ the set of color clusters.
 Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $U$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}_{\rm rem}$ and a digraph collection $\mathcal{J}^{\prime}$ satisfying (i)-(v) of Lemma 2.2.
@@
 Furthermore, $Lm\leq|U|\leq Mm+\epsilon_{0}n$ and $Mm\leq|\mathcal{C}_{\rm rem}|\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$.
-Thus we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$.
+Thus, we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$.
 Based on Lemma 2.4, we obtain that, for each vertex $i\in[L]$, there are at least $(1-d^{1/4})M$ colors $j\in[M]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$; for each color $j\in[M]$, there are at least $(1-d^{1/4})L$ vertices $i\in[L]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$.
@@
 Clearly, $U=\bigcup_{i\in[s]}V(P^{i})$.
-It follows that $|\mathcal{C}_{\rm rem}\setminus{\rm col}(\bigcup_{i\in[s]}P^{i})|=s$.
+It follows that $\big|\mathcal{C}_{\rm rem}\setminus{\rm col}(\bigcup_{i\in[s]}P^{i})\big|=s$.
 Next, we do the following operation for each $i\in[s]$ in turn.
@@
 This is feasible since Claim 4.4 implies that there are at least $2^{-9}\beta n$ choices for $Q_{i}$, of which at most $s\leq 2^{-10}\beta n$ have been used.
-Now, $(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})$ is totally absorbable, so there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j_{i}}$-absorbing path of $(v_{j_{i}},v_{j_{i}}^{\prime})$ inside ${C}$.
+Now, $(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})$ is totally absorbable, so there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j_{i}}$-absorbing path of $(v_{j_{i}},v_{j_{i}}^{\prime})$ inside C{C}.
 Then choose one of them, say $S_{i}:=x_{1}^{i}x_{2}^{i}x_{3}^{i}x_{4}^{i}$, whose vertices have not been previously chosen since $s<2^{-10}\beta n<2^{-10}\lambda^{2}n$, and whose colors are $b_{1}^{i},b_{2}^{i},b_{3}^{i}$ in order.
@@
 For each one, there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j}$-absorbing paths of $(v_{j},v_{j})$ and disjoint rainbow directed $c_{j}^{\prime}$-absorbing paths of $(v_{j}^{\prime},v_{j}^{\prime})$ inside $C$.
-Since $\beta<\lambda^{2}$, there are two such rainbow directed rainbow paths $T_{i},T_{i}^{\prime}$ for each $(c_{j},v_{j})$ and $(c_{j}^{\prime},v_{j}^{\prime})$, which are disjoint and whose vertices have not previously been chosen.
+Since $\beta<\lambda^{2}$, there are two such rainbow directed paths $T_{i},T_{i}^{\prime}$ for each $(c_{j},v_{j})$ and $(c_{j}^{\prime},v_{j}^{\prime})$, which are disjoint and whose vertices have not previously been chosen.
 At the end of this process, we have a collection $\{S_{i}:i\in[s]\},\ \{T_{i}:i\in I\},\ \{T_{i}^{\prime}:i\in I\}$ of disjoint rainbow directed paths inside $C$ satisfying
@@
 For each $i\in[s]$, we replace $S_{i}$ by $x_{1}^{i}x_{2}^{i}v_{j_{i}}x_{i}P^{i}y_{i}v_{j_{i}}^{\prime}x_{3}^{i}x_{4}^{i}$ with colors $b_{1}^{i},c_{j_{i}},a_{i}$ followed by the colors inherited from $P^{i}$, followed by $c_{j_{i}}^{\prime},b_{2}^{i},b_{3}^{i}$.
-That is, we have replaced $S_{i}$ by a directed rainbow path with the same endpoints, vertices $V(S_{i})\cup V(P^{i})\cup\{v_{j_{i}},v_{j_{i}}^{\prime}\}$, and colors ${\rm col}(S_{i})\cup{\rm col}(P^{i})\cup\{c_{j_{i}},c_{j_{i}}^{\prime},a_{i}\}$.
+That is, we have replaced $S_{i}$ by a rainbow directed path with the same endpoints, vertices $V(S_{i})\cup V(P^{i})\cup\{v_{j_{i}},v_{j_{i}}^{\prime}\}$, and colors ${\rm col}(S_{i})\cup{\rm col}(P^{i})\cup\{c_{j_{i}},c_{j_{i}}^{\prime},a_{i}\}$.
 For each $i\in I$, we replace $T_{i}:=y_{1}^{i}y_{2}^{i}y_{3}^{i}y_{4}^{i}$ by $y_{1}^{i}y_{2}^{i}v_{i}y_{3}^{i}y_{4}^{i}$ where colors are inherited except ${\rm col}(y_{2}^{i}v_{i})=c_{i}$ and ${\rm col}(v_{i}y_{3}^{i})={\rm col}(y_{2}^{i}y_{3}^{i})$.
@@
 ${\rm col}(T_{i}^{\prime})\cup\{c_{i}^{\prime}\}$).
-Thus we have obtained a directed rainbow cycle using vertices $V(C)\cup\bigcup_{i\in[s]}V(P^{i})\cup\{v_{i},v_{i}^{\prime}:i\in[r]\}=V$ and colors ${\rm col}(C)\cup\bigcup_{i\in[s]}{\rm col}(P^{i})\cup\{c_{i},c_{i}^{\prime}:i\in[r]\}=[n]$ where each color is used at most once (and hence exactly once).
+Thus we have obtained a rainbow directed cycle using vertices $V(C)\cup\bigcup_{i\in[s]}V(P^{i})\cup\{v_{i},v_{i}^{\prime}:i\in[r]\}=V$ and colors ${\rm col}(C)\cup\bigcup_{i\in[s]}{\rm col}(P^{i})\cup\{c_{i},c_{i}^{\prime}:i\in[r]\}=[n]$ where each color is used at most once (and hence exactly once).
 That is, $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
@@
 5 Extremal case
-Choose constants $\mu,\alpha,\gamma,\epsilon,\delta$ such that
-$0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1\ \text{and}\ \delta\ll\min\left\{\zeta,\frac{1}{2}-\zeta-\epsilon\right\},$
-where $\zeta$ is defined in Lemma 3.3.
-Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$.
+In this section, we consider the extremal case, in which most digraphs in the collection are $(\epsilon,\mathrm{ECk})$-extremal for some $k\in[3]$.
+The following result follows directly from Lemma 4.3.
+Lemma 5.1.
+Assume that $0<\frac{1}{n}\ll 8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1$ and $m=(1-4\sqrt{\delta})n$.
+Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$.
+If $\mathcal{D}$ contains no transversal directed Hamilton cycles, then by swapping indices if necessary, one of the following holds:
+  1. (B1)
+    either $D_{i}$ is $(\epsilon,{\rm EC1})$-extremal or $(\epsilon,{\rm EC2})$-extremal and it admits an $\epsilon$-characteristic partition $(\tilde{A}_{i},\tilde{B}_{i},\tilde{L}_{i})$ for each $i\in[m]$,
+  2. (B2)
+    or $D_{i}$ is $(\epsilon,{\rm EC3})$-extremal and it admits an $\epsilon$-characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each $i\in[m]$.
+In (B1), for every $i\in[m]$, $D_{i}$ also admits a partition $(A_{i},B_{i},L_{i})$ satisfying
+  • (C1)
+    if $i=1$ then $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$; for $i\geq 2$,
+  • (C2)
+    if $D_{i}$ is $(\epsilon,K_{\lceil\frac{n}{2}\rceil}\cup K_{\lfloor\frac{n}{2}\rfloor})$-extremal, then for $Y\in\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{D_{i}}(v,Y_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
+  • (C3)
+    if $D_{i}$ is $(\epsilon,K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor})$-extremal, then for $\{Y,Z\}=\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{D_{i}}(v,Z_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
+  • (C4)
+    subject to (C2)-(C3), the partition $(A_{i},B_{i},L_{i})$ is chosen such that $|A_{i}\cup B_{i}|$ is maximized for each $i\in[2,m]$.
+Furthermore, for every $i\in[m]$ and $Y\in\{A,B\}$, we have
+$|L_{i}|\leq 2\epsilon n\ \ \text{and}\ \ |{Y}_{1}{\triangle}Y_{i}|<2\delta n.$
+Proof.
+Choose constants $\alpha,\gamma,\epsilon,\delta$ such that
+$0<\frac{1}{n}\ll\alpha\ll\gamma,8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1.$
+Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$.
 Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles.
-By Lemma 4.3, we know $\mathcal{D}$ is not $(\gamma,\alpha)$-strongly stable.
-Hence there are at least $(1-\gamma)n$ digraphs in $\mathcal{D}$ such that each of them is $\alpha$-extremal.
+By Lemma 4.3, we know that $\mathcal{D}$ is not $(\gamma,\alpha)$-strongly stable.
+Hence, there are at least $(1-\gamma)n$ digraphs in $\mathcal{D}$ such that each of them is $\alpha$-extremal.
 Without loss of generality, assume that $D_{i}$ is $\alpha$-extremal for each $i\in[(1-\gamma)n]$.
-Note that every $\alpha$-extremal digraph is also $\epsilon$-extremal, since $\alpha\ll\epsilon$.
+Note that every $\alpha$-extremal digraph is also $\epsilon$-extremal, since $\alpha<\epsilon$.
 Denote
+$\tilde{\mathcal{C}}_{1}$ $:=\{i\in[n]:D_{i}\ \text{is}\ (\epsilon,{\rm EC}1)\text{-extremal}\},$
+$\tilde{\mathcal{C}}_{2}$ $:=\{i\in[n]:D_{i}\ \text{is}\ (\epsilon,{\rm EC}2)\text{-extremal}\},$
+$\tilde{\mathcal{C}}_{3}$ $:=\{i\in[n]:D_{i}\ \text{is}\ (\epsilon,{\rm EC}3)\text{-extremal}\}.$
 […]
-[…]
-[…]
-Hence $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|+|\tilde{\mathcal{C}}_{3}|\geq(1-\gamma)n$.
-By Lemma 3.3, we know that $D_{i}$ has a characteristic partition $(\tilde{A_{i}},\tilde{B_{i}},\tilde{L_{i}})$ for each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, or a characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each $i\in\tilde{\mathcal{C}}_{3}$.
-Under the above characteristic partition, applying Lemma 4.3 yields that $\mathcal{D}$ is not $(\epsilon,\delta)$-weakly stable.
-For $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{1}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing.
-For $i\in\tilde{\mathcal{C}}_{3}$, define $I^{2}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing.
-For $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{3}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing.
+Under the above characteristic partition, applying Lemma 4.3 again yields that $\mathcal{D}$ is not $(\epsilon,\delta)$-weakly stable.
+For each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{1}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing.
+For each $i\in\tilde{\mathcal{C}}_{3}$, define $I^{2}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing.
+For each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{3}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing.
 Then we have
@@
 Hence there exists a color set, say $I^{k}$ (where $I^{k}\subseteq\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ if $k=1$ and $I^{k}\subseteq\tilde{\mathcal{C}}_{3}$ if $k\in\{2,3\}$), with size at most $\sqrt{\delta}n$ such that the following hold:
-  1. (B1)
-    if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\Delta\tilde{A}_{i}|<\delta n$ and $|\tilde{B}_{1}\Delta\tilde{B}_{i}|<\delta n$ for all $i\in(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$,
-  2. (B2)
-    if $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels, for all $k\in[4]$ and all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{2}$ we have $|W_{1}^{k}\Delta W_{i}^{k}|<\delta n$, in particular $|C_{1}^{k}\Delta C_{i}^{k}|<2\delta n$,
-  3. (B3)
-    if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\Delta W_{i}^{1}|<\delta n$ and either $|\tilde{A}_{1}\Delta W_{i}^{2}|<\delta n$ or $|\tilde{A}_{1}\Delta W_{i}^{4}|<\delta n$ for all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$.
-Note that if ${\bf(B3)}$ holds, then for each $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$, there is an $i_{0}\in\{2,4\}$ such that
-$2\delta n>|\tilde{A}_{1}\Delta W_{i}^{1}|+|\tilde{A}_{1}\Delta W_{i}^{i_{0}}|\geq|W_{i}^{1}\Delta W_{i}^{i_{0}}|>2\delta n,$
-a contradiction.
-This implies that exactly one of $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$ holds.
-Based on (B1)-(B2), we proceed our proof by considering the following two theorems.
+  • if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\triangle\tilde{A}_{i}|<\delta n$ and $|\tilde{B}_{1}\triangle\tilde{B}_{i}|<\delta n$ for all $i\in(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$,
+  • if $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels, for all $k\in[4]$ and all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{2}$ we have $|W_{1}^{k}\triangle W_{i}^{k}|<\delta n$, in particular $|C_{1}^{k}\triangle C_{i}^{k}|<2\delta n$,
+  • if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\triangle W_{i}^{1}|<\delta n$ and either $|\tilde{A}_{1}\triangle W_{i}^{2}|<\delta n$ or $|\tilde{A}_{1}\triangle W_{i}^{4}|<\delta n$ for all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$.
+Note that if last itemize holds, then for each $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$, there is an $i_{0}\in\{2,4\}$ such that
+$2\delta n>|\tilde{A}_{1}\triangle W_{i}^{1}|+|\tilde{A}_{1}\triangle W_{i}^{i_{0}}|\geq|W_{i}^{1}\triangle W_{i}^{i_{0}}|>2\delta n,$
+a contradiction. This implies that exactly one of $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$ holds.
+[…]
+  • if $D_{i}$ is $(\epsilon,{\rm EC}1)$-extremal, then for $Y\in\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{i}^{+}(v,Y_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
+  • if $D_{i}$ is $(\epsilon,{\rm EC}2)$-extremal, then for $Y\in\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{i}^{+}(v,Z_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
+  • subject to the above two conditions, the partition $(A_{i},B_{i},L_{i})$ is chosen such that $|A_{i}\cup B_{i}|$ is maximized for each $i\in[2,m]$.
+Denote $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$.
+Hence for every $i\in[m]$ and $Y\in\{A,B\}$, we have $|Y_{1}\triangle Y_{i}|\leq|\tilde{Y_{1}}\triangle\tilde{Y_{i}}|+|\tilde{L}_{i}|\leq\delta n+2{\epsilon}n<2\delta n$, as desired.
+∎
+Based on Lemma 5.1, we proceed with the proof by considering the following two theorems.
 Theorem 5.2.
-Assume that $\delta$ is defined in (5).
+Assume that $0<\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll 1$ and $m=(1-4\sqrt{\delta})n$.
 Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$.
-If $|\tilde{\mathcal{C}}_{3}|<2\sqrt{\delta}n$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
+If $D_{i}$ is $(\epsilon,{\rm EC1})$-extremal or $(\epsilon,{\rm EC2})$-extremal for each $i\in[m]$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
 Theorem 5.3.
-Assume that $\delta$ is defined in (5).
+Assume that $0<\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll 1$ and $m=(1-4\sqrt{\delta})n$.
 Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$.
-If $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
-In the following of this section, we prove the above two theorems respectively.
-5.1 Proof of Theorem 5.1
-In this subsection, we give the proof of Theorem 5.1.
-Proof of Theorem 5.1.
-Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles.
-Let $\epsilon,\delta$ be defined in (5), and choose $\eta$ with $\delta\ll\eta\ll 1$.
-Denote $[m]:=(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$ and $\mathcal{C}_{\rm bad}:=[m+1,n]$.
-By adding colors to $\mathcal{C}_{\rm bad}$ if necessary we may assume $m=(1-4\sqrt{\delta})n$.
-Recall that $\epsilon\ll\delta$.
-Hence for each $i\in[2,m]$, there exists a new partition $(A_{i},B_{i},L_{i})$ of $D_{i}$ such that
-  • (C1)
+If $D_{i}$ is $(\epsilon,{\rm EC3})$-extremal for each $i\in[m]$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
+In the following two sections, we prove the above two theorems respectively.
+6 Proof of Theorem 5.2
+In this section we prove Theorem 5.2.
+Following the outline presented in Section 1.4, we first identify several types of “bad” vertices and cover them by short rainbow directed paths whose endpoints possess additional properties.
+Next we handle the “bad” colors by choosing matchings.
+We then connect all these short rainbow directed paths via a connecting lemma.
+Finally, by the transversal blow-up lemma, we construct a long transversal directed path, which can be closed to form the desired transversal directed cycle.
+For the sake of clarity in presenting the following lemmas, we summarize some properties of the digraph collection $\mathcal{D}$.
+  1. $(\dagger)$
+    Let $0<\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll\eta\ll 1$ and $m=(1-4\sqrt{\delta})n$.
+Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta(\mathcal{D})\geq\frac{n}{2}$.
+We assume the following conditions hold.
+    1. (a)
+      For every $i\in[m]$, $D_{i}$ is either $(\epsilon,{\rm EC1})$-extremal or $(\epsilon,{\rm EC2})$-extremal and it admits a partition $({A_{i}},{B_{i}},{L_{i}})$ satisfying (C1)-(C4).
+We extend $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$.
+    2. (b)
+      Define
+      $\mathcal{C}_{1}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC1})\textrm{-extremal}\},$
+      $\mathcal{C}_{2}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC2})\textrm{-extremal}\},\ \ \text{and}\ \ \mathcal{C}_{\rm bad}=\mathcal{C}\setminus(\mathcal{C}_{1}\cup\mathcal{C}_{2}).$
+      Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise.
+    3. (c)
 […]
-  • (C2)
+    4. (d)
+      Define
+      $X:=\{x\in V:x\not\in A_{i}\cup B_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in\hat{\mathcal{C}}\}$.
 […]
-  • (C3)
-    subject to (C1)-(C2), each vertex in $V$ lies in as many as possible $A_{i}\cup B_{i}$ for $i\in[2,m]$.
-Denote $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$.
-Expand $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$.
-For simplicity, we assume $Y\in\{A,B\}$ and $Z\in\{A,B\}\setminus\{Y\}$ in the following.
-Let $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ be a partition of $[m]$ with
-[…]
-[…]
-Hence for $i\in[2,m]$, we have $|Y_{1}\Delta Y_{i}|\leq|\tilde{Y_{1}}\Delta\tilde{Y_{i}}|+|\tilde{L}_{i}|\leq\delta n+2{\epsilon}n<2\delta n$.
-Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise.
-Clearly, $|\hat{\mathcal{C}}|\geq(1-4\sqrt{\delta}-\eta)n$.
-Define
-$X:=$$\{x\in V:x\not\in A_{i}\cup B_{i}\leavevmode\nobreak\ \textrm{forat least}\leavevmode\nobreak\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in\hat{\mathcal{C}}\},$
-$X_{Y}:=$$\{x\in Y\setminus X:x\not\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 10\sqrt{\delta}|\hat{\mathcal{C}}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in\hat{\mathcal{C}}\}.$
-Recall that $|L_{i}|\leq 2{\epsilon}n$ and $|Y_{1}\Delta Y_{i}|<2\delta n$ for each $i\in\hat{\mathcal{C}}$. Hence $6\sqrt{\delta}|\hat{\mathcal{C}}||X|\leq 2{\epsilon}n|\hat{\mathcal{C}}|$ and
-$10\sqrt{\delta}|\hat{\mathcal{C}}||X_{Y}|\leq\sum_{i\in\hat{\mathcal{C}}}|Y\setminus Y_{i}|\leq\sum_{i\in\hat{\mathcal{C}}}(|Y\setminus Y_{1}|+|Y_{1}\Delta Y_{i}|)\leq(2\delta+2{\epsilon})n|\hat{\mathcal{C}}|.$
-It follows that $|X|<\frac{1}{3}\sqrt{\epsilon}n$ and $|X_{Y}|<\frac{1}{4}\sqrt{\delta}n$. For $k\in[2]$, define
-$X_{Y}^{k}=\left\{\begin{array}[]{ll}\{x\in X_{Y}:x\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 3\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{k}}\},&\textrm{if $|\mathcal{C}_{k}|\geq\eta n$,}\\[5.0pt] \emptyset,&\textrm{otherwise.}\end{array}\right.$
-For a vertex $x\in X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})$, we know $x\in Y_{i}$ for at most $6\sqrt{\delta}n$ colors $i\in\hat{\mathcal{C}}$ and $x\in A_{i}\cup B_{i}$ for at least $(1-6\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
-Thus, $x\in Y\setminus Y_{i}$ for at least $|\hat{\mathcal{C}}|-6\sqrt{\delta}n$ colors $i\in\hat{\mathcal{C}}$ and $x\in Z_{i}$ for at least $(1-6\sqrt{\delta})|\hat{\mathcal{C}}|-6\sqrt{\delta}n\geq(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
-Hence
-$|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|(|\hat{\mathcal{C}}|-6\sqrt{\delta}n)\leq\sum_{i\in\hat{\mathcal{C}}}|Y\setminus Y_{i}|\leq(2\delta+2\epsilon)n|\hat{\mathcal{C}}|.$
-It follows that $|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|<4\delta n$.
-Move vertices in $X_{A}\setminus(\bigcup_{k\in[2]}X_{A}^{k})$ to $B$ and vertices in $X_{B}\setminus(\bigcup_{k\in[2]}X_{B}^{k})$ to $A$.
-Let $X^{\prime}$ be a subset of $X$ consisting of vertices $x$ such that $d_{i}^{+}(x,Y_{i})\geq(1-3\sqrt{\delta})n$ (resp.
-$d_{i}^{-}(x,Y_{i})\geq(1-3\sqrt{\delta})n$) for at least $(1-3\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$ and $d_{i}^{-}(x,Z_{i})\geq\frac{5}{2}\sqrt{\delta}n$ (resp.
-$d_{i}^{+}(x,Z_{i})\geq\frac{5}{2}\sqrt{\delta}n$) for at least $3\sqrt{\delta}|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
-Delete vertices in $X^{\prime}$ from $A\cup B$.
-Without loss of generality, assume that $|B|-|A|=r$ with $0\leq r\leq 8{\delta}n+\sqrt{\epsilon}n$.
-Define $V_{\rm bad}:=\bigcup_{k\in[2]}(X^{k}_{A}\cup X^{k}_{B})\cup(X\setminus X^{\prime})$.
-Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+\frac{1}{2}\sqrt{\delta})n$.
-Moreover, each vertex in $Y\setminus V_{\rm bad}$ lies in $Y_{i}$ for at least $(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
-The subsequent result can be used to connect two disjoint short rainbow directed paths into a single short rainbow directed path.
-Claim 5.3 (Connecting tool).
+The next lemma shows that any vertex contained in $L_{i}$ for many colors $i\in\mathcal{C}$ can be covered by a rainbow directed path $P_{3}$ with endpoints not in $V_{\rm bad}$.
+To enhance flexibility, we select a rainbow star $S_{5}$ (instead of a simple directed $P_{3}$), which allows each endpoint of the resulting path to be replaced by an alternative available vertex when needed.
+Lemma 6.1.
+Suppose that $(\dagger)$6 holds and $\{Y,Z\}=\{A,B\}$.
+Let $P$ be a rainbow directed path in $\mathcal{D}$ with length at most $\frac{7}{3}\sqrt{\delta}n$.
+Assume $X\setminus(X^{\prime}\cup V(P))=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$.
+If $x_{i}\in X\cap Y$, then there exist colors $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{j}\setminus({\rm col}(P)\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and vertices $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},x_{i}^{4}\in W\setminus(V(P)\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$, where
+  • $j=2$ and $W=Z$ if $|\mathcal{C}_{1}|<\eta n$,
+  • $j=1$ and $W=Y$ if $|\mathcal{C}_{2}|<\eta n$,
+  • either $j=1$ and $W=Y$, or $j=2$ and $W=Z$ otherwise.
+Proof.
+We only prove the case that $|\mathcal{C}_{1}|<\eta n$, and the other cases can be proved by similar arguments.
+Suppose there exists an $i_{0}\in[s]$ such that Lemma 6.1 holds for all $i\in[i_{0}-1]$ but does not hold for $i_{0}$.
+Assume for convenience that $x_{i_{0}}\in X\cap A$.
+Denote $\mathcal{C}_{2}^{\prime}:=\mathcal{C}_{2}\setminus({\rm col}(P)\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i_{0}-1]\})$.
+Let
+$\mathcal{C}_{2}^{1}:=\big\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{+}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P)|\big\}\ \text{and}$
+$\mathcal{C}_{2}^{2}:=\big\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{-}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P)|\big\}.$
+Hence either $|\mathcal{C}_{2}^{1}|\leq 3$ or $|\mathcal{C}_{2}^{2}|\leq 3$.
+Without loss of generality, assume that $|\mathcal{C}_{2}^{1}|\leq 3$.
+Thus, for each $j\in\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}$, we have
+$d_{j}^{+}(x_{i_{0}},A_{j})$ $\geq d_{j}^{+}(x_{i_{0}},A)-|A\setminus A_{1}|-|A_{1}\triangle A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
+$\geq\delta^{0}(D_{j})-d_{j}^{+}(x_{i_{0}},B)-|A\setminus A_{1}|-|A_{1}\triangle A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
+$\geq\frac{n}{2}-\left(4(i_{0}-1)+4+|V(P)|\right)-2{\epsilon}n-2\delta n-4\delta n-\sqrt{\epsilon}n$
+$\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n.$
+Furthermore, $|\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}|\geq(1-3\sqrt{\delta})|\mathcal{C}_{2}|$.
+Notice that $x_{i_{0}}\notin X^{\prime}$.
+Then, for at least $(1-3\sqrt{\delta})|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, $d_{j}^{-}(x_{i_{0}},B_{j})<\frac{5}{2}\sqrt{\delta}n$, and consequently, $d_{j}^{-}(x_{i_{0}},A_{j})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for the same set of colors.
+By (C4), we know that $x_{i_{0}}\in A_{j}\cup B_{j}$ for all but at most $6\sqrt{\delta}|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, which leads to $x_{i_{0}}\not\in X$, a contradiction.
+∎
+Next, we give an application of the transversal blow-up lemma (see [10]) for embedding transversal undirected Hamilton paths inside very dense bipartite graph collections, which can be proved by minor modifications to the proof of [9, Lemma 6.1], and we omit the proof here.
+We say that $P$ is an undirected subgraph of a digraph collection $\mathcal{D}=\{{D_{1}},\ldots,D_{n}\}$ if, for every edge $uv\in E(P)\cap E(D_{i})$, the reverse arc $vu\in E(D_{i})$.
+Lemma 6.2.
+Suppose $(\dagger)$6 holds and $\{W,Z\}=\{A,B\}$.
+Let $W^{*}\subseteq W\setminus V_{\rm bad}$ and $Z^{*}\subseteq Z\setminus V_{\rm bad}$, where $|W^{*}|,|Z^{*}|\geq\eta n$, $W^{*}\cap Z^{*}=\emptyset$ and $|W^{*}|-|Z^{*}|=t\in\{0,1\}$.
+Let $T^{*}=Z^{*}$ if $t=0$ and $T^{*}=W^{*}$ if $t=1$.
+Let $\mathcal{C^{*}}\subseteq\mathcal{C}$ satisfy $|\mathcal{C^{*}}|=|W^{*}|+|Z^{*}|-1$, where $\mathcal{C^{*}}\subseteq\mathcal{C}_{1}$ if $W=Z$ and $\mathcal{C^{*}}\subseteq\mathcal{C}_{2}$ if $W\neq Z$.
+Let $W^{-}\subseteq W^{*}$ and $T^{+}\subseteq T^{*}$ with $|W^{-}|,|T^{+}|\geq\frac{\eta n}{8}$.
+Then there is a transversal undirected Hamilton path in $\{D_{i}^{\pm}[W^{*},Z^{*}]:i\in\mathcal{C}^{*}\}$ starting at $W^{-}$ and ending at $T^{+}$.
+The subsequent result can be used to connect two disjoint short rainbow paths into a single short rainbow path.
+Lemma 6.3 (Connecting tool).
+Suppose that $(\dagger)$6 holds and $\{Y,Z\}=\{A,B\}$.
 Assume $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ are two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s},v_{1}\notin V_{\rm bad}\cup X^{\prime}$ and $s+t\leq 5\eta n$.
@@
     If $u_{s},v_{1}\in Y$ and $|\mathcal{C}_{1}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{1}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
-Proof of Claim 5.3.
+Proof.
 We only give the proof of (i), the other two statements can be proved by similar discussions, whose procedures are omitted.
-Note that $u_{s},v_{1}\not\in V_{\rm bad}\cup X^{\prime}$.
-Then $u_{s}$ (resp.
-$v_{1}$) lies in $Y_{i}$ (resp.
-$Z_{i}$) for at least $(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
-Since $|\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, we obtain that $u_{s}$ (resp.
-$v_{1}$) lies in $Y_{i}$ (resp.
-$Z_{i}$) for at least $\sqrt{\delta}n$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$.
-Hence there are two distinct colors $c_{1},c_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$ such that $u_{s}\in Y_{c_{1}}$ and $v_{1}\in Z_{c_{2}}$.
-It is routine to check that
-$|N_{{c_{1}}}^{+}(u_{s})\cap(Z\setminus(V_{\rm bad}\cup X^{\prime}))|\geq$ $|N_{{c_{1}}}^{+}(u_{s})\cap Z_{c_{1}}|-|Z_{1}\Delta Z_{c_{1}}|-|X\cup X_{A}\cup X_{B}|$
+[…]
+$|N_{{c_{1}}}^{+}(u_{s})\cap(Z\setminus(V_{\rm bad}\cup X^{\prime}))|\geq$ $|N_{{c_{1}}}^{+}(u_{s})\cap Z_{c_{1}}|-|Z_{1}\triangle Z_{c_{1}}|-|X\cup X_{A}\cup X_{B}|$
 $\geq$ $\left(\frac{1}{2}-3\sqrt{\delta}\right)n-\left(2\delta+\sqrt{\epsilon}+\frac{1}{2}\sqrt{\delta}\right)n\geq\left(\frac{1}{2}-4\sqrt{\delta}\right)n.$
@@
 ∎
+To prove Theorem 5.2, we need the following lemma.
+This establishes that every digraph collection $\mathcal{D}$ contains a transversal directed Hamilton cycle if it admits an almost balanced partition, with the small parts inducing an empty graph for almost all colors.
+We first state a preliminary result from [11].
+Lemma 6.4 ([11]).
+Let $\mathcal{C}$ be a set of colors, and $\mathcal{G}=\{G_{i}[Y,B]:i\in\mathcal{C}\}$ be a collection of bipartite graphs with the common bipartition $Y\cup B$ such that $7|Y|<|B|\leq\frac{3}{5}|\mathcal{C}|$.
+If $\sum_{i\in\mathcal{C}}|E(D_{i}[Y,B])|\geq t|B||\mathcal{C}|$ for some integer $t$ with $1\leq t\leq|Y|$, then $\mathcal{G}$ contains $t$ disjoint rainbow stars on $5$ vertices, and each of them has its center in $Y$ and other vertices in $B$.
+Lemma 6.5.
+Assume $0<\frac{1}{n}\ll\delta\ll 1$ and $0\leq\gamma\leq 3{\delta}$.
+Let $\mathcal{C}$ be a set of $n$ colors, and let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs with the common vertex set $V$ of size $n$ such that $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$.
+Let $A\cup B$ be a partition of $V$ with $|A|=\left\lceil\frac{n}{2}\right\rceil+\gamma n$, and let $\mathcal{C}^{\prime}\cup\mathcal{C}^{\prime\prime}$ be a partition of $\mathcal{C}$ with $|\mathcal{C}^{\prime\prime}|\leq\delta n$.
+Assume that $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$.
+Then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
+Proof.
+Define $Y_{1}:=\{v\in A:d_{i}^{+}(v,B)\leq(1-\delta^{\frac{1}{4}})|B|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$, $Y_{2}:=\{v\in A:d_{i}^{-}(v,B)\leq(1-\delta^{\frac{1}{4}})|B|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$.
+Since $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$, it is routine to check that
+$\left\lceil\frac{n}{2}\right\rceil|B||\mathcal{C}^{\prime}|$ $\leq\sum_{i\in\mathcal{C}^{\prime}}|E({D_{i}}[A,B])|$
+$\leq|Y_{1}|(1-\delta^{\frac{1}{4}})|B|\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|+|Y_{1}||B|(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|+(|A|-|Y_{1}|)|B||\mathcal{C}^{\prime}|$
+$=(|A|-\delta^{\frac{1}{2}}|Y_{1}|)|B||\mathcal{C}^{\prime}|.$
+It implies that $|Y_{1}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$.
+Similarly, $|Y_{2}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$.
+For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp.
+$v\in Y_{2}\setminus Y_{1}$), we have $d_{i}^{-}(v,B)>(1-\delta^{\frac{1}{4}})|B|$ (resp.
+$d_{i}^{+}(v,B)>(1-\delta^{\frac{1}{4}})|B|$) for at least $(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|$ colors $i\in\mathcal{C}^{\prime}$.
+Notice that Lemma 6.2 and Lemma 6.3 (i)-(ii) hold by setting $V_{\rm bad}:=Y_{1}\cup Y_{2}$ and $\mathcal{C}_{2}=\mathcal{C}^{\prime}$.
+Since $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$, we have $|E(D_{i}[A,B])|,|E(D_{i}[B,A])|\geq|B|\left\lceil\frac{n}{2}\right\rceil$ for each $i\in\mathcal{C}^{\prime}$, i.e., there are at most $|B|(\left\lceil\frac{n}{2}\right\rceil+\gamma n)-|B|\left\lceil\frac{n}{2}\right\rceil=|B|\gamma n$ non-edges from $A$ to $B$ (resp., from $B$ to $A$) in $D_{i}$.
+Hence
+$\sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[Y_{1},B])|\geq(|Y_{1}|-\gamma n)|B||\mathcal{C}^{\prime}|\ \ \text{and}\ \ \sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[B,Y_{2}])|\geq(|Y_{2}|-\gamma n)|B||\mathcal{C}^{\prime}|.$
+According to the sizes of $|Y_{1}|$ and $|Y_{2}|$, we proceed by considering the following three cases.
+Case 1. $|Y_{1}|,|Y_{2}|\geq\gamma n$.
+In view of Lemma 6.4, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ copies inside $\{D_{i}[Y_{1},B]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ copies inside $\{D_{i}[B,Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$.
+Recall that each vertex in $Y_{1}$ (resp.
+$Y_{2}$) has at least $\left\lceil\frac{n}{2}\right\rceil-(1-\delta^{\frac{1}{4}})|B|>4|Y|$ out-neighbors (resp.
+in-neighbors) in $A$ for at least $\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|>\delta^{\frac{1}{4}}(1-\delta)n>4|Y|$ digraphs $D_{i}$ with $i\in\mathcal{C}^{\prime}$.
+Therefore, by using colors in $\mathcal{C}^{\prime}$,
+  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by disjoint rainbow directed $P_{3}$ copies with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $B$,
+  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies starting at $B$ and ending at $A$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$,
+  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies starting at $A$ and ending at $B$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$,
+  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $A$.
+For each $y\in Y_{1}\cup Y_{2}$, let $P_{y}:=y^{1}yy^{2}$ denote the rainbow directed $P_{3}$ with center $y$, where $P_{y}$ has colors $c_{y}^{1}$ and $c_{y}^{2}$.
+Let $\mathbf{P}=\{P_{y}:y\in Y_{1}\cup Y_{2}\}$.
+By using Lemma 6.3, one may connect all of the above rainbow directed $P_{3}$ copies into a single rainbow directed path $P^{1}$ that starts at $A$ and ends at $B$.
+Then
+$|V(P^{1})\cap A|-|V(P^{1})\cap B|=|Y_{1}^{\prime}\setminus Y_{2}^{\prime}|+|Y_{2}^{\prime}\setminus Y_{1}^{\prime}|+2|Y_{1}^{\prime}\cap Y_{2}^{\prime}|=|Y_{1}^{\prime}|+|Y_{2}^{\prime}|=2\gamma n.$
+Hence, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|+\sigma$, where $\sigma=0$ if $n$ is odd and $\sigma=1$ otherwise.
+It is routine to check that there exists a rainbow matching inside $\{D_{i}^{\pm}[A\setminus V(P^{1}),B\setminus V(P^{1})]:i\in\mathcal{C}^{\prime\prime}\}$, say $M$, such that $D_{j}[A\setminus V(P^{1}\cup M),B\setminus V(P^{1}\cup M)]\neq\emptyset$ for each $j\in{\rm col}(M)$.
+Furthermore, $D_{j}$ is $(19\sqrt{\delta},{\rm EC}1)$-extremal for all $j\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$.
+Subcase 1.1. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have the same parity.
+We greedily choose two disjoint rainbow directed paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\}$ with lengths $\lceil\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rceil$ and $\lfloor\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rfloor$ respectively, such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$.
+Applying Lemma 6.3 (i), we successively connect $P^{1}$, all rainbow edges in $M$, $P_{A}$ and $P_{B}$ into a single rainbow path $P^{2}$ whose endpoints lie in different parts.
+Clearly, $|A\setminus V(P^{2})|=|B\setminus V(P^{2})|$.
+Together with Lemma 6.2, we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
+Subcase 1.2. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have different parity.
+Notice that $|B|=\lfloor\frac{n}{2}\rfloor-\gamma n$.
+Hence for each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $B$ has at least $\gamma n+1$ in-neighbors in $A$.
+Therefore, each vertex in $B$ has at least one in-neighbor outside $Y_{1}^{\prime}$.
+We first assume $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|\geq 1$.
+Choose an unused vertex $y\in B$ and a color $c\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$.
+Assume that ${xy}\in E(D_{c}[A,B])$ with $x\notin Y_{2}^{\prime}$.
+If $x\in A\setminus(Y_{1}\cup Y_{2})$, then $x$ can be chosen outside $V(P^{1}\cup M)$.
+Using Lemma 6.3 (i)-(ii), we connect $xy$ and $P^{1}$ into a single rainbow directed path $P$.
+If $x\in(Y_{1}\cup Y_{2})\setminus Y_{1}^{\prime}$, then we replace the end vertex of $P_{x}$ by $y$.
+In either situation the problem reduces to Subcase 1.1, yielding a transversal directed Hamilton cycle in $\mathcal{D}$, as desired.
+Now, we consider that $n$ is odd and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|=0$.
+For each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $A$ has at least one in-neighbor outside $Y_{1}^{\prime}$.
+Choose an unused vertex $y\in A\setminus(Y_{1}\cup Y_{2})$ and a color $c\in\mathcal{C}$.
+Then there exists a vertex $x\in A\setminus Y_{1}^{\prime}$ such that ${xy}\in E(D_{c}[A])$.
+By a similar discussion as above, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
+Case 2. At least one of $Y_{1}$ and $Y_{2}$ has size less than $\gamma n$.
+In this case, we consider only $|Y_{1}|\geq\gamma n$ and $|Y_{2}|<\gamma n$; the remaining two cases are analogous.
+Following the argument of Case 1, we know that all vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed copies of $P_{3}$ with centers in $Y_{1}$ and endpoints in $A\setminus(Y_{1}\cup Y_{2})$.
+Furthermore, each vertex in $Y_{2}\setminus Y_{1}$ can be covered by a rainbow directed $P_{3}$ with its center in $Y_{2}\setminus Y_{1}$, starting from $A\setminus(Y_{1}\cup Y_{2})$ and using colors in $\mathcal{C}^{\prime}$.
+(In this process, for the choice of the other endpoint of each rainbow directed $P_{3}$, we prefer using vertices in $A$ over vertices in $B$.) By using unused colors in $\mathcal{C}$ and unused vertices in $A\setminus Y$, one may extend those rainbow directed $P_{3}$ as long as possible or choose all other disjoint rainbow directed paths.
+Let $\mathbf{P}=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow directed paths in the above, where each $Q_{i}$ has at least one endpoint in $A\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $A$.
+We next consider the following two possible cases.
+Subcase 2.1. $|B|\geq|A|-(s_{1}+\cdots+s_{t})-\sigma$.
+In this subcase, there exists a set $\mathbf{P}^{\prime}=\{Q_{1}^{\prime},Q_{2}^{\prime},\ldots,Q_{\ell}^{\prime}\}$ such that $|B|=|A|-(s_{1}^{\prime}+\cdots+s_{\ell}^{\prime})-\sigma$, the endpoints of each $Q_{i}^{\prime}$ lie outside $Y$ and $Y_{2}\cup Y_{1}^{\prime}\subseteq V(Q_{1}^{\prime}\cup\ldots\cup Q_{\ell}^{\prime})$, where $s_{i}^{\prime}=|E(Q_{i}^{\prime}[A])|$ for each $i\in[\ell]$.
+By Lemma 6.3 (i)-(ii), we connect all rainbow directed paths in $\mathbf{P}^{\prime}$ into a single rainbow directed path $P^{1}$, whose endpoints are in different parts.
+Since $|A|-|B|=2\gamma n+\sigma$, we have $\ell\leq s_{1}^{\prime}+\cdots+s_{\ell}^{\prime}=2\gamma n$.
+Therefore, $|E(P^{1})|\leq 8\gamma n$ and $|A\setminus V(P^{1})|-\sigma=|B\setminus V(P^{1})|$.
+Recall that all vertices in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ can also be covered by a set of disjoint rainbow directed copies of $P_{3}$ with centers in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ and both endpoints in $B$.
+Applying Lemma 6.3 (i) again to connect those rainbow directed paths with $P^{1}$, we get a rainbow directed path $P^{2}$ with length at most $8\gamma n+12\sqrt{\delta}n$, and with endpoints in different parts.
+Choose a maximum rainbow matching inside $\{D_{i}^{\pm}[A\setminus V(P^{2}),B\setminus V(P^{2})]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{2})\}$, say $M$, such that $D_{j}^{\pm}[A\setminus V(P^{2}\cup M),B\setminus V(P^{2}\cup M)]\neq\emptyset$ for each $j\in{\rm col}(M)$.
+Clearly, $D_{j}$ is $(13\sqrt{\delta},{\rm EC}1)$-extremal for all $j\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)$.
+In fact, when constructing $P^{1}-B-Y$, we may prioritize the use of colors from $\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)$ over other colors.
+Hence either ${\rm col}(P^{1}-B-Y)\subseteq\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$ or $\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\subseteq{\rm col}(P^{1}-B-Y)$.
+Based on Lemma 6.3 (i), one may connect $P^{1}$ and all edges of $M$ into a single rainbow directed path $P^{2}$, whose endpoints are in different parts and length is at most $8\gamma n+4{\delta}n$.
+Next, we are to choose two disjoint rainbow paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)\}$ such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$ respectively, whose lengths are determined by the parity of $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$.
+Clearly, if $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have the same parity, then by Lemma 6.2, we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
+Hence $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have different parity.
+By an argument analogous to that in the proof of Step 4 in Case 1, we can obtain the desired transversal Hamilton cycle in $\mathcal{D}$.
+Subcase 2.2. $|B|<|A|-(s_{1}+\cdots+s_{t})-\sigma$.
+In this case, $\sum_{i=1}^{t}s_{i}<2\gamma n$.
+Let $w$ be an arbitrary vertex in $A\setminus V(\mathbf{P})$ and $c_{1},c_{2}$ be two colors in $\mathcal{C}\setminus{\rm col}(\mathbf{P})$.
+Clearly, in $D_{c_{1}}$ and $D_{c_{2}}$, $w$ has at least $\left\lceil\frac{n}{2}\right\rceil-(\lfloor\frac{n}{2}\rfloor-\gamma n)=\gamma n+\sigma$ out-neighbors (resp.
+in-neighbors) in $V(\mathbf{P})$ and it cannot be adjacent to the start vertex of $Q_{i}[A]$ (resp.
+from the end vertex of $Q_{i}[A]$) for all $i\in[t]$ (by the maximality of $\mathbf{P}$).
+For convenience, assume that $s_{0}=0$ and $V(Q_{i})\cap A:=\{v_{s_{1}+\cdots+s_{i-1}+i},\ldots,v_{s_{1}+\cdots+s_{i}+i}\}$ for each $i\in[t]$.
+Define
+$I_{1}:=\left\{i:v_{i}\in N_{{c_{1}}}^{-}(w)\right\}\ \ \text{and}\ \ I_{2}:=\left\{i:v_{i+1}\in N_{{c_{2}}}^{+}(w)\right\}.$
+Hence $I_{1},I_{2}\subseteq\bigcup_{i\in[t]}\left[\sum_{j\in[i-1]}s_{j}+i,\sum_{j\in[i]}s_{j}+i-1\right]$.
+Therefore, $I_{1}\cap I_{2}\neq\emptyset$, which contradicts the maximality of $\mathbf{P}$.
+∎
+Now, we are ready to give the proof of Theorem 5.2.
+Proof of Theorem 5.2.
+Choose constant $\epsilon,\delta,\eta$ such that
+$\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll\eta\ll 1,$
+let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$.
+Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles.
+By Lemma 5.1, for each $i\in[m]$, $D_{i}$ has a characteristic partition $({A_{i}},{B_{i}},{L_{i}})$ satisfying (C1)-(C4).
+Moreover, for every $i\in[m]$ and $Y\in\{A,B\}$, we have
+$|L_{i}|\leq 2\epsilon n\ \text{and}\ |{Y}_{1}{\triangle}Y_{i}|<2\delta n.$ (2)
+Expand $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$.
+For simplicity, we assume $Y\in\{A,B\}$ and $Z\in\{A,B\}\setminus\{Y\}$ in the following.
+Let $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ be a partition of $[m]$ with
+$\mathcal{C}_{1}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC}1){\text{-extremal}}\},$
+$\mathcal{C}_{2}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC}2){\text{-extremal}}\}.$
+Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise.
+Clearly, $|\hat{\mathcal{C}}|\geq(1-4\sqrt{\delta}-\eta)n$.
+Define
+$X:=$ $\{x\in V:x\not\in A_{i}\cup B_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in\hat{\mathcal{C}}\}$,
+$X_{Y}:=$ $\{x\in Y\setminus X:x\not\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 10\sqrt{\delta}|\hat{\mathcal{C}}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in\hat{\mathcal{C}}\}$.
+[…]
+[…]
+It follows that $|X|<\frac{1}{3}\sqrt{\epsilon}n$ and $|X_{Y}|<\frac{1}{4}\sqrt{\delta}n$. For $k\in[2]$, define
+$X_{Y}^{k}:=\left\{\begin{array}[]{ll}\{x\in X_{Y}:x\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 3\sqrt{\delta}n\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}_{k}}\},&\textrm{if $|\mathcal{C}_{k}|\geq\eta n$,}\\[5.0pt] \emptyset,&\textrm{otherwise.}\end{array}\right$.
+[…]
+[…]
+It follows that $|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|<4\delta n$.
+Then we move vertices in $X_{A}\setminus(\bigcup_{k\in[2]}X_{A}^{k})$ to $B$ and vertices in $X_{B}\setminus(\bigcup_{k\in[2]}X_{B}^{k})$ to $A$.
+[…]
 Next, based on the sizes of $|\mathcal{C}_{1}|$ and $|\mathcal{C}_{2}|$, we distinguish our proof into the following three cases.
@@
 In this case, each vertex in $Y\setminus(X\cup X_{Y}^{2})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{2}|$ colors $i\in\mathcal{C}_{2}$.
-Recall that $|B|-|A|=r$ with $0\leq r\leq(8{\delta}+\sqrt{\epsilon})n$.
+Recall that |B||A|=r|B|-|A|=r with $0\leq r\leq(8{\delta}+\sqrt{\epsilon})n$.
 The proof is divided into four steps.
@@
 Assume that $\{Q_{1},\ldots,Q_{t}\}$ is a set of disjoint maximal rainbow directed paths in $\mathcal{D}[B]$.
-If $|E(Q_{1})|+\cdots+|E(Q_{t})|\geq|B|-|A|-\sigma$, then there must exist a set of disjoint rainbow directed paths, say $\{Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}\}$, such that $|E(Q_{1}^{\prime})|+\cdots+|E(Q_{t^{\prime}}^{\prime})|=|B|-|A|-\sigma$.
+On the one hand, suppose that $|E(Q_{1})|+\cdots+|E(Q_{t})|\geq|B|-|A|-\sigma$.
+Then there must exist a set of disjoint rainbow directed paths, say $\{Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}\}$, such that $|E(Q_{1}^{\prime})|+\cdots+|E(Q_{t^{\prime}}^{\prime})|=|B|-|A|-\sigma$.
 That is, $|A|-t^{\prime}=|B|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-\sigma$.
-Using Claim 5.3 (ii), we connect $Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}$ into a single rainbow directed path $P^{1}$ with length at most $27\delta n$, starting at $A$ and ending at $B$.
-(In fact, by using Claim 5.3 (ii), we only get a rainbow path with two endpoints in $B$.
+Using Lemma 6.3 (ii), we connect $Q_{1}^{\prime},\ldots,Q_{t^{\prime}}^{\prime}$ into a single rainbow directed path $P^{1}$ with length at most $27\delta n$, starting at $A$ and ending at $B$.
+(In fact, by using Lemma 6.3 (ii), we only get a rainbow path with two endpoints in $B$.
 But we usually want to find a rainbow directed path with endpoints in different parts.
@@
 Therefore, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|-\sigma$, as desired.
-In what follows, it suffices to consider $|E(Q_{1})|+\cdots+|E(Q_{t})|<|B|-|A|-\sigma=r-2s-\sigma$.
-Now, we move all vertices of $Q_{1},\ldots,Q_{t}$ to $A$.
+On the other hand, suppose that $|E(Q_{1})|+\cdots+|E(Q_{t})|<|B|-|A|-\sigma=r-2s-\sigma$.
+In this case, we move all vertices of $Q_{1},\ldots,Q_{t}$ to $A$.
 Let $\mathcal{C}^{\prime}:=\mathcal{C}\setminus{\rm col}(\bigcup_{i\in[t]}Q_{i})$.
@@
 It follows from $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$ that $\left\lceil\frac{n}{2}\right\rceil\leq|A|+|X^{\prime}|\leq\frac{n}{2}+19{\delta}n$.
-Now, by Lemma 5.14 (which is presented in the final of this subsection), we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
+Now, by Lemma 6.5, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
 In this step, if we only move vertices in $V_{\rm bad}$ from $B$ to $A$, then let $P^{1}$ be a null digraph (i.e., no vertices); otherwise, assume that $P^{1}$ is a rainbow directed path obtained in the above, whose length is at most $27\delta n$ and endpoints are $u_{1}\in A\setminus V_{\rm bad},\,v_{1}\in B\setminus V_{\rm bad}$.
 Hence $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|-\sigma$.
-Step 2. Construct a series of disjoint rainbow directed rainbow $P_{3}$ such that all centers of them are exactly all vertices in $V_{\rm bad}\cup X^{\prime}$.
+Step 2. Construct a sequence of disjoint rainbow directed $P_{3}$ copies such that all centers of them are exactly all vertices in $V_{\rm bad}\cup X^{\prime}$.
 Recall that $|X_{A}^{2}\cup X_{B}^{2}|\leq|X_{A}\cup X_{B}|<\frac{1}{2}\sqrt{\delta}n$.
@@
 If $v\in Y$, then $v$ belongs to $Y_{i}$ for at least $(3\sqrt{\delta}-27\delta)n>4|V_{\rm bad}\cup X^{\prime}|$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{1})$.
-Hence, using colors in $\mathcal{C}_{2}\setminus{\rm col}(P^{1})$, we can greedily choose $|X_{A}^{2}\cup X_{B}^{2}|$ disjoint rainbow directed $P_{3}$, such that their centers correspond to all vertices in $X_{A}^{2}\cup X_{B}^{2}$.
+Hence, using colors in $\mathcal{C}_{2}\setminus{\rm col}(P^{1})$, we can greedily choose $|X_{A}^{2}\cup X_{B}^{2}|$ disjoint rainbow directed $P_{3}$ copies, such that their centers correspond to all vertices in $X_{A}^{2}\cup X_{B}^{2}$.
 Moreover, if $v\in(X_{A}^{2}\cup X_{B}^{2})\cap Y$, then the rainbow directed $P_{3}$ with center $v$ has endpoints in $Z\setminus(V_{\rm bad}\cup X^{\prime}\cup V(P^{1}))$.
-For vertices in $X^{\prime}$, we know from its definition that by avoiding vertices and colors in $P^{1}$ and all rainbow directed paths chosen for vertices in $X_{A}^{2}\cup X_{B}^{2}$, there exists a set of disjoint rainbow directed $P_{3}$ with centers in $X^{\prime}$ and endpoints in different parts.
-By using Claim 5.3 (i)-(ii), we connect $P^{1}$, all rainbow directed $P_{3}$ with centers in $X_{A}^{2}\cup X_{B}^{2}\cup X^{\prime}$ (as obtained above) into a single rainbow directed path.
+For vertices in $X^{\prime}$, we know from its definition that by avoiding vertices and colors in $P^{1}$ and all rainbow directed paths chosen for vertices in $X_{A}^{2}\cup X_{B}^{2}$, there exists a set of disjoint rainbow directed $P_{3}$ copies with centers in $X^{\prime}$ and endpoints in different parts.
+By using Lemma 6.3 (i)-(ii), we connect $P^{1}$, all rainbow directed $P_{3}$ copies with centers in $X_{A}^{2}\cup X_{B}^{2}\cup X^{\prime}$ (as obtained above) into a single rainbow directed path.
 This results in a rainbow directed path $P^{2}=u_{1}\ldots v_{2}$ with $v_{2}\in B\setminus V_{\rm bad}$ and length at most
@@
 Notice that the position (in $A$ or $B$) of vertices in $X\setminus X^{\prime}$ does not affect our proof, so if $v\in X$ is a vertex moved from $B$ to $A$, then we only need consider $v\in A$.
-Claim 5.4.
-Assume $X\setminus V(P^{2})=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$.
-If $x_{i}\in X\cap Y$, then there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{2}\setminus({\rm col}(P^{2})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},x_{i}^{4}\in Z\setminus(V(P^{2})\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$.
-Proof of Claim 5.4.
-Suppose there exists an $i_{0}\in[s]$ such that Claim 5.4 holds for all $i\in[i_{0}-1]$ but does not hold for $i_{0}$.
-Assume for convenience that $x_{i_{0}}\in X\cap A$.
-Denote $\mathcal{C}_{2}^{\prime}:=\mathcal{C}_{2}\setminus({\rm col}(P^{2})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i_{0}-1]\})$.
-Let
-$\mathcal{C}_{2}^{1}:=\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{+}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P^{2})|\}\ \text{and}$
-$\mathcal{C}_{2}^{2}:=\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{-}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P^{2})|\}.$
-Hence either $|\mathcal{C}_{2}^{1}|\leq 3$ or $|\mathcal{C}_{2}^{2}|\leq 3$.
-Without loss of generality, assume that $|\mathcal{C}_{2}^{1}|\leq 3$.
-Thus, for each $j\in\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}$, we have
-$d_{j}^{+}(x_{i_{0}},A_{j})$ $\geq d_{j}^{+}(x_{i_{0}},A)-|A\setminus A_{1}|-|A_{1}\Delta A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
-$\geq\delta^{0}(D_{j})-d_{j}^{+}(x_{i_{0}},B)-|A\setminus A_{1}|-|A_{1}\Delta A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
-$\geq\frac{n}{2}-\left(4(i_{0}-1)+4+|V(P^{2})|\right)-2{\epsilon}n-2\delta n-4\delta n-\sqrt{\epsilon}n$
-$\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n.$
-Furthermore, $|\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}|\geq(1-3\sqrt{\delta})|\mathcal{C}_{2}|$.
-Notice that $x_{i_{0}}\notin X^{\prime}$.
-Then, for at least $(1-3\sqrt{\delta})|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, $d_{j}^{-}(x_{i_{0}},B_{j})<\frac{5}{2}\sqrt{\delta}n$, and consequently, $d_{j}^{-}(x_{i_{0}},A_{j})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for the same set of colors.
-By (C3), we know that $x_{i_{0}}\in A_{j}\cup B_{j}$ for all but at most $6\sqrt{\delta}|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, which leads to $x_{i_{0}}\not\in X$, a contradiction.
-∎
-By Claim 5.4, for each $x_{i}\in(X\setminus X^{\prime})\cap Y$, there exists a rainbow directed path $Q_{x_{i}}=x_{i}^{1}x_{i}x_{i}^{2}$ with ${\rm col}(Q_{x_{i}})=\{c_{i}^{1},c_{i}^{2}\}\subseteq\mathcal{C}_{2}$ and $x_{i}^{1},x_{i}^{2}\in Z\setminus V_{\rm bad}$.
+By Lemma 6.1, for each $x_{i}\in(X\setminus X^{\prime})\cap Y$, there exists a rainbow directed path $Q_{x_{i}}=x_{i}^{1}x_{i}x_{i}^{2}$ with ${\rm col}(Q_{x_{i}})=\{c_{i}^{1},c_{i}^{2}\}\subseteq\mathcal{C}_{2}$ and $x_{i}^{1},x_{i}^{2}\in Z\setminus V_{\rm bad}$.
 Moreover, each color or endpoint in $Q_{x_{i}}$ can be replaced by an unused color or vertex (if needed).
-Using Claim 5.3 (i)-(ii), one may connect $P^{2}$ and all rainbow directed paths in $\{Q_{x_{i}}:x_{i}\in X\setminus X^{\prime}\}$ into a single rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{3}\in B\setminus V_{\rm bad}$, whose length is at most $1027\delta n+2\sqrt{\delta}n+5\sqrt{\epsilon}n+$ 10 .
+Using Lemma 6.3 (i)-(ii), one may connect $P^{2}$ and all rainbow directed paths in $\{Q_{x_{i}}:x_{i}\in X\setminus X^{\prime}\}$ into a single rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{3}\in B\setminus V_{\rm bad}$, whose length is at most $27\delta n+2\sqrt{\delta}n+5\sqrt{\epsilon}n+10.$
 Step 3. Select a rainbow matching with colors in $\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{3})$.
-Choose a maximum rainbow matching in $\{D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]:i\in\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{3})\}$, say $\tilde{M}$.
-If there exists an edge with color $j$ in $\tilde{M}$ such that $D_{j}^{\pm}[A\setminus V(P^{3}\cup\tilde{M}),B\setminus V(P^{3}\cup\tilde{M})]$ contains no $2$-matching, then delete all such edges from $\tilde{M}$ and denote the resulting rainbow directed matching by $M$.
-Connect $P^{3}$ and all rainbow directed edges in $M$ by Claim 5.3 (i), we obtain a rainbow directed path $P^{4}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $v_{4}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{3})|+16\sqrt{\delta}n\leq 18\sqrt{\delta}n+5\sqrt{\epsilon}n+27\delta n+10$.
+[…]
 […]
@@
 By the construction of $P^{4}$, we know that if $w\in V(P^{4})$ is a vertex not in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X$, then it is possible to avoid $w$ when constructing $P^{4}$; if $i\in{\rm col}(P^{4})$ is a color not in ${\rm col}(P^{1}-A)\cup{\rm col}(M)$, then it is possible to avoid $i$ when constructing $P^{4}$.
-That is, we can assume that $w\not\in V(P^{4})$ unless $w\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i\not\in{\rm col}(P^{4})$ unless $i\in{\rm col}(P^{1}-A)\cup{\rm col}(M)$.
-Furthermore, in the process of choosing rainbow directed paths inside $\mathcal{D}[B]$ (in Step 1), one may use colors from $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$ before other colors and avoid vertices in $X^{\prime}\cup V_{\rm bad}\cup V(M)$.
+In other words, we may assume that $w\not\in V(P^{4})$ unless $w\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i\not\in{\rm col}(P^{4})$ unless $i\in{\rm col}(P^{1}-A)\cup{\rm col}(M)$.
+Furthermore, while selecting rainbow directed paths inside $\mathcal{D}[B]$ (in Step 1), one may give priority to using colors from $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$ and avoid vertices in $X^{\prime}\cup V_{\rm bad}\cup V(M)$.
 This implies that either $\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}\subseteq{\rm col}(P^{1})$ or ${\rm col}(P^{1}-A)\subseteq\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime}$.
 Denote $\tilde{\mathcal{C}}:=(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P^{1})$.
-Notice that for each color $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{4})$, every vertex $v\in A\setminus V(P^{4})$ has at least $(\frac{1}{2}-3\sqrt{\delta})n-|V(P^{4})|-|B\Delta B_{i}|$ out-neighbors (resp.
+Notice that for each color $i\in\mathcal{C}_{2}\setminus{\rm col}(P^{4})$, every vertex $v\in A\setminus V(P^{4})$ has at least $(\frac{1}{2}-3\sqrt{\delta})n-|V(P^{4})|-|B\triangle B_{i}|$ out-neighbors (resp.
 in-neighbors) in $B\setminus V(P^{4})$ in $D_{i}$.
@@
 Therefore, by replacing every pair of arcs with the same endpoints and opposite directions by an undirected edge, the digraph $D_{i}^{\pm}[A\setminus V(P^{4}),B\setminus V(P^{4})]$ can be viewed as an undirected graph $G_{i}$ with vertex set $(A\cup B)\setminus V(P^{4})$ and minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$.
-Next, we give an application of the transversal blow-up lemma (see [9]) for embedding transversal undirected Hamilton paths inside very dense bipartite graph collections, which can be proved by minor modifications to the proof of [10, Lemma 6.1], and we omit the proof here.
-Claim 5.5.
-Assume $W,Z\in\{A,B\}$.
-Let $W^{*}\subseteq W\setminus V_{\rm bad}$ and $Z^{*}\subseteq Z\setminus V_{\rm bad}$, where $|W^{*}|,|Z^{*}|\geq\eta n$, $W^{*}\cap Z^{*}=\emptyset$ and $|W^{*}|-|Z^{*}|=t\in\{0,1\}$.
-Let $T^{*}=Z^{*}$ if $t=0$ and $T^{*}=W^{*}$ if $t=1$.
-Let $\mathcal{C^{*}}\subseteq\mathcal{C}$ satisfy $|\mathcal{C^{*}}|=|W^{*}|+|Z^{*}|-1$, where $\mathcal{C^{*}}\subseteq\mathcal{C}_{1}$ if $W=Z$ and $\mathcal{C^{*}}\subseteq\mathcal{C}_{2}$ if $W\neq Z$.
-Let $W^{-}\subseteq W^{*}$ and $T^{+}\subseteq T^{*}$ with $|W^{-}|,|T^{+}|\geq\frac{\eta n}{8}$.
-Then there is a transversal undirected Hamilton path in $\{G_{i}[W^{*},Z^{*}]:i\in\mathcal{C}^{*}\}$ starting at $W^{-}$ and ending at $T^{+}$.
+The next claim employs Lemma 6.2 to construct a transversal directed Hamilton cycle.
 Claim 6.6.
 Let $Q=x_{1}\ldots x_{s}$ be a rainbow directed path inside $\mathcal{D}$ with $s\leq 19\sqrt{\delta}n$, $V_{\rm bad}\subseteq V(Q)$, $(\mathcal{C}_{1}\cup\mathcal{C}_{\rm bad})\setminus\tilde{\mathcal{C}}\subseteq{\rm col}(Q)$, $x_{1}\in A\setminus V_{\rm bad}$ and $x_{s}\in B\setminus V_{\rm bad}$.
-If $|\tilde{\mathcal{C}}|-||B\setminus V(Q)|-|A\setminus V(Q)||$ is a nonnegative even integer, then $\mathcal{D}$ has a transversal directed Hamilton cycle.
+If $|\tilde{\mathcal{C}}|-\big||B\setminus V(Q)|-|A\setminus V(Q)|\big|$ is a nonnegative even integer, then $\mathcal{D}$ has a transversal directed Hamilton cycle.
 Proof of Claim 6.6.
-Using colors in $\tilde{\mathcal{C}}$, one may greedily choose two disjoint rainbow directed paths $P^{5}$ and $P^{6}$ in $\mathcal{D}[A\setminus V(Q)]$ and $\mathcal{D}[B\setminus V(Q)]$ with lengths $\frac{|\tilde{\mathcal{C}}|-(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$ and $\frac{|\tilde{\mathcal{C}}|+(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$ respectively, whose endpoints are $u_{5},v_{5}\in A\setminus V_{\rm bad}$ and $u_{6},v_{6}\in B\setminus V_{\rm bad}$.
-(Note that if a rainbow path has length 00, then it contains one vertex.) Applying Claim 5.3 (i)-(ii) to connect $Q,P^{5}$ and $P^{6}$ in turn, we get a rainbow directed path $P$ with endpoints $x_{1}\in A\setminus V_{\rm bad}$ and $v_{6}\in B\setminus V_{\rm bad}$.
+Using colors in $\tilde{\mathcal{C}}$, one may greedily choose two disjoint rainbow directed paths $P^{5}$ and $P^{6}$ in $\mathcal{D}[A\setminus V(Q)]$ and $\mathcal{D}[B\setminus V(Q)]$ with lengths
+$\frac{|\tilde{\mathcal{C}}|-(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}\ \ \text{and}\ \ \frac{|\tilde{\mathcal{C}}|+(|B\setminus V(Q)|-|A\setminus V(Q)|)}{2}$
+respectively, whose endpoints are $u_{5},v_{5}\in A\setminus V_{\rm bad}$ and $u_{6},v_{6}\in B\setminus V_{\rm bad}$.
+(Note that if a rainbow path has length 0, then it contains one vertex.) Applying Lemma 6.3 (i)-(ii) to connect $Q,P^{5}$ and $P^{6}$ in turn, we get a rainbow directed path $P$ with endpoints $x_{1}\in A\setminus V_{\rm bad}$ and $v_{6}\in B\setminus V_{\rm bad}$.
 Clearly, $|V(P)|\leq 19\sqrt{\delta}n+4\sqrt{\delta}n+\eta n+4$ and $|A\setminus V(P)|=|B\setminus V(P)|$.
@@
 It is routine to check that $|W^{*}|=|T^{*}|$, $|\mathcal{C}^{*}|=|W^{*}|+|T^{*}|-1$,
-$|W^{-}|$ $\geq d_{G_{i_{2}}}(v_{6},A)-|V(P)|\geq d_{G_{i_{2}}}(v_{6},A_{i_{2}})-|A_{1}\Delta A_{i_{2}}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|-|V(P)|$
+$|W^{-}|$ $\geq d_{G_{i_{2}}}(v_{6},A)-|V(P)|\geq d_{G_{i_{2}}}(v_{6},A_{i_{2}})-|A_{1}\triangle A_{i_{2}}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|-|V(P)|$
 $\geq\left(\frac{1}{2}-45\sqrt{\delta}\right)n-2\delta n-4{\delta}n-\sqrt{\epsilon}n-|V(P)|>\left(\frac{1}{2}-2\eta\right)n,$
 and similarly $|T^{+}|>\left(\frac{1}{2}-2\eta\right)n$.
-Applying Claim 5.5 yields that $\{G_{i}[W^{*},T^{*}]:i\in\mathcal{C}^{*}\}$ contains a transversal undirected path $P^{\prime}$ starting at $v^{\prime}\in W^{-}$ and ending at $u^{\prime}\in T^{+}$.
+Applying Lemma 6.2 yields that $\{G_{i}[W^{*},T^{*}]:i\in\mathcal{C}^{*}\}$ contains a transversal undirected path $P^{\prime}$ starting at $v^{\prime}\in W^{-}$ and ending at $u^{\prime}\in T^{+}$.
 Thus, $x_{1}Pv_{6}v^{\prime}P^{\prime}u^{\prime}x_{1}$ is a transversal directed Hamilton cycle inside $\mathcal{D}$.
 ∎
-If $n$ and $|\tilde{\mathcal{C}}|$ have the same parity, then in view of Claim 5.6, there exists a transversal directed Hamilton cycle inside $\mathcal{D}$, a contradiction.
-Hence, it suffices to consider that $n$ and $|\tilde{\mathcal{C}}|$ have the different parity.
-Therefore, either $|\tilde{\mathcal{C}}|-(|B\setminus V(P^{4})|-|A\setminus V(P^{4})|)$ is a positive odd integer or $|\tilde{\mathcal{C}}|=0$ and $n$ is odd.
-Based on the definition of $X^{\prime}$, move vertices in $X^{\prime}$ to $Y$ if it is covered by a rainbow directed path from $Z$ to $Y$ (in Step 2).
+[…]
+Based on the definition of $X^{\prime}$, we move vertices in $X^{\prime}$ to $Y$ if it is covered by a rainbow directed path from $Z$ to $Y$ (in Step 2).
+The following claim considers the local structure of $D_{i}$.
 Claim 6.7.
@@
 Recall that each of $w_{1}$ and $w_{2}$ can be chosen outside $V(P^{4})$ unless it is in $(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$, and $i_{0}$ can be chosen outside ${\rm col}(P^{4})$ unless $i_{0}\in{\rm col}(M)$.
-If $i_{0}\in{\rm col}(M)$, then let $\tilde{P}^{4}$ be a rainbow directed path obtained by deleting the edge with color $i_{0}$ from $P^{4}$ and connecting its two endpoints by Claim 5.3 (i).
+If $i_{0}\in{\rm col}(M)$, then let $\tilde{P}^{4}$ be a rainbow directed path obtained by deleting the edge with color $i_{0}$ from $P^{4}$ and connecting its two endpoints by Lemma 6.3 (i).
 It is routine to check that $|B\setminus V(\tilde{P}^{4})|-|A\setminus V(\tilde{P}^{4})|=|B\setminus V({P}^{4})|-|A\setminus V({P}^{4})|$.
-Let $P^{4}:=\tilde{P}^{4}$ if $i_{0}\in{\rm col}(M)$.
-Therefore, it suffices to consider $i_{0}\not\in{\rm col}(P^{4})$.
-$\bullet$ $w_{1},w_{2}\not\in V_{\rm bad}$.
+Let $P^{4}:=\tilde{P}^{4}$ if $i_{0}\in{\rm col}(M)$, which reduces our proof to the case where $i_{0}\not\in{\rm col}(P^{4})$.
+We consider the following several cases.
+$\bullet$ Suppose $w_{1},w_{2}\not\in V_{\rm bad}$.
 Then $w_{1},w_{2}\not\in V(P^{4})$.
-Connect $w_{1}$ with the endpoint $v_{4}$ of $P^{4}$ by Claim 5.3 (ii).
+Connect $w_{1}$ with the endpoint $v_{4}$ of $P^{4}$ by Lemma 6.3 (ii).
 Denote the resulting rainbow directed path by $Q_{1}$.
-$\bullet$ $w_{1}\not\in V_{\rm bad}$ and $w_{2}\in V_{\rm bad}$.
+$\bullet$ Suppose $w_{1}\not\in V_{\rm bad}$ and $w_{2}\in V_{\rm bad}$.
 Hence $w_{1}\not\in V(P^{4})$ and there exists a rainbow subpath $u_{1}P^{4}w_{2}^{1}w_{2}^{2}w_{2}$ of $P^{4}$ with $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$, $w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$.
-Replace $w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$, and connect $w_{1}$ with the vertex in $N_{P^{4}}^{-}(w_{2}^{1})$ by Claim 5.3 (i).
+Replace $w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$, and connect $w_{1}$ with the vertex in $N_{P^{4}}^{-}(w_{2}^{1})$ by Lemma 6.3 (i).
 Let $Q_{2}$ be the resulting rainbow directed path.
 The case for $w_{1}\in V_{\rm bad}$ and $w_{2}\not\in V_{\rm bad}$ is similar.
-$\bullet$ $w_{1},w_{2}\in V_{\rm bad}$.
-By rearranging, one may assume that $w_{1}$ and $w_{2}$ is connected by a rainbow directed path $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ in $P^{4}$, where $w_{1}^{1},w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$ and $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$.
-Replace $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$.
-Denote the resulting rainbow directed path by $Q_{3}$.
-In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})||$ is a nonnegative even integer.
+$\bullet$ Suppose $w_{1},w_{2}\in V_{\rm bad}$.
+By rearranging, one may assume that $w_{1}$ and $w_{2}$ are connected by a rainbow directed path $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ in $P^{4}$, where $w_{1}^{1},w_{2}^{2}\in A\setminus(V_{\rm bad}\cup X^{\prime})$ and $w_{2}^{1}\in B\setminus(V_{\rm bad}\cup X^{\prime})$.
+Replace $w_{1}w_{1}^{1}w_{2}^{1}w_{2}^{2}w_{2}$ with $w_{1}w_{2}$ in $P^{4}$; and denote the resulting rainbow directed path by $Q_{3}$.
+In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-\big||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})|\big|$ is a nonnegative even integer.
 It follows from Claim 6.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
 Hence $z_{1}\notin V(P^{4})$.
-$\bullet$ $z_{2}\not\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$.
+Next, we consider several cases.
+$\bullet$ Suppose $z_{2}\not\in(V(P^{1})\cap B)\cup V_{\rm bad}\cup X^{\prime}$.
 Then $z_{2}\not\in V(P^{4})$.
-Connect $z_{2}$ with the endpoint $u_{1}$ of $P^{4}$ by Claim 5.3 (i).
-Let $Q_{1}$ be the resulting rainbow directed path, and $\tilde{\mathcal{C}}_{1}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
-$\bullet$ $z_{2}\in V(P^{1})\cap B$.
+Using Lemma 6.3 (i), connect $z_{2}$ to the endpoint $u_{1}$ of $P^{4}$.
+Denote the resulting rainbow directed path by $Q_{1}$, and set $\tilde{\mathcal{C}}_{1}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
+$\bullet$ Suppose $z_{2}\in V(P^{1})\cap B$.
 Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{\prime}z_{2}P^{4}v_{4}$, where $z_{2}^{\prime}z_{2}\in E(P^{1})\cap E(D_{i_{2}})$ for some $i_{2}\in\mathcal{C}$.
@@
 Hence $z_{2}^{\prime}\not\in V_{\rm bad}$.
-Connect $u_{1}P^{1}z_{2}^{\prime}$ with $z_{1}z_{2}P^{4}v_{4}$ into a single rainbow directed path by Claim 5.3 (i)-(ii).
-Denote the resulting rainbow directed path by $Q_{2}$.
+Apply Lemma 6.3 (i)-(ii) to connect $u_{1}P^{1}z_{2}^{\prime}$ and $z_{1}z_{2}P^{4}v_{4}$ into a single rainbow directed path, which we denote by $Q_{2}$.
 Let $\tilde{\mathcal{C}}_{2}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$ if $z_{2}^{\prime}\in A$, and $\tilde{\mathcal{C}}_{2}:=(\tilde{\mathcal{C}}\setminus\{i_{1}\})\cup\{i_{2}\}$ if $z_{2}^{\prime}\in B$.
-$\bullet$ $z_{2}\in V_{\rm bad}\cup X^{\prime}$.
+$\bullet$ Suppose $z_{2}\in V_{\rm bad}\cup X^{\prime}$.
 Then $P^{4}$ can be written as $u_{1}P^{4}z_{2}^{1}z_{2}P^{4}v_{4}$, where $z_{2}^{1}\in A\setminus V_{\rm bad}$.
-Connect $u_{1}P^{4}z_{2}^{1}$ with $z_{1}z_{2}z_{2}^{2}P^{4}v_{4}$ into a single rainbow directed path by Claim 5.3 (ii).
-Let $Q_{3}$ be the resulting rainbow directed path and $\tilde{\mathcal{C}}_{3}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
-In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})||$ is a nonnegative even integer.
+By Lemma 6.3 (ii), we connect $u_{1}P^{4}z_{2}^{1}$ with $z_{1}z_{2}z_{2}^{2}P^{4}v_{4}$ into a single rainbow directed path, say $Q_{3}$.
+Set $\tilde{\mathcal{C}}_{3}:=\tilde{\mathcal{C}}\setminus\{i_{1}\}$.
+In each of the above cases, we get a rainbow directed path $Q_{i}$ ($i\in[3]$) such that $|\tilde{\mathcal{C}}|-\big||B\setminus V(Q_{i})|-|A\setminus V(Q_{i})|\big|$ is a nonnegative even integer.
 It follows from Claim 6.6 that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
 Notice that $|B\setminus(X^{\prime}\cup V(P^{1}))|\geq\frac{n}{2}-28\delta n$ and $|{\rm col}(P^{1}-A)|\leq 27\delta n$.
-Together with Lemma 5.14 (see the final of this subsection), we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
-If $|\tilde{\mathcal{C}}|\geq 1$, then based on Claim 5.7 (ii) and the fact that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we obtain $|A|,|B|\geq\frac{n}{2}+1$, a contradiction.
+Together with Lemma 6.5 (see the end of this subsection), we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
+If $|\tilde{\mathcal{C}}|\geq 1$, then based on Claim 6.7 (ii) and the fact that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we obtain $|A|,|B|\geq\frac{n+1}{2}$, a contradiction.
 Case 2. $|\mathcal{C}_{2}|<\eta n$.
 In this case, each vertex in $Y\setminus(X\cup X_{Y}^{1})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{1}|$ colors $i\in\mathcal{C}_{1}$.
-By a similar discussion as in Case 1, we know that there are $|V_{\rm bad}|$ disjoint rainbow directed $P_{3}$ with centers in $V_{\rm bad}$ and both endpoints in the same part.
+Following the argument from Step 2 of Case 1, we know that there are $|V_{\rm bad}|$ disjoint rainbow directed $P_{3}$ copies, each having its center in $V_{\rm bad}$ and both endpoints in the same part.
 We now consider vertices in $X^{\prime}$.
-Based on the definition of $X^{\prime}$, we move vertices in $X^{\prime}$ to $Y$ if they are covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1).
+Based on the definition of $X^{\prime}$, we move a vertex in $X^{\prime}$ to $Y$ whenever it is covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1).
 In order to connect those rainbow directed $P_{3}$ with centers in $X^{\prime}$ into a single rainbow directed path, we balance the number of rainbow directed paths from $A$ to $B$ and those from $B$ to $A$.
 Claim 6.8.
-There exists a rainbow directed path $P^{0}$ such that its endpoints are not in $V_{\rm bad}\cup X^{\prime}$, $X^{\prime}\subseteq V(P^{0})$ and ${\rm col}(P^{0})\subseteq\mathcal{C}_{1}$, and its length is at most $8\sqrt{\epsilon}n$.
-Furthermore, such $P^{0}$ can be chosen starting at $A$ and ending at $B$ or two endpoints in the same part.
+There exists a rainbow directed path $P^{0}$ with length at most $8\sqrt{\epsilon}n$ such that $X^{\prime}\subseteq V(P^{0})$, ${\rm col}(P^{0})\subseteq\mathcal{C}_{1}$, and its endpoints lie outside $V_{\rm bad}\cup X^{\prime}$.
+Furthermore, $P^{0}$ can be chosen to either start at $A$ and end at $B$, or have both endpoints in the same part.
 Proof of Claim 6.8.
-If $|X^{\prime}|=0$, then by the minimum semi-degree condition, our result holds obviously.
-Without loss of generality, assume that $|X^{\prime}|\geq 1$ and $|A\cap X^{\prime}|-|B\cap X^{\prime}|\geq 0$.
-We first select all disjoint non-trivial maximum rainbow directed paths, say $Q_{1},\ldots,Q_{t}$, inside $\{D_{i}[A\cap X^{\prime}]:i\in\mathcal{C}_{1}\}$.
-Let $t^{\prime}:=|V(Q_{1})|+\cdots+|V(Q_{t})|$ and $Q_{i}:=x_{i}\ldots y_{i}$ for each $i\in[t]$.
-Next, choose a maximum rainbow directed matching, denoted by $\tilde{M}$, inside $\{D_{i}[A,B]:i\in\mathcal{C}_{1}\setminus{\rm col}(Q_{1}\cup Q_{2}\cup\ldots\cup Q_{t})\}$ avoiding vertices in $\bigcup_{i\in[t]}V(Q_{i}-y_{i})$.
-Partition $\tilde{M}$ to $M_{1}\cup M_{2}$ such that $V(M_{2})\cap A=V(\tilde{M})\cap A\cap X^{\prime}$.
-Hence, each vertex in $V(M_{2})\cap A$ or each $Q_{i}$ with $y_{i}\in V(M_{2})\cap A$ can be covered by a rainbow directed path with length at most $|X^{\prime}|+2$ and both endpoints in $B\setminus(V_{\rm bad}\cup X^{\prime})$.
-Denote $\hat{X}:=(A\cap X^{\prime})\setminus V\left(\bigcup_{i\in[t]}(Q_{i}-y_{i})\cup\tilde{M}\right)$.
-If $|\hat{X}|=0$ and $|B\cap X^{\prime}|\geq 1$, then the desired rainbow directed path can be obtained immediately by Claim 5.3 (iii) (note that if there is a rainbow directed path with some endpoint in $V_{\rm bad}$, then it can be extended by at most two vertices so that both of its endpoints remain in the corresponding part and are no longer in $V_{\rm bad}$).
-Next, we consider $|\hat{X}|\geq 1$.
+[…]
+Assume that $|\hat{X}|\geq 1$.
 We are to prove that $|E(M_{1})|\geq|\hat{X}|+1$.
-Suppose that $|E(M_{1})|\leq|\hat{X}|\leq\sqrt{\epsilon}n$.
+Suppose for contradiction that $|E(M_{1})|\leq|\hat{X}|\leq\sqrt{\epsilon}n$.
 Together with the definition of $X^{\prime}$, one has $B\cap X^{\prime}\subseteq V(\tilde{M})\cap B$ and $\tilde{M}$ can be chosen such that each matching edge incident to a vertex in $B\cap X^{\prime}$ is not incident to vertices in $A\cap X^{\prime}$.
-Choose $u_{1}\in\hat{X}$, $u_{2}\in B\setminus V(\tilde{M})$ and two colors $c_{1},c_{2}\in\mathcal{C}_{1}\setminus{\rm col}(\tilde{M}\cup Q_{1}\cup\ldots\cup Q_{t})$.
-We consider $N_{c_{1}}^{+}(u_{1})$ and $N_{c_{2}}^{-}(u_{2})$ respectively.
-It is easy to see that $u_{1}$ has at most $t^{\prime}-t$ out-neighbors inside $A\cap X^{\prime}$ and $N_{c_{1}}^{+}(u_{1},B)=N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))$.
-Hence
-$\left\lceil\frac{n}{2}\right\rceil\leq|N_{c_{1}}^{+}(u_{1})|\leq|A\setminus X^{\prime}|+(t^{\prime}-t)+|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|.$ (2)
+[…]
+[…]
 Similarly, we have
@@
 $|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|+|N_{c_{2}}^{-}(u_{2},A\cap V(\tilde{M}))|\geq|\hat{X}|+|E(M_{2})|+1.$
-Notice that, if there exists an edge $w_{1}w_{2}\in E(\tilde{M})$ such that $u_{1}w_{1}\in E(D_{c_{1}})$ and $w_{2}u_{2}\in E(D_{c_{2}})$, then we will find a large rainbow matching, a contradiction.
-Hence $|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|+|N_{c_{2}}^{-}(u_{2},A\cap V(\tilde{M}))|\leq|E(\tilde{M})|$.
-It follows that $|E(\tilde{M})|\geq|\hat{X}|+|E(M_{2})|+1$, which implies $|E(M_{1})|\geq|\hat{X}|+1$.
-Thus, together with Claim 5.3 (iii), the desired rainbow directed path $P^{0}$ can be obtained.
-Finally, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|=0$.
+[…]
+Now, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|=0$.
 If $|E(M_{1})|\geq 1$, then we are done.
@@
 Using a similar argument as (3) and (4), there is a vertex in $A\setminus(V_{\rm bad}\cup X^{\prime})$ that is adjacent to some vertex in $\{x_{1},\ldots,x_{t}\}\cup((A\cap X^{\prime})\setminus V(Q_{1}\cup\ldots\cup Q_{t}))$.
-Together with Claim 5.3 (iii), the desired rainbow directed path can be obtained immediately.
-∎
-By a similar discussion as Claim 5.4, we get the following claim.
-Claim 5.9.
-Assume $X\setminus X^{\prime}=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$.
-If $x_{i}\in(X\setminus X^{\prime})\cap Y$, then there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{1}\setminus({\rm col}(P^{0})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},\,x_{i}^{4}\in Y\setminus(V(P^{0})\cup V_{\rm bad}\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$.
-By Claim 5.9 and a similar discussion as Steps 1-3 in Case 1, the following hold.
+Together with Lemma 6.3 (iii), the desired rainbow directed path can be obtained immediately.
+Finally, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|\geq 1$.
+In this case, the desired rainbow directed path also follows immediately from part (iii) of Lemma 6.3.
+(Note that if there is a rainbow directed path with some endpoint in $V_{\rm bad}$, then it can be extended by at most two vertices so that both of its endpoints remain in the corresponding part and are no longer in $V_{\rm bad}$.) ∎
+By Lemma 6.1 and a similar discussion as Steps 1-3 in Case 1, the following hold.
   1. (D1)
-    In $\{D_{i}[A\setminus V(P^{0})]\cup D_{i}[B\setminus V(P^{0})]:i\in\mathcal{C}_{1}\setminus{\rm col}(P^{0})\}$, there are $|V_{\rm bad}\setminus V(P^{0})|$ disjoint rainbow directed $P_{3}$ with centers in $V_{\rm bad}\setminus V(P^{0})$ and both endpoints in the same part.
-Let $\mathbf{P}$ be a set consisting of those rainbow directed $P_{3}$.
+    In $\{D_{i}[A\setminus V(P^{0})]\cup D_{i}[B\setminus V(P^{0})]:i\in\mathcal{C}_{1}\setminus{\rm col}(P^{0})\}$, there are $|V_{\rm bad}\setminus V(P^{0})|$ disjoint rainbow directed $P_{3}$ copies, each with its center in $V_{\rm bad}\setminus V(P^{0})$ and both endpoints in the same part.
+Let $\mathbf{P}$ be the set of these rainbow directed $P_{3}$ copies.
   2. (D2)
-[…]
+    For colors in $\mathcal{C}_{\rm bad}$, there exists a maximal rainbow matching, say $M$, inside $\{D_{i}[A\setminus V(\mathbf{P}\cup P^{0})]\cup D_{i}[B\setminus V(\mathbf{P}\cup P^{0})]:i\in\mathcal{C}_{\rm bad}\}$ such that for each edge in $M[Y]$ with color $j$ we have $D_{j}[Y\setminus V(\mathbf{P}\cup P^{0}\cup{M})]$ contains a $2$-matching.
+Denote $\mathcal{C}_{\rm bad}^{\prime}:=\mathcal{C}_{\rm bad}\setminus{\rm col}(M)$.
+Then $D_{j}$ is $(19\sqrt{\delta},{\rm EC}2)$-extremal for all $j\in\mathcal{C}_{\rm bad}^{\prime}$.
   3. (D3)
     Let $q:=|\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime}|$.
-In the digraph collection $\{D_{i}^{\pm}[A,B]:i\in\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\}$, there exists a rainbow directed path $Q$ of length $q-1$ or $q$ that avoids vertices in $\mathbf{P}\cup P^{0}\cup M$.
+In the digraph collection $\{D_{i}^{\pm}[A,B]:i\in\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\}$, there exists a rainbow directed path $Q$ of length q1q-1 or $q$ that avoids vertices in $\mathbf{P}\cup P^{0}\cup M$.
   4. (D4)
     We can assume that a vertex $v\notin V(\mathbf{P}\cup M\cup Q)$ unless $v\in V_{\rm bad}$, and a color $c\notin{\rm col}(\mathbf{P})$ if $c\in\mathcal{C}_{1}$.
-Next, we give the following claim.
+Next, we give a claim that will be used later to obtain transversal directed Hamilton cycles.
 Claim 6.9.
@@
 $|N_{{i}}^{+}(z,I_{j})|$ $\geq d_{i}^{+}(z,A\setminus V(Q_{1}\cup Q_{2}))-|I_{3-j}|$
-$\geq d_{i}^{+}(z,A_{i})-|A_{1}\Delta A_{i}|-|X_{A}\setminus X_{A}^{1}|-|X^{\prime}|-|V(Q_{1}\cup Q_{2})|-|I_{3-j}|$
+$\geq d_{i}^{+}(z,A_{i})-|A_{1}\triangle A_{i}|-|X_{A}\setminus X_{A}^{1}|-|X^{\prime}|-|V(Q_{1}\cup Q_{2})|-|I_{3-j}|$
 $\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n-\left(2\delta+4\delta+4\eta+\sqrt{\epsilon}\right)n-\left\lceil\frac{n+r}{4}\right\rceil>\frac{n}{4}-5\eta n.$
-Similarly, we have $|N_{{i}}^{-}(z,I_{j})|\geq\frac{n}{4}-5\eta n$. It implies that
+Similarly, we have $|N_{{i}}^{-}(z,I_{j})|\geq\frac{n}{4}-5\eta n$. This implies that
 $|N_{{i}}^{+}(z,I_{j})\cap N_{{i}}^{-}(z,I_{j})|\geq 2\left(\frac{n}{4}-5\eta n\right)-|I_{j}|\geq\frac{n}{4}-10\eta n.$
-Hence, for each $i\in\mathcal{\mathcal{C}}_{a}\cup\{i_{1},i_{2}\}$, one may view $D_{i}^{\pm}[I_{1},I_{2}]$ as an undirected graph $G_{i}[I_{1},I_{2}]$ with minimum degree at least $\frac{n}{4}-10\eta n$, where two vertices are adjacent in $G_{i}[I_{1},I_{2}]$ if and only if there are two opposite arcs between them.
-By Claim 5.5, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with two endpoints $x_{a}\in N_{G_{i_{1}}}(z_{1},I_{1})$ and $y_{a}\in N_{G_{i_{2}}}(z_{1}^{\prime},I)$ (here $y_{a}\in I_{2}$ if and only if $|I_{1}|=|I_{2}|$).
+Hence, for each color $i\in\mathcal{\mathcal{C}}_{a}\cup\{i_{1},i_{2}\}$, one may view $D_{i}^{\pm}[I_{1},I_{2}]$ as an undirected graph $G_{i}[I_{1},I_{2}]$ with minimum degree at least $\frac{n}{4}-10\eta n$, where two vertices are adjacent in $G_{i}[I_{1},I_{2}]$ if and only if there are two opposite arcs between them.
+By Lemma 6.2, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with two endpoints $x_{a}\in N_{G_{i_{1}}}(z_{1},I_{1})$ and $y_{a}\in N_{G_{i_{2}}}(z_{1}^{\prime},I)$ (here $y_{a}\in I_{2}$ if and only if $|I_{1}|=|I_{2}|$).
 Similarly, there exists a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with two endpoints $x_{b}\in N_{G_{i_{3}}}(z_{s},J_{1})$ and $y_{b}\in N_{G_{i_{4}}}(z_{t}^{\prime},J)$ (here $y_{b}\in J_{2}$ if and only if $|J_{1}|=|J_{2}|$).
@@
 Subcase 2.1. $q$ is even and $q\geq 2$.
-Using Claim 5.3 (iii), we connect $P^{0}$ (with endpoints in the same part), all rainbow directed paths in $\mathbf{P},M$ and $Q$ (from $A$ to $B$ with length $q-1$) into a single rainbow directed path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in A\setminus(V_{\rm bad}\cup X^{\prime})$, $y\in B\setminus(V_{\rm bad}\cup X^{\prime})$.
-Next, choose an edge $u_{q}v_{q}$ from $D_{c}[B\setminus V(P),A\setminus V(P)]$, where $c\in(\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P)$.
+Using Lemma 6.3 (iii), we connect $P^{0}$ (with endpoints in the same part), all rainbow directed paths in $\mathbf{P},$ all edges in $M$ and $Q$ (from $A$ to $B$ with length q1q-1) into a single rainbow directed path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in A\setminus(V_{\rm bad}\cup X^{\prime})$, $y\in B\setminus(V_{\rm bad}\cup X^{\prime})$.
+Choose $c\in(\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P)$ and let $u_{q}v_{q}$ be an edge in $D_{c}[B\setminus V(P),A\setminus V(P)]$.
 Consequently, $P$ and $u_{q}v_{q}$ are two disjoint rainbow directed paths satisfying all conditions in Claim 6.9.
@@
 Subcase 2.2. $q$ is odd.
-Using Claim 5.3 (iii), we connect all rainbow paths in $\mathbf{P},M$ and $Q$ (from $B$ to $A$ with length $q$) into a single rainbow directed path $P$, whose length is at most $(3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in B\setminus V_{\rm bad}$, $y\in A\setminus V_{\rm bad}$.
-Consequently, $P$ and $P^{0}$ (from $A$ to $B$) are two disjoint rainbow paths satisfying all conditions in Claim 5.10.
-Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
+Using Lemma 6.3 (iii), we connect all rainbow paths in $\mathbf{P},$ all edges in $M$ and $Q$ (from $B$ to $A$ with length $q$) into a single rainbow directed path $P$, whose length is at most $(3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in B\setminus V_{\rm bad}$, $y\in A\setminus V_{\rm bad}$.
+Therefore, $P$ and $P^{0}$ (from $A$ to $B$) are two disjoint rainbow paths satisfying all conditions in Claim 6.9.
+It follows that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
 Subcase 2.3. $q=0$.
@@
 Delete $u$ from $X^{\prime}$ when constructing $P^{0}$.
-Based on the definition of $X^{\prime}$, there exists a rainbow directed $P_{3}$ from $B$ to $A$ with center $u$ that avoids vertices in $P^{0}$.
-Applying Claim 5.3 (iii) again to connect $P^{0}$ (from $A$ to $B$), all rainbow directed paths in $\mathbf{P},M$ into a single rainbow path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$, with endpoints in different parts.
+Based on the definition of $X^{\prime}$, there exists a rainbow directed copy of $P_{3}$ from $B$ to $A$ with center $u$, whose vertices are disjoint from $P^{0}$.
+Applying Lemma 6.3 (iii) again to connect $P^{0}$ (from $A$ to $B$), all rainbow directed paths in $\mathbf{P},M$ into a single rainbow path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$, with endpoints in different parts.
 Thus, $P$ and $P_{x}$ are two disjoint rainbow paths satisfying all conditions in Claim 6.9.
@@
   • if $v_{i}\in X_{Y}^{2}$, then $c_{4i-4+j}\in\mathcal{C}_{2}$, $v_{i}^{j}\in Z\cap Z_{c_{4i-4+j}}$, ${v_{i}^{j}v_{i}}\in E(D_{c_{4i-4+j}})$ for all $j\in[2]$ and ${v_{i}v_{i}^{j}}\in E(D_{c_{4i-4+j}})$ for all $j\in\{3,4\}$.
-Therefore, there exists a set of disjoint rainbow directed path $\mathbf{P}^{0}=\{v_{i}^{1}v_{i}v_{i}^{2}:i\in[t]\}$ with ${\rm col}(v_{i}^{1}v_{i}v_{i}^{2})=\{c_{4i-3},c_{4i-2}\}$ for each $i\in[t]$.
-Based on the definition of $X^{\prime}$, move vertices in $X^{\prime}$ to $Y$ if they are covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1).
-Thus, by avoiding all vertices and colors in $\mathbf{P}^{0}$, there exists a set of disjoint rainbow directed paths $\mathbf{P^{1}}=\{v^{1}vv^{2}:v\in X^{\prime}\}$ with $v^{1}\in Z\setminus(V_{\rm bad}\cup X^{\prime})$ and $v^{2}\in Y\setminus(V_{\rm bad}\cup X^{\prime})$ if $v\in Y\cap X^{\prime}$.
-Furthermore, at least one of the colors of ${v^{1}v}$ and ${vv^{2}}$ can be chosen from both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$.
+Therefore, there exists a set of disjoint rainbow directed paths $\mathbf{P}^{0}=\{v_{i}^{1}v_{i}v_{i}^{2}:i\in[t]\}$ with ${\rm col}(v_{i}^{1}v_{i}v_{i}^{2})=\{c_{4i-3},c_{4i-2}\}$ for each $i\in[t]$.
+Based on the definition of $X^{\prime}$, we move a vertex from $X^{\prime}$ to $Y$ whenever it is covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1).
+Avoiding all vertices and colors appearing in $\mathbf{P}^{0}$, we then obtain a family of disjoint rainbow directed paths $\mathbf{P^{1}}=\{v^{1}vv^{2}:v\in X^{\prime}\}$, where for each vertex $v\in Y\cap X^{\prime}$ we have $v^{1}\in Z\setminus(V_{\rm bad}\cup X^{\prime})$ and $v^{2}\in Y\setminus(V_{\rm bad}\cup X^{\prime})$.
+Furthermore, for every such path at least one of the colors of $v^{1}v$} and $vv^{2}$} can be chosen from both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$.
 Without loss of generality, assume that $|A|\leq|B|$.
-Connect all rainbow directed paths in $\mathbf{P}^{0}\cup\mathbf{P}^{1}$ by using colors in $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and Claim 5.3 (i)-(iii), one may obtain a rainbow directed path $P^{1}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{1}\in B\setminus V_{\rm bad}$, whose length is at most $4t+4|X^{\prime}|\leq 2\sqrt{\delta}n+4|X^{\prime}|$.
-By a similar discussion as Claim 5.4, we obtain the following claim.
-Claim 5.11.
-Assume $X\setminus X^{\prime}=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$.
-Then for each $i\in[s]$, there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in(\mathcal{C}_{1}\cup\mathcal{C}_{2})\setminus({\rm col}(P^{1})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},\,x_{i}^{4}\in V\setminus(V(P^{1})\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that: if $x_{i}\in Y$, then ${x_{i}^{j}x_{i}}\in E(D_{c_{i}^{j}})$ for all $j\in[2]$, and ${x_{i}x_{i}^{j}}\in E(D_{c_{i}^{j}})$ for all $j\in\{3,4\}$, and either $x_{i}^{j}\in Y$ and $c_{i}^{j}\in\mathcal{C}_{1}$ for all $j\in[4]$, or $x_{i}^{j}\in Z$ and $c_{i}^{j}\in\mathcal{C}_{2}$ for all $j\in[4]$.
-By Claim 5.11, there exists a set of disjoint rainbow directed path $\mathbf{P}^{2}=\{x_{i}^{1}x_{i}x_{i}^{2}:i\in[s]\}$, each of which has colors $c_{i}^{1},c_{i}^{2}$ for $i\in[s]$.
-Connect all these rainbow directed paths and $P^{1}$ using Claim 5.3 (i)-(ii), and we obtain a rainbow directed path $P^{2}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{2}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{1})|+4|X\setminus X^{\prime}|+2\leq(2\sqrt{\delta}+4\sqrt{\epsilon})n+2$.
-[…]
+Using colors in $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and applying Lemma 6.3 (i)-(iii), we connect all rainbow directed paths in $\mathbf{P}^{0}\cup\mathbf{P}^{1}$ into a single rainbow directed path $P^{1}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{1}\in B\setminus V_{\rm bad}$, whose length is at most $4t+4|X^{\prime}|\leq 2\sqrt{\delta}n+4|X^{\prime}|$.
+By Lemma 6.1, there exists a set of disjoint rainbow directed paths $\mathbf{P}^{2}=\{x_{i}^{1}x_{i}x_{i}^{2}:i\in[s]\}$, each of which has colors $c_{i}^{1},c_{i}^{2}$ for $i\in[s]$.
+Applying Lemma 6.3 (i)-(ii), we connect all these rainbow directed paths together with $P^{1}$ into a single rainbow directed path $P^{2}$, which has endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{2}\in B\setminus V_{\rm bad}$ and length at most $|E(P^{1})|+4|X\setminus X^{\prime}|+2\leq(2\sqrt{\delta}+4\sqrt{\epsilon})n+2$.
+Choose a maximal rainbow matching, say $M$, inside $\{D_{i}[A\setminus V(P^{2})]\cup D_{i}[B\setminus V(P^{2})]:i\in\mathcal{C}_{\rm bad}\}$ such that for each edge in $M[Y]$ with color $j$, we have $D_{j}[Y\setminus V(P^{2}\cup M)]$ contains a $2$-matching.
+Denote $\mathcal{C}_{\rm bad}^{\prime}:=\mathcal{C}_{\rm bad}\setminus{\rm col}(M)$.
+Note that $D_{j}$ is $(19\sqrt{\delta},{\rm EC}2)$-extremal for all $j\in\mathcal{C}_{\rm bad}^{\prime}$.
+Hence, $\{D_{i}[A\setminus V(P^{2}\cup M),B\setminus V(P^{2}\cup M)]:i\in\mathcal{C}_{\rm bad}^{\prime}\}$ contains a transversal matching, say $M^{\prime}$.
+By using Lemma 6.3 (i)-(iii) to connect $P^{2}$ and all rainbow edges in $M\cup M^{\prime}$, we obtain a rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{3}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{2})|+16\sqrt{\delta}n\leq(18\sqrt{\delta}+4\sqrt{\epsilon})n+2$.
 Since $u_{1},w_{3}\notin V_{\rm bad}$, there exist $i_{0},i_{1}\in\mathcal{C}_{1}\setminus{\rm col}(P^{3})$ such that $u_{1}\in A_{i_{0}}$ and $w_{3}\in B_{i_{1}}$.
@@
 Similar to the above two cases, we know that each digraph $D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{2}$ can be viewed as an undirected graph with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$; each digraph $D_{i}[A\setminus V(P^{3})]\cup D_{i}[B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{1}$ can be viewed as the union of two disjoint undirected graphs with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$.
-In what follows, we proceed by considering the parity of $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$.
-Subcase 3.1. $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is odd.
-Split $A\setminus V(P^{3})$ into $I_{1}$ and $I_{2}$ such that $|I_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|+1}{2}$, and split $B\setminus V(P^{3})$ into $J_{1}$ and $J_{2}$ such that $|J_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|+1}{2}$.
-Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus{\rm col}(P^{3})\}$.
-By Claim 5.5, there exists a transversal undirected path $Q$ inside $\mathcal{D}_{1}$ with endpoints $x\in I_{1}$ and $y\in J_{1}$.
-Since $x,y\not\in V_{\rm bad}$, there exist $j_{0},j_{1}\in\mathcal{C}_{1}$ such that $x\in A_{j_{0}}$ and $y\in B_{j_{1}}$.
-Suppose that $|I_{2}|=|A\setminus V(P^{3})|-|I_{1}|=k_{1}$ and $|J_{2}|=|B\setminus V(P^{3})|-|J_{1}|=k_{2}$.
-We further split $I_{2}$ into $I_{2}^{\prime}$ and $I_{2}^{\prime\prime}$ such that $|I_{2}^{\prime}|=\lceil\frac{k_{1}}{2}\rceil$, split $J_{2}$ into $J_{2}^{\prime}$ and $J_{2}^{\prime\prime}$ such that $|J_{2}^{\prime}|=\lceil\frac{k_{2}}{2}\rceil$.
-Then split the color set $\mathcal{C}_{1}\setminus({\rm col}(P^{3})\cup\{j_{0},j_{1}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$.
-Let $\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}$ and $\mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}$.
-Applying Claim 5.5 again yields that there is a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with endpoints $x_{a}\in N_{{i_{0}}}^{-}(u_{1},I_{2}^{\prime}),\,y_{a}\in N_{{j_{0}}}^{+}(x,I_{2})$; and a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with endpoints $x_{b}\in N_{{i_{1}}}^{+}(w_{3},J_{2}^{\prime}),\,y_{b}\in N_{{j_{1}}}^{-}(y,J_{2})$.
-Thus, $u_{1}P^{3}w_{3}x_{b}Q_{b}y_{b}yQxy_{a}Q_{a}x_{a}u_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$, a contradiction.
-Subcase 3.2. $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even.
-In this subcase, we first prove the following claim.
+We first prove the following claim.
 Claim 6.10.
@@
 Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{4}\})\}$.
-By Claim 5.5, there exists a transversal undirected path $Q$ inside $\mathcal{D}_{1}$ with endpoints $z_{1}^{\prime\prime}\in N_{{\ell_{4}}}^{+}(z_{1}^{\prime})\cap J_{1}$ and $y\in I_{1}$.
+By Lemma 6.2, $\mathcal{D}_{1}$ contains a transversal undirected path $Q$ with endpoints $z_{1}^{\prime\prime}\in N_{{\ell_{4}}}^{+}(z_{1}^{\prime})\cap J_{1}$ and $y\in I_{1}$.
 Since $y\not\in V_{\rm bad}$, there exists $\ell_{5}\in\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3}\})$ such that $y\in A_{\ell_{5}}$.
@@
 Split the color set $\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3},\ell_{5}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$.
-Let $\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}$ and $\mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}$.
-By Claim 5.5, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with endpoints $x_{a}\in N_{{\ell_{1}}}^{-}(z_{1})\cap I_{2}^{\prime},\,y_{a}\in N_{{\ell_{5}}}^{+}(y)\cap I_{2}$; and a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with endpoints $x_{b}\in N_{{\ell_{2}}}^{+}(z_{k})\cap J_{2}^{\prime},\,y_{b}\in N_{{\ell_{3}}}^{-}(z_{t}^{\prime})\cap J_{2}$.
-Hence, $z_{1}Q_{1}z_{k}x_{b}Q_{b}y_{b}z_{t}^{\prime}Q_{2}z_{1}^{\prime}z_{1}^{\prime\prime}Qyy_{a}Q_{a}x_{a}z_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$.
-∎
+Let
+$\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}\ \ \text{and}\ \ \mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}.$
+Applying Lemma 6.2 again yields that
+  • $\mathcal{D}_{a}$ contains a transversal undirected path $Q_{a}$ with endpoints $x_{a}\in N_{{\ell_{1}}}^{-}(z_{1})\cap I_{2}^{\prime},\,y_{a}\in N_{{\ell_{5}}}^{+}(y)\cap I_{2}$;
+  • $\mathcal{D}_{b}$ contains a transversal undirected path $Q_{b}$ with endpoints $x_{b}\in N_{{\ell_{2}}}^{+}(z_{k})\cap J_{2}^{\prime},\,y_{b}\in N_{{\ell_{3}}}^{-}(z_{t}^{\prime})\cap J_{2}$.
+Hence, $z_{1}Q_{1}z_{k}x_{b}Q_{b}y_{b}z_{t}^{\prime}Q_{2}z_{1}^{\prime}z_{1}^{\prime\prime}Qyy_{a}Q_{a}x_{a}z_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$. ∎
+Assume that $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is odd.
+Notice that there exists an edge $z_{1}^{\prime}z_{2}^{\prime}\in E(D_{i_{2}}[A\setminus V(P^{3}),B\setminus V(P^{3})])$ for some $i_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P^{3})$.
+This edge together with the path $P^{3}$ form two rainbow directed paths that satisfy all conditions in Claim 6.10.
+It follows that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
+Hence $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even.
 Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$ and $|A|\leq|B|$, each vertex in $A$ has an in-neighbor in $B$ in each $D_{i}$ with $i\in\mathcal{C}_{1}\cup{\rm col}(M)$.
-Suppose that there exists an edge $y_{2}y_{1}\in E(D_{i_{2}}[B,A])$ for some $i_{2}\in\mathcal{C}_{1}\cup{\rm col}(M)$, where $y_{1}\in A\setminus(V(P^{3})\cup V_{\rm bad})$ and $y_{2}\in B$.
-If $i_{2}\in{\rm col}(M)$, then assume ${uv}\in E(M[A])\cap E(D_{i_{2}})$.
-By Claim 5.3 (i)-(ii), we connect $P^{2}$, all rainbow edges in $M-uv$ and all rainbow edges in $M^{\prime}$ in turn.
-This results a rainbow directed path $\tilde{P}^{3}$ with endpoints $\tilde{u}_{1}\in A\setminus V_{\rm bad}$ and $\tilde{w}_{3}\in B\setminus V_{\rm bad}$.
+Suppose that there exists an edge $y_{2}y_{1}\in E(D_{i_{3}}[B,A])$ for some $i_{3}\in\mathcal{C}_{1}\cup{\rm col}(M)$, where $y_{1}\in A\setminus(V(P^{3})\cup V_{\rm bad})$ and $y_{2}\in B$.
+If $i_{3}\in{\rm col}(M)$, then assume ${uv}\in E(M[A])\cap E(D_{i_{3}})$.
+Applying Lemma 6.3 (i)-(ii), we successively connect $P^{2}$, all rainbow edges in MuvM-uv and all rainbow edges in $M^{\prime}$.
+This results in a rainbow directed path $\tilde{P}^{3}$ with endpoints $\tilde{u}_{1}\in A\setminus V_{\rm bad}$ and $\tilde{w}_{3}\in B\setminus V_{\rm bad}$.
 Therefore, $P^{3}$ can be written as $uvv^{\prime}\tilde{u}_{1}\tilde{P}^{3}\tilde{w}_{3}$, where the colors of $vv^{\prime}$} and $v^{\prime}\tilde{u}_{1}$} are in $\mathcal{C}_{2}$.
 Since $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even, one has $|\mathcal{C}_{2}\setminus{\rm col}(\tilde{P}^{3})|$ is even.
-Hence one may let $P^{3}:=\tilde{P}^{3}$ if $i_{2}\in{\rm col}(M)$ and further assume $i_{2}\not\in{\rm col}(P^{3})$.
+Hence one may let $P^{3}:=\tilde{P}^{3}$ if $i_{3}\in{\rm col}(M)$ and further assume $i_{3}\not\in{\rm col}(P^{3})$.
 Notice that $y_{2}\not\in V(P^{3})$ unless it is in $V_{\rm bad}\cup X^{\prime}$.
@@
 Hence, there exists a rainbow directed $\hat{P^{3}}$ starting at $A$ and ending at $B$ such that $P^{3}$ is obtained by connecting $y_{2}^{1}y_{2}y_{2}^{2}$ and $\hat{P^{3}}$ if $y_{2}^{1}\in A$, and $P^{3}$ is obtained by connecting $\hat{P^{3}}$ and $y_{2}^{1}y_{2}y_{2}^{2}$ otherwise.
-If $y\in X^{\prime}$, then by adjusting the selections of $y_{2}^{1},y_{2}^{2}$ and $j_{1},j_{2}$, one may get a rainbow directed path $P^{3}$ such that $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd.
-This can be reduced to Subcase 3.1.
-If $y\in V_{\rm bad}$, then $|\mathcal{C}_{2}\setminus{\rm col}(\hat{P^{3}})|$ is even.
-Let $\hat{P}:=y_{2}^{1}y_{2}y_{1}$ if $y_{2}^{1}\in B$, and let $\hat{P}:=y^{\prime}y_{2}^{1}y_{2}y_{1}$ with $y^{\prime}y_{2}^{1}\in E(D_{c^{\prime}})$ for unused color $c^{\prime}\in\mathcal{C}_{2}$ and unused vertex $y^{\prime}\in B$ if $y_{2}^{1}\in A$.
-Consequently, $\hat{P^{3}}$ and $\hat{P}$ are two rainbow directed paths satisfying all conditions in Claim 5.12.
-Hence, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
-∎
-In order to complete the proof of Theorem 5.1, we need to show the existence of transversal directed Hamilton cycles in Step 1 of Case 1. We first state a preliminary result from [11].
-Lemma 5.13 ([11]).
-Let $\mathcal{C}$ be a set of colors, and $\mathcal{G}=\{G_{i}[Y,B]:i\in\mathcal{C}\}$ be a collection of bipartite graphs with the common bipartition $Y\cup B$ such that $7|Y|<|B|\leq\frac{3}{5}|\mathcal{C}|$.
-If $\sum_{i\in\mathcal{C}}|E(D_{i}[Y,B])|\geq t|B||\mathcal{C}|$ for some integer $t$ with $1\leq t\leq|Y|$, then $\mathcal{G}$ contains $t$ disjoint rainbow star on $5$ vertices, each with the center in $Y$ and other vertices in $B$.
-Lemma 5.14.
-Assume $0<\frac{1}{n}\ll\delta\ll 1$ and $0\leq\gamma\leq 3{\delta}$.
-Let $\mathcal{C}$ be a set of $n$ colors, and let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs with the common vertex set $V$ of size $n$ such that $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$.
-Let $A\cup B$ be a partition of $V$ with $|A|=\left\lceil\frac{n}{2}\right\rceil+\gamma n$, and let $\mathcal{C}^{\prime}\cup\mathcal{C}^{\prime\prime}$ be a partition of $\mathcal{C}$ with $|\mathcal{C}^{\prime\prime}|\leq\delta n$.
-Assume that $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$.
-Then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
-Proof.
-Define $Y_{1}:=\{v\in A:d_{i}^{+}(v,B)\leq(1-\delta^{\frac{1}{4}})|B|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$, $Y_{2}:=\{v\in A:d_{i}^{-}(v,B)\leq(1-\delta^{\frac{1}{4}})|B|\ \textrm{for at least}\ \delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|\ \textrm{colors}\ i\in\mathcal{C}^{\prime}\}$.
-Since $D_{i}[B]=\emptyset$ for all $i\in\mathcal{C}^{\prime}$, it is routine to check that
-$\left\lceil\frac{n}{2}\right\rceil|B||\mathcal{C}^{\prime}|$ $\leq\sum_{i\in\mathcal{C}^{\prime}}|E({D_{i}}[A,B])|$
-$\leq|Y_{1}|(1-\delta^{\frac{1}{4}})|B|\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|+|Y_{1}||B|(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|+(|A|-|Y_{1}|)|B||\mathcal{C}^{\prime}|$
-$=(|A|-\delta^{\frac{1}{2}}|Y_{1}|)|B||\mathcal{C}^{\prime}|.$ (4)
-It implies that $|Y_{1}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$.
-Similarly, $|Y_{2}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$.
-For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp.
-$v\in Y_{2}\setminus Y_{1}$), we have $d_{i}^{-}(v,B)>(1-\delta^{\frac{1}{4}})|B|$ (resp.
-$d_{i}^{+}(v,B)>(1-\delta^{\frac{1}{4}})|B|$) for at least $(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|$ colors $i\in\mathcal{C}^{\prime}$.
-Notice that Claim 5.5 and Claim 5.3 (i)-(ii) hold by setting $V_{\rm bad}:=Y_{1}\cup Y_{2}$ and $\mathcal{C}_{2}=\mathcal{C}^{\prime}$.
-Since $\delta^{0}(\mathcal{D})\geq\left\lceil\frac{n}{2}\right\rceil$, we have $|E(D_{i}[A,B])|,|E(D_{i}[B,A])|\geq|B|\left\lceil\frac{n}{2}\right\rceil$ for each $i\in\mathcal{C}^{\prime}$, i.e., there are at most $|B|(\left\lceil\frac{n}{2}\right\rceil+\gamma n)-|B|\left\lceil\frac{n}{2}\right\rceil=|B|\gamma n$ non-edges from $A$ to $B$ (resp., from $B$ to $A$) in $D_{i}$.
-Hence
-$\sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[Y_{1},B])|\geq(|Y_{1}|-\gamma n)|B||\mathcal{C}^{\prime}|\ \text{and}\ \sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[B,Y_{2}])|\geq(|Y_{2}|-\gamma n)|B||\mathcal{C}^{\prime}|.$
-According to the sizes of $|Y_{1}|$ and $|Y_{2}|$, we proceed by considering the following three cases.
-Case 1. $|Y_{1}|,|Y_{2}|\geq\gamma n$.
-In view of Lemma 5.13, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[Y_{1},B]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[B,Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$.
-Recall that each vertex in $Y_{1}$ (resp.
-$Y_{2}$) has at least $\left\lceil\frac{n}{2}\right\rceil-(1-\delta^{\frac{1}{4}})|B|>4|Y|$ out-neighbors (resp.
-in-neighbors) in $A$ for at least $\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|>\delta^{\frac{1}{4}}(1-\delta)n>4|Y|$ digraphs $D_{i}$ with $i\in\mathcal{C}^{\prime}$.
-Therefore, by using colors in $\mathcal{C}^{\prime}$,
-  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by disjoint rainbow directed $P_{3}$ with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $B$,
-  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ starting at $B$ and ending at $A$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$,
-  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ starting at $A$ and ending at $B$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$,
-  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $A$.
-For each $y\in Y_{1}\cup Y_{2}$, denote the rainbow directed $P_{3}$ with center $y$ by $P_{y}:=y^{1}yy^{2}$ with colors $c_{y}^{1}$ and $c_{y}^{2}$.
-Let $\mathbf{P}=\{P_{y}:y\in Y_{1}\cup Y_{2}\}$.
-By using Claim 5.3, one may connect all of the above rainbow directed $P_{3}$ into a single rainbow directed path $P^{1}$ starting at $A$ and ending at $B$.
-Then
-$|V(P^{1})\cap A|-|V(P^{1})\cap B|=|Y_{1}^{\prime}\setminus Y_{2}^{\prime}|+|Y_{2}^{\prime}\setminus Y_{1}^{\prime}|+2|Y_{1}^{\prime}\cap Y_{2}^{\prime}|=|Y_{1}^{\prime}|+|Y_{2}^{\prime}|=2\gamma n.$
-Hence, $|A\setminus V(P^{1})|=|B\setminus V(P^{1})|+\sigma$, where $\sigma=0$ if $n$ is odd and $\sigma=1$ otherwise.
+If $y_{2}\in X^{\prime}$, then by adjusting the selections of $y_{2}^{1},y_{2}^{2}$ and $j_{1},j_{2}$, one may get a rainbow directed path $P^{3}$ such that $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd.
+This situation therefore reduces to the case where $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd.
 […]
-Subcase 1.1. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have the same parity.
-Greedily choose two disjoint rainbow directed paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\}$ with lengths $\lceil\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rceil$ and $\lfloor\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rfloor$ respectively, such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$.
-Applying Claim 5.3 (i) to connect $P^{1}$, all rainbow edges in $M$, $P_{A}$ and $P_{B}$ in turn, we get a single rainbow path $P^{2}$ with endpoints in different parts.
-Clearly, $|A\setminus V(P^{2})|=|B\setminus V(P^{2})|$.
-Together with Claim 5.5, we know $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
-Subcase 1.2. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have the different parity.
-Notice that $|B|=\lfloor\frac{n}{2}\rfloor-\gamma n$.
-Hence for each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $B$ has at least $\gamma n+1$ in-neighbors in $A$.
-Therefore, each vertex in $B$ has at least one in-neighbor outside $Y_{1}^{\prime}$.
-We first consider $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|\geq 1$.
-Choose an unused vertex $y\in B$ and a color $c\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$.
-Suppose that ${xy}\in E(D_{c}[A,B])$ with $x\notin Y_{2}^{\prime}$.
-$\bullet$ $x\in A\setminus(Y_{1}\cup Y_{2})$.
-Then $x$ can be chosen outside $V(P^{1}\cup M)$.
-By Claim 5.3 (i)-(ii), one may connect $xy$ and $P^{1}$ into a single rainbow directed path $P$.
-Applying a similar discussion as Subcase 1.1, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
-$\bullet$ $x\in(Y_{1}\cup Y_{2})\setminus Y_{1}^{\prime}$.
-Then replace the end vertex of $P_{x}$ by $y$.
-Applying a similar discussion as Subcase 1.1, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
-Therefore, it suffices to consider $n$ is odd and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|=0$.
-For each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $A$ has at least one in-neighbor outside $Y_{1}^{\prime}$.
-Choose an unused vertex $y\in A\setminus(Y_{1}\cup Y_{2})$ and a color $c\in\mathcal{C}$.
-Then there exists a vertex $x\in A\setminus Y_{1}^{\prime}$ such that ${xy}\in E(D_{c}[A])$.
-By a similar discussion as above, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
-Case 2. At least one of $Y_{1}$ and $Y_{2}$ has size less than $\gamma n$.
-In this case, we consider only $|Y_{1}|\geq\gamma n$ and $|Y_{2}|<\gamma n$; the other two cases can be discussed similarly.
-Based on the proof of Case 1, we know that all vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{1}$ and endpoints in $A\setminus(Y_{1}\cup Y_{2})$.
-Furthermore, each vertex in $Y_{2}\setminus Y_{1}$ can be covered by a rainbow directed $P_{3}$ with center in $Y_{2}\setminus Y_{1}$ and start vertex in $A\setminus(Y_{1}\cup Y_{2})$ using colors in $\mathcal{C}^{\prime}$.
-(In this process, for the choice of the other endpoint of each rainbow directed $P_{3}$, we prefer using vertices in $A$ over vertices in $B$.) By using unused colors in $\mathcal{C}$ and unused vertices in $A\setminus Y$, one may extend those rainbow directed $P_{3}$ as long as possible or choose all other disjoint rainbow directed paths.
-Let $\mathbf{P}=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow directed paths in the above, where each $Q_{i}$ has at least one endpoint in $A\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $A$.
-Subcase 2.1. $|B|\geq|A|-(s_{1}+\cdots+s_{t})-\sigma$.
-In this subcase, there exists a set $\mathbf{P}^{\prime}=\{Q_{1}^{\prime},Q_{2}^{\prime},\ldots,Q_{\ell}^{\prime}\}$ such that $|B|=|A|-(s_{1}^{\prime}+\cdots+s_{\ell}^{\prime})-\sigma$, the endpoints of $Q_{i}^{\prime}$ are not in $Y$ for all $i\in[\ell]$ and $Y_{2}\cup Y_{1}^{\prime}\subseteq V(Q_{1}^{\prime}\cup\ldots\cup Q_{\ell}^{\prime})$, where $s_{i}^{\prime}:=|E(Q_{i}^{\prime}[A])|$ for each $i\in[\ell]$.
-By Claim 5.3 (i)-(ii), we connect all rainbow directed paths in $\mathbf{P}^{\prime}$ into a single rainbow directed path $P^{1}$, whose endpoints are in different parts.
-Since $|A|-|B|=2\gamma n+\sigma$, we have $\ell\leq s_{1}^{\prime}+\cdots+s_{\ell}^{\prime}=2\gamma n$.
-Therefore, $|E(P^{1})|\leq 8\gamma n$ and $|A\setminus V(P^{1})|-\sigma=|B\setminus V(P^{1})|$.
-Recall that all vertices in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ can also be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ and both endpoints in $B$.
-Applying Lemma 5.3 (i) again to connect those rainbow directed paths with $P^{1}$, we get a rainbow directed path $P^{2}$ with length at most $8\gamma n+12\sqrt{\delta}n$, and with endpoints in different parts.
-[…]
-Based on Claim 5.3 (i), one may connect $P^{1}$ and all rainbow directed edges in $M$ into a single rainbow directed path $P^{2}$, whose endpoints are in different parts.
-Clearly, $|E(P^{2})|\leq 8\gamma n+4{\delta}n$.
-Next, we are to choose two disjoint rainbow paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)\}$ such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$ respectively, whose lengths are determined by the parity of $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$.
-Clearly, if $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have the same parity, then by Claim 5.5, we know $\mathcal{G}$ contains a transversal Hamilton cycle, a contradiction.
-Hence $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have different parity.
-By similar arguments as the proof of Step 4 in Case 1, we can prove our result.
-Subcase 2.2. $|B|<|A|-(s_{1}+\cdots+s_{t})-\sigma$.
-In this case, $\sum_{i=1}^{t}s_{i}<2\gamma n$.
-Let $w$ be an arbitrary vertex in $A\setminus V(\mathbf{P})$ and $c_{1},c_{2}$ be two colors in $\mathcal{C}\setminus{\rm col}(\mathbf{P})$.
-Clearly, in $D_{c_{1}}$ and $D_{c_{2}}$, $w$ has at least $\left\lceil\frac{n}{2}\right\rceil-(\lfloor\frac{n}{2}\rfloor-\gamma n)=\gamma n+\sigma$ out-neighbors (resp.
-in-neighbors) in $V(\mathbf{P})$ and it cannot be adjacent to the start vertex of $Q_{i}[A]$ (resp.
-from the end vertex of $Q_{i}[A]$) for all $i\in[t]$ (by the maximality of $\mathbf{P}$).
-For convenience, assume that $s_{0}=0$ and $V(Q_{i})\cap A:=\{v_{s_{1}+\cdots+s_{i-1}+i},\ldots,v_{s_{1}+\cdots+s_{i}+i}\}$ for each $i\in[t]$.
-Define
-$I_{1}:=\left\{i:v_{i}\in N_{{c_{1}}}^{-}(w)\right\}\ \text{and}\ I_{2}:=\left\{i:v_{i+1}\in N_{{c_{2}}}^{+}(w)\right\}.$
-Hence $I_{1},I_{2}\subseteq\bigcup_{i\in[t]}\left[\sum_{j\in[i-1]}s_{j}+i,\sum_{j\in[i]}s_{j}+i-1\right]$.
-Therefore, $I_{1}\cap I_{2}\neq\emptyset$, which contradicts the maximality of $\mathbf{P}$.
-∎
-5.2 Proof of Theorem 5.2
-In this subsection, we give the proof of Theorem 5.2.
+7 Proof of Theorem 5.3
+In this section we present the proof of Theorem 5.3.
+Since the overall strategy parallels that of Theorem 5.2, the argument itself is relatively short; therefore we do not separate out preliminary claims but instead give a complete, self-contained proof of Theorem 5.3.
 Proof of Theorem 5.3.
-Recall that if $|\mathcal{C}_{3}|\geq 2\sqrt{\delta}n$, then $|\mathcal{C}_{1}|+|\mathcal{C}_{2}|<2\sqrt{\delta}n$.
-Denote $[m]:=\mathcal{C}_{3}\setminus I^{2}$ and $\mathcal{C}_{\rm bad}:=[m+1,n]$.
-By adding colors to $\mathcal{C}_{\rm bad}$ if necessary, we may assume $m=(1-4\sqrt{\delta})n$.
-By (B2), we have $|C_{1}^{k}\Delta C_{i}^{k}|\leq 2\delta n$ for all $k\in[4]$ and all $i\in[m]$.
+Choose constant $\epsilon,\delta$ such that $\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll 1,$ let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$.
+Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles.
+For each $i\in[m]$, since $D_{i}$ is $(\epsilon,{\rm EC3})$-extremal, one may assume that $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ is a $\epsilon$-characteristic partition of $D_{i}$.
+By Lemma 5.1, one has $|C_{1}^{k}\triangle C_{i}^{k}|\leq 2\delta n$ for all $k\in[4]$ and all $i\in[m]$.
 Expand $C_{1}^{1}\cup C_{1}^{2}\cup C_{1}^{3}\cup C_{1}^{4}$ into a partition $C^{1}\cup C^{2}\cup C^{3}\cup C^{4}$ of $V$ such that $|C^{1}|-|C^{3}|\leq 1$ and $|C^{2}|=|C^{4}|$.
-Hence, for $i\in[2,m]$ and $k\in[4]$, we have $|C^{k}\Delta C_{i}^{k}|\leq 2\delta n+2{\epsilon}n<3\delta n$.
+Hence, for $i\in[2,m]$ and $k\in[4]$, we have $|C^{k}\triangle C_{i}^{k}|\leq 2\delta n+2{\epsilon}n<3\delta n$.
 Now, we define the set of bad vertices. Let $(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}$. Define
-$X:=\left\{x\in V:x\in L_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 3\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
-$X_{YZ}:=\left\{x\in(Y\cup Z)\setminus X:x\not\in Y_{i}\cup Z_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 10\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
-$X_{YZ}^{\prime}:=\left\{x\in X_{YZ}:x\in W_{i}\leavevmode\nobreak\ \textrm{for some $W\in\{Y,Z\}$ with at least}\leavevmode\nobreak\ \frac{3}{2}\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{3}}\right\}.$
-Recall that for all $i\in{\mathcal{C}}_{3}$, we have $|L_{i}|\leq 2{\epsilon}n$ and $|(Y_{1}\cup Z_{1})\Delta(Y_{i}\cup Z_{i})|<4\delta n$. Hence
+$X:=\left\{x\in V:x\in L_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 3\sqrt{\delta}|\mathcal{C}_{3}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}}_{3}\right\},$
+$X_{YZ}:=\left\{x\in(Y\cup Z)\setminus X:x\not\in Y_{i}\cup Z_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 10\sqrt{\delta}|\mathcal{C}_{3}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}}_{3}\right\},$
+$X_{YZ}^{\prime}:=\left\{x\in X_{YZ}:x\in W_{i}\penalty 10000\ \textrm{for some $W\in\{Y,Z\}$ with at least}\penalty 10000\ \frac{3}{2}\sqrt{\delta}n\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}_{3}}\right\}.$
+Recall that for all $i\in{\mathcal{C}}_{3}$, we have $|L_{i}|\leq 2{\epsilon}n$ and $|(Y_{1}\cup Z_{1})\triangle(Y_{i}\cup Z_{i})|<4\delta n$. Hence
 $3\sqrt{\delta}|{\mathcal{C}}_{3}||X|\leq 2{\epsilon}n|{\mathcal{C}}_{3}|\ \ \text{and}\ \ 10\sqrt{\delta}|{\mathcal{C}}_{3}||X_{YZ}|\leq\sum_{i\in\mathcal{C}_{3}}|(Y\cup Z)\setminus(Y_{i}\cup Z_{i})|\leq(4\delta+2{\epsilon})n|{\mathcal{C}}_{3}|.$
@@
 $|X_{YZ}\setminus X_{YZ}^{\prime}|(1-4\sqrt{\delta})|{\mathcal{C}}_{3}|\leq$ $\sum_{i\in{\mathcal{C}}_{3}}|(Y\cup Z)\setminus(Y_{i}\cup Z_{i})|$
-$\leq$ $\sum_{i\in\mathcal{C}_{3}}\left(|(Y\cup Z)\setminus(Y_{1}\cup Z_{1})|+|(Y_{1}\cup Z_{1})\Delta(Y_{i}\cup Z_{i})|\right)$
+$\leq$ $\sum_{i\in\mathcal{C}_{3}}\left(|(Y\cup Z)\setminus(Y_{1}\cup Z_{1})|+|(Y_{1}\cup Z_{1})\triangle(Y_{i}\cup Z_{i})|\right)$
 $\leq$ $(4\delta+2\epsilon)n|{\mathcal{C}}_{3}|.$
@@
 Let $\{Y,Z\}\in\left\{\{C^{1},C^{3}\},\{C^{2},C^{4}\}\right\}$. Define
-$X_{Y}:=$$\left\{x\in Y\setminus(X\cup X_{YZ}^{\prime}):x\not\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 20\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
-$X_{Y}^{\prime}:=$$\left\{x\in X_{Y}:x\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 6\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{3}}\right\}.$
+$X_{Y}:=$ $\left\{x\in Y\setminus(X\cup X_{YZ}^{\prime}):x\not\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 20\sqrt{\delta}|\mathcal{C}_{3}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}}_{3}\right\},$
+$X_{Y}^{\prime}:=$ $\left\{x\in X_{Y}:x\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 6\sqrt{\delta}n\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}_{3}}\right\}.$
 By a similar discussion as above, we know that $|X_{Y}|<\frac{1}{2}\sqrt{\delta}n$ and each vertex of $X_{Y}\setminus X_{Y}^{\prime}$ lies in $Z_{i}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in\mathcal{C}_{3}$.
 Furthermore, $|X_{Y}\setminus X_{Y}^{\prime}|\leq 4\delta n$.
-Then move vertices in $X_{Y}\setminus X_{Y}^{\prime}$ to $Z$.
+We then move vertices in $X_{Y}\setminus X_{Y}^{\prime}$ to $Z$.
 Hence $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n$.
-Define $V_{\rm bad}:=X\cup(\bigcup_{(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}}X_{YZ}^{\prime})\cup(\bigcup_{k\in[4]}X_{C^{k}}^{\prime})$.
+Define $V_{\rm bad}:=X\cup\big(\bigcup_{(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}}X_{YZ}^{\prime}\big)\cup\big(\bigcup_{k\in[4]}X_{C^{k}}^{\prime}\big)$.
 Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+3\sqrt{\delta})n$.
 Moreover, each vertex in $C^{k}\setminus V_{\rm bad}$ lies in $C_{i}^{k}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in{\mathcal{C}}_{3}$.
-The subsequent result can be used to connect two disjoint short rainbow directed paths into a single short rainbow directed path.
-The proof follows a similar argument to that of Claim 5.3, and thus we omit it.
+The subsequent result allows us to connect two disjoint short rainbow directed paths into a single short rainbow directed path.
+Its proof follows the same lines as that of Lemma 6.3, and is therefore omitted.
 Claim 7.1 (Connecting tool).
 Let $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ be two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s}\in Y\setminus V_{\rm bad}$, $v_{1}\in Z\setminus V_{\rm bad}$, where $Y,Z\in\{C^{1},C^{2},C^{3},C^{4}\}$.
-Assume that $|\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)|>21\sqrt{\delta}n$ and $|C^{i}\setminus(V(P\cup Q)\cup V_{\rm bad})|>25\delta n$ if $w_{1}$ or $w_{1}^{\prime}$ is chosen from $C^{i}$ for some $i\in[4]$.
+Assume that $|\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)|>21\sqrt{\delta}n$, and that for all $i\in[4]$, if $C^{i}$ is a candidate set for selecting $w_{1}$ or $w_{1}^{\prime}$, then $|C^{i}\setminus(V(P\cup Q)\cup V_{\rm bad})|>25\delta n$.
   1. (i)
     If $(Y,Z)\in\{(C^{1},C^{2}),(C^{3},C^{4})\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
-Similar result holds for $(Y,Z)\in\{(C^{2},C^{3}),(C^{4},C^{1})\}$.
+A similar result holds for $(Y,Z)\in\{(C^{2},C^{3}),(C^{4},C^{1})\}$.
   2. (ii)
@@
     If $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(1,3),(1,4),(4,3)\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in C^{2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
-Similar result holds for $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(3,1),(3,2),(2,1)\}$.
+A similar result holds for $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(3,1),(3,2),(2,1)\}$.
   5. (v)
-    If $(Y,Z)=(C^{i},C^{i})$ with $i\in\{2,4\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and two vertices $w_{1}\in C^{i+2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
+    If $(Y,Z)=(C^{i},C^{i})$ with $i\in\{2,4\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in C^{i+2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
 Delete vertices in $X$ from $\bigcup_{k\in[4]}C^{k}$ and choose a vertex $x\in X$.
@@
   1. (a)
-    If $|N_{i}^{+}(x,C^{1}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{3}$ with center in $x$ and both endpoints in $C^{1}$.
-Move such $x$ to $C^{1}$.
+    If $|N_{i}^{+}(x,C^{1}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed copy of $P_{3}$ centered at $x$ with both endpoints in $C^{1}$.
+We then move such $x$ to $C^{1}$.
   2. (b)
     Notice that each vertex in $C^{1}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ in-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$.
-If $|N_{i}^{+}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{4}$ with center in $x$ and endpoints in $C^{4},C^{2}$ respectively.
-Move such $x$ to $C^{1}$.
+If $|N_{i}^{+}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed copy of $P_{4}$ centered at $x$ with endpoints in $C^{4},C^{2}$ respectively.
+We then move such $x$ to $C^{1}$.
     Suppose (a)-(b) do not hold.
 Then for at least $(1-400\delta)|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|C^{3}\cup C^{4}|-400\delta n-|V_{\rm bad}|$ out-neighbors in $(C^{3}\cup C^{4})\setminus V_{\rm bad}$.
-Notice that each vertex in $C^{3}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ out-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$.
+Observe that every vertex in $C^{3}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ out-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$.
   3. (c)
-    If $|N_{i}^{-}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{3}$ or $P_{4}$ with center in $x$ and endpoints in $C^{2},C^{4}$ respectively.
-Move such $x$ to $C^{3}$.
-Similar result holds when $|N_{i}^{-}(x,C^{3}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, move such $x$ to $C^{3}$.
+    If $|N_{i}^{-}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then we can find a rainbow directed copy of $P_{3}$ or $P_{4}$ centered at $x$ with endpoints in $C^{2},C^{4}$ respectively; we then move such $x$ to $C^{3}$.
+The same conclusion holds when $|N_{i}^{-}(x,C^{3}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$; in that case we also move $x$ to $C^{3}$.
   4. (d)
@@
 Then in each $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|(C^{1}\cup C^{4})\setminus V_{\rm bad}|-400\delta n-|V_{\rm bad}|$ in-neighbors inside $(C^{1}\cup C^{4})\setminus V_{\rm bad}$ for all but at most $(1-400\delta)|\mathcal{C}_{3}|$ colors $i\in\mathcal{C}_{3}$.
-Move such $x$ to $C^{2}$.
-Applying the above operation for $j\in\{2,3,4\}$ in turn.
-Denote by $X^{\prime}$ the subset of $X$ consisting of vertices from item (d).
-Hence, for vertices in $X\setminus X^{\prime}$, there exists a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $X\setminus X^{\prime}$, each having both endpoints in $C^{k}$ with $k\in\{1,3\}$, or one endpoint in $C^{2}$ and the other in $C^{4}$.
+We then move $x$ to $C^{2}$.
+We perform the above operation successively for $j=2,3,4$.
+Let $X^{\prime}$ denote the subset of $X$ formed by the vertices described in item (d).
+Hence, for the vertices in $X\setminus X^{\prime}$, there exists a set of disjoint rainbow directed copies of $P_{3}$ or $P_{4}$ centered at those vertices, where each copy either has both endpoints in the same part $C^{k}$ with $k\in\{1,3\}$, or has one endpoint in $C^{2}$ and the other in $C^{4}$.
 Moreover, $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n+2\sqrt{\epsilon}n$.
-By the definition of $V_{\rm bad}\setminus X$, for vertices in $(V_{\rm bad}\setminus X)\cap Y$, we can find a set of disjoint rainbow directed $P_{3}$ with centers in $(V_{\rm bad}\setminus X)\cap Y$ and endpoints in $Y\setminus V_{\rm bad}$ (if $Y\in\{C^{1},C^{3}\}$) or in $(C^{2}\cup C^{4})\setminus(Y\cup V_{\rm bad})$ (if $Y\in\{C^{2},C^{4}\}$).
+By the definition of $V_{\rm bad}\setminus X$, for vertices in $(V_{\rm bad}\setminus X)\cap Y$, we can find a set of disjoint rainbow directed copies of $P_{3}$ with centers in $(V_{\rm bad}\setminus X)\cap Y$ and endpoints in $Y\setminus V_{\rm bad}$ (if $Y\in\{C^{1},C^{3}\}$) or in $(C^{2}\cup C^{4})\setminus(Y\cup V_{\rm bad})$ (if $Y\in\{C^{2},C^{4}\}$).
 Denote the rainbow directed path with center $x$ obtained above by $P_{x}$ for each $x\in V_{\rm bad}\setminus X^{\prime}$.
@@
 Choose a set of disjoint maximal rainbow directed paths in $\mathcal{D}[C^{2}]\cup\mathcal{D}[C^{2},C^{1}]$, denote such a set by $\{Q_{1},\ldots,Q_{t}\}$.
-Assume $\{Q_{1}^{\prime},\ldots,Q_{k}^{\prime}\}$ is a set of disjoint rainbow directed $P_{3}$ in $\mathcal{D}[C^{2},C^{3}]$ with centers in $C^{2}$ and endpoints in $C^{3}$.
-If $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$, then there must exist an integer $k^{\prime}\in[k]$ and a set of disjoint rainbow directed subpaths of $\{Q_{1},\ldots,Q_{t}\}$, say $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}}\}$, such that $|E(\tilde{Q}_{1})|+\cdots+|E(\tilde{Q}_{t^{\prime}})|+k^{\prime}=|C^{2}|-|C^{4}|-\sigma$.
+Assume that $\{Q_{1}^{\prime},\ldots,Q_{k}^{\prime}\}$ is a set of disjoint rainbow directed copies of $P_{3}$ in $\mathcal{D}[C^{2},C^{3}]$ with centers in $C^{2}$ and endpoints in $C^{3}$.
+On the one hand, we consider that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$.
+Then there exist an integer $k^{\prime}\in[k]$ and a set of disjoint rainbow directed subpaths of $\{Q_{1},\ldots,Q_{t}\}$, say $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}}\}$, such that $|E(\tilde{Q}_{1})|+\cdots+|E(\tilde{Q}_{t^{\prime}})|+k^{\prime}=|C^{2}|-|C^{4}|-\sigma$.
 That is, $|C^{4}|-t^{\prime}=|C^{2}|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-k^{\prime}-\sigma$.
 If the end vertex of $\tilde{Q_{i}}$ with $i\in[t^{\prime}]$ (resp.
-the start vertex of $Q_{i}^{\prime}$ with $i\in[k^{\prime}]$) lies in $V_{\rm bad}\setminus X^{\prime}$, then extend it by a vertex not in $V_{\rm bad}$.
-Using Claim 5.15 to connect the resulting $\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}},Q_{1}^{\prime},\ldots,Q_{k^{\prime}}^{\prime}$ into a single rainbow path $P^{0}$ with endpoints in $C^{2}$ and $C^{4}$ respectively, whose length is at most $190\delta n$.
+the start vertex of $Q_{i}^{\prime}$ with $i\in[k^{\prime}]$) lies in $V_{\rm bad}\setminus X^{\prime}$, then we extend it by a vertex not in $V_{\rm bad}$.
+Applying Claim 7.1, we then connect the modified paths $\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}},Q_{1}^{\prime},\ldots,Q_{k^{\prime}}^{\prime}$ into a single rainbow path $P^{0}$ with endpoints in $C^{2}$ and $C^{4}$ respectively, whose length is at most $190\delta n$.
 Therefore, $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-\sigma$, as desired.
-In each of the above cases, the remaining vertices in $C^{2}$ and $C^{4}$ are balanced.
+In each of the above cases, that is, either $|C^{2}|-|C^{4}|-\sigma\leq s_{1}+2s_{2}$, or $|C^{2}|-|C^{4}|-\sigma>s_{1}+2s_{2}$ and $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$, the number of the remaining vertices in $C^{2}$ and $C^{4}$ are balanced.
 Since $|E(P^{0})|\leq 190\delta n$, by avoiding vertices in $V(P^{0})$, one may choose a set of disjoint rainbow directed paths $\mathbf{P}^{0}=\{P_{x}:x\in V_{\rm bad}\setminus(X^{\prime}\cup P^{0})\}$.
 For colors in $\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{0})$, we select a maximal rainbow matching $M$ inside $\mathcal{D}^{\pm}[C^{2},C^{4}]\cup\mathcal{D}[C^{1}]\cup\mathcal{D}[C^{3}]\cup(\bigcup_{i\in[4]}\mathcal{D}[C^{i},C^{i+1}])$ ensuring it avoids vertices in $V(P^{0}\cup\mathbf{P}^{0})$.
-If there exists a color $c\in\mathcal{C}\setminus({\rm col}(P^{0}\cup M))$, then $D_{c}^{\pm}[C^{1},C^{3}]$ is almost an undirected complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost undirected complete graphs.
-For colors in $\mathcal{C}\setminus({\rm col}(P^{0}\cup M))$,we can choose a set of disjoint rainbow directed paths, say $\mathbf{P}^{1}$, with centers in $C^{1}$ and endpoints inside $C^{3}$, ensuring that at most one color in $\mathcal{C}_{\rm bad}$ is remains unused.
+Notice that if there exists a color $c\in{\mathcal{C}_{\rm bad}}\setminus({\rm col}(P^{0}\cup M))$, then $D_{c}^{\pm}[C^{1},C^{3}]$ is almost an undirected complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost undirected complete graphs.
+Consequently, for colors in ${\mathcal{C}_{\rm bad}}\setminus({\rm col}(P^{0}\cup M))$, we can choose a set of disjoint rainbow directed paths, say $\mathbf{P}^{1}$, with centers in $C^{1}$ and endpoints inside $C^{3}$, ensuring that at most one color in $\mathcal{C}_{\rm bad}$ remains unused.
 Let $\mathbf{P}:=\mathbf{P}^{0}\cup\mathbf{P}^{1}\cup M\cup\{P^{0}\}$.
@@
 We proceed by considering the following two cases.
-$\bullet$ $\sigma=0$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|$.
+$\bullet$ Suppose that $\sigma=0$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|$.
 We first assume that $\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})=\emptyset$.
-By applying Claim 5.5 inside the digraph collection $\{D_{i}[C^{1}]\cup D_{i}[C^{3}]:i\in\mathcal{C}_{3}\setminus\{i_{1},i_{2},i_{3},i_{4}\}\}$ and avoiding vertices and colors in $Q_{0}\cup Q_{1}\cup Q_{2}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, where $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$.
-Connecting $Q_{1},Q_{2}$ by using Claim 5.15 (iv), one gets a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.
-Applying Claim 5.5 to the digraph collection $\{D_{i}^{\pm}[C^{2}\setminus V(P\cup Q_{0}),C^{4}\setminus V(P\cup Q_{0})]:i\in\mathcal{C}_{3}\setminus({\rm col}(P\cup Q_{0})\cup\{i_{1},i_{2},c_{1},c_{2}\})\}$, there exist two disjoint rainbow directed paths $P^{3}=x_{3}\ldots y_{3}$ and $P^{4}=x_{4}\ldots y_{4}$ with $x_{3}\in N_{{c_{1}}}^{+}(y_{2},C^{4}\setminus V_{\rm bad})$, $y_{3}\in N_{{i_{1}}}^{-}(u_{0},C^{4}\setminus V_{\rm bad})$, $x_{4}\in N_{{i_{2}}}^{+}(v_{0},C^{2}\setminus V_{\rm bad})$ and $y_{4}\in N_{{c_{2}}}^{-}(x_{1},C^{4}\setminus V_{\rm bad})$, respectively.
+By applying Lemma 6.2 to the digraph collection $\{D_{i}[C^{1}]\cup D_{i}[C^{3}]:i\in\mathcal{C}_{3}\setminus\{i_{1},i_{2},i_{3},i_{4}\}\}$, while avoiding vertices and colors in $Q_{0}\cup Q_{1}\cup Q_{2}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$.
+Here $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$.
+We then connect $Q_{1},Q_{2}$ via Claim 7.1 (iv) to form a rainbow directed path $P=x_{1}\ldots y_{2}$, which satisfies $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.
+Applying Lemma 6.2 to the digraph collection $\{D_{i}^{\pm}[C^{2}\setminus V(P\cup Q_{0}),C^{4}\setminus V(P\cup Q_{0})]:i\in\mathcal{C}_{3}\setminus({\rm col}(P\cup Q_{0})\cup\{i_{1},i_{2},c_{1},c_{2}\})\}$, we obtain two disjoint rainbow directed paths $P^{3}=x_{3}\ldots y_{3}$ and $P^{4}=x_{4}\ldots y_{4}$, where $x_{3}\in N_{{c_{1}}}^{+}(y_{2},C^{4}\setminus V_{\rm bad})$, $y_{3}\in N_{{i_{1}}}^{-}(u_{0},C^{4}\setminus V_{\rm bad})$, $x_{4}\in N_{{i_{2}}}^{+}(v_{0},C^{2}\setminus V_{\rm bad})$ and $y_{4}\in N_{{c_{2}}}^{-}(x_{1},C^{4}\setminus V_{\rm bad})$, respectively.
 Hence $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
 Now, assume that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$.
-Together with the condition $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, there exists an edge $uv\in E(D_{c}[C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{3}\cup C^{4}])$ if $|C^{1}|+|C^{2}|\leq|C^{3}|+|C^{4}|$, and $vu\in E(D_{c}[C^{4}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{1}\cup C^{2}])$ otherwise.
-By a similar discussion as above and considering whether the vertex $v$ belongs to $V_{\rm bad}$ or not, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
-$\bullet$ $\sigma=1$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-1$.
+Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, there exists an edge $uv\in E(D_{c}[C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{3}\cup C^{4}])$ if $|C^{1}|+|C^{2}|\leq|C^{3}|+|C^{4}|$, and $vu\in E(D_{c}[C^{4}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{1}\cup C^{2}])$ otherwise.
+By an argument analogous to that above, and considering whether $v\in V_{\rm bad}$, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
+$\bullet$ Suppose that $\sigma=1$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-1$.
 We first consider that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$.
-Choose a directed edge, say ${xy}$, in $E(D_{c}[C^{3}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3}),C^{1}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3})])$.
-By applying Claim 5.5 to the digraph collection $\{D_{i}[C^{1}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{x\})]\cup D_{i}[C^{3}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{y\})]:i\in\mathcal{C}_{3}\setminus({\rm col}(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{i_{1},i_{2},i_{3},i_{4},c\})\}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, with $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$.
-By connecting $Q_{1},{xy},Q_{3}$ in turn using Claim 5.15 (iv), we obtain a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.
+Choose a directed edge, say xy{xy}, in $E(D_{c}[C^{3}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3}),C^{1}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3})])$.
+By applying Lemma 6.2 to the digraph collection $\{D_{i}[C^{1}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{x\})]\cup D_{i}[C^{3}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{y\})]:i\in\mathcal{C}_{3}\setminus({\rm col}(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{i_{1},i_{2},i_{3},i_{4},c\})\}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, where $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$.
+By sequentially connecting $Q_{1},{xy},Q_{3}$ via Claim 7.1 (iv), we obtain a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.
 Similar to the previous discussion, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
 By a similar argument as above, and considering whether the vertex $v$ belongs to $V_{\rm bad}$ or not, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
-In what follows, it suffices to consider the case that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k<|C^{2}|-|C^{4}|-\sigma=r-s_{1}-2s_{2}-\sigma$.
-Move all vertices inside $Q_{1},\ldots,Q_{t}$, all vertices in $(V(Q_{1}^{\prime})\cup\ldots\cup V(Q_{k}^{\prime}))\cap C^{2}$ along with their in-neighbors in $C^{3}$ to $C^{4}$.
+On the other hand, we consider that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k<|C^{2}|-|C^{4}|-\sigma=r-s_{1}-2s_{2}-\sigma$.
+In this case, we move all vertices inside $Q_{1},\ldots,Q_{t}$, all vertices in $(V(Q_{1}^{\prime})\cup\ldots\cup V(Q_{k}^{\prime}))\cap C^{2}$ along with their in-neighbors in $C^{3}$ to $C^{4}$.
 Denote the resulting partition by $\tilde{C}^{1}\cup\tilde{C}^{2}\cup\tilde{C}^{3}\cup\tilde{C}^{4}$.
@@
 which implies that $|\tilde{C}^{2}|\leq|\tilde{C}^{4}|$.
-We complete our proof by the subsequent lemma (i.e., Lemma 5.16), which leads to a contradiction by showing that $\mathcal{D}$ contains a transversal directed Hamilton cycle.
+We complete our proof by using the subsequent lemma (i.e., Lemma 7.2), which leads to a contradiction by showing that $\mathcal{D}$ contains a transversal directed Hamilton cycle.
 ∎
@@
 $\leq|C^{2}||C^{3}||\mathcal{C}^{\prime}|+|Y_{2}|(1-\delta^{\frac{1}{4}})|C^{2}|\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|+|Y_{2}||C^{2}|(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|+(|C^{4}|-|Y_{2}|)|C^{2}||\mathcal{C}^{\prime}|$
-$=(|C^{4}|-\delta^{\frac{1}{2}}|Y_{2}|+|C^{3}|)|C^{2}||\mathcal{C}^{\prime}|.$ (5)
-Thus, $|Y_{2}|\leq 3\sqrt{\delta}n$. Similarly, $|Y_{1}|\leq 3\sqrt{\delta}n$. Notice that Claim 5.5 and Claim 5.15 hold by setting $V_{\rm bad}:=Y$ and $\mathcal{C}_{3}:=\mathcal{C}^{\prime}$.
+$=(|C^{4}|-\delta^{\frac{1}{2}}|Y_{2}|+|C^{3}|)|C^{2}||\mathcal{C}^{\prime}|.$
+Thus, $|Y_{2}|\leq 3\sqrt{\delta}n$. Similarly, $|Y_{1}|\leq 3\sqrt{\delta}n$. Notice that Lemma 6.2 and Claim 7.1 hold by setting $V_{\rm bad}:=Y$ and $\mathcal{C}_{3}:=\mathcal{C}^{\prime}$.
 For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp.
@@
 This implies that for each $i\in\mathcal{C}^{\prime}$, there are at most $|C^{2}|(\left\lceil\frac{n}{2}\right\rceil+\gamma_{1}n)-|C^{2}|\left\lceil\frac{n}{2}\right\rceil=|C^{2}|\gamma_{1}n$ non-edges in $D_{i}[C^{2},C^{3}\cup C^{4}]$ and at most $|C^{2}|\gamma_{2}n$ non-edges in $D_{i}[C^{1}\cup C^{4},C^{2}]$.
-Therefore, Hence
+Therefore,
 $\sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[Y_{1},C^{2}])|\geq(|Y_{1}|-\gamma_{1}n)|C^{2}||\mathcal{C}^{\prime}|\ \ \text{and}\ \ \sum_{i\in\mathcal{C}^{\prime}}|E(D_{i}[C^{2},Y_{2}])|\geq(|Y_{2}|-\gamma_{2}n)|C^{2}||\mathcal{C}^{\prime}|.$
-We proceed our proof by considering the values of $|Y_{1}|$ and $|Y_{2}|$.
+We proceed with the proof by considering the values of $|Y_{1}|$ and $|Y_{2}|$.
 Case 1. $|Y_{1}|\geq\gamma_{1}n,\ |Y_{2}|\geq\gamma_{2}n$.
-In view of Lemma 5.13, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma_{1}n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ inside $\{D_{i}[Y_{1},C^{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma_{2}n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[C^{2},Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$.
+In view of Lemma 6.4, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma_{1}n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies inside $\{D_{i}[Y_{1},C^{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma_{2}n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies inside $\{D_{i}[C^{2},Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$.
 Therefore, by using colors in $\mathcal{C}^{\prime}$,
-  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $C^{2}$,
-  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$, starting at $C^{4}$ and ending at $C^{2}$,
-  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$, starting at $C^{2}$ and ending at $C^{4}$,
-  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$, $P_{4}$ or $P_{5}$ with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $C^{4}$.
+  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by a set of disjoint rainbow directed $P_{3}$ copies with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $C^{2}$,
+  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ copies with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$, starting at $C^{4}$ and ending at $C^{2}$,
+  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ copies with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$, starting at $C^{2}$ and ending at $C^{4}$,
+  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$, $P_{4}$ or $P_{5}$ copies with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $C^{4}$.
 For each $y\in Y$, denote the rainbow directed path with center $y$ by $P_{y}=y^{1}*y*y^{2}$ with colors in $\mathcal{C}_{3}$. Let $\mathbf{P}:=\{P_{y}:y\in Y\}$.
 It is routine to verify that there exists a rainbow matching inside $D_{i}^{\pm}[C^{2}\setminus V(\mathbf{P}),C^{4}\setminus V(\mathbf{P})]\cup D_{i}[C^{1}]\cup D_{i}[C^{3}]\cup(\bigcup_{k\in[4]}D_{i}[C^{k},C^{k+1}]):i\in\mathcal{C}^{\prime\prime}\}$, say $M$.
-For all but at most one unused colors in $\mathcal{C}_{\rm bad}$, we can find a rainbow directed $P_{3}$ with center in $Y$ and endpoints in $Z$, where $\{Y,Z\}=\{C^{1},C^{3}\}$.
-If some color $c\in\mathcal{C}_{\rm bad}$ remains unused, then $D_{c}[C^{1},C^{3}]$ is nearly a complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost complete graphs.
+For all but at most one unused color in $\mathcal{C}^{\prime\prime}$, we can find a rainbow directed copy of $P_{3}$ with center in $Y$ and endpoints in $Z$, where $\{Y,Z\}=\{C^{1},C^{3}\}$.
+If some color $c\in\mathcal{C}^{\prime\prime}$ remains unused, then $D_{c}[C^{1},C^{3}]$ is nearly a complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost complete graphs.
 Then, by a similar argument to that in Theorem 6.5, $\mathcal{D}$ contains a transversal directed Hamilton cycle.
@@
 We only consider the case that $|Y_{1}|\geq\gamma_{1}n$ and $|Y_{2}|<\gamma_{2}n$; other cases can be discussed similarly.
-Notice that vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed paths with centers in $Y_{1}$ and ending at $C^{4}\setminus(Y_{1}\cup Y_{2})$.
+Notice that vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed paths with centers in $Y_{1}$ and ending in $C^{4}\setminus(Y_{1}\cup Y_{2})$.
 In particular, those rainbow directed paths can be chosen such that none of them contains edges inside $\mathcal{D}[C^{4},C^{3}]$ or $\mathcal{D}[C^{1},C^{4}]$.
-Furthermore, vertices in $Y_{2}\setminus Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{2}\setminus Y_{1}$ and starting at $C^{4}\setminus Y$, using colors in $\mathcal{C}^{\prime}$.
-In the digraph collection $\mathcal{D}[C^{4}\setminus Y]$, extend those directed rainbow paths or choose other disjoint directed rainbow directed paths into a set of disjoint maximal rainbow directed paths.
+Furthermore, vertices in $Y_{2}\setminus Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ copies with centers in $Y_{2}\setminus Y_{1}$, starting at $C^{4}\setminus Y$, and using colors in $\mathcal{C}^{\prime}$.
+In the digraph collection $\mathcal{D}[C^{4}\setminus Y]$, extend those rainbow directed paths or choose other disjoint rainbow directed paths into a set of disjoint maximal rainbow directed paths.
 Let $\mathbf{P}:=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow paths in the above, each of which has at least one endpoint in $C^{4}\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $C^{4}$.
-Let $\{e_{1},e_{2},\ldots,e_{r}\}$ be a set consisting of all disjoint rainbow edges in $\mathcal{D}[C^{4}\setminus Y,C^{3}]$, and let $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{k}\}$ be a set of disjoint maximal directed rainbow $P_{3}$ in $\mathcal{D}^{\pm}[C^{4}\setminus Y,C^{1}]$ with centers in $C^{4}$ and endpoints in $C^{1}$, each of which is chosen while avoiding previously used vertices and colors.
+Let $\{e_{1},e_{2},\ldots,e_{r}\}$ be a set consisting of all disjoint rainbow edges in $\mathcal{D}[C^{4}\setminus Y,C^{3}]$, and let $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{k}\}$ be a set of disjoint maximal rainbow directed $P_{3}$ in $\mathcal{D}^{\pm}[C^{4}\setminus Y,C^{1}]$ with centers in $C^{4}$ and endpoints in $C^{1}$, each of which is chosen while avoiding previously used vertices and colors.
 Subcase 2.1. $|C^{2}|\geq|C^{4}|-k-r-(s_{1}+\cdots+s_{t})-\sigma$.
@@
 By the maximality of $\mathbf{P}$, in $D_{c_{1}}$ (resp.
-$D_{c_{2}}$), $w$ cannot be adjacent to the start vertices of $Q_{i}[C^{4}]$ (resp.
-adjacent from the end vertices of $Q_{i}[C^{4}]$) for all $i\in[t]$.
+$D_{c_{2}}$), $w$ cannot be adjacent to the starting vertices of $Q_{i}[C^{4}]$ (resp.
+cannot be adjacent from the ending vertices of $Q_{i}[C^{4}]$) for all $i\in[t]$.
 Define
@@
 8 Concluding remarks
-In this paper, we generalize Ghouila-Houri’s theorem [19] (Theorem 1.3) to transversal version, which solves a problem proposed by Chakraborti, Kim, Lee, Seo [6].
-Thereby, the transversal version of Dirac’s theorem can be deduced, which was obtained by Joos and Kim [21].
-An anti-directed cycle is a digraph in which the underlying graph forms a cycle and no pair of consecutive edges forms a directed path.
-DeBiasio and Molla [17] proved the the anti-directed Hamilton cycle is guaranteed to appear in a digraph $D$ if $\delta^{0}(D)\geq\frac{n}{2}+1$.
-DeBiasio, Kühn, Molla, Osthus and Taylor [16] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one.
-It would be interesting to generalize the above two results to transversal version.
+In this paper, we establish a transversal analogue of Ghouila-Houri’s theorem [19] (Theorem 1.3), thereby resolving a problem proposed by Chakraborti, Kim, Lee, and Seo [6] for all sufficiently large $n$.
+As a consequence, we recover the transversal version of Dirac’s theorem, previously obtained by Joos and Kim [21].
+An anti-directed cycle is a digraph in which the underlying graph forms a cycle, and no pair of consecutive edges forms a directed path.
+DeBiasio and Molla [17] proved that the anti-directed Hamilton cycle is guaranteed to appear in a digraph $D$ if $\delta^{0}(D)\geq\frac{n}{2}+1$.
+DeBiasio, Kühn, Molla, Osthus, and Taylor [16] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one.
+It would be interesting to generalize the above two results to a transversal version.
 An oriented graph is a digraph with no cycle of length two.
-Keevash, Kühn, and Osthus [23] proved the oriented version of Dirac’s theorem, which states that every $n$-vertex digraph with minimum semidegree at least $\frac{3n-4}{8}$ contains a directed Hamilton cycle.
+Keevash, Kühn, and Osthus [23] proved the oriented version of Dirac’s theorem, which states that every $n$-vertex digraph with minimum semi-degree at least $\frac{3n-4}{8}$ contains a directed Hamilton cycle.
 It would be interesting to consider a transversal version of this theorem.
-Question 6.1.
+Question 8.1.
 Let $\mathcal{D}=\left\{D_{1},\ldots,D_{n}\right\}$ be a collection of oriented graphs with common vertex set $V$ of size $n.$ If $\delta^{0}(\mathcal{D})\geq\lceil\frac{3n-4}{8}\rceil$, does $\mathcal{D}$ contain a transversal directed Hamilton cycle?
@@
 To overcome this difficulty, one has to introduce new stable conditions and build an absorption structure using Lemma 3.4.
-It seems that new ideas are needed to prove the tight bound in Question 6.1.
+It seems that new ideas are needed to prove the tight bound in Question 8.1.
+9 Acknowledgment
+We would like to express our gratitude to the anonymous reviewers for their valuable comments that greatly improved the presentation of this paper.
 Appendix A Regularity for digraph collections
@@
 Before proceeding with the proof, we first make some preparations, including some necessary definitions and key theorems.
-A $k$-uniform hypergraph (or $k$-graph) $\mathscr{F}=(V,E)$ consists of a vertex set $V$ and an edge set $E$ which is a family of $k$-element subsets of $V$, i.e., $E\subseteq\binom{V}{k}.$ For a vertex $v\in V$, the degree of $v$ in $\mathscr{F}$, denoted by $\deg_{\mathscr{F}}(v)$, is the number of edges containing $v$.
+A $k$-uniform hypergraph (or $k$-graph) $\mathscr{F}=(V,E)$ consists of a vertex set $V$ and an edge set $E$ which is a family of $k$-element subsets of $V$, i.e., $E\subseteq\binom{V}{k}$.
+For a vertex $v\in V$, the degree of $v$ in $\mathscr{F}$, denoted by $\deg_{\mathscr{F}}(v)$, is the number of edges containing $v$.
 Given a $k$-graph $\mathscr{F}$, and $k$ disjoint sets $V_{1},V_{2},\ldots,V_{k}\subseteq V(\mathscr{F})$, define $\mathscr{F}[V_{1},V_{2},\ldots,V_{k}]$ to be the subhypergraph of $\mathscr{F}$ with edge set $\left\{\{v_{1},v_{2},\ldots,v_{k}\}\in E(\mathscr{F}):v_{i}\in V_{i}\ {\text{for all}}\ i\in[k]\right\}$.
@@
   • for each $i\in[s]$ and all $k$-tuples $\{i_{1},i_{2},\ldots,i_{k}\}\in\binom{[L]}{k}$, we have that $(U_{i_{1}},U_{i_{2}},\ldots,U_{i_{k}})$ is either empty or $(\epsilon,d)$-regular in $\mathscr{F}^{\prime}_{i}$.
-Since the proof of Theorem A.4 closely follows the proof of Theorem A.3, requiring only a repeated application of the cleaning operation on $s$ hypergraphs, we omit it here for brevity.
-Let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs on a common vertex set $[n]$. Based on $\mathcal{D}$, we construct the auxiliary $4$-graph $\mathscr{H}$ as follows:
-$V(\mathscr{H})=[n]\cup\mathcal{C}\cup S_{1}\cup S_{2},\ \text{where}\ |S_{1}|=|S_{2}|=n,$
-$E(\mathscr{H})=\left\{\{i,j,c,x\}:i,j\in[n]\ \text{with}\ ij\in E(D_{c})\ {\text{for some}}\ c\in\mathcal{C}\ \text{and}\ x\in S_{1}\ \text{if}\ i<j,\ x\in S_{2}\ \text{if}\ i>j\right\}.$
+Since the proof of Theorem A.4 closely follows that of Theorem A.3, requiring only a repeated application of the cleaning operation on $s$ hypergraphs, we omit it here for brevity.
+Let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs on a common vertex set $[n]$.
+Based on $\mathcal{D}$, we construct the auxiliary $4$-graph $\mathscr{H}$ as follows.
+The vertex set is $V(\mathscr{H})=[n]\cup\mathcal{C}\cup S_{1}\cup S_{2}$, where $S_{1}$ and $S_{2}$ are new sets with $|S_{1}|=|S_{2}|=n$, and the four sets $[n]$, $\mathcal{C}$, $S_{1}$, $S_{2}$ are pairwise disjoint.
+The edge set of $\mathscr{H}$ is defined by
+${E(\mathscr{H})=\left\{\{i,j,c,x\}:i,j\in[n],\ c\in\mathcal{C},\ ij\in E(D_{c});\ x\in S_{1}\ \text{if}\ i<j\ \text{and}\ x\in S_{2}\ \text{otherwise}\right\}}.$
 Now, we prove Lemma 2.2.
@@
 Let $n_{1}$ be derived from Theorem A.3 by applying it with the parameters $(\epsilon^{\prime},L_{1},4)$.
-Then, $n_{2}$ is obtained from Theorem A.4 using the parameters $(\epsilon^{\prime},2,L_{0},n_{1},3)$.
-Let $n_{0}:=\frac{4n_{2}}{\alpha}$.
-By increasing $n_{0}$, we may assume that $L_{1}\ll n_{1}\ll n_{2}\ll n_{0}$.
+Next, let $n_{2}$ be obtained from Theorem A.4 using the parameters $(\epsilon^{\prime},2,L_{0},n_{1},3)$.
+Define $n_{0}:=\frac{4n_{2}}{\alpha}$.
+By increasing $n_{0}$ if necessary, we may assume that $L_{1}\ll n_{1}\ll n_{2}\ll n_{0}$.
 Altogether,
-$0<\frac{1}{n_{0}}\ll\frac{1}{n_{2}}\ll\frac{1}{n_{1}}\ll\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1$
+$0<\frac{1}{n_{0}}\ll\frac{1}{n_{2}}\ll\frac{1}{n_{1}}\ll\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1.$
 Let $n\geq n_{0}$ be an integer and $\mathcal{D}$ be a digraph collection on vertex set $V:=[n]$ and with color set $\mathcal{C}$, where $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$.
@@
 $\epsilon^{\prime}|U|^{3}+Km^{\prime 2}|U|\leq\epsilon^{\prime}|U|^{3}+m^{\prime}|U|^{2}\leq\epsilon^{\prime}|U|^{3}+\frac{|U|^{3}}{L_{1}}\leq 2\epsilon^{\prime}|U|^{3}.$
-Combining $(c)$, we have
-${\rm deg}_{\mathscr{H}^{\prime}}(v)\geq{\rm deg}_{\mathscr{H}}(v)-(d+3\epsilon^{\prime})|U|^{3}$ (6)
-for all $v\in U$.
-Partition each cluster $U_{i}$, $i\in[K]$, into $\frac{1}{\alpha}$ subclusters of size at most $m_{0}$ so that all but at most four subclusters of $U_{i}$ have size exactly $m_{0}$ and the property that they lie entirely within one of $V$, $\mathcal{C}$, $S_{1}$ and $S_{2}$.
+Combining $(c)$, for each $v\in U$ we have
+${\rm deg}_{\mathscr{H}^{\prime}}(v)\geq{\rm deg}_{\mathscr{H}}(v)-(d+3\epsilon^{\prime})|U|^{3}.$
+Partition each cluster $U_{i}$, $i\in[K]$, into $\frac{1}{\alpha}$ subclusters of size at most $m_{0}$ so that all but at most $4$ subclusters of $U_{i}$ have size exactly $m_{0}$ and the property that they lie entirely within one of $V$, $\mathcal{C}$, $S_{1}$ and $S_{2}$.
 If a subcluster does not have this property, then add it to $U_{0}$.
@@
 $\left\{uv:u<v,\ x\in S_{1}\ \text{and}\ \{u,v,c,x\}\in\mathscr{H}^{\prime}\right\}\cup\left\{vu:u<v,\ y\in S_{2}\ \text{and}\ \{u,v,c,y\}\in\mathscr{H}^{\prime}\right\}.$
-We claim that the properties of the Lemma 2.2 (i)-(iv) are satisfied.
+We claim that the properties of Lemma 2.2 (i)-(iv) are satisfied.
 Property (i). Note that $m^{\prime}K\leq|U|$. Since $\epsilon^{\prime}\ll\epsilon$, one has
@@
 Also,
-$L+M\leq\frac{2K}{\alpha}\leq\frac{2n_{1}}{\alpha}\leq\frac{n_{0}}{2},$ (8)
+$L+M\leq\frac{2K}{\alpha}\leq\frac{2n_{1}}{\alpha}\leq\frac{n_{0}}{2},$
 proving the required upper bound for both $L$ and $M$. Furthermore, $m_{0}(L+M)\leq n+|\mathcal{C}|$ and
@@
 for all $v\in V$, and similarly for $\sum_{c\in\mathcal{C}}d^{-}_{D_{c}^{\prime}}(v)$.
-Additionally, observe that $e(D_{c})=\frac{{\rm deg}_{\mathscr{H}}(c)}{n}$ for all $c\in\mathcal{C}$.
-It follows that
-$e(D_{c}^{\prime})\geq\frac{{\rm deg}_{\mathscr{H}^{\prime}}(c)}{n}\geq\frac{{\rm deg}_{\mathscr{H}}(c)}{n}-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}=e(D_{c})-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}$ (12)
-for all $c\in\mathcal{C}$.
+Additionally, observe that $e(D_{c})=\frac{{\rm deg}_{\mathscr{H}}(c)}{n}$ for each $c\in\mathcal{C}$.
+It follows that for each $c\in\mathcal{C}$,
+$e(D_{c}^{\prime})\geq\frac{{\rm deg}_{\mathscr{H}^{\prime}}(c)}{n}\geq\frac{{\rm deg}_{\mathscr{H}}(c)}{n}-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}=e(D_{c})-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}.$ (9)
 Property (iv). It follows by construction.
@@
 $E(\mathscr{H}_{1})=\left\{\{u,v,c\}:uv\in D^{\prime}_{c}[V_{i},V_{j}],u,v\in V\setminus V_{0},c\in\mathcal{C}\setminus\mathcal{C}_{0},1\leq i<j\leq L\right\}$, $E(\mathscr{H}_{2})=\left\{\{u,v,c\}:uv\in D^{\prime}_{c}[V_{j},V_{i}],u,v\in V\setminus V_{0},c\in\mathcal{C}\setminus\mathcal{C}_{0},1\leq i<j\leq L\right\}$.
-Notice that we can view partition $P=(V_{1},\ldots,V_{L},\mathcal{C}_{1},\ldots,\mathcal{C}_{M})$ as comprising two different partitions $P_{1}$, $P_{2}$ of hypergraphs $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ (both partitions divide $V$ and $\mathcal{C}$ in the same way, but the sets of edges among the partition sets are different).
-By increasing $n_{2}$, we may assume that $n_{2}\gg N(\epsilon^{\prime},2,L_{0},L+M,3)$ (which is defined in Theorem A.4).
-Consequently, $n_{0}\gg N(\epsilon^{\prime},2,L_{0},L+M,3)$ as well.
-Applying Theorem A.4, we can find a refinement partition of $P$, say $U_{0}^{\prime},U_{1}^{\prime},\ldots,U_{K^{\prime}}^{\prime}$, and spanning subhypergraphs $\mathscr{H}_{1}^{\prime}\subseteq\mathscr{H}_{1}$ and $\mathscr{H}_{2}^{\prime}\subseteq\mathscr{H}_{2}$ such that the following hold:
+[…]
   • (a’)
@@
 B.1 Proof of Lemma 2.7
+In this subsection we prove Lemma 2.7, which provides a characteristic partition of each balanced bipartite graph whose minimum degree is just below $\frac{n}{2}$, allowing a tiny fraction of vertices that may have even smaller degrees.
 Proof of Lemma 2.7.
-We first add vertices with small degree to $C_{1}\cup C_{2}$.
+We first add vertices with degree at most $\left(\frac{1}{2}-\epsilon^{5}\right)n$ to $C_{1}\cup C_{2}$.
 Let $\mu=\epsilon^{5}$.
@@
 Let $C_{1}=V_{1}\setminus(A_{1}\cup B_{1})$ and $C_{2}=V_{2}\setminus(A_{2}\cup B_{2})$.
-Hence, $(A_{1},B_{1},C_{1},A_{2},B_{2},C_{2})$ is the desired partition.
+Hence, $(A_{1},B_{1},C_{1},A_{2},B_{2},C_{2})$ is the desired partition of $G$.
 ∎
 B.2 Proof of Theorem 2.10
-In this subsection, we prove Theorem 2.10. Before proceeding with the proof, we first present some key lemmas that will be used in the proof.
+In this subsection we prove Theorem 2.10, giving a sufficient condition for the existence of a transversal perfect matching in bipartite graph collections.
+Before proceeding with the proof, we first present some key lemmas that will be used in the proof.
 Assume that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$.
@@
 Lemma B.1.
-Suppose that $0<\frac{1}{n}\ll\mu^{\prime}\ll\alpha\ll\gamma,\epsilon,\ll\delta\ll 1$.
+Suppose that $0<\frac{1}{n}\ll\mu^{\prime}\ll\alpha\ll\gamma,{\epsilon}\ll\delta\ll 1$.
 Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$.
@@
 For the sake of contradiction, suppose that $\mathcal{G}$ is not $\mu^{\prime}$-nice.
-Then there is $A\subseteq V_{1}$ of size $\lfloor\frac{n}{2}\rfloor$ and $B\subseteq V_{2}$ of size $\lfloor\frac{n}{2}\rfloor$ such that $e_{\mathcal{G}}(A,B)\leq\mu^{\prime}n^{3}$.
+Then there are sets $A\subseteq V_{1}$ and $B\subseteq V_{2}$, each of size $\lfloor\frac{n}{2}\rfloor$, such that $e_{\mathcal{G}}(A,B)\leq\mu^{\prime}n^{3}$.
 Therefore, for all but at most $\sqrt{\mu^{\prime}}n$ colors $i\in[n]$, we have $e_{G_{i}}(A,B)<\sqrt{\mu^{\prime}}n^{2}$, implying that such $G_{i}$ is not $\sqrt{\mu^{\prime}}$-nice.
-This further implies that $\mathcal{G}$ is not $(1-\sqrt{\mu^{\prime}},\sqrt{\mu^{\prime}})$-strong stable and hence not $(\gamma,\alpha)$-strong stable, since $\mu^{\prime}\ll\gamma,\alpha$.
+This further implies that $\mathcal{G}$ is not $(1-\sqrt{\mu^{\prime}},\sqrt{\mu^{\prime}})$-strongly stable and hence not $(\gamma,\alpha)$-strongly stable, since $\mu^{\prime}\ll\gamma,\alpha$.
 Without loss of generality, assume that for each $i\in[(1-\sqrt{\mu^{\prime}})n]$, $G_{i}$ is not $\sqrt{\mu^{\prime}}$-nice.
@@
 It follows that, after possible relabeling $A_{1}^{i},B_{1}^{i},A_{2}^{i},B_{2}^{i}$, for all but at most $\delta^{2}n$ colors $i\in[(1-\sqrt{\mu^{\prime}})n]$ we have
-$|A\Delta A_{1}^{i}|\leq\frac{\delta n}{2}\,\,\,\text{and}\,\,\,|B\Delta B_{2}^{i}|\leq\frac{\delta n}{2}.$
+$|A\triangle A_{1}^{i}|\leq\frac{\delta n}{2}\ \ \text{and}\ \ |B\triangle B_{2}^{i}|\leq\frac{\delta n}{2}.$
 Otherwise, $e_{\mathcal{G}}(A,B)\geq\frac{\delta^{3}n^{3}}{8}>\mu^{\prime}n^{3}$ by $\mu^{\prime}\ll\epsilon\ll\delta$, a contradiction.
-Delete such colors from $[(1-\sqrt{\mu^{\prime}})n]$ and let $\mathcal{C}$ be the remaining color set.
+Delete such colors from $(1-\sqrt{\mu^{\prime}})n]$ and let $\mathcal{C}$ be the remaining color set.
 Thus, for every $i,j\in\mathcal{C}$ the graphs $G_{i}$ and $G_{j}$ are not $\delta$-crossing.
@@
 Thus, it suffices to show that for all $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}]\cup\mathcal{M}[N_{2},N_{2}^{+}])$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$.
-If so,
+If this holds, then
 $e_{\mathcal{G}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}(n-2\mu n)+2\mu n^{3}<6\mu n^{3},$
-a contradiction to $\mathcal{G}$ being $6\mu$-nice.
+contradicting the assumption that $\mathcal{G}$ is $6\mu$-nice.
 Next, we consider the case $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}])$; the proof for the case $c\in{\rm col}(\mathcal{M}[N_{2},N_{2}^{+}])$ is similar and is omitted.
@@
 Thus,
-$|N_{3}^{+}\Delta N_{1}^{+}|=|N_{3}^{+}\cup N_{1}^{+}|-|N_{3}^{+}\cap N_{1}^{+}|\leq(n-|N_{2}|)-(|N_{1}|+|N_{3}|-(n-|N_{2}|))\leq 4\mu n.$ (13)
-On the other hand, for the maximality of matching $\mathcal{M}^{\prime}$, we have $e_{G_{c}}(N_{3}^{+},N_{2}^{+})=0$.
+$|N_{3}^{+}\triangle N_{1}^{+}|=|N_{3}^{+}\cup N_{1}^{+}|-|N_{3}^{+}\cap N_{1}^{+}|\leq(n-|N_{2}|)-(|N_{1}|+|N_{3}|-(n-|N_{2}|))\leq 4\mu n.$ (10)
+On the other hand, by the maximality of $\mathcal{M}^{\prime}$, we have $e_{G_{c}}(N_{3}^{+},N_{2}^{+})=0$.
 Thus, together with (10), we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$, as desired.
@@
 $(A_{1}^{i},B_{1}^{i},C_{1}^{i},A_{2}^{i},B_{2}^{i},C_{2}^{i})$ of $G_{i}$.
-We define a vertex $v\in V$ is $i$-good if either $G_{i}$ is $\epsilon$-extremal and $v\in V\setminus(C^{1}_{i}\cup C^{2}_{i})$ or $G_{i}$ is not $\epsilon$-extremal and $d_{G_{i}}(v)\geq(\frac{1}{2}-\epsilon^{3})n$.
+A vertex $v\in V$ is called $i$-good if either $G_{i}$ is $\epsilon$-extremal and $v\in V\setminus(C^{1}_{i}\cup C^{2}_{i})$ or $G_{i}$ is not $\epsilon$-extremal and $d_{G_{i}}(v)\geq(\frac{1}{2}-\epsilon^{3})n$.
 Note that for each $i\in[n]$ and $j\in\{1,2\}$, there are at most $2\epsilon n$ vertices in $V_{j}$ that are not $i$-good.
@@
 Given $\delta,\delta^{\prime},\gamma,\gamma^{\prime}\geq 0$, a rainbow matching $\mathcal{M}$ is an absorbing matching with parameters $(\delta,\delta^{\prime},\gamma,\gamma^{\prime})$ if $|E(\mathcal{M})|\leq\gamma n$ and there exists a color set $\mathcal{C}$ of size at least $\delta n$ such that given any color $c\in\mathcal{C}$ and any $c$-good vertex $v\in V_{2}$, for all but at most $\delta^{\prime}n$ vertices $u\in V_{1}$, there are at least $\gamma^{\prime}n$ $c$-absorbing edges of $(u,v)$ within $\mathcal{M}$.
-$u$$v$$w_{1}$$w_{2}$$c$
+[…]
 Figure 5: $w_{1}w_{2}$ is a $c$-absorbing edge of $(u,v)$.
@@
 Proof.
-Without loss of generality, we assume that $G_{1},\ldots,G_{\gamma n}$ are $\alpha$-nice. For $i\in[\lambda n]$, define
+Without loss of generality, assume that $G_{1},\ldots,G_{\gamma n}$ are $\alpha$-nice. For each $i\in[\lambda n]$, define
 $\mathcal{F}_{i}:=\left\{(u,v)\in V_{1}\times V_{2}:i\in L(uv)\right\}.$
 The degree condition of $G_{i}$ implies that $|\mathcal{F}_{i}|=e(G_{i})\geq(n-dn)\left(\frac{1}{2}-\mu\right)n>\frac{1}{4}n^{2}.$ Fix a color $c\in[n]$ and two vertices $u\in V_{1},v\in V_{2}$ such that $v$ is $c$-good, let
-$Z_{i}(c,uv):=\{(w,w^{\prime})\in\mathcal{F}_{i}:ww^{\prime}\,\,\text{is a}\,\,c\text{-absorbing edge of}\,\,(u,v)\}\ \text{and}\ Z(c,uv):=\bigcup_{i\in[\lambda n]}Z_{i}(c,uv).$
+$Z_{i}(c,uv):=\big\{(w,w^{\prime})\in\mathcal{F}_{i}:ww^{\prime}\,\,\text{is a}\,\,c\text{-absorbing edge of}\,\,(u,v)\big\}\ \ \text{and}\ \ Z(c,uv):=\bigcup_{i\in[\lambda n]}Z_{i}(c,uv).$
 Since $G_{i}$ is $\alpha$-nice, $e(G_{i}[N_{G_{i}}(u),N_{G_{c}}(v)])\geq\frac{\alpha n^{2}}{2}$ for all but at most $dn$ indices $i\in[\gamma n]$.
@@
 Applying Lemma 3.4 with
-[…]
+$\mathbf{H}:=\{\mathcal{F}_{i}:i\in[\lambda n]\},\,\,\,\mathbf{Z}:=\{Z(c,uv):c\in[n],(u,v)\in V_{1}\times V_{2}\,\,\text{and}\,\,v\,\,\text{is}\,\,c\textrm{-good}\},$
 and parameters $t:=\lambda n$, $\epsilon:=\frac{\alpha}{2}$, to obtain a rainbow matching $\mathcal{M}$ inside $\mathbf{H}$ of size at least $(1-\frac{\alpha^{2}}{16})\lambda n$ such that $|E(Z(c,uv))\cap E(\mathcal{M})|\geq\frac{\alpha^{2}\lambda n}{16}$ for all $Z(c,uv)\in\mathbf{Z}$.
@@
 Lemma B.5.
-Let $10<\frac{1}{n}\ll d\ll\lambda,\mu\ll\epsilon\ll\delta<$ 1.
+Let $0<\frac{1}{n}\ll d\ll\lambda,\mu\ll\epsilon\ll\delta<1$.
 Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$ such that the following conditions hold:
@@
 For any vertex $v\in V$, let
-[…]
+$\mathcal{C}_{v}:=\{i\in[n]:v\ \text{is}\ i\text{-good}\}.$
 Since $e(C_{\mathcal{G}}^{\epsilon,\delta})\geq\delta n^{2}$, there exists a subgraph $H$ of $C_{\mathcal{G}}^{\epsilon,\delta}$ such that $|V(H)|\geq\delta n$ and $\delta(H)\geq\delta n$.
@@
     for every $i\in V(H)$ and $u\in\overline{T_{i}}$, we have $|\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})|\geq\frac{\delta\kappa n}{4}$.
-Fix such a $\mathcal{U}$ and let $\overline{\mathcal{U}}:=V(H)\setminus\mathcal{U}$. By relabeling colors, one may assume $\mathcal{U}=:[t]$. Clearly,
+Fix such a set $\mathcal{U}$ and let $\overline{\mathcal{U}}:=V(H)\setminus\mathcal{U}$. By relabeling colors, one may assume $\mathcal{U}=:[t]$. Clearly,
 $|\overline{\mathcal{U}}|=|V(H)|-|\mathcal{U}|\geq\frac{\delta n}{2},\ \ \text{and}\ \ \delta(H[\overline{\mathcal{U}}])\geq\delta(H)-|\mathcal{U}|\geq\frac{\delta n}{2}.$
 Define the multi-graph collection
-[…]
+$\mathbf{Z}:=\left\{Z(i,uv):=\bigcup_{j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})}Z_{j}(i,uv):i\in\overline{\mathcal{U}},u\in\overline{T_{i}}\ \text{and}\ v\ \text{is}\ i\text{-good}\right\}.$
 Given $i$ and $u$, the number of choices for $j$ is at least $\frac{\delta\kappa n}{4}$ by (ii).
@@
 $\bullet$ For vertices in $V_{2}$, there are at most $2\epsilon n+dn$ vertices $w$ that are either not $c_{i}^{2}$-good or $d_{G_{c}}(w)<\left(\frac{1}{2}-\mu\right)n$.
-Thus, we have at least $n-2\epsilon n-dn-2\geq\frac{n}{2}$ choices for $w_{i}^{2^{\prime}}$ such that $w_{i}^{2^{\prime}}$ is $c_{i}^{2}$-good and $d_{G_{c}}(w)\geq\left(\frac{1}{2}-\mu\right)n$.
+Thus, we have at least $n-2\epsilon n-dn-2\geq\frac{n}{2}$ choices for $w_{i}^{2^{\prime}}$ such that $w_{i}^{2^{\prime}}$ is $c_{i}^{2}$-good and $d_{G_{c}}(w_{i}^{2^{\prime}})\geq\left(\frac{1}{2}-\mu\right)n$.
 $\bullet$ Fix $w_{i}^{1^{\prime}}$ and $w_{i}^{2^{\prime}}$.
@@
 Let $\{c_{i}^{1},c_{i}^{2}:i\in[2\beta n]\}\subseteq\mathcal{C}$ be a set of distinct colors. Define
-$\mathcal{F}_{i}:=\{(w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})\in V_{1}\times V_{2}\times V_{1}\times V_{2}:(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})\,\,\text{and}\,\,(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})\,\,\,\text{are absorbable}\}.$
+$\mathcal{F}_{i}:=\{(w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})\in V_{1}\times V_{2}\times V_{1}\times V_{2}:(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})\penalty 10000\ \text{and}\penalty 10000\ (c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})\penalty 10000\ \text{are absorbable}\}.$
 Since for each color $c\in\mathcal{C}$, there are at most $2\epsilon n$ vertices in each part that are not $c$-good.
@@
 Thus, all vertices in $U\setminus V(\mathcal{M}_{1})$ and all colors in $\mathcal{C}_{\rm rem}\setminus{\rm col}(\mathcal{M}_{1})$ are contained in a rainbow matching of $\mathcal{G}$.
-At the end of this process, there remains a set $I\subseteq[r]$ such that the $q_{i}$ with $i\in[I]$ are precisely the 6-tuples that were not chosen.
+At the end of this process, there remains a set $I\subseteq[r]$ such that the $q_{i}$ with $i\in I$ are precisely the 6-tuples that were not chosen.
 Recall that $(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})$ and $(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})$ are absorbable for each $i\in I$.